A Complete Bibliography of Publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences

halieutiques et aquatiques (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 August 2024 Version 1.05

Title word cross-reference

10^{3} [1033]. 10^{4} [1033]. > [1054]. ^{+} [2258, 1919, 1992, 2570, 389, 2167]. ^{14} [151, 97, 2378, 1919, 707]. ^{15} [1833, 2378]. ^{2+} [2167]. ^{210} [1349, 1401]. ^{228} [2442]. ^{232} [1865]. ^{238} [1865]. ^{60} [1610]. _{1} [321]. _{2} [1019, 1946, 1234, 2570]. _{3} [2387]. _{3}^{−} [1234]. _{6} [1847]. a [313, 344, 1075, 631, 278, 1229, 782, 1144, 2202, 1251]. α [2372, 1269]. β [1449, 2317]. χ^{2} [147, 645]. : [1547]. δ [1136]. μ [1054]. n [1082, 399]. p, p′ [299]. r [438]. Σ [1116]. T^{2} [1278]. × [364, 346].

-Aminolevulinic [1136]. -ATPase [2258, 1919, 1992]. -butyltin [1082]. -DDT [299]. -hexyl [399]. -Methyltestosterone [2372, 1269].

0-Age [1611].

120-kHz [87]. 1254 [1521, 1894]. 13/Carbon [98]. 17 [1449, 2317]. 1970s [1306]. 1973 [498].

2 [2125]. 2-Methylisoborneol [1926, 185]. 20E [752]. 223 [974]. 227 [65].

303 [2352].

4 [2125]. 4-D [2125]. 450 [865].

50-Year [2594].

70-kHz [2066, 2605]. 73 [628, 1831]. 77^{o} [739].

Îles [2047, 2196, 870, 319, 1767, 1862]. Îles-de-la-Madeleine [870].

A. [1485, 2550, 362]. Abalone [1133]. Aberdeenshire [1435]. Aberrant [1682]. Ability [551, 1347, 1768, 692, 2265, 1012, 369, 629]. Abiogenic [2264]. Abiotic [1451, 1875, 1150, 2316, 406]. Abnormalities [1482]. Abondance [2142, 2550]. Aboveground [1905]. Abra [1810]. Abramis [1415]. Absence [422, 2281, 755]. Absorbance [896]. Absorption [2596, 1379, 2117, 421]. Abundance [623, 2019, 1882, 1134, 1854, 1770, 2262, 2510, 1764, 970, 2211, 1076, 127, 1382, 1850, 2194, 2192, 1672, 1958, 228, 1447, 728, 819, 2102, 1954, 2452, 236, 1393, 2260, 345, 2320, 1867, 668, 2464, 407, 245, 420, 427, 2209, 571, 1555, 1565, 671, 1848, 1081, 2049, 941, 1653, 1914, 2591, 1057, 1503, 2483, 627, 1333, 1237, 1407, 565, 1434, 1869, 2180, 466, 1368, 554, 1454, 1370, 1403, 1750]. Abundances [2614, 1238, 2382]. acanthias [1471]. Acartia [1581]. Accelerated [2029]. Accelerates [2375]. Access [1978, 1498, 936]. Accident [851]. Acclimate [2254, 133]. Acclimated [715]. Acclimation [405, 1829, 1830, 1708, 1532, 2279, 2280]. Acclimation-Induced [1532]. Acclimitization [2144]. Accounting [1122]. Accumulate [1683]. Accumulated [1013]. Accumulation [1257, 1601, 32, 1249, 2140, 1349, 278, 97, 110, 822, 2219, 299, 981, 1354, 2297, 2573, 1082, 2382, 1421, 2042, 1196, 1588, 1339, 545, 1211, 478]. Accuracy [2124, 2177]. Acephate [186]. Acetone [1203]. Acetylcholinesterase [186]. Acetylene [63]. Acid [1186, 2253, 1030, 1337, 2507, 2596, 2394, 1294, 2392, 2234, 1168, 1573, 197, 22, 1863, 1293, 1615, 2283, 2284, 2121, 1494, 1465, 2332, 1510, 1695, 1136, 1018, 1414, 2233, 564, 2285, 2199, 1628, 2278, 2279, 2280]. Acid-Exposed [1695]. Acid-Insoluble [1186]. Acid-Rain [1168]. Acid-Stressed [2234, 2233]. Acid/Aluminum [2278, 2279, 2280]. Acid/Base [2392]. Acidic [2386, 1460, 2396, 2392, 1883, 1164, 2008, 1580, 1283, 2601, 1942, 755, 1735, 2391, 1751, 275]. Acidification [1456, 1601, 2276, 1020, 1589, 1332, 1497, 1557, 2468, 2519, 1726, 973, 693, 2010, 1223, 2018, 1580, 948, 1731, 49, 974, 2455, 2385, 755, 1515, 1989, 48, 51, 1539, 818]. Acidified [2397, 759, 1951, 1119, 1150, 2410, 282, 1105, 1754, 1469, 1908]. Acidifying [2121]. Acidity [2389, 2100, 1221, 1994, 1434, 1936]. Acids [2473, 9, 307, 995, 1425, 25]. Acipenser [1411, 2204]. Acoustic [2037, 1723, 2052, 1886, 2165, 1673, 2605, 435, 627, 1438, 2003]. Acoustical [1966]. Acoustically [283]. Acquired [1576]. Across [2348]. Act [498]. Action [2021]. Activable [303, 456]. Activation [186]. Active [900]. Actively [2585, 577]. Activities [1037, 2244, 865, 473]. Activity [646, 1953, 2585, 746, 1894, 1542, 187, 76, 1620, 1700, 1297, 1959, 1527, 182, 1992, 2368, 2075, 131, 1695, 1136, 1006, 2449, 389]. Actual [2489]. aculeatus [2400, 2142, 1807, 2043, 1162]. Acute [1337, 1340, 573, 2125, 999, 1849, 715, 1897, 340, 1694, 2167]. Acyl [2473]. Adams [2004]. Adaptability [593]. Adaptation [861, 851, 1102, 1622, 1493, 265]. Adaptations [854]. Adapted [2606]. Adaptive [1851, 811, 2481, 1807, 828, 328, 329, 378, 2450, 2119]. Added [52]. Addition [150, 869, 60, 2391, 720]. Additions [1964]. Address [1757]. Adenosine [826, 389, 550]. Adenovirus [1411]. Adenovirus-Like [1411]. Adenylate [599]. Adhuc [540]. Adipose [984]. Adirondack [2010, 2397, 856, 2395, 2388, 504]. Adjacent [1249, 347]. adjusting [265]. adjustment [1327]. Adjuvant [1406]. Adjuvant-Dependent [1406]. Administered [762, 2306]. Administration [382]. administré [2306]. Adsorption [869, 1691, 2087, 312]. Adult [232, 219, 1882, 2108, 2276, 823, 1362, 220, 1607, 2230, 1742, 1954, 23, 221, 1602, 971, 2580, 2283, 2284, 675, 846, 437, 2030, 570, 1503, 2226, 2584, 392, 2278, 2279, 2280, 1611]. Adults [2200, 2492, 365, 644]. Advance [1816]. Advanced [389]. Advantage [2066, 700]. aeglefinus [1820, 2428, 1474, 1821, 2540, 1594]. Aequorea [2532]. Aeration [910, 1036]. Aerococcus [1862]. Aeromonas [464, 551, 726, 1406, 1182, 2463, 1576, 268]. aestivalis [745, 1485, 2550, 362]. Affected [1865]. Affecting [1797, 2110, 547, 2147, 2338, 1380, 1958, 782, 1777, 1148, 84, 176, 2562, 1123]. Affects [173, 427]. Affinities [505, 1345]. African [2343]. After [1009, 1690, 938, 2285, 817, 2333, 2421, 2389, 1088, 1087, 1845, 221, 997, 1838, 227, 1523, 2393]. Ag [432]. Against [1406, 551]. Age [177, 1851, 418, 562, 1471, 1538, 1788, 918, 2428, 1367, 1973, 2055, 681, 327, 1004, 1309, 2458, 317, 34, 1350, 2220, 680, 893, 2206, 1721, 1849, 915, 2574, 1467, 2442, 1513, 828, 1378, 773, 345, 794, 1101, 1925, 1924, 68, 2464, 718, 1446, 1536, 1823, 594, 920, 977, 1364, 225, 2588, 2623, 499, 1583, 921, 2354, 831, 612, 769, 1151, 824, 1064, 2528, 1590, 1514, 1891, 774, 1611, 2472]. Age- [2428]. Age-0 [2574, 2464]. Age-at-First-Capture [1513]. Age-Dependent [920]. Age-Structured [1788, 1004, 34, 1467, 1513, 828, 1378, 824, 1514, 1891, 893]. Ageing [1699]. Agent [2095, 1841]. Agents [898, 250]. Ages [1243]. Agglutinin [1406]. Agglutinins [1576]. Aggregating [980, 2355]. Aggregation [1999].

Aggregations [2511]. Aggregative [1403]. Aggression [1842]. Aggressive [1459]. Aging [567]. Aglantha [200]. Agnatha [215]. Agonistic [1569, 2119]. aguabonita [897]. Aid [2358, 699]. Aided [2201]. Aiming [1856]. Air [1998]. Air-Breathing [1998]. Al [2506]. ALA [1136]. ALA-D [1136]. alalunga [2461, 1440]. Alaska [1825, 1335, 580, 1051, 2194, 142, 159, 1560, 2431, 501, 444, 2471, 1153, 582, 2195, 270, 396, 889, 2564, 1446, 1035, 2528, 2612]. Alaskan [1349, 2201, 1910, 1311, 1814, 1692]. alaskensis [476]. alba [1810]. albacares [1722]. Albacore [2461, 1440]. Alberta [187, 1808, 1117, 2363, 2478, 2479, 1975, 2160]. albula [1702]. Aldosterone [1627]. Alevin [1282]. Alevins [133, 1987, 733, 544, 1283, 935, 390]. Alewife [1607, 1447, 728, 2550, 167, 362, 1667]. Alewives [23, 1485]. Alga [1835, 1082, 1957, 1469, 1211, 1965]. Algae [895, 592, 32, 2510, 2123, 1164, 63, 1342, 578, 1890, 315, 1127, 1, 431, 596, 1518, 1975, 2158, 2578, 1176, 311, 589, 743]. Algal [150, 55, 1635, 1542, 1730, 2319, 857, 1214, 2331, 429, 1643, 1646, 2527, 2390, 1551, 474, 1237, 2016, 1434, 2587, 1809, 409, 917]. Algorithms [546]. alimentaire [1621, 2306, 265]. Alimentary [1291]. alimentation [1621]. Alkaline [1542, 1297, 1381, 2075, 474]. Alkalinity [1340, 2594, 2336, 1688, 2562, 1106, 2419, 2460]. Allele [2058, 1637]. Allelic [489, 513]. Allocapnia [1727]. Allocation [1366, 395, 394, 1512, 1355, 914]. Allochthonous [1197]. Allochthony [309]. Allometric [624, 2128, 2600, 2418]. Allometry [2600]. Allopatric [719, 2246, 512]. Allotting [804]. Allowing [1522]. Allozyme [2058, 2608]. Along [584, 780, 1547, 2605, 959, 942, 2614, 292]. Alosa [2535, 952, 1159, 1452, 1451, 1558, 395, 394, 1607, 1447, 728, 23, 745, 1485, 2550, 167, 362, 1667, 1579]. alosoides [2458]. alpinus [2474, 1703, 2146, 2451, 661, 2439, 320, 2448, 1863, 2183, 712, 2273, 291, 2271, 912, 1623]. Alsek [281]. Alta [1729]. Alteration [468, 938, 1018]. Alterations [2183, 642]. Altered [2125, 2014, 79]. Alternative [209, 2004, 2198, 1848, 1593, 1531, 1927, 377, 2167]. Alters [1508, 573]. Aluminum [2277, 1661, 1439, 2013, 1987, 1464, 1105, 1529, 2281, 2283, 2284, 1487, 1494, 1532, 148, 2168, 1751, 2285, 1956, 2282, 2278, 2279, 2280]. Aluminum-Contaminated [1751]. alutus [2101]. Ambient [156, 1670, 2607, 1679, 1235]. Ambystoma [2013, 2174]. Amelioration [1010]. America [2494, 744, 539, 2616, 1931, 475, 1885, 1956, 1637, 1345]. American [1290, 1271, 1779, 1680, 2370, 620, 233, 364, 2535, 1292, 946, 978, 1276, 1597, 2582, 2456, 1990, 952, 1159, 1452, 1451, 1558, 1269, 1179, 1784, 126, 1768, 1758, 2244, 1791, 395, 394, 672, 780, 1662, 101, 1700, 1580, 1660, 1959, 617, 1579, 2561, 658, 2189, 2188, 846, 648, 1785, 966, 1289, 1488, 1291, 1782, 466]. americana [1312, 369, 629]. americanus [1290, 1271, 1779, 1681, 1680, 2370, 620, 1781, 1955, 1292, 946, 978, 1276, 1597, 1771, 1770, 2582, 2055, 1990, 2013, 1005, 324, 1179, 1294, 1784, 124, 1768, 1758, 1113, 1791, 865, 1174, 2069, 672, 780, 2141, 1662, 1767, 2047, 2196, 572, 591, 2434, 2558, 1293, 1959, 617, 1862, 2561, 870, 2189, 2188, 846, 2322, 1772, 648, 1778, 1785, 1425, 1967, 1289, 1488, 1782, 430, 466]. Amictic [10]. Amino [9, 307, 1573, 995, 2332]. Aminocarb [2125]. Aminolevulinic [1136]. Aminonaphthylamidase [13]. Ammocoete [214]. Ammocoetes [1191, 723, 226, 208, 1590]. Ammodytes [2595, 397, 812, 2152]. Ammonia [106, 614, 1662, 1111, 788, 175, 428, 429, 181, 2606, 419, 436, 2449]. Ammonia-Caused [788]. Ammonium [2056, 564]. Amnicola [2460]. Among [2175, 2413, 2414, 2547, 1542, 508, 1341, 1266, 787, 2043, 350, 1512, 1656, 268, 2243, 1495, 1496, 1279, 2104, 1254, 2500, 1141, 514, 1910, 2337, 2373, 566, 291, 1807, 1909, 2503, 2379, 2382, 2555, 2203, 1034]. Among-Population [2413, 2414, 2547]. Amount [2462]. Amphibian [1439]. Amphibians [231]. Amphipod [2568, 438, 1999, 1875, 2166, 797, 1539, 2545, 2167]. Amphipoda [426, 2311, 1951, 1950, 1, 1463, 2293, 2567]. Ampoule [1846]. anadromes [714]. Anadromous [2182, 233, 2407, 2474, 1866, 1584, 1887, 2097, 2146, 2371, 847, 320, 1316, 654, 1416, 1288, 1503, 19, 498, 2404, 714]. Anaerobic [156, 2406]. Analog [2375, 1977]. Analogues [762, 2054]. Analyses [1898, 2422, 2425, 1122, 409]. Analysing [1583]. Analysis [923, 1668, 2415, 1520, 1137, 316, 1654, 603, 2161, 2065, 313, 2575, 95, 94, 2055, 2559, 1366, 366, 690, 2433, 1350, 2563, 1741, 1726, 2358, 376, 893, 1969, 1857, 876, 1847, 2138, 1188, 2102, 2434, 1720, 1493, 1112, 1988, 1070, 721, 1178, 1500, 57, 1961, 1105, 2614, 62, 2526, 1800, 2025, 1603, 1926, 1039, 925, 1455, 1921, 2209, 1578, 1429, 2505, 2579, 2154, 184, 1932, 424, 924, 1568, 1802, 1392, 1916, 12, 172, 1124, 1303, 2377, 602, 1938, 2293, 824, 1869, 2350, 595, 2557, 1252, 1353, 369, 629, 830, 1970, 812, 1664, 1922, 2613, 2089]. Analytab [752]. Analytical [769]. Analyzing [174, 680, 1858, 1522]. Anaplopoma [2108]. anatomy [881]. Ancestry [987]. Anchovy [2093]. and/or [1894]. Anesthesia [997]. Angiosperm [2215]. Angiosperms [2032]. Angled [1538]. Angler [1531, 1816]. Angling [2617, 1006]. Anguilla [2456, 1269, 1307, 1700, 1415, 658, 1738]. anguillarum [1184, 430]. anguille [1307, 658]. Animal [619, 2412, 145, 1245]. Animals [737]. Anion [2432, 340]. Anion-Exchange [2432]. Anisakidae [1666]. Anisakis [1666]. Anisogammaridae [1463]. Ankistrodesmus [1082, 1211]. Anne [187, 2479]. Annelida [36, 157]. Anniversary [1041]. Announcements [302, 355, 385, 403, 732, 810, 991]. Announcements/Avis [302, 355, 385, 403, 732, 810, 991]. Annual [527, 66, 1174, 986, 2530, 742, 815, 1581, 68, 321, 966, 2427, 1991, 2619]. annularis [1736]. Annuli [918]. Annulus [1471]. Anoplopoma [562, 1000, 1617, 659]. Anoxic [1049, 1104, 1570, 2145]. Antarctic [2624, 875]. Antennal [1482]. Anthracene [797]. Anthropogenic [1726, 829, 1221, 2242]. Anthropogenically [1951, 1120]. Antibacterial [1066]. Antibody [92, 1844, 1967]. Antifreeze [1113, 2133, 1630]. Antigen [1841]. Antigens [551]. Antimicrobial [2495]. Antimicrobic [421]. Any [7]. Apatite [80]. Aperture [409]. Apex [1722]. API [752]. API-20E [752]. Apicomplexa [1983]. Appalachian [2318, 78, 332]. Apparatus [1976, 663]. Apparent [1338, 1505, 2324]. Appetite [2232]. Application [11, 592, 1723, 998, 621, 1607, 1831, 1961, 2310, 345, 500, 279, 2505, 1510, 2349, 844, 1769, 2028, 595, 2130, 1922]. Applications [495, 1212, 536, 1265, 2475, 1515]. Applied [2585, 2559, 730, 1921, 2154, 38, 1167]. Appraisal [2000]. Approach [2050, 1839, 1558, 2526, 345, 805, 2604, 135, 1653, 172, 1003, 1124, 1006, 381, 2493, 1941, 1669, 663]. Approaches [2422]. Appropriate [1378]. Approximations [2514, 1800]. Aquacultural [1486]. Aquaculture [2274, 539, 2307]. Aquatic [1497, 951, 1075, 2262, 2321, 557, 1336, 2021, 2409, 737, 2469, 1480, 1737, 231, 332, 2190, 643, 27, 1420, 407, 1613, 2536, 1904, 98, 2029, 2030, 1014, 149, 412, 326, 1058]. Aqueous [2498, 458, 1105]. arc [1749, 265]. arc-en-ciel [1749, 265]. Arches [1663]. Archipelago [75]. Archoplites [548]. Arctic [2474, 1470, 1703, 1074, 1349, 580, 2146, 1051, 2112, 2201, 2451, 28, 911, 569, 1677, 1311, 1814, 661, 2439, 320, 2448, 878, 1863, 2183, 712, 2273, 1393, 291, 396, 2617, 1897, 889, 2271, 912, 1623, 588, 1333, 1725, 2549, 1314, 1176, 1348, 1649]. arcticus [2112, 1897, 588]. Area [603, 1399, 1887, 67, 584, 1382, 2192, 59, 289, 190, 561, 2025, 448, 2316, 2330, 1565, 1358, 1196, 178, 1370, 66, 1262, 1089, 48, 42, 56]. Areal [2517]. Areas [1335, 2437, 1407, 1024, 1463, 1828, 1448]. arenaceodentata [36, 157]. arenaria [336]. argent [2306]. Argopecten [2024]. argus [1775, 1777, 1760, 1774]. Aroclor [1521, 1894]. Aromatic [2322]. Aromatics [391]. Array [2003]. Arsenate [2304, 2303]. Arsenic [895, 55, 778, 1586, 359]. artedii [1436, 1732, 2373, 1602, 654, 1392, 1916, 602, 2546]. Artemia [1425]. Articles [1629]. Artifact [2098]. Artifacts [2447]. Artificial [344, 1634, 2302, 501, 470, 49]. Artificially [2454, 897, 1488]. Asbestos [1508]. Aschers. [744]. Ascorbic [2596, 22, 1293]. Ash [1186, 1121]. Asia [537]. Asiatic [1508]. Aspects [1290, 672, 519, 1199, 2225]. Assay [1847, 1708, 1841, 149, 1400, 1967, 21]. Assemblage [2318, 2034]. Assemblages [2482, 1557, 2563, 2025, 2197, 2527, 2363, 2016, 1686, 1524]. Assess [2536, 2209, 1400, 750]. Assessed [105]. Assessing [1254, 1715, 345, 1418, 522]. Assessment [1789, 74, 2255, 860, 1868, 1100, 444, 1966, 1952, 539, 1148, 262, 123, 1615, 1201, 368, 2419, 1359, 2150, 608, 379]. Assigned [2477]. Assimilation [2078, 2077, 569, 2007, 945, 2176, 757]. Associated [1205, 527, 657, 1411, 2151, 601, 1867, 953, 305, 195, 1213, 1050]. Associates [541]. Association [1290, 1384, 1, 1572]. Associations [455, 1034, 743, 851]. Assuming [556]. Assumption [769]. Assumptions [1789, 1023]. Astro [1379]. Astro-turf [1379]. Atlantic [376, 2310, 283, 177, 1187, 1538, 1584, 1037, 2247, 848, 1854, 2161, 558, 1820, 2050, 94, 1557, 1367, 724, 1973, 1563, 6, 677, 762, 2341, 2538, 839, 1310, 133, 719, 1453, 1107, 1317, 2092, 2133, 2441, 2454, 2131, 1143, 2461, 2248, 859, 1379, 2405, 780, 1108, 1746, 1060, 1729, 2193, 2291, 2439, 2208, 1592, 343, 1945, 799, 567, 1622, 2448, 446, 1902, 1322, 776, 2200, 1598, 190, 699, 740, 975, 1228, 1821, 1845, 2470, 1344, 1283, 1994, 2601, 1327, 2132, 639, 1517, 109, 617, 1435, 1316, 1992, 1489, 1666, 1983, 969, 1921, 2159, 585, 1832, 2367]. Atlantic [2368, 1444, 727, 399, 1181, 1656, 1180, 2590, 2465, 747, 1232, 1078, 268, 1429, 2064, 2197, 1859, 103, 853, 1140, 2015, 496, 1364, 2623, 1278, 274, 1619, 328, 329, 843, 1072, 1932, 871, 1568, 2236, 2591, 1735, 1418, 1057, 1503, 1816, 497, 890, 1326, 2099, 921, 1123, 1163, 2354, 195, 1694, 563, 1139, 1358, 1359, 581, 2149, 1151, 25, 494, 2199, 1637, 2290, 1241, 2404, 2113, 1321, 818, 466, 1026, 271, 2449, 2426, 390, 2569, 1923, 2152]. atlantiques [1327]. Atmosphere [2235, 2236]. Atmospheric [1671, 856, 2313, 1391]. Atomic [2498]. ATP [1214]. ATPase [2258, 76, 1919, 1992]. atratulus [1055]. Atrazine [1739]. atromaculatus [1055]. Attached [32, 746]. Attaching [1312]. Attachment [1279]. Attack [1459, 240]. Attacked [1596]. Attacking [1032]. Attained [84]. Attempts [133]. Attenuation [2215, 1460]. Attraction [2330]. Atypical [268]. auratus [2358, 1240, 445, 767]. aurea [1230, 2173]. auritus [2523]. Australian [2358]. Autocorrelated [2581]. Automated [614]. Automatic [543]. automatique [543]. Autoradiography [57]. Autotrophic [1803, 1964]. autour [1767]. Autumn [2477, 1072]. Autumn-Spawning [2477]. Autumnal [843]. autumnalis [2201]. Auxiliary [1350]. Availability [678, 2262, 1417, 1849, 2359, 199, 1104, 1391, 83]. Available [1372, 1389, 2432, 2525, 2588]. Avant [43]. Avant-propos [43]. Average [317, 499, 774]. Averaging [174]. Aversion [689]. Avirulent [551]. Avis [302, 355, 385, 403, 732, 810, 991]. Avoidance [542, 868, 2417, 601, 1002, 391, 392]. Avoiding [2344]. ayresi [232, 1882]. azteca [1951, 1950, 1875, 797, 1539, 2293].

B [1847, 850]. B.C [764, 1472]. Babine [1023, 861, 1625, 644, 1903]. Bacillus [1709, 2577]. Back [2584]. Back-Calculated [2584]. Background [1318, 1566, 48]. Backscattering [1886, 1438]. Backward [546]. Bacteria [2554, 2467, 746, 995, 1544, 195, 194]. Bacterial [1134, 1542, 1659, 2134, 350, 363, 1201, 1841, 1400, 2587]. Bacterioplankton [1690]. Baffin [682, 1019]. bairdi [708, 308]. Baited [584, 2540]. Balaena [1074]. Balaenoptera [1517]. Balance [2277, 1222]. Balances [1883]. Balanus [888]. balteatus [2011]. balthica [911, 432]. Bank [1429, 2252, 2426, 1820, 1861, 656, 2141, 673, 404, 1594, 2424, 397, 812]. Banks [2310, 1]. Barrage [1801]. Barriers [253]. Barrow [1399]. Base [2392, 261, 203, 1510, 1414, 1628, 2278]. Based [1953, 2429, 603, 1458, 1362, 1306, 1868, 2192, 1662, 960, 2580, 785, 321, 645, 2489, 1753, 2444, 2445, 1006, 850, 1886, 2425, 912]. basée [1307]. Baseline [815]. Bases [2295]. Basin [626, 333, 1852, 309, 2042, 492, 909, 515, 1566]. Basis [2456, 485, 328, 2576]. Basket [416]. Bass [2486, 1004, 1609, 388, 1648, 1184, 1172, 2408, 1230, 506, 2366, 19, 2610, 1596]. Batch [869, 1257, 1585]. Bathylagid [1387]. bathymetric [1267]. bathymétrique [1267]. batrachus [411]. Bay [1098, 2194, 672, 1620, 1447, 2204, 2166, 776, 236, 582, 1693, 815, 1581, 1035, 372, 237, 269, 746, 946, 1770, 682, 909, 1371, 1395, 622, 161, 1019, 568, 471, 2320, 245, 654, 2446, 1238, 1421, 2377, 1745, 840, 1343, 1313, 679, 1801]. Bayer [628, 1831]. Bayesian [2194]. Bays [288, 143]. Be [2200, 2519]. Beach [1974]. Beak [35]. Beaked [925, 1455, 404, 747]. Beam [1723, 2238, 2052, 2237, 2270]. Bear [281]. Bearing [1730]. Bearings [2425]. Beaufort [580]. Bed [196, 1046, 2211, 613, 1389, 669]. Bedrock [2587, 899]. Beds [2184]. Beef [335]. Been [7]. Before [1690, 817, 1088, 1087]. Behavior [1242, 620, 2108, 1422, 1866, 980, 346, 510, 1025, 558, 1005, 529, 719, 811, 597, 288, 868, 1794, 745, 705, 1959, 200, 211, 763, 1489, 2347, 1312, 1483, 648, 735, 2599, 590, 1069, 1769, 2531, 750, 1569, 2119, 19, 936, 325]. Behavioral [1453, 1055, 1746, 487, 1863, 2183, 566]. Behaviour [1717, 1976, 2341, 1634, 2551, 1984, 2296, 2545]. Behavioural [2558]. below [2113]. Beluga [2020, 2539, 2450]. Belugas [1333]. Benefit [366, 1024, 1498]. Benefits [259]. Benthic [1456, 725, 2123, 2353, 934, 1263, 2118, 1659, 127, 1389, 1390, 417, 792, 469, 2297, 2124, 1430, 2132, 1693, 552, 858, 1, 1200, 1198, 1646, 2527, 453, 840, 1216, 1801, 2426]. Benthos [982, 884, 1437, 1587, 427]. Benzene [97]. Benzo [278]. Bering [2059, 142, 444, 1954, 308, 987, 376]. Berried [1770]. Bertalanffy [2300, 915, 121, 1706, 1593, 16, 31]. Better [892]. Between [725, 1445, 2126, 1764, 2110, 1366, 823, 67, 764, 450, 1389, 2018, 1702, 1672, 1822, 782, 1474, 349, 112, 1416, 183, 644, 1238, 1, 2011, 1778, 329, 965, 1728, 2293, 1215, 466, 2240, 702, 1370, 1790, 1866, 2328, 1134, 1260, 714, 2456, 1973, 464, 1610, 2563, 1856, 2351, 1125, 2081, 2525, 261, 362, 2214, 2503, 1983, 870, 2406, 1421, 1330, 2531, 1196, 1013, 2270, 1028, 1604, 415]. Beverton [1069]. Bias [2156, 434, 2110, 2462, 2111, 163, 162, 1264]. Biased [1243, 2584]. bicolor [1381]. Biennial [2471]. bifasciatum [1652]. Bight [656]. Biliary [1919]. bilinearis [202, 866, 2574]. Binary [2167]. Binational [1941]. Binding [2244]. Bio [2520]. Bio-Socio-Economic [2520]. Bioaccumulation [2459, 624, 1464, 922, 290, 1011, 1014, 1425, 1012, 1965, 1525]. Bioassay [1892, 312]. Bioassays [423]. Bioavailability [2118, 1136]. Biochemical [1305, 1431, 2026, 1820, 376, 2371, 892, 865, 142, 159, 1118, 1143, 2216, 1863, 805, 2035, 641]. Biochemistry [1830]. Bioconcentration [578, 1201]. Bioconversion [1425]. Bioeconomic [1366, 876]. Bioenergefics [1254]. Bioenergetic [2489, 1785, 1058]. Bioenergetics [1520, 1866, 2585, 258, 1722, 129]. Biogenic [1421]. Biogeochemical [1553, 2242]. Biogeographic [1515]. Biological [2429, 91, 465, 847, 618, 882, 228, 1071, 975, 1017, 1489, 2295, 1041, 913, 798, 1104, 1438, 111, 1927, 1486, 949, 750, 1869, 178, 498, 2576, 83, 812, 2612]. Biologically [2432, 1124]. Biology [1290, 1466, 1534, 1851, 232, 233, 220, 1387, 1761, 1047]. Bioluminescence [1400]. Biomass [72, 165, 983, 2517, 2321, 2593, 1336, 759, 1419, 2211, 1480, 1737, 2032, 1905, 1659, 957, 1250, 559, 1068, 440, 568, 1889, 447, 1295, 365, 349, 427, 1555, 1144, 2088, 2082, 1740, 768, 1934, 2259, 972, 1518, 2016, 1330, 1975, 1935, 2203, 1599]. Biomonitoring [2513]. BIONESS [93]. Biota [801, 1332, 1497, 27]. Biotic [1451, 1875, 1150, 2295, 2150, 406]. Biotransformation [797, 1749]. Biphenyl [913]. Biphenyls [1481]. Birth [317]. Bisazir [458]. Bivalve [2554, 278, 2586, 704, 432]. Bivalves [1810]. Bivalvia [612, 1567]. Bivariate [2600]. Black [2522, 2176, 2294, 2577]. Blackflies [2338]. Blackfly [1709, 1655]. Blackman [737]. Blacknose [1055]. Blackwater [1659, 1993]. Blade [1250]. Blanc [1490]. Bleached [2272, 386]. Bloater [1872]. Blood [2253, 1902, 699, 997, 813, 214, 2292, 1695, 1414, 553, 2278, 2279, 2280]. Bloom [150, 1043, 1044, 1432, 1469]. Bloom-Forming [1469]. Blooms [2319, 2331, 1551]. Blossom [1035]. Blubber [1912, 2212]. Blue [2510, 1029, 2319, 63, 2471, 2331, 2009, 2501, 1890, 429, 1518, 61, 1975, 743]. Blue-Green [2510, 2319, 63, 2331, 1518, 743, 429, 1975]. Blueback [745, 1485, 2550, 362]. Bluefin [1364]. Bluegill [779, 2014, 1648, 745, 2094, 2599, 276]. Boats [1192]. Bodies [2473]. Body [646, 1286, 1851, 2429, 1141, 901, 1154, 2267, 2458, 1609, 2232, 764, 2234, 1268, 566, 2402, 1900, 1517, 1666, 2464, 2588, 328, 432, 1225, 1495, 1032, 1596, 1644, 702, 1864]. Body-Lead [764]. Bog [1956]. Bone [693, 2410]. Bongo [77]. Book [18, 40, 100, 119, 137, 154, 189, 286, 338, 354, 384, 402, 425, 443, 461, 555, 575, 616, 632, 649, 667, 684, 701, 753, 771, 790, 809, 842, 863, 886, 907, 931, 967, 990, 1016, 1042, 1061, 1115, 1132, 1149, 1170, 1189, 1207, 1227, 1273, 1300, 1328, 1377, 1398, 1413, 1427, 1442, 1479, 1502, 1533, 1561, 1591, 1616, 1641, 1665, 1685, 1712]. Bootstrap [2209]. Boreal [1098, 2357, 277, 749]. borealis [1778]. Boreogadus [580]. Boron [293]. both [915]. Bothnian [236, 237]. Bottlenose [1275, 927, 928, 1304, 1627]. Bottom [1691, 145, 1566, 1011, 651, 640]. Boundary [1390, 1216]. Bounds [1513, 756]. Bovine [1269, 2375]. Bowhead [1074]. Box [750, 1940]. Boxes [926]. Brachionus [97, 10]. Brachydanio [1978]. Brachyura [1572]. Brackish [1873]. Brain [529, 215, 186]. brama [1415]. Branchial [2440]. Break [433]. Breakdown [1092]. Bream [1415]. Breathing [1998]. Breeding [147, 2005, 1325]. Breton [2062]. breviturrita [1906]. Brevoortia [1418]. Briefly [2183]. Bristol [2194, 582, 1238]. Britain [494]. British [1953, 947, 1305, 1431, 1466, 1286, 1534, 861, 1141, 1263, 618, 339, 1852, 2151, 2530, 800, 883, 1674, 880, 1387, 867, 1536, 2257, 1145, 2438, 2556, 1463, 2383, 1657, 817, 1604, 2584, 1917, 1903, 623, 2162, 370, 1851, 2026, 1048, 288, 1690, 606, 1253, 1287, 2400, 2616, 1837, 191, 852, 1705, 422, 500, 451, 1578, 1190, 1964, 71, 1238, 1446, 2532, 723, 1001, 831, 972, 2349, 2205, 1330, 1495, 1496, 1933, 1374, 2612]. Brood [1783]. Broods [1930]. Broodstocks [2337]. Brook [646, 310, 2276, 2277, 551, 726, 1744, 2356, 2396, 450, 1709, 502, 1119, 2234, 2389, 2299, 1984, 1499, 2214, 1316, 2281, 1924, 948, 2283, 2284, 181, 1235, 1523, 1057, 2391, 2395, 1751, 1537, 1754, 1893, 2233, 1947, 2285, 2282, 504, 390, 2278, 2279, 2280]. Brown [128, 1744, 450, 1614, 640, 1822, 2230, 2302, 567, 2448, 566, 2379, 2489, 8, 2035, 2243, 1241]. Browns [2141, 2424, 2252, 2426]. Brownwater [1883]. Brunswick [2430, 1693, 1581]. Bubble [2060, 2362]. Buccinum [2330]. Bucket [2469]. Budget [60, 937, 28, 1086]. Budgets [441, 1971, 1748, 577, 2419]. Buffer [1692, 140]. Buffered [1164, 1434]. Bufo [2013]. Bulk [995]. Bulk-Stored [995]. Bullhead [566]. Buoyancy [84, 1327, 1327]. Buoys [127]. Burdens [829, 2412, 2402]. Burrowing [2412]. Burst [1208, 1496]. Butyltin [2459, 1082].

C [151, 342, 652, 97, 1129, 2378, 1919, 2202, 1632, 707, 2113]. C. [1436, 708, 1602, 654, 308]. côte [870]. côtiers [2225]. CaCO [2387]. Caddisflies [687, 742]. Caddisfly [1826]. Cadmium [2229, 1601, 2403, 1990, 1689, 775, 829, 1222, 2219, 1240, 54, 448, 922, 410, 1679, 2120, 1481, 2293, 1828, 1718, 56, 478]. Cadmium-Stressed [54]. Cage [591, 1199]. Caging [145]. Calanoid [900, 1618, 2221]. Calanoida [901, 1154, 1811, 325]. calanoides [1154, 1811]. Calanus [299, 2063]. Calcareous [1590]. Calcite [2397, 2394, 2051, 2393]. Calcium [2386, 2594, 2392, 28, 1268, 50, 2283, 2284, 1235, 2033]. Calculated [983, 2001, 1039, 2584]. Calculating [1608]. Calculation [1476]. Calculations [937]. Calculator [351]. California [89, 2104, 802, 1025, 1788, 1362, 1605, 1961, 279, 668, 420, 1765, 784, 917]. Calipers [351]. Callorhinus [2534, 2533, 774]. Cambarus [1752]. Campbell [1674, 1963, 738]. camtschatica [2471]. Can [1079, 147, 927, 2421, 2200, 2519, 2614]. Canada [2297, 2103, 1277, 2507, 232, 625, 2108, 1093, 793, 1771, 1557, 1898, 1719, 1051, 1206, 2373, 713, 1283, 2132, 1000, 1387, 396, 1808, 1911, 758, 2623, 723, 1117, 2578, 1448]. canadense [711, 1261]. Canadian [569, 978, 2201, 64, 2323, 507, 1071, 75, 1410, 1393, 62, 1549, 5, 1921, 1041, 2590, 1932, 122, 497, 1333, 840, 1725, 2149, 2549, 1448, 587, 1599]. Canal [814, 1636]. Cancer [802, 1025, 1788, 1362, 613, 591, 1605, 279, 668, 2561, 1578, 1778, 1244, 2556, 2555]. Cannibalism [2059, 1575]. Cannulation [1078, 2280]. Canola [795, 1645]. Canonical [2563]. Canopy [887]. Cap [1146, 1490]. Capacities [2087]. Capacity [1436, 2184, 1306, 698, 905, 2292]. Cape [1801, 1490, 2062, 2595, 1761]. capelan [677, 1621]. Capelin [94, 2572, 1974, 465, 314, 657, 621, 919, 1174, 995, 1209, 1157, 323, 1368, 677, 1621]. capitata [545]. Capitella [545]. capricornutum [1297]. Captive [2450, 590]. Captivity [2592]. Capture [2592, 1614, 1513, 1101]. Captured [1776, 343]. Capturing [351, 548]. Carassius [1240, 445, 767]. Carbohydrases [361]. Carbohydrate [2129, 1020, 1589, 1372, 2076, 1893]. Carbon [60, 2575, 2078, 2077, 295, 1461, 717, 2357, 1323, 2178, 704, 1126, 945, 2352, 2362, 1846, 749, 1658, 98, 1035, 1348, 1649]. Carbon-12 [98]. Carbon-13 [98]. Carbon-13/Carbon-12 [98]. Carbonate [2386, 2392]. Carcasses [1396, 2524]. Carcharhinid [2134]. Carcharhinus [1367].

Carcinus [2033]. Cardiac [2367]. Carlo [2111, 164]. Carnation [1263, 606, 1852, 2151, 2530, 1145, 2438, 817]. Carnitines [2473]. Carnivorous [2221]. Carolina [128]. Carotenoid [1385]. Carotenoids [845]. Carp [2274, 670, 1576, 382, 2005]. carpio [2274, 670, 1576, 382]. Carrying [905]. Cascade [2315, 304]. Cascading [1410]. Case [1650, 2467, 737, 2274, 2040]. Casein [1662]. Casein- [1662]. castelnaui [2557]. Castle [89, 784]. Catalytic [1967]. Catch [1538, 1350, 1856, 680, 273, 2299, 1247, 1485, 1605, 721, 2526, 1378, 2369, 2172, 637, 1424, 1522, 1932, 924, 2236, 2483, 1303, 546, 280, 2555, 1024]. Catch-Age [1350]. Catch-Per-Unit-Effort [637, 1424, 1538]. Catch/Effort [924]. Catchability [202, 1452, 2561, 1370]. Catches [336, 2050, 1380, 2434, 794, 1937, 1763, 496, 1057]. Catching [1253, 1522]. Catchment [1363]. Catchments [1105]. Categorical [1802]. Catfish [1110, 2105, 2499, 1998, 115, 2381, 652, 615, 997, 1926, 788, 411, 421, 1230, 1507, 436]. Cation [2338, 1082, 1491, 162]. Cation-Exchange [162]. Cations [1140, 2167]. catodon [283]. Catostomidae [1136]. Catostomus [1601, 775, 693, 2410, 1849, 203, 2250, 2536, 1936]. Caught [1614, 1435, 2465]. caurinus [1837, 2241]. Causal [2095]. Causative [1841]. Cause [1406, 1160, 1907, 422, 1280, 185, 2614]. Caused [2156, 1946, 788, 972, 271]. Causes [1452, 1519, 2010, 1218, 111]. Causing [268]. Cave [1700]. Cave-Spring [1700]. Cavity [1666, 1943]. Cayuga [622]. CCA [2563]. Cd [203, 1103, 1899, 2167]. Cedar [185]. Cell [115, 1894, 1009, 2234, 139, 445, 727, 2539, 411, 1381, 431, 2031, 2034, 152]. Cells [1943, 2233]. Cellular [2517, 642, 531, 1284]. Censored [2485]. Census [1962, 1477]. Central [1961, 2010, 2248, 2018, 1966, 2313, 1105, 1221, 673, 2562, 941, 2460, 1813, 2293, 961]. Centrarchidae [2523]. Centrarchids [2191]. Centrifugation [353]. Centrifuge [266]. Ceratium [2224]. Ceratomyxa [941]. Cereal [950]. cerebralis [132, 422, 266, 647]. Cerelose [686]. Certain [417]. certains [2225]. Cesium [1722]. Cestoda [861]. Cetacean [2624]. Cetaceans [1969]. Chain [9, 1865, 1339]. Chains [326]. chalcogramma [623, 2059, 142, 188, 2416]. Chamber [596]. Chameleons [488]. Change [7, 1618, 438, 289, 1420, 2312, 1532, 1099, 702]. Changes [2482, 2105, 1205, 1422, 1357, 1538, 357, 2267, 2484, 2458, 952, 1224, 2386, 2594, 1129, 76, 1880, 2014, 2392, 866, 1913, 1732, 1585, 228, 2102, 1259, 289, 1388, 1500, 1638, 62, 1329, 1838, 1840, 1992, 1528, 245, 227, 2368, 1881, 2446, 1578, 1200, 2305, 143, 247, 499, 1072, 1870, 8, 1827, 229, 1213, 972, 2606, 2450, 2285, 178, 1302, 311, 834]. Changing [483, 2172]. Channel [1110, 2105, 2499, 996, 115, 2381, 652, 615, 997, 1926, 788, 905, 411, 421, 1230, 1507, 2383, 436]. Channelized [1740]. Channels [821, 2302]. Chaoborus [2560, 2314]. Char [2474, 1703, 2246, 2439, 320, 2448, 2265, 1863, 2183, 291, 520, 2271, 912, 1623, 912]. Characteristic [453]. Characteristics [1110, 2592, 2211, 1310, 1063, 2410, 1087, 228, 512, 1743, 878, 75, 139, 752, 188, 1195, 1675, 997, 1316, 2408, 813, 2420, 2143, 749, 1647, 695, 496, 641, 2479, 1933, 194]. Characterization [1056, 898, 1733, 2190, 2463, 2083, 2310, 2605]. Characters [603, 2161, 94, 847, 2373, 2043]. Charge [481, 599]. Charr [2473, 346, 2146, 2451, 661, 712, 2273, 2214]. Chemical [910, 1781, 533, 2387, 2118, 2421, 1150, 70, 2336, 1988, 2573, 982, 777, 528, 2025, 2408, 227, 2259, 111, 46, 2450, 1642, 669, 2306]. Chemically [1206]. Chemicals [1457, 2117, 849]. Chemistry [2486, 2253, 2386, 2392, 1052, 53, 1902, 1283, 813, 1577, 1391, 1642, 1383]. Chemoreceptors [2570]. Chemosensory [324]. chez [1267, 1525, 478, 1749, 1810, 265]. chimique [1307]. chimiques [2306]. China [538]. Chinook [2258, 1335, 1141, 1699, 1844, 796, 76, 1910, 1721, 2494, 1073, 733, 544, 1961, 1900, 1674, 2251, 2347, 703, 1360, 1142, 1282, 1569, 2119, 1611]. Chionoecetes [1716, 1267, 1310, 2335, 1954, 2062, 308]. Chironomid [161, 2573, 1423, 1302]. Chironomidae [792, 1407, 1034, 1093, 122, 1094, 1482, 2227, 2513, 2179, 2180]. Chironomus [161, 2402, 1482, 2227]. Chlamydomonas [869, 1257, 1585]. Chloride [1234, 1381, 162, 1588]. Chlorinated [1683, 1120]. Chlorination [386]. Chlorination-Stage [386]. Chlorine [1492, 542, 388, 1184, 2183, 788]. Chloroform [1251]. Chloroguaiacol [1201]. Chlorophyceae [869, 1257, 1297]. Chlorophyll [11, 1134, 1385, 313, 95, 1606, 344, 1075, 631, 1229, 782, 287, 2503, 1731, 1911, 1144, 2504, 758, 2202, 784, 1352, 1203, 1251]. Chloroveratrole [1201]. Choice [2230, 888, 1909]. Chondrus [2184, 2185, 1535]. Chorion [2015]. Chromarod [1904]. Chromatographic [628]. Chromatography [2498, 313, 940, 1474, 1172, 1904]. Chromosome [36]. Chronic [356, 1372, 1222, 22, 30, 1988, 290, 1537]. Chronology [1116, 2442]. Chrysochromulina [1906]. Chrysophrys [2358]. Chrysophyceae [2256]. Chrysophycean [2519, 2008]. Chrysophyte [2563]. Chrysophytes [2256]. Chrysotile [1508]. Chub [1055]. Chukchi [1158]. Chum [821, 996, 814, 1305, 1466, 1286, 1534, 1851, 2026, 2175, 2500, 1842, 1850, 1331, 1571, 653, 2530, 593, 1102, 781, 107, 2163, 938, 2438, 1194, 1028, 2275]. Churchill [1095, 1088, 1087, 1084]. Cichlid [877]. Cichlids [2173]. ciel [1749, 265]. Ciliates [671]. Ciliophora [2085]. Circulation [179, 2310, 2424, 1314]. Circulatory [1912]. Circulus [2000]. Circumneutral [2121, 1942, 1908]. Cisco [1436, 2201, 2373, 318, 654, 1392, 1916, 2546]. Ciscoes [513, 294]. Cladocera [1320, 636, 1064, 1864]. Cladophora [2088]. Clam [336, 28, 444]. Clams [1508, 1079, 2336, 854]. Clarias [1998, 411]. Clarity [2552, 266]. clarki [413, 1260, 1887, 1346, 2246, 1260, 1887, 1346, 2246, 1259, 2044, 2315, 451, 2334, 1454]. Class [128, 952, 1159, 2223, 1612, 1157, 2185, 19, 629, 369]. Classes [1702, 1904]. Classification [209, 951, 847, 2124, 240, 960, 404, 2164]. Classifications [2177]. Clay [648]. Clean [1828]. Cleaning [1083]. Clear [1985, 78, 304, 1692, 917]. Clear-cut [1985, 304, 1692]. Clear-cutting [78]. Clearing [929]. Climate [1821]. Climates [1622]. Climatic [1558, 2223, 1371, 1123]. Cloned [1733]. Closed [2453, 1024]. Closer [2066]. Clupea [729, 2053, 2430, 6, 677, 1621, 1118, 339, 2198, 1958, 1996, 2138, 2139, 776, 740, 1209, 2609, 1489, 1666, 1983, 2477, 1065, 585, 399, 1180, 2611, 2257, 1583, 1156, 706, 831, 563, 1139, 1358, 1359, 590, 2205, 1995, 1604, 2576, 138, 1026, 271]. clupeaformis [1205, 1436, 1096, 511, 1097, 1732, 33, 933, 512, 105, 1266, 604, 1602, 68, 1262, 654, 1695, 256, 243]. Clupeid [1741]. Cluster [2209, 554]. Clutch [204, 1864]. Co [2295, 2377, 1019, 2570, 1610]. Co-distribution [2295]. Co-occurrences [2377]. Coal [1834, 757]. Coal-Derived [757]. Coarse [1046, 2170]. Coast [232, 625, 780, 744, 1000, 2361, 466, 870, 2108, 1852, 713, 1966, 320, 2565, 2354, 1368]. Coastal [196, 2272, 2333, 2104, 1260, 1873, 896, 2430, 580, 1552, 1719, 2218, 657, 1850, 1690, 606, 2457, 306, 1287, 2400, 2245, 155, 787, 191, 852, 1867, 2565, 1805, 451, 1964, 1823, 2048, 1001, 2591, 1827, 1728, 972, 1967, 1330, 2028, 1933, 1219, 881]. Cobalt [2118]. Cobalt-60 [2118]. Cobequid [372]. Cod [1820, 2428, 580, 1051, 893, 1878, 2208, 2200, 190, 1474, 1821, 639, 2540, 617, 1921, 1444, 727, 274, 1619, 2236, 2591, 2427, 2199, 1604, 2130, 1448]. Cods [1325]. Coefficient [1370]. Coefficients [69]. Coelenterates [14]. Coelomic [1844]. Coenagrionidae [2019]. Coenocline [942]. Coexistence [1366]. Coexisting [1541, 180]. cognatus [708, 1311]. Cohabitants [1448]. Coho [1308, 1422, 872, 1985, 583, 342, 1396, 2524, 823, 2375, 1155, 2232, 2218, 1977, 2345, 1331, 1571, 606, 1742, 2151, 2530, 2245, 2491, 2265, 2616, 176, 2508, 1838, 391, 2548, 904, 1565, 1805, 2266, 1483, 891, 2505, 696, 695, 1208, 2181, 1870, 352, 1145, 2438, 1804, 2384, 611, 102, 2531, 2383, 1495, 1496, 2528, 817, 2054, 830, 1917, 335, 1038]. Cohort [2415, 1209, 1800]. Cohorts [2557]. Cold [2498, 542, 567, 1622, 135, 1036]. Cold-Water [542, 1036, 135]. Coldwater [2383]. Coliform [2554]. Collapse [1096, 2468, 1721]. Collected [1485]. Collecting [2213, 2190]. Collection [2469, 806]. Colloidal [807, 1059, 940]. Colombia [606, 2556, 1917]. Colonization [741, 2215, 1217, 2032, 1048, 470, 125, 1587, 1166, 944]. Colonized [1408]. Colony [2159]. Coloured [1945]. Columbia [1387, 723, 1657, 1953, 947, 2162, 370, 1305, 1431, 1466, 1286, 1534, 1851, 2026, 861, 1141, 1263, 1048, 618, 288, 339, 1690, 1852, 1253, 2151, 2530, 1287, 2400, 800, 883, 2616, 1674, 880, 1837, 852, 1705, 422, 500, 451, 1578, 1190, 1964, 867, 71, 1238, 1446, 1536, 2532, 1001, 2257, 831, 1145, 2438, 972, 1463, 2349, 2205, 1330, 2383, 1495, 1496, 1933, 817, 1374, 1604, 2584, 1903, 2612]. Columbian [623, 191]. Column [620, 1204, 1229, 2621, 1908]. Combinations [47]. Combined [1957]. Combining [637]. comme [658]. comment [2603]. comment/lettre [2603]. commentaire [2603]. commentaires [118, 170, 301, 337, 383, 460, 574, 598, 666, 683, 716, 731, 789, 808, 827, 906, 930, 1015, 1114, 1131, 1169, 1246, 1272, 1299, 1397, 1412, 1640, 1684, 1711, 1755, 1817, 1876, 1901, 1949, 2023, 2041, 2136, 2268, 2309, 2374, 2443, 2490, 136]. Commentary [1148]. Comments [209, 1051, 2143, 1375, 118, 170, 301, 337, 383, 460, 574, 598, 666, 683, 716, 731, 789, 808, 827, 906, 930, 1015, 1114, 1131, 1169, 1246, 1272, 1299, 1397, 1412, 1640, 1684, 1711, 1755, 1817, 1876, 1901, 1949, 2023, 2041, 2136, 2268, 2309, 2374, 2443, 2490, 136]. comments/Lettres [118, 170, 301, 337, 383, 460, 574, 598, 666, 683, 716, 731, 789, 808, 827, 906, 930, 1015, 1114, 1131, 1169, 1246, 1272, 1299, 1397, 1412, 1640, 1684, 1711, 1755, 1817, 1876, 1901, 1949, 2023, 2041, 2136, 2268, 2309, 2374, 2443, 2490, 136]. Commercial [2127, 144, 273, 819, 1926, 1763, 1099, 185]. Commercially [1783]. commersoni [1601, 775, 693, 2410, 1849, 203, 2250, 2536, 1936]. Commission [206]. Common [14, 2274, 670, 25, 1403, 1929, 1930]. Commonly [2600]. communauté [677]. communautés [899]. Communication [1781, 528, 254]. Communities [1456, 2303, 2517, 1021, 2397, 2593, 1336, 2326, 1839, 174, 2027, 1880, 2081, 2018, 692, 1690, 1677, 145, 1256, 1547, 2343, 2214, 54, 1840, 1487, 1942, 2082, 778, 2314, 475, 2391, 453, 1018, 1245, 152, 1302, 275, 56, 899]. Community [1279, 2304, 2547, 1741, 640, 2622, 1648, 2544, 473, 470, 975, 568, 448, 922, 634, 2617, 1144, 1896, 959, 2455, 2579, 1200, 2390, 517, 2016, 2587, 2034, 2566, 832, 2298, 2173, 81, 1582, 677]. Company [2204]. Comparable [2206]. Comparaison [2502]. Comparative [1534, 993, 1347, 956, 1542, 502, 1493, 1675, 773, 2380, 2191, 2610, 1327]. Compared [2448, 2584]. comparée [1327]. Comparing [418, 316, 631, 1706, 1937, 1914]. Comparison [2249, 1335, 2326, 2052, 2371, 63, 1143, 388, 1356, 2264, 2373, 1369, 704, 745, 2471, 1598, 1529, 2616, 1450, 1040, 318, 1581, 68, 1655, 1968, 1673, 2210, 571, 2595, 1432, 2064, 630, 1089, 2489, 2202, 637, 1352, 924, 1751, 280, 2157, 1203, 2084, 2549, 2608, 587, 2190, 2502]. Comparisons [1654, 603, 1185, 2543, 2139, 121]. Compartments [1103]. Competing [1512, 1103]. Competition [1366, 450, 2351, 1281, 1672, 1822, 2230, 566, 634, 1643, 1778, 1551, 2109]. Competitive [725, 1634]. complanata [1196]. Complement [36]. Complementarity [375]. Complex [657, 1922]. Complexation [1957]. Complexing [898]. Complexity [638, 2548, 2545]. Component [2172, 2347]. Components [114, 1161, 911, 1952, 982, 1499, 2172, 1938, 553]. Composite [2481]. Composition [196, 1790, 2162, 150, 1538, 336, 477, 1820, 2593, 327, 998, 2458, 1609, 1726, 1698, 1063, 693, 856, 307, 2496, 1987, 1677, 1485, 445, 1344, 2025, 794, 1888, 214, 1731, 785, 2143, 2455, 1515, 1646, 2382, 1827, 1213, 972, 2528, 2199, 669, 1644, 2152, 2279, 2613, 1960]. Compositional [311]. Compounds [1835, 97, 1546, 1391, 589, 1965]. Comprehensive [864]. comprendre [2225]. Compte [575, 931]. Comptes [18, 40, 100, 119, 137, 154, 189, 286, 338, 354, 384, 402, 425, 443, 461, 555, 616, 632, 649, 667, 684, 701, 753, 771, 790, 809, 842, 863, 886, 907, 967, 990, 1016, 1042, 1061, 1115, 1132, 1149, 1170, 1189, 1207, 1227, 1273, 1300, 1328, 1377, 1398, 1413, 1427, 1442, 1479, 1502, 1533, 1561, 1591, 1616, 1641, 1665, 1685, 1712]. Computer [1342, 2287, 776, 351, 192]. Computer-Controlled [1342]. Computer-Driven [2287]. Computerized [2417]. Concentration [2476, 1134, 168, 1670, 2522, 764, 1977, 2338, 1268, 674, 1752, 1978, 266, 468, 635, 184, 38, 462, 1235, 1363, 1284, 720, 1864]. Concentrations [1871, 2182, 414, 1037, 423, 801, 13, 2126, 344, 1075, 1990, 628, 457, 2485, 722, 2410, 1819, 62, 203, 2408, 2562, 1487, 2121, 1722, 1678, 2621, 2293, 432, 564, 2350, 419, 599, 1588]. Concept [495, 491, 479, 480, 486, 490, 690, 1188, 501, 2526, 521, 500, 523, 1375, 17]. Conceptual [1701, 2288, 2577]. Concomitant [2368]. Condensed [2493]. Condition [2552, 2148, 2586, 1517, 669]. Conditional [2453]. Conditioning [1884]. Conditions [1948, 929, 2001, 1159, 1835, 1734, 2111, 2330, 1867, 778, 1304, 659, 29, 19, 21]. Conductivity [2594, 1806, 2546]. Cone [2008]. confervicolus [1, 1463]. Confidence [37]. Confinement [108, 2097]. Conflicting [1681]. Confounding [788]. Congener [2471]. Congeners [913]. Conifer [1092]. Coniferous [1704]. Connecticut [952, 1159, 1452, 1451]. Conowingo [123]. Consecutive [1782]. Consequence [2622]. Consequences [1452, 2144, 395, 1739, 1019, 107, 1181, 517, 949]. Conservation [1776, 520, 493]. Conservative [1019]. Consideration [769]. Considerations [2249, 1785]. Constance [2223]. Constant [26, 1369, 445, 2099]. Constants [1033]. Constituents [386, 2450]. Constraint [1631]. Constraints [1512, 2275]. constricta [735]. Constrictions [1379]. Consumed [2413, 2462, 447]. Consumption [1670, 198, 1433, 1672, 670, 1450, 711, 1381, 462, 1426]. Container [2002]. Containing [2403, 2345, 2250]. Contaminant [801, 2456, 1405, 981, 2339, 2007, 1921, 2536, 399]. Contaminants [321, 798, 149]. Contaminated [981, 1683, 2342, 1120, 113, 1400, 1751, 1828, 275]. Contamination [604, 841, 2306]. Content [1833, 1030, 2586, 1475, 2497, 2616, 2615, 709, 2076, 1982, 545, 714]. Contents [1722]. Contiguous [1051, 884, 2379]. Continental [2310, 840, 1657, 618, 2511]. Continuous [872, 32, 2325, 1552, 139, 1547, 57, 579, 1751, 56]. Continuous-Flow [32]. Continuously [1083, 1067]. Continuum [690, 1375, 17]. Contrasting [2224, 2203]. Contribution [2333, 1174, 2412, 2132, 2093, 2082, 1054, 73, 2349, 2051]. Contributions [2419, 2096, 294]. Control [1205, 532, 230, 529, 535, 973, 1709, 2481, 2319, 231, 252, 228, 2287, 1831, 253, 2327, 260, 471, 262, 1506, 1992, 250, 227, 873, 530, 229, 222, 1486, 243, 259, 371, 1611]. Controlled [1342, 1510, 928]. Controlling [1492, 1389, 2331, 24, 2621, 739]. Controls [1158]. Controversies [1999]. Conundrum [486]. Conventional [1356]. Conversion [619, 983, 801, 1269, 1250]. Cook [159]. Cooling [806]. Cooperative [485]. Copepod [1404, 1548, 1618, 656, 717, 2221, 299, 1710, 296, 2425, 1270]. Copepoda [901, 1154, 1811, 325]. Copepods [900, 72, 165, 1509, 433, 1581, 850, 204, 939]. Copper [2258, 1249, 1990, 1689, 405, 1372, 898, 1430, 1688, 1829, 1830, 2250, 2120, 2599, 1067, 743]. Copper-Complexing [898]. Coral [1652]. Core [1865, 837, 1050]. Coregonid [488]. Coregoninae [513, 294]. Coregonines [1732]. Coregonus [1205, 1436, 1096, 511, 2223, 2201, 892, 1097, 1732, 1702, 33, 2373, 933, 512, 105, 1266, 604, 1602, 239, 318, 68, 1262, 654, 2095, 1872, 1392, 1916, 1695, 602, 256, 243, 267, 2546]. Corer [2158]. Corophium [614]. Corrected [629]. Correction [2412, 1111, 1167, 2483]. Corrections [41, 171, 322, 576, 685, 791, 908, 1062, 1171, 1301, 1428, 1562, 1713, 1877, 2137, 2269, 2399, 2537, 41, 171, 322, 576, 685, 791, 908, 1713, 1877, 2137, 2399, 2537]. Correlated [1634]. Correlation [2467, 1973, 1720, 27, 1420, 432]. Correlations [2452, 1819, 1238, 2293, 1637]. Correspondence [2563, 2240]. Corticosteroid [436]. Corticosteroids [1172]. Cortisol [108, 1838, 2606, 1627]. Corynebacterium [92]. Cossuridae [370]. Cost [2144, 2618, 2124, 1208, 1024]. Cost-Benefit [1024]. Costs [549, 259]. Cottus [725, 477, 708, 1311]. Coulter [409]. Count [1673]. Counter [1402]. Counterimmunoelectrophoresis [1009]. Counterparts [1185]. Counting [820, 1070]. Counts [1357, 351]. Coupling [2235, 2236]. Covariability [1076]. Covariance [2161, 925, 1455]. Cover [1948, 2321, 1737, 1984, 2548, 1813, 1454]. Covered [1277, 24, 691]. Cowlitz [1142]. CPUE [1653]. Crab [802, 1025, 1788, 1079, 1716, 1310, 613, 1662, 1759, 2471, 1605, 2442, 279, 668, 2062, 1578, 2555, 308, 2033]. crabe [1267]. Crabs [1362, 2335, 1954, 591, 2561, 1778, 1572, 1244, 2556, 192, 1267]. Crappie [1736]. Crassostrea [126, 2244, 874, 2306, 669]. Crayfish [2486, 964, 2469, 2468, 1382, 309, 1624, 1889, 50, 601, 763, 1628]. Creek [1263, 1055, 640, 1852, 2151, 2530, 2327, 2022, 1145, 2438, 817, 942, 606, 281]. crenatus [888]. Cresol [944]. creuses [408]. crispus [2184, 2185, 1535]. cristata [1854, 2177]. cristatus [2221]. Criteria [1996, 498]. Criterion [1284]. Critical [1550, 642, 1860]. Critique [714, 1351, 714]. croissance [2502]. Crop [1445, 840]. Cross [1308, 1238, 1027]. Cross-Sectional [1308]. Crosses [874]. Croton [1978]. Crustacea [1787, 2293, 2567, 1864]. Crustacean [1419, 2586, 1068, 1173, 1090, 1780, 475, 275, 1599]. Crustaceans [2442, 2580, 1783]. Cryogenic [2498]. Cryptochironomus [2227]. crysoleucas [2512, 2452]. Crystallization [1682]. Ctenophore [1526]. Cu [203, 1899, 432, 2167]. Cultivation [1898, 2492]. Cultural [1580, 755]. Culture [1404, 869, 1257, 1509, 1009, 1585, 1342, 518, 139, 57, 536, 537, 579, 408, 1926, 2332, 1535, 381, 1686, 2034, 401]. Cultured [1411, 1528]. Cultures [726, 1552, 857, 1547]. Culturing [56]. Cumberland [909]. Cumulative [1528]. cupida [2338]. Current [900, 1436, 2522, 1698, 348, 2422, 84, 2548, 331, 1646, 1569, 494]. Currents [763, 2235, 2236, 2071, 141]. Curve [1476, 915, 1593]. Curves [815, 121, 1706, 1655, 1669]. cut [1985, 304, 1692]. Cutthroat [413, 1260, 1887, 1346, 2246, 1259, 2044, 2315, 451, 2334, 1454]. cutting [78]. Cyanide [715, 710, 1749]. Cyanobacterial [2083]. cyanure [1749]. Cycle [762, 2112, 560, 1676, 2471, 1736, 2031, 2383]. Cycles [2050, 1605, 279, 2242, 966, 939, 1991]. Cyclic [26, 76]. Cyclical [200]. Cycling [1833, 1541, 445, 1549, 1348]. Cyclopoid [734, 1080]. cygnus [1776, 1764, 1763]. cylindraceum [654]. Cylindrical [556, 300]. Cymathere [1047]. cynoglossus [603, 393]. Cyprid [888]. Cyprinid [2207]. Cyprinidae [1978]. Cyprinids [2480]. Cyprinus [2274, 670, 1576, 382]. Cystidicola [987, 894, 838]. Cystophora [1854, 2177]. Cysts [1409, 902, 679]. Cytochrome [865, 1967]. Cytogenetic [157]. cytological [2306]. cytologiques [2306]. Cytology [2234].

d [1862, 2125, 686, 1136]. D-Glucose [686]. Dace [725, 1055]. Daily [869, 2106, 650, 952, 527, 2059, 288, 1296, 703, 581, 193, 702]. Dajidae [476]. d’ajustement [1327]. Dakota [2174]. Dam [2407]. Damage [2175, 788]. d’Amérique [658]. Dams [741]. dans [543, 1490, 1267, 1621, 1862]. Daphnia [2403, 898, 307, 636, 1464, 1415, 468, 607, 697, 757, 1339]. Data [146, 1308, 336, 533, 970, 2001, 2485, 2466, 147, 892, 680, 2206, 273, 1857, 2543, 1070, 915, 1952, 2260, 2526, 1378, 2369, 773, 341, 2580, 833, 2289, 2209, 2605, 1494, 637, 38, 1522, 924, 1802, 2483, 48, 1303, 1869, 595, 608, 1264, 1891, 1982, 1026]. Dating [110, 822]. Dauphin [2552, 2186]. Davis [840]. Day [1474, 2099]. Daylength [2078]. Daytime [2353]. DCMU [2080, 131]. DCMU-Enhanced [2080, 131]. DDE [1116]. DDE/ [1116]. DDT [1116, 299]. Dead [2255, 955]. Death [221]. d’eau [1490]. débit [1125]. Debris [2333, 1396, 1092, 2181, 1704]. Decade [332]. Decapod [2580]. Decapoda [873, 1572]. Decapods [476]. Decay [1865]. Dechlorination [1354]. decipiens [2492]. Decision [2124, 2187]. Decline [838, 877, 974, 1523, 602, 248, 2534, 1664]. Declines [1519]. Decomposing [995]. Decomposition [1836, 1334, 944]. Decreased [593, 50]. Decreasing [1374]. Dee [1435]. Deep [631, 1880, 179, 1019, 396, 130, 784, 694]. Deep-Chlorophyll [784]. Deep-sea [694]. Deep-water [396]. Deepwater [925]. Defense [325]. Deferentia [1681]. Deficiency [1847, 58, 2073, 2076]. Definable [486]. Defining [308]. Definition [1758, 892, 740, 485, 498]. Deformation [900]. Deformities [1482, 2227, 2513]. Degradation [55]. Degraded [999]. Degrees [2345]. Dehydratase [1136]. Dehydroascorbic [1294]. Dehydrogenase [1632, 21]. Delay [1868, 1353]. Delay-Difference [1868, 1353]. Delayed [1652]. d’élevage [1327]. Delineation [1455]. Delphinapterus [2592, 2020, 2539, 1333, 1745, 2450]. Demand [1546]. Demersal [1242, 820, 1210]. Demic [2376]. Demographic [1181, 812]. Demography [438]. Demonstration [1152]. Dendrograms [2210, 2613]. Denil [1473]. Denitrification [65, 82]. Dense [857]. Densities [820, 1478, 1055]. Density [1790, 1079, 1509, 964, 1452, 1558, 613, 2223, 2551, 893, 2521, 469, 1612, 887, 2138, 1853, 2400, 2550, 781, 440, 2195, 1437, 1937, 298, 1239, 1823, 2435, 1423, 1916, 706, 2528, 1826, 138]. Density-Dependent [1790, 1079, 1452, 1558, 2223, 1612, 2195, 298, 1239, 1823, 893]. Density-Independent [298]. Dentin [90]. Dentinal [918]. denudation [899, 899]. Depend [147]. Dependence [1721, 2435, 2160]. Dependent [1790, 1079, 1406, 1452, 1558, 2223, 865, 1612, 2195, 296, 298, 1239, 1823, 1127, 920, 2597, 2235, 942, 893]. d’éperlan [677, 1621]. Depletion [1277, 7, 1907, 1265, 24, 183, 1553, 29, 2160]. Deployment [1192]. Deposit [2118, 2007]. Deposit-Feeders [2118]. Deposition [1671, 1046, 1168, 2313, 1018]. Deposits [90]. Depress [636]. Depression [1986, 277]. Depressions [1946]. Depth [1286, 1998, 1336, 2215, 339, 837, 1434, 1038]. Depths [344, 2032]. Depuration [797]. Deriso [1353]. DerisoCatch [2472]. DerisoCatch-at-Age [2472]. Derived [283, 70, 815, 757, 554, 1135]. Describe [489, 836]. Description [2035, 208]. Descriptions [1093, 723]. Descriptive [2167]. Desert [2319]. Design [93, 831, 1472, 1192, 2180]. Designations [2152]. Designing [891]. Designs [1212, 2153]. Desorption [312]. Destructive [466]. Detect [1409, 927, 369, 629]. Detected [1943]. Detecting [1844, 1420, 2581, 292]. Detection [592, 2498, 92, 1346, 2260, 210, 435, 1904, 2480]. Deteriorating [996]. Determinants [310, 1875, 2299, 1253]. Determination [414, 1471, 2498, 423, 681, 1563, 1847, 134, 343, 604, 2442, 639, 1172, 1101, 182, 1846, 35, 1364, 80, 1933, 608, 550, 543]. Determinations [418, 1385, 68]. Determine [2482, 562, 1145, 2557, 1590]. Determined [1399, 2210, 1429, 612, 2427]. Determining [106, 628, 709, 831, 1024, 61, 1454]. Deterministic [993, 1915]. Detritus [934, 1263, 125, 545]. Developed [1620]. Developing [1067, 835]. Development [1955, 2019, 996, 1471, 1187, 1825, 1511, 1973, 1618, 762, 1051, 465, 1206, 1996, 348, 1341, 251, 470, 2124, 2339, 975, 330, 411, 853, 1766, 1491, 2527, 2029, 2030, 2257, 1785, 1504, 2495, 2480, 2578, 2569]. Developmental [1466, 1534, 1851, 388, 2616, 2615, 468]. Developments [1376]. Deviations [1543]. Devices [2355]. DHP [399]. Di- [399]. diacetylmonoxime [543, 543]. Diacyclops [1548]. Diagnostically [813]. Dialkyllead [1965]. dialysantes [408]. dialyse [408]. Dialysis [1529]. Diameter [1996]. Diaptomid [939]. Diaptomus [901, 1635, 358, 204]. Diatom [2482, 1671, 2507, 2517, 1730, 1166, 148, 2332, 1343, 841, 152]. Diatom-Inferred [1730]. Diatoms [1404, 2122, 2147]. Diel [1242, 1811, 187, 859, 1620, 1700, 1959, 2031, 1696, 2224]. Diet [310, 2105, 2401, 173, 1165, 845, 2389, 1311, 686, 1945, 1902, 270, 773, 456, 412, 803, 1621, 265, 1621]. Dietary [2499, 2061, 2381, 1990, 1618, 1894, 2440, 1269, 1372, 1268, 765, 22, 1293, 1505, 2117, 1425, 2096, 1013]. Diets [79, 2345, 288, 795, 1662, 981, 1645, 836, 390]. Difference [1868, 1449, 1353]. Differences [1254, 626, 1141, 104, 395, 394, 2081, 1341, 566, 1527, 2271, 2292, 2379, 2058, 2419, 2531, 2243, 1917]. Different [1871, 115, 1161, 2456, 1855, 551, 1462, 2345, 1702, 1729, 955, 53, 1622, 1493, 715, 710, 617, 630, 506, 778, 736, 2555, 1304, 2427, 1686, 1028, 2113]. Differential [1674, 331, 276, 2089]. Differentially [2508, 2227]. Differentiating [350, 135]. Differentiation [2104, 376, 1974, 816, 2216, 1506, 1455, 267, 294, 1109, 1345]. Differently [1708]. Differing [872, 2218, 2292, 2527]. Difficulties [567]. Diffuse [1678]. Diffusion [69, 2015]. Diflubenzuron [1404, 692, 79]. Digestibilité [265]. Digestibility [1186, 1645, 639, 2324, 265]. Digestion [1311, 686, 617]. Digestive [2155, 2596, 182, 361, 1044]. digitale [200]. Diked [1747, 1748]. Dilution [1686]. Dimensional [1155, 1727]. Dimethyl [95, 1203]. Dimethylformamide [1203]. Dimictic [1450]. Dimilin [1404]. Dimorphism [984]. Dinoflagellate [2589, 2320, 2224, 401, 1066, 271, 679]. d’intensité [1490]. Dioxide [2078, 2077]. Diphyllobothrium [1703]. Diptera [122, 1094, 1482, 2179, 2180, 1093, 2522, 1655, 2029, 2030, 2227, 2513]. Dipteran [67]. Direct [92, 1741, 299, 148, 80, 1586]. Directly [937]. Discharge [1839, 2131, 1150, 1636]. Discharges [123]. Discount [689]. Discovered [444]. Discovery [1650]. Discrete [509, 2325, 740]. Discreteness [521, 1933, 494]. Discretion [1878]. Discriminant [925, 1455]. Discrimination [1307, 362]. Discussion [198]. Disease [770, 652, 1411, 132, 2060, 2362, 422, 1564, 363, 1841, 305, 195, 194, 1007]. Diseased [1244]. Diseases [1206]. dispar [1933]. Dispersal [2108, 894, 2582, 676, 2141, 1853, 1236, 1885]. Dispersed [1854, 1318]. Dispersion [2430, 2522, 1288, 2424]. Disposition [2372]. Dissipation [1255]. Dissolved [2129, 1833, 929, 295, 2357, 1806, 1546, 1747, 1126, 1318, 2307, 826, 1846, 2114, 768, 1035, 419, 1997, 2306]. Dissolved/Dispersed [1318]. dissoute [2306]. Distance [779, 1631]. Distinct [498]. Distinction [2456, 714]. Distinguishing [1474, 1983, 925, 714]. Distribution [283, 2129, 2229, 894, 1292, 2300, 1399, 746, 2067, 1336, 1005, 1895, 2147, 347, 1698, 127, 1951, 2323, 339, 1702, 2544, 1108, 1958, 1700, 2246, 433, 70, 728, 1831, 2515, 2448, 1636, 236, 2620, 1393, 503, 191, 2320, 2508, 763, 1044, 224, 2289, 571, 671, 2602, 1081, 1793, 478, 2305, 837, 1727, 941, 2181, 1572, 1392, 2212, 229, 769, 1333, 565, 958, 1138, 1539, 1696, 1036, 25, 2085, 902, 1657, 817, 83, 1454, 677, 1767, 2295]. Distributions [2568, 1811, 1049, 682, 820, 2098, 75, 867, 2236, 2591, 453, 1358, 2427, 1064]. District [2204]. Disturbances [2272, 2153]. ditremum [1703]. Diurnal [202]. Diurnally [445]. Diver [2469, 706]. Divergence [1807, 2043, 2567, 2404]. Divergent [1053]. Diverse [326]. Diversion [1096, 1095, 1086, 1395, 1090, 1091, 1088, 1087, 1084]. Diversity [759, 2337, 360, 517, 2566, 1374]. Divisions [713]. DMSO [95]. DNA [2535, 1137, 2328, 2169, 1733, 2461, 2337, 2222, 2463, 305, 2257, 2608, 1970]. DNA-Level [2222]. Do [2254, 1546, 1293, 1304, 1833]. Does [1542, 1687]. Dogfish [1471]. Dolly [1346, 2246, 1259, 2265]. dolomieui [2486, 1609, 2408]. Dolphin [1275]. Dolphins [927, 928, 1304, 1627]. Domain [1891]. Domestic [502]. Dominance [1890]. Dominant [1445, 1734]. Domoic [2507, 2332]. Don’t [416]. d’origine [658]. Dorsal [1471, 1699]. Dosage [543]. Dosages [1925]. Dose [2543, 1715, 1714, 2393]. Double [12]. Double-Tagging [12]. Doubleday [2472]. down [2135]. Downstream [985, 2538, 328]. Dragonflies [1717]. Dragraked [2184].

Drainage [1852, 2419, 281]. Drainages [1728]. Draining [1985, 2042]. Drawdown [1027]. Dreissena [2553, 1810]. Drift [1274, 1918, 1224, 1361]. Drifting [2545]. Drilling [620, 324, 1179]. Driven [2287]. droebachiensis [980, 1244, 25, 466]. Drug [334]. Drugs [421]. Drum [835]. Dry [1864]. Dual [1723, 2052, 2007, 1078]. Dual-Beam [1723, 2052]. dubius [397, 812]. Due [434, 567, 926, 2186, 2380]. dulcicoles [714]. Dummy [560]. d’un [1490, 2502, 1307]. Dunaliella [359]. Dungeness [802, 1025, 1788, 1362, 1605, 279, 668, 1578, 2556, 2555]. Duoculture [2439]. Durable [2017]. Duration [2238, 2356, 2237, 2180]. Duration-in-Beam [2238, 2237]. Durations [1754]. During [1948, 869, 558, 65, 628, 2594, 395, 394, 1379, 653, 670, 330, 1667, 974, 1257, 1589, 1618, 1043, 2341, 2010, 1222, 2219, 1732, 1052, 2338, 1088, 1087, 2352, 1838, 1044, 2562, 2368, 1491, 1870, 2483, 2480, 2024, 1929]. Dwarf [661, 712, 318]. Dwelling [1384, 1454, 1750]. Dynamic [1854, 1017, 2526, 2475]. Dynamics [923, 2129, 1650, 1308, 2517, 802, 1074, 1408, 993, 2520, 1918, 2220, 1719, 1461, 1879, 1880, 1791, 2357, 976, 657, 909, 78, 2264, 933, 1247, 1980, 2525, 161, 1547, 2261, 2609, 2352, 1624, 1564, 54, 1840, 2565, 1808, 71, 2048, 1679, 1793, 2114, 2115, 784, 2029, 1235, 871, 113, 2109, 1194, 1993, 1540, 1554, 436, 1935, 942, 2130, 1923, 2024].

Early [1146, 821, 814, 1005, 314, 610, 653, 2506, 1341, 604, 705, 2616, 2615, 1000, 1280, 1239, 2093, 1151, 2480, 1321, 1611]. Earthy [1160]. Earthy-Flavor [1160]. East [2595]. Eastern [1771, 2005, 2507, 626, 793, 2430, 508, 2059, 569, 728, 654, 1931, 2623, 475, 1745, 308, 1158, 1929, 1930]. Eastmain [1395]. eau [543]. écailles [714]. Ecdysial [873]. Ecdysis [697]. échelle [88]. Echo [2238, 820, 1070, 1886, 1937, 1916, 554]. Echosounder [2052, 1905]. ECOLOG [820]. Ecological [2429, 580, 1526, 1168, 687, 2553, 1739, 2417, 2214, 1420, 1623, 1753, 1802, 2243, 226]. Ecology [269, 1796, 2423, 127, 1457, 672, 117, 2047, 2142, 23, 2574, 234, 2315, 2191, 212, 700, 588, 238, 323, 1261]. Economic [114, 2249, 2520, 144]. Economically [623]. Economics [114]. Economy [1801]. Ecophysiology [1588]. Ecoregions [2240]. Ecosystem [1116, 2403, 1376, 1147, 1658, 2385, 1724]. Ecosystems [619, 957, 367, 1697, 2536, 2135, 113, 1510, 1809, 1704, 1482, 2240, 81]. Ecotoxicological [2303, 1839]. Ecotypes [2373]. Edaphic [1267]. édaphique [1267]. Eddy [1560]. Edible [1047]. Editorial [524, 754, 992, 1818, 524, 754, 992]. Editorial/Éditorial [524, 754, 992]. edulis [1898, 1029, 1390, 2009, 2501, 1525, 1899]. Edward [2507]. Edwardsiella [1507, 2495]. Eel [1415, 658]. eelpouts [396]. Eels [2456, 1700, 1738]. Effect [1456, 1789, 1337, 1466, 741, 1998, 1165, 72, 983, 2403, 1025, 779, 55, 1670, 1204, 2215, 1521, 1990, 1894, 2002, 2228, 2345, 2521, 657, 856, 1150, 1672, 1268, 643, 1086, 2230, 733, 765, 688, 1354, 1853, 160, 1978, 2301, 2362, 794, 1101, 2172, 365, 1840, 552, 2464, 2176, 2271, 2283, 2284, 1731, 2, 1401, 1678, 778, 2168, 1334, 2623, 293, 73, 2314, 2236, 1583, 31, 1285, 474, 1194, 1258, 2177, 944, 372, 409, 1596, 1091, 2569]. Effected [77]. Effective [1385, 825, 2330, 1804]. Effectiveness [633, 300]. Effects [678, 2229, 2105, 623, 2340, 910, 620, 2253, 2019, 2473, 2258, 2303, 2277, 165, 1788, 1589, 2397, 1332, 198, 82, 2522, 1439, 1834, 1159, 1451, 1558, 1689, 230, 1842, 1269, 1372, 2125, 438, 1741, 457, 2223, 2346, 116, 730, 893, 866, 2497, 2069, 231, 909, 1168, 1987, 78, 2018, 1560, 1939, 692, 999, 1585, 674, 2138, 101, 544, 1447, 728, 22, 30, 2151, 2012, 1484, 2103, 2142, 2100, 2550, 2339, 1902, 1747, 1365, 2111, 2171, 786, 1344, 982, 1660, 261, 922, 2571, 2281, 1312, 1420, 788, 49, 304, 1692, 858, 1487, 1647, 2121, 1201, 2189, 140, 798, 2005]. Effects [298, 1090, 846, 1964, 1234, 290, 1892, 225, 1423, 2120, 2029, 1481, 1778, 2390, 935, 2591, 340, 1586, 2599, 890, 2099, 352, 2354, 697, 2391, 51, 46, 1694, 2438, 2366, 972, 2384, 1518, 2016, 2493, 2034, 2566, 1504, 1537, 564, 2528, 2039, 2054, 276, 2282, 379, 939, 936, 1991, 2207, 1801, 720, 743, 834, 389, 56, 2306]. Effets [2306]. Efficacy [2412]. Efficiencies [875, 1008, 2207]. Efficiency [1226, 619, 983, 801, 2618, 621, 1987, 47, 321, 2117, 1198, 2395, 1192, 2393]. Effluent [32, 570, 1121]. Effluents [2272, 386, 45, 565]. Efflux [2234]. Effort [1790, 549, 1538, 802, 273, 1369, 2526, 1378, 2369, 37, 1512, 637, 1424, 1522, 924, 1303, 280, 804]. Egg [1955, 1466, 1851, 1597, 1770, 1458, 465, 1174, 1849, 2400, 2492, 1705, 2420, 926, 2095, 2188, 1962, 1679, 1282, 2063, 1537, 1402, 1995, 1064, 835, 1488, 1291]. Egg-Laying [2492]. Egg-Mass [1962]. Eggers [2106]. Eggs [1439, 2057, 1174, 1097, 197, 339, 1108, 1060, 544, 781, 1474, 2607, 1283, 2616, 2615, 805, 798, 1140, 1679, 2532, 416, 2286, 2199, 2113, 204]. Ehrbg. [10]. EIBS [1900]. Eight [2618, 763]. Eimeria [1983]. Eimeridae [1983]. ELA [47]. Electric [1402]. Electrode [1679]. Electrolyte [1589, 1222, 1679, 29]. Electrolytes [2234]. Electron [470, 2083]. Electrophoresis [105]. Electrophoretic [1310, 1063, 512, 969, 2143, 1241, 1917]. Eleginus [1051]. Element [2358, 2412, 1819, 303, 456]. Elemental [693, 785, 1213, 1960]. Elements [2573, 387, 1045, 1590]. élevage [1862]. Elevated [2250]. Elimination [2554, 2372, 467]. Elliott [2106]. Elliptio [1196]. Elvers [1269]. Embayment [816]. Embolism [2060]. Embryo [2359, 853]. Embryonic [1955, 2569]. Embryos [133, 935]. Emerged [1569, 2119]. Emergence [67, 1974, 314, 859, 1523, 2179]. Emergent [1813]. Emerging [122, 1094, 2180]. Emigration [1985, 2405, 2530]. Emphasis [2323, 1975]. Empirical [1134, 557, 1635, 2211, 717, 1951, 876, 559, 1068, 349, 2294, 1934, 1989, 966, 1330, 2237, 1599]. Empiricism [700]. emploi [408]. Employed [1385]. Employing [2231]. Enchytraeid [1993]. Enclosed [2027]. Enclosure [60, 2597, 2017, 56]. Enclosures [796, 1739, 591, 2571, 1386, 1138, 1012, 1824, 2231]. Encounter [556, 1167]. Endangered [498]. Endemic [291]. Endocrine [532]. Endocrinology [536, 1780]. Endonuclease [1137]. Energetic [1955, 2619, 2620, 2275]. Energetics [179, 711, 844, 659]. Energy [1871, 1174, 1389, 395, 394, 1371, 161, 577, 1837, 10, 2250, 709, 599, 835, 1644, 335]. England [497]. English [1226, 2339, 786, 2621, 1823, 2048, 1481, 1010, 1011, 1014, 1012, 1013].

Engraulis [2093]. Enhanced [905, 2080, 131]. Enhancement [1795, 197, 1023, 1486]. Enriched [1166]. Enrichment [2122, 2262, 2123, 1552, 950, 1964, 1423, 1827, 1553, 841]. Enteric [770]. Enteromorpha [1835]. Entire [2379]. Entrained [1636, 806, 756]. Entrainment [300]. entre [714, 1125, 870]. Entry [1153, 582]. Environment [882, 2343, 2540, 2266, 412, 1628, 1960, 899, 88]. Environmental [1779, 2156, 1788, 1020, 650, 547, 527, 2437, 2563, 438, 2006, 2466, 2244, 1875, 621, 1223, 909, 1100, 2515, 2297, 153, 2620, 2260, 1344, 1148, 330, 1867, 2295, 1647, 321, 798, 2094, 1536, 496, 2363, 368, 890, 965, 1694, 972, 1158, 2293, 371, 294, 1066, 19, 1738, 2308, 2578, 2545, 954]. Environments [104, 2543, 955, 2191, 2502]. Enzyme [1141, 1847, 182, 1841]. Enzyme-linked [1841]. Enzymes [1037, 1044, 1633]. Eocene [2624]. Eogammarus [1, 1463]. eperlanus [2046, 1415]. Ephemera [2186]. Ephemeral [854]. Ephemeroptera [985, 2412, 2186]. Epicarida [476]. epidemic [1862]. épidémie [1862]. Epidemiology [1206]. Epifauna [127]. Epigenetic [989]. Epilimnetic [1541, 2496, 855, 2571, 1022, 1225]. Epilithic [2587]. Epipelic [2390, 1434]. Epiphytic [1443]. Epischura [1618, 325, 1270]. Episodic [1994, 1751]. Epithelia [289]. Epizootic [1031, 268, 1783]. Epontic [2158]. Equation [86, 31, 546, 1353]. Equations [2514, 721, 1800, 2418]. Equatorial [618]. Equilibrium [1467, 326, 1669, 1935]. Erie [626, 7, 660, 610, 815, 2382]. Erosion [1085]. Error [1520, 2110, 2267, 970, 756, 163, 1613, 184]. Errors [940, 2614, 380, 1932, 707, 61, 379]. Erythrocyte [1128, 1136]. Erythrocytic [1900, 107, 274, 1619]. Erythropoiesis [1240]. Escape [1984]. Escapement [560, 1938]. Escapements [377]. Esocids [1254, 2380, 2610]. Esox [1098, 2436, 2552, 187, 2148, 203, 1885, 1013, 1477]. espèce [714]. Especially [190]. Essay [1355]. Essential [2573]. established [2304]. Establishment [411, 2587]. Esterase [2146]. Esterase-2 [2146]. Esters [1226]. Estimate [283, 557, 1905, 1306, 986, 273, 2421, 280, 257, 1426]. Estimated [2206, 2614, 668, 2015]. Estimates [434, 848, 327, 760, 2002, 613, 717, 1250, 440, 2111, 380, 2165, 1218, 756, 37, 68, 1437, 1673, 1937, 2143, 427, 1722, 437, 2202, 1198, 1916, 2483, 1745, 2157, 256, 2177, 581, 850, 164, 554]. Estimating [146, 1308, 1833, 2162, 2106, 970, 2453, 317, 1480, 86, 584, 66, 1063, 63, 2192, 2378, 1356, 577, 1952, 2111, 794, 1040, 341, 1663, 1581, 571, 645, 424, 1003, 280, 192, 2533, 1477, 2153, 2130, 1750, 426]. Estimation [2156, 1243, 918, 1854, 336, 2037, 91, 2267, 1309, 34, 1443, 2485, 398, 1806, 2192, 2432, 2462, 2550, 915, 2369, 1295, 1888, 2580, 1655, 2128, 2577, 953, 2237, 1264, 554, 1007, 1864]. Estimator [2238, 2000]. Estimators [998]. Estradiol [1521, 1449, 2317]. Estradiol-17 [1449, 2317]. Estradiol-lnduced [1521]. estuaire [6, 677, 1621]. estuarien [2502, 88]. Estuaries [676, 704, 2251, 2275]. Estuarine [2538, 1835, 438, 999, 1102, 1360, 1586, 599, 942, 2502, 88]. Estuarine-Dependent [942]. Estuary [558, 1552, 932, 127, 2431, 1674, 1220, 1288, 1827, 1463, 323, 677, 1621, 74, 6, 819, 561, 2360, 1963, 1416, 1]. état [478]. Etching [2336]. Ethanol [1059, 866]. Ethoxycoumarin [646]. ethylase [646]. Ethylation [2498]. Etomidate [997]. Eubothrium [861]. Eulamellibranches [1810]. Euphausia [875, 478]. Euphausiacée [478]. Euphausiid [2511]. Euphausiids [1248, 568, 87]. Euphorbiaceae [1978]. Europe [234]. European [2046, 1898, 236, 2442, 239, 238, 237, 1738]. Eurytemora [1581]. Eurythermality [597]. Eutrophic [910, 161, 1549, 1245, 2566, 917]. Eutrophication [610, 2352, 45]. Evacuation [2155, 572, 1077, 1722, 2416]. Evaluated [2319, 1967]. Evaluating [681, 2481, 246, 447, 1872, 2205]. Evaluation [1186, 562, 2106, 1606, 1786, 356, 2057, 1361, 2331, 2263, 1172, 1774, 1065, 709, 1381, 758, 158, 1510, 2109, 2447, 162, 1192, 2237, 1975, 2411, 1824, 2231]. Evasive [276, 548]. Event [2559]. Events [1324]. Evidence [2376, 1680, 14, 2340, 2036, 364, 2328, 838, 509, 1887, 1699, 1405, 2133, 865, 2319, 618, 215, 2018, 1281, 610, 289, 2619, 740, 1195, 1449, 2347, 2367, 1230, 1678, 1772, 2045, 328, 329, 474, 2241, 1374, 818, 1903, 2279, 2280]. Evolution [533, 490, 215, 1633, 416, 1130, 44, 1630, 1135, 1135]. Evolutionary [989, 1632, 2567, 525]. évolutives [2225]. Examination [203, 467]. Examined [2563]. Example [1788, 831]. Examples [128]. excavata [2320, 902]. excavate [679]. Excess [1401]. Excessive [1486]. Exchange [60, 2432, 2525, 673, 1416, 162, 1215]. Exchanges [2253, 1628]. Exclusion [1607, 1230]. Excretion [2440, 614, 1662, 670, 1919, 181, 2449]. Exercise [2182, 872, 2012]. Existence [1810]. Exotic [504]. Expected [1025]. Experience [719, 1100]. Experienced [2555]. Experiment [980, 1614, 59, 1567, 2251, 2385]. Experimental [1456, 150, 2106, 32, 2122, 55, 2326, 2468, 1741, 126, 1648, 1281, 2138, 2139, 53, 1101, 1564, 448, 2342, 99, 1262, 468, 1964, 891, 2455, 1619, 2365, 1653, 113, 1510, 48, 1507, 944, 2173, 2153, 1626, 2231, 2306, 478, 67, 66, 1382, 59, 2352, 1089, 44, 42, 56]. expérimentale [2306, 478]. Experimentally [438, 33, 2230, 1259, 282, 1481, 553]. expérimentaux [43]. Experiments [434, 2002, 1858, 1986, 1361, 2597, 1183, 12, 1386, 1510, 1330, 841, 463, 2180]. Explain [1951]. Exploitation [1650, 1319, 1174, 1721, 819, 2556]. Exploited [1719, 33, 933]. Exploratory [1355]. Exploring [38]. Exponents [2600]. Export [2325, 1808, 749]. Exposed [1030, 775, 615, 2183, 445, 642, 2536, 1481, 1695, 1067, 2606, 1414, 1754, 436, 757, 2017, 1588]. Exposure [2253, 1337, 1894, 2353, 2356, 356, 1222, 2219, 1987, 30, 1676, 1752, 2362, 2281, 1897, 2283, 2284, 290, 1751, 1537, 1893, 2285, 2278]. Exposures [620]. Expression [1151]. Extended [2592, 1223, 2219]. Extension [2442, 1252]. Extent [1111]. External [639, 362, 1180, 938, 1498]. Extinction [2407]. Extracellular [464, 1182, 401]. Extract [1978]. Extracted [726]. Extraction [1385, 95, 1214, 1529, 353, 1203, 1251]. Extrapolation [2543]. Extrathyroidal [457]. Extreme [1622]. Extrusion [2188]. Eyed [935].

F [2202]. f. [161]. Fabr. [1267]. Face [816, 377]. facteurs [1621]. Factor [2499, 405, 1298, 1539]. Factors [646, 1797, 1606, 2510, 2110, 2572, 901, 1320, 2522, 547, 1451, 1558, 1895, 2147, 2223, 66, 127, 465, 1389, 2338, 1380, 780, 674, 1958, 2515, 782, 888, 1777, 2297, 2103, 84, 176, 24, 2316, 689, 2562, 224, 703, 2188, 2446, 962, 1896, 2621, 2259, 2363, 1066, 739, 1738, 2308, 2578, 587, 1626, 1621]. Facultative [745]. FAD [2355]. falcatus [1082, 1211]. Fall [1229, 2373, 740, 2534]. Fall-Spawning [2373]. Family [1800, 1804]. farionis [987]. farm [1862]. farmed [1327]. Farming [1199, 1285]. fasciatus [1598, 925, 1455, 404, 747]. Fast [1529]. Fast-Oxine [1529]. Fat [1517, 29, 1864]. Fatally [2589]. Fate [151, 2524, 412, 326, 2393]. Fates [52]. Fathead [2100, 2417, 84, 1663, 849, 1032]. Fatty [2473, 1425, 25, 2199]. Fauna [492, 837]. Faunal [2323]. Feasibility [929]. Features [362, 1358]. February [2245]. Fecal [173]. Fecundity [1310, 2139, 1073, 211, 1705, 1000, 2283, 2250, 2029, 1586, 1995]. Fed [845, 1662]. Feed [1186]. Feeders [2118]. Feeding [900, 1242, 1422, 202, 2436, 2267, 2583, 1741, 613, 2232, 2059, 796, 220, 2551, 288, 2221, 1987, 1379, 299, 1620, 2138, 733, 2142, 23, 2448, 2007, 2574, 1693, 2317, 634, 1489, 2420, 213, 1173, 1, 1008, 588, 1617, 590, 581, 323, 2173, 1261, 1644, 325, 1270]. Female [1955, 1466, 1584, 357, 1362, 1742, 10, 2317, 2189, 1488, 297, 1038]. Females [1954, 2289]. Fenitrothion [2125, 186]. Fenvalerate [2027]. FEP [2423]. Ferrous [1049]. ferruginea [1861]. Fertile [2334]. Fertilization [2116, 1287, 2331, 2571, 1262, 1045, 1166, 1330, 1488, 1782]. Fertilized [1642, 2179, 2364]. Fertilizer [1045]. Fiber [414, 765, 2202]. Fibers [1508]. fibres [408]. Fibrils [807, 1059]. Fibroblastic [139]. Fickian [676]. Fidelity [1579]. Field [1681, 1550, 1717, 1820, 2001, 1155, 556, 1168, 332, 1281, 1356, 2462, 1361, 2251, 642, 1494, 321, 435, 2597, 938, 131, 2207]. Fifth [1041]. Filamentous [194]. Films [1304]. Filter [1741, 1083, 2173]. Filter-Feeding [1741, 2173]. Filtering [1320, 1891]. Filters [2202, 1514]. Filtration [2476, 2065, 2027, 1844]. fimbria [562, 2108, 1000, 1617, 659]. Fin [984, 1699, 1517, 68, 195, 2243, 1032, 442]. Fin-Ray [68]. Financial [1285, 1099]. Findings [1969]. Fine [1046, 2035, 292, 1704, 300]. Fine-spotted [2035]. Fingerling [108]. Fingerlings [2245, 904]. Fingernail [854]. Finite [2539, 1183]. Finland [236, 235, 237]. finmarchicus [299, 2063]. Finned [1296]. Fire [46]. First [317, 332, 1379, 1513, 1667, 1983, 748, 48, 1151, 2290, 390, 774]. First- [1983]. First-Generation [2290]. First-Order [332]. Fish [923, 2107, 495, 2272, 2038, 2340, 492, 1030, 1520, 414, 1601, 316, 1866, 532, 1095, 2106, 2413, 2414, 2547, 483, 1175, 779, 2037, 156, 2050, 1373, 1511, 1811, 2559, 2436, 2127, 1834, 964, 2353, 2437, 230, 2433, 2002, 529, 820, 2052, 535, 1634, 2000, 2006, 2358, 1185, 2466, 1113, 676, 932, 1076, 1478, 2098, 2586, 2318, 1027, 640, 986, 1206, 877, 2192, 559, 538, 999, 734, 1080, 887, 2462, 473, 1369, 1915, 1247, 2355, 688, 2506, 160, 518, 622, 878, 1819, 1388, 1636, 1384, 139, 1410, 1365, 1228, 132, 485, 536, 537, 1210, 2614, 1683, 806, 1638]. Fish [350, 528, 330, 828, 454, 1172, 521, 2165, 836, 1693, 1603, 1527, 1053, 1394, 262, 345, 447, 1040, 1101, 1329, 1387, 2316, 2571, 351, 634, 2617, 1629, 689, 303, 456, 711, 1468, 805, 1937, 833, 1201, 953, 2446, 321, 913, 798, 2135, 1736, 298, 2322, 1848, 2064, 2049, 2532, 135, 305, 2588, 755, 1515, 1103, 1989, 1632, 1643, 73, 1932, 2363, 914, 2164, 1392, 2483, 517, 493, 2600, 1386, 965, 1928, 2444, 2445, 1473, 831, 1630, 2366, 586, 412, 1018, 1245, 565, 534, 452, 750, 1036, 1869, 1402, 2003, 595, 522, 1524, 1498, 141, 599, 835]. Fish [1652, 942, 442, 1007, 2344, 185, 1824, 2231, 714, 851]. Fish-Aggregating [2355]. Fished [584, 2330]. Fisheries [472, 1650, 114, 549, 2276, 2429, 2050, 169, 993, 994, 1366, 144, 366, 1310, 1461, 698, 1063, 2481, 398, 1981, 2222, 1212, 876, 1338, 1980, 374, 153, 1701, 2422, 2518, 519, 2132, 1513, 521, 500, 668, 1615, 1512, 2295, 1559, 879, 71, 920, 1792, 1728, 2349, 1769, 2028, 1252, 235, 2308, 1664, 2612, 883, 2423]. Fishermen [1279, 1776, 582, 2187]. Fishery [1096, 486, 802, 1025, 1764, 993, 2520, 128, 206, 2228, 893, 2299, 2204, 2301, 1760, 2172, 279, 1888, 2062, 2590, 1761, 1298, 694, 1278, 1355, 73, 1418, 1816, 2354, 864, 1928, 602, 1133, 523, 1069, 824, 1099, 804]. Fishes [947, 725, 1319, 2585, 2061, 201, 2055, 542, 1406, 1347, 527, 375, 662, 811, 1856, 2346, 1850, 597, 1281, 1672, 487, 752, 577, 1236, 488, 5, 96, 1002, 2288, 85, 199, 2117, 1147, 1740, 655, 2377, 29, 883]. Fishing [1790, 664, 730, 2355, 2614, 200, 1512, 1429, 1167, 1816, 2483, 1133, 1024, 237, 1448]. Fishless [2396]. Fishway [1825]. Fishways [1473]. Fitness [1701, 102, 525]. Fitting [1706, 158]. Five [1164, 1573, 512, 622, 2615, 1693, 203, 1936, 2167]. Fixation [968, 64, 63, 674, 2447, 743, 917]. Fixed [1638]. Fixing [2057]. Fjord [1248, 2619, 2620]. Fjords [884]. Flagfish [101, 1676, 1580]. Flame [1904]. flaring [1422]. Flatfish [2289, 416, 748]. flavescens [2376, 2068, 1449, 1506, 2317, 1077, 1884, 2464, 2292, 1962, 2, 1200, 2045, 2305, 2597, 570, 1392, 1231, 1696, 2479]. Flavor [1160, 185]. Fleet [1253, 1247, 1980]. Fleets [1366]. Flemish [1146]. Flesh [1409]. Fleshed [2494]. Flex [949]. Flexibacteria [135, 305]. Flexibility [2436, 811]. Flexible [86]. Flies [2522]. Flin [203]. Floating [1813]. Floating-Leaved [1813]. Flocculation [168]. Flock [1403]. Flon [203]. Flooded [1095]. Flooding [1747, 2039]. Floodplain [1852]. Floods [2346, 2323]. Flora [9]. Florida [963, 2126, 1021, 1075, 2262, 2321, 2510, 1775, 1777, 1760, 1774]. floridae [101, 1676, 1580]. flottabilité [1327]. Flounder [603, 650, 2055, 1861, 393, 865, 1174, 572, 2434, 617, 2322, 594, 1425, 1967, 430]. Flounders [783]. Flow [2476, 2104, 32, 2575, 1887, 1389, 816, 469, 1484, 75, 806, 10, 1351, 1658, 1917, 1058, 1125]. Flow-through [1058]. Flowing [2409, 1768]. Fluctuating [2543, 2031]. Fluctuations [1159, 1759, 1867, 1962, 2225]. Fluid [1844, 2607, 1943]. Fluids [1179]. Fluorescence [2498, 2080, 131, 2270]. Fluorescent [92, 1844]. Flushed [166, 1284]. fluviatilis [236, 239, 238, 237]. Flux [2217, 2496, 1150, 1126, 62, 1414, 1258, 1225, 449]. Fluxes [2277, 1883, 2261, 2120, 1843, 2279]. Fly [1614, 2176, 2294, 2577, 1121]. Fly-Caught [1614]. Focal [2175, 2500]. Follicle [1499]. Followed [2498]. Following [1456, 1224, 2386, 2392, 1913, 332, 1831, 1752, 593, 1493, 904, 2395, 1383, 297, 1422, 1096, 1862, 1200]. Fonction [86]. fontinalis [646, 310, 2276, 346, 2277, 551, 726, 1744, 2356, 2396, 450, 1709, 502, 1119, 2234, 2389, 2299, 1984, 1499, 2214, 1316, 2281, 948, 2283, 2284, 181, 1235, 1523, 1057, 2391, 2395, 1751, 1537, 1754, 1893, 2233, 1947, 2285, 2282, 504, 390, 2278, 2279, 2280]. Food [2046, 1717, 151, 1319, 2413, 983, 2522, 1861, 375, 1269, 1784, 9, 1390, 2497, 2462, 661, 1849, 1259, 2452, 566, 1517, 1693, 1963, 1527, 447, 1040, 1884, 71, 98, 1158, 581, 1504, 326, 757, 1826, 1339, 2207, 1454, 2306]. Food-Chain [1339]. Food-Induced [2046]. Food-Limited [1390]. Forage [1388, 261]. Foraging [2585, 1462, 811, 2512, 1346, 258, 1959, 2163, 1483, 2197, 2599, 2523]. Forbes [1618, 1270]. Force [261, 260, 262]. Forced [2235, 2236, 2591, 1869]. Forcing [1788, 2619, 2620]. Forecast [2475]. Forecasting [2194, 1940, 718, 1163, 1006, 452, 2205]. Forecasts [2249, 2604, 2598]. Foreign [2244]. Foreign-Particle [2244]. Forest [2357, 1852, 277, 749, 1045, 46, 2438, 1704]. Forested [46]. Foreword [42]. Fork [1663]. Form [437, 1032, 1596, 2306]. Formalin [866, 674]. Formation [2333, 1440, 2624]. forme [2306]. Formed [2103]. formes [478]. Forming [1469, 2089]. Forms [1105, 1151, 2371, 478]. Formulation [1709]. Fortnightly [1371, 1867]. Forward [546]. Fossil [2482, 1019]. Found [2600, 208]. Four [2554, 1226, 2127, 1883, 977, 2243, 1935, 1032, 1596]. Fourier [1654]. Fourteen [1819]. Fowlkes [2210]. Foxe [1566]. Fraction [1033, 1059]. Fractionation [2118]. Fractions [630]. Framework [2288]. Fraser [1, 2162, 1534, 1882, 560, 561, 1536, 1860, 2516]. Frazer [2195]. Free [423, 746, 307, 1573, 1078]. Free-Living [746]. Free-Swimming [1078]. Freeze [2200, 837]. Freeze-Core [837]. Freezing [866, 1491, 2611]. French [543, 899, 1490, 2502, 714, 6, 1267, 677, 1621, 1125, 1135, 1767, 761, 1327, 2306, 1862, 658, 870, 88, 1525, 478, 1749, 265]. Frequencies [489, 2058, 1637]. Frequency [1141, 1972, 986, 1607, 1440, 2235, 172, 2243, 1891]. Frequency-Dependent [2235]. Fresh [1134, 2097, 611]. Freshwater [2529, 895, 625, 72, 165, 505, 156, 1618, 1462, 1419, 1542, 2027, 1850, 1164, 395, 63, 1573, 1281, 1672, 2544, 538, 473, 903, 1739, 1395, 593, 1902, 1430, 387, 1807, 1900, 1236, 1638, 1002, 1697, 1111, 2135, 1896, 2542, 143, 113, 1285, 51, 431, 61, 1957, 1539, 1809, 2034, 1196, 29, 2298, 1482, 1906, 311, 1211, 1965, 2167, 714]. Freshwaters [2140, 1104]. Frond [2185]. Front [2066]. Frontal [292]. Frustules [148]. Fry [1709, 197, 1052, 859, 606, 2302, 1853, 593, 1102, 365, 2315, 926, 1145, 2241, 1754, 1569, 2119, 702, 1997, 2530]. Fuel [1019, 1481]. Full [2090]. Fullness [1173]. fulvescens [2204]. Function [2156, 1839, 560, 876, 2100, 1940, 2135, 16, 1014, 1893, 1302]. Functional [1716, 2335, 2102, 1720, 2111, 1121, 1264, 881]. Functions [91]. Fundulus [1733, 2606]. Fundy [746, 946, 1511, 1770, 909, 776, 568, 2320, 2377, 372, 679, 1801]. Funnel [122]. Further [1568]. Furunculosis [1406, 99]. fuscum [2338]. fuscum/mixtum [2338]. Future [539]. Fyke [1057].

Gadid [1051]. Gadus [1820, 2428, 1878, 2208, 2200, 190, 1474, 1821, 639, 2540, 617, 1921, 1444, 727, 274, 1619, 2236, 2591, 2427, 2199, 1325, 1604, 2130, 1448]. gaffkemia [1862]. gaffkémie [1862]. gairdneri [1459, 1668, 2107, 1186, 2254, 2253, 108, 1337, 510, 1165, 13, 1340, 1020, 1589, 1260, 1521, 1894, 2372, 2356, 1689, 2440, 133, 405, 1372, 2125, 1719, 457, 1614, 1160, 1222, 1223, 2219, 197, 897, 795, 1268, 1847, 1822, 1746, 686, 981, 765, 1645, 22, 30, 1354, 2012, 1853, 160, 955, 289, 1493, 1919, 186, 715, 710, 2359, 1688, 1829, 1830, 1505, 1195, 1675, 2060, 2362, 1708, 1312, 2342, 99, 813, 1944, 1494, 467, 913, 451, 331, 1532, 1190, 410, 181, 1234, 2334, 2324, 2168, 1749, 2011, 2120, 2361, 738, 462, 862, 1523, 340, 1067, 1213, 2447, 1414, 2096, 2348, 419]. gairdneri [1588, 2226, 319, 751, 389, 21, 265]. galilaea [2173]. galilaeus [877]. Galilee [877]. Gammaracanthus [2568]. Gammaridae [1]. Gammarus [438, 2167, 1789, 2442]. Gape [2038]. Gas [2498, 60, 1715, 1474, 2171, 2060, 2362, 1280]. Gases [1414, 2278]. Gasoline [2022]. Gaspé [6]. gaspésienne [6]. Gasterosteus [2400, 2142, 1807, 2043, 1162]. Gastric [572, 1077, 1722, 2416]. Gastropoda [2460]. Gauntlet [2349]. Gear [2462, 2465, 1167, 2427]. Gel [940]. Gene [2104, 1887, 2133, 816, 1630, 1917, 1135]. Genera [2568]. General [2423, 680, 1980, 1715, 1500, 1467, 1303, 1714, 1928, 381]. Generalist [1633]. Generalized [2453, 1868, 1178, 1706]. Generalizing [1252]. Generating [2604]. Generation [299, 2290]. generosa [612]. Genetic [1305, 1431, 2026, 2104, 1887, 508, 438, 147, 376, 730, 2092, 2371, 104, 816, 1913, 1910, 142, 159, 1878, 2553, 1729, 2216, 1701, 291, 740, 519, 515, 2044, 1444, 2595, 1190, 157, 506, 2361, 8, 493, 1694, 1885, 2101, 2035, 1194, 522, 294, 2290, 2404, 1664, 1345, 1135]. Genetics [491, 1118, 1143, 484, 1629, 805]. génétique [1135]. géniques [1135]. Genotypic [911]. Genus [2328, 2402, 2595]. Geochemical [1071, 1339]. Geochemistry [1012]. Geodock [612]. Geographic [1292, 2456, 2494, 2448, 2289, 637, 328]. Geographical [901, 658]. géographique [658]. Geography [492]. Geometry [2457]. Geophysically [2591]. Geophysically-Forced [2591]. George [1890]. Georges [672, 1620, 673, 1820, 1861, 656, 1594]. Georgia [883, 880, 1387, 2257, 14, 1882, 882, 772, 881, 884, 879, 1472, 885]. Geosmin [1160, 185]. Geothermal [982]. GF [2202]. GF/C [2202]. GF/F [2202]. gibberum [2340, 1864]. gibbosus [180, 2094, 2523, 954]. gibhosus [276]. gigas [2306, 669]. Gill [2258, 1422, 76, 289, 488, 1329, 1992, 788, 1167, 2285, 194, 389, 550]. Gill-flaring [1422]. Gills [1226, 410, 1234, 2447]. Glacial [2564, 2568]. Glaciated [475]. Glacier [2044, 1826]. Glass [2190, 2202, 312]. Glass-Fiber [2202]. Gluconeogenesis [2233]. Glucose [2258, 686]. Glutamate [21]. Glycine [707]. Glycol [2476]. Glyptocephalus [603, 393]. Gmelin [1979]. Godbout [1538]. Goitred [2616]. Golden [2512, 897, 2452]. Goldeye [2458]. Goldfish [1240, 445, 767]. golfe [1267, 1767]. Gonad [1902]. Gonadal [1187, 762, 529, 2479]. Gonadosomatic [2477]. Gonadotropin [1977, 2054]. Gonadotropin-Releasing [1977]. Gonadotropin-Releaslng [2054]. Gonadotropins [533]. Gonads [531]. Gonyaulax [2320, 902, 679]. gorbuscha [1431, 1286, 1913, 288, 1560, 1331, 2488, 705, 2163, 365, 2005, 1194, 1860]. Gorge [2327]. Governing [901]. gracilis [1051]. Gradient [1547, 959, 1582]. Gradients [111, 1725, 832]. Graduate [541]. Graebn. [744]. Grain [950]. Grand [404, 812, 1771, 397]. Granulomatosis [2046]. Graphical [597, 1858]. Gravel [196, 821, 1046, 950, 648, 2438]. Gravels [20]. Grayling [1897, 588]. Graylings [2112]. Grazing [1635, 656, 1942, 2566, 466]. Great [492, 2263, 175, 223, 510, 2169, 592, 838, 483, 206, 1913, 231, 2553, 487, 512, 2216, 153, 241, 240, 485, 705, 2616, 578, 579, 262, 211, 224, 427, 225, 1553, 259, 294, 208]. Greater [2417]. greeleyi [1924]. Green [895, 2510, 1835, 2319, 63, 2331, 1082, 1890, 1518, 1957, 743, 429, 1975, 622, 245]. Greenland [2161, 376, 699, 1081, 1623, 2354]. Grey [1912, 793, 1031]. Grilse [1435]. groenlandica [848, 918, 1309, 1031, 871, 1568]. groenlandicus [357]. Gross [321]. Ground [584, 1491, 2532]. Groundfish [1650, 2515]. Groundfishes [651]. Groundwater [2357, 1401]. Group [91, 2040]. Groups [847, 1983, 2549]. Growing [1269]. Grown [1585, 1686, 669]. Growth [1668, 2107, 1681, 1797, 1404, 1550, 177, 821, 996, 2182, 1030, 869, 1257, 1508, 1601, 1187, 2474, 316, 1998, 26, 2413, 2414, 2547, 983, 2122, 2517, 1025, 115, 342, 1776, 55, 1820, 2300, 1161, 650, 2428, 946, 1873, 1367, 1990, 464, 726, 128, 737, 2068, 1451, 1687, 1552, 1842, 1269, 664, 1029, 719, 2375, 1634, 2000, 86, 1185, 1107, 1317, 76, 1879, 1248, 116, 730, 2218, 2345, 2521, 1382, 2206, 621, 1174, 502, 1390, 1625, 636, 1164, 2389, 1612, 1379, 2138, 33, 1729, 1447, 765, 1567, 688, 2012, 2208, 2047, 2196, 1853, 1777, 1296, 373, 1945, 915, 1344, 387]. Growth [1710, 503, 773, 1837, 1449, 2317, 874, 2501, 1667, 2508, 341, 815, 296, 68, 2464, 711, 121, 1706, 2420, 585, 420, 2128, 1968, 2271, 2283, 49, 2250, 1647, 703, 1360, 2189, 321, 1803, 2094, 607, 2266, 1429, 1239, 1823, 290, 977, 2489, 1892, 2588, 1423, 1593, 2365, 2257, 1235, 914, 1523, 31, 1008, 2099, 431, 655, 172, 439, 2444, 2395, 2366, 1067, 1535, 612, 1213, 563, 707, 474, 1486, 1463, 1504, 1537, 1754, 2233, 2479, 2557, 2546, 1652, 1028, 936, 138, 2207, 193, 702, 397, 743, 917, 2502, 1985]. Growth.of [1617]. grypus [1912, 793, 1031]. Guidelines [518]. Gulf [2184, 1267, 376, 1767, 2196, 2235, 323, 1321, 441, 2428, 1511, 127, 142, 780, 2141, 306, 2609, 1318, 2066, 1441, 1358, 319]. Gull [248]. Gulland [158]. gulosus [725]. Gun [2469]. Gupeidae [1579]. Guppies [2365, 936]. Guppy [290]. Gustatory [2344, 2570]. Gut [9, 2412, 1173].

H [2570]. H. [1143]. Habitat [201, 2127, 375, 662, 1744, 1702, 2192, 887, 1746, 2230, 661, 2457, 145, 2142, 1963, 2548, 2315, 304, 2564, 843, 1463, 1219, 1454]. Habitats [2515, 854, 407]. Habits [613, 2059, 1158, 581]. Haddock [1820, 2428, 1474, 1821, 2540, 1594]. Haematractidium [109]. Haemolymph [1628]. Haemolysin [1182]. Hake [202, 4, 866, 2574, 1941, 976]. Half [913]. Half-lives [913]. Halibut [2161, 376, 2092, 1143, 1108, 699, 341, 2590, 2465, 1424, 2149]. Halichoerus [1912, 793, 1031]. Hamilton [2342]. Hampshire [2008, 1434]. Handline [694]. Handling [623, 108, 1337]. Harbor [1031]. Harbour [558, 2342]. Hard [2348]. Hardness [1340, 1688]. Hardwater [855]. hareng [6, 677, 1621]. harengus [729, 6, 677, 1621, 339, 2198, 1958, 1996, 2138, 2139, 776, 740, 1666, 1983, 2477, 1065, 585, 399, 1180, 2611, 2257, 1156, 706, 831, 563, 1139, 590, 2205, 1995, 1604, 2576, 1026, 271, 2053, 2430, 6, 677, 1621, 776, 740, 1209, 2609, 1489, 1666, 1983, 2477, 585, 399, 1180, 1583, 563, 1139, 1358, 1359, 138, 1026, 271]. Harp [357, 848, 918, 1309, 1031, 871, 1568, 1577, 1363]. Harpacticoid [1404, 1509, 717]. Harvest [1366, 2198, 1369, 1516, 2349, 2598, 297]. Harvested [1467]. Harvesting [34, 2301, 2185, 2438, 255]. Harvests [266]. Hatcheries [1492, 373]. Hatchery [177, 1985, 823, 719, 1857, 1675, 1674, 2508, 971, 363, 1805, 1483, 268, 1142, 8, 195, 1213, 1804, 2290]. Hatchery-Produced [1483]. Hatchery-Raised [1857]. Hatchery-Reared [177, 1674, 195, 1213]. Hatching [2448, 2359, 103, 2286]. Hawaii [1762, 1285]. Hawaiian [2435, 694]. HCB [467]. HCO [1234]. Head [2348, 1944]. Headwater [2325, 1895, 2313, 2114, 2115, 655, 2388]. Headwaters [947]. Healing [1275]. Health [2536]. Heart [2368, 2033]. Heat [441, 289, 1505, 1991]. Heat-Stressed [289]. Heated [565]. Heavy [592, 2326, 51, 1018, 1957, 2549, 81, 658]. Heavy-Metal [81]. Hecate [893, 867, 1604]. Held [1902, 188, 611]. Helicolenus [969]. Hematocrit [2258, 29]. Hematologic [2450]. Hematological [1223, 813]. Hematology [2278]. Hematopoietic [2476, 675, 926, 1812, 1943, 1944]. Hemitripterus [1113]. Hemocytes [2244]. Hemoglobin [445]. Hemoglobins [2451]. Henneguya [1056, 1409]. Hepatic [646, 865, 186, 738]. Hepatocytes [1468]. Herbivore [173]. Herbivores [2298]. Herbivorous [271]. herdmani [1581]. Here [260]. Hereditary [2347]. Herklotsichthys [2557]. Herring [729, 1882, 1654, 2053, 2430, 2345, 1118, 339, 2198, 674, 772, 1958, 1996, 2138, 2139, 713, 745, 1485, 2550, 776, 740, 1209, 2609, 1602, 362, 1489, 1666, 1983, 2477, 1065, 585, 399, 1180, 2532, 2611, 1583, 1156, 706, 831, 1472, 602, 563, 1139, 1358, 1359, 590, 2205, 1995, 2557, 1604, 2576, 138, 1026, 271, 1923, 6, 677, 1621]. Herrings [2257]. heteroclitus [1733, 2606]. Heterogeneity [2550, 2086]. Heterogeneous [1980]. Heterotrophic [2593, 2082]. Hexachlorobenzene [467]. Hexachlorobiphenyl [578, 579]. Hexagenia [2412, 2186]. hexyl [399]. Hg [203, 1103, 1899]. High [313, 1509, 628, 438, 2213, 1365, 1393, 1172, 2605, 1753, 1333, 1725, 83]. High-Density [1509]. High-Performance [628]. High-Pressure [313]. High-Resolution [2605]. High-Speed [2213]. Highly [398, 1127, 1906]. Hills [643]. Hiodon [2458]. Hippoglossoides [617, 2161, 376, 699]. Hippoglossus [2092, 1143, 1108, 2590, 2465, 1424, 2149]. hirundinella [2224]. hispida [1399, 270, 2212, 2549]. Histological [892, 1528]. Histology [727]. Histopathoiogy [2060]. Historical [2482, 509, 1421]. Histories [1154, 1429, 2557]. History [1146, 2036, 1882, 2535, 510, 1141, 2008, 395, 1073, 444, 512, 1266, 2273, 883, 705, 1675, 1000, 2425, 420, 654, 1516, 979, 833, 1544, 1130, 1326, 222, 2226, 1321, 1982]. Hjort [1324]. Hokkaido [2175]. Holopedium [2340, 1864]. Holophryxus [476]. Holt [1069]. homard [1767, 870]. homards [1862]. homari [1290, 1292, 1597, 1862, 1289, 1291]. Homarus [1290, 1271, 1779, 1681, 1680, 2370, 620, 1781, 1955, 1789, 1292, 946, 978, 1276, 1597, 1771, 1770, 2582, 1990, 1005, 324, 1179, 1294, 1784, 124, 1768, 1758, 1791, 2069, 672, 780, 2141, 1662, 1767, 2047, 2196, 591, 2558, 1293, 1959, 2442, 1862, 2561, 870, 873, 2189, 2188, 846, 1772, 648, 1778, 1785, 1289, 1488, 1782, 466]. Homeostatic [878]. Homing [2441, 487, 713, 2347, 1772, 1142, 2058, 1026]. Homology [305]. Homozygous [331]. Hood [814]. Hooded [1854, 2177]. Hook [694]. Hooking [2263]. Hooks [584, 2540]. Hoplonemertea [1289]. Horizontal [1811, 1323, 433, 673, 1767]. horizontale [1767]. Hormonal [529, 1449]. Hormone [1668, 2107, 2182, 1269, 1977, 2616, 2615, 2266, 611]. Hormone-induced [611]. Hormone-Sterilized [2266]. Hormones [1506, 1992, 2054]. Host [604]. Hot [647]. Hotelling [1278]. Hovvatn [1383]. Howe [1953]. Hoxie [2327]. hoyi [2311, 2497, 2166, 1872]. HPLC [1847]. huître [2306]. Hudson [1371, 471, 654, 1745, 1313, 269, 2204, 1566, 143, 949, 369, 629]. hudsonica [1581]. Human [1457, 1298]. Humic [1541, 1986, 277]. Humpback [2589, 284, 1081]. Humus [1971]. Huron [2593, 222, 511, 933, 1447, 2166, 287, 243]. Hyalella [1951, 1950, 1875, 797, 1539, 2293]. hyalina [1415]. Hybrid [346, 502, 504]. Hybridization [364, 2328, 413, 1260, 1733]. Hybridized [897]. Hybrids [988, 1151, 1241]. Hydraulic [2457]. Hydroacoustic [440, 2164, 2237]. Hydrocarbon [2132]. Hydrocarbons [636, 1683, 846, 2322, 392]. Hydrodynamic [2527]. Hydrodynamics [763, 1869]. Hydroelectric [1205, 2312]. Hydroextraction [168]. Hydrogen [2013, 2509, 2356]. Hydrographic [1159, 1453, 465, 1157]. Hydrographical [618]. Hydrography [5, 1657]. Hydrologic [973, 2170]. Hydrology [1971, 2388]. Hydrolysis [2134]. Hydromedusa [2532]. Hydromedusae [200]. Hydromedusan [1242]. hydrophila [1576]. Hydrophobic [2190]. Hydropsyche [1218]. Hydropsychid [687, 1826]. Hydropsychidae [1218]. Hynes [986]. Hyolimnion [1836]. Hypereutrophic [2331, 2202]. Hypersaline [1725]. Hypertrophic [968, 428, 429]. Hypochlorite [106]. Hypolimnetic [910, 1433, 57, 1036]. Hypolimnion [198, 1049, 1920, 1570, 45]. Hyposaline [2191]. hypothèse [2225]. Hypotheses [377]. Hypothesis [1951, 289, 258, 1324, 621]. Hypoxia [2324, 1696, 936]. Hypoxic [599].

iavaretus [2095]. Ice [1277, 1948, 909, 1371, 691, 1907, 188, 471, 24, 596, 1991, 2158, 2578]. Ice-Covered [1277, 24, 691]. Iceland [2273]. Icelandic [921, 1123, 1163, 2354]. Ichthyofauna [5]. Ichthyomyzon [1924]. Ichthyophthiriasis [751]. Ichthyoplankton [633, 2213, 756, 2289, 1220, 1050, 300]. Icta [1507]. ictaluri [1507, 2495]. Ictalurus [1110, 2105, 2499, 2381, 79, 652, 615, 566, 997, 1926, 788, 411, 421, 1230, 436]. Idaho [2022]. Identification [1550, 1305, 1431, 2026, 2247, 609, 1458, 1009, 1051, 159, 516, 1191, 1598, 752, 515, 387, 2044, 600, 969, 2477, 805, 785, 268, 493, 1979, 1241, 1922, 2612]. Identifications [2358]. Identified [399]. Identify [1128, 1278]. Identifying [1302, 1514]. Identity [708]. if [1699]. II [2414, 2397, 1011, 1930]. III [2547, 1012]. illecebrosus [1296, 35]. Illex [1296, 35]. Image [2053]. Imaging [1409]. Immersion [458, 751]. Immigration [696]. Immobilization [334]. Immune [1165, 1838]. Immunity [1406]. Immunization [551, 751]. Immunochemical [2239]. Immunocompetence [2345]. Immunoenzyme [350]. Immunofluorescent [1291]. Immunogen [751]. Immunogenic [726]. Immunogenicity [551]. Immunosorbent [1841]. Impact [1882, 2401, 2326, 656, 2027, 1974, 640, 1100, 1739, 1676, 1256, 1148, 1603, 2214, 2185, 368, 2487, 522, 818]. Impacted [2095]. Impacts [1648, 2151, 1615, 1145, 226, 2173]. Impairs [861]. Impassable [2407]. Implantation [1838]. Implicated [185]. Implication [533, 711, 1594]. Implications [2247, 514, 366, 1310, 1519, 1580, 1759, 291, 945, 2617, 68, 2590, 644, 506, 2579, 98, 497, 1728, 254, 817, 323, 1339]. Import [2170]. Importance [899, 74, 2585, 1075, 557, 2217, 934, 1530, 299, 145, 1102, 2297, 1221, 836, 2251, 2446, 2094, 1326, 1216, 917, 899]. Important [623, 66, 813, 1448, 1783]. Impoundment [1096, 1098, 1417, 1086, 1088, 1087, 2103, 1084, 1090, 1091]. Impoundments [2138]. Imprecise [1914]. Imprinting [2441, 487, 2367]. Improve [929]. Improved [596, 1402, 1058, 1864]. Improvements [1800, 635]. Improving [1776]. In-Lake [2419]. In-line [1083]. Inactivation [647]. Incidence [2005, 1007, 1862]. Including [1800]. Inclusion [1900]. Incomplete [146]. Incorporating [893]. Incorporation [359]. Increase [1183]. Increased [2230, 765, 1945, 57, 245, 988, 419]. Increases [2182, 1095, 1946, 1994, 266, 1735]. Increasing [549, 1101]. Increment [915, 1440, 1505, 1429]. Incremental [1351]. Increments [650, 1250, 1296, 1710, 703, 702]. Incubated [1379]. Incubation [197, 1060, 926, 1943, 2095, 2438, 2113, 56]. Incubators [821, 781]. Independent [2098, 298, 1722]. Independently [2210]. Index [2586, 2550, 638, 787, 2477, 1867, 1653, 131, 1503, 2418, 1324, 707, 2150, 81, 1864, 1490]. Indexing [1357, 1482]. Indian [1093, 1096, 1098, 1092, 1097, 1986, 1086, 1100, 1088, 1087, 1085, 1084, 1090, 122, 1094, 1099, 1091, 71]. Indiana [855]. Indicated [2204, 713, 1500, 1136, 1026]. indicateurs [658]. Indicator [1186, 987, 2580, 738]. Indicators [2256, 58, 75, 639, 1489, 948, 599, 658]. indice [1490]. Indices [1764, 2001, 645, 1914, 1057, 669]. Indigenous [504]. Indirect [2340, 1586, 1741]. Individual [1520, 1025, 1247, 31, 1957, 804].

Individuals [1812, 226]. Induced [2046, 2105, 2258, 2304, 2014, 865, 1206, 289, 1329, 1532, 988, 256, 1967, 178, 257, 611]. Induction [2342]. Industrial [1839, 698, 849]. Industry [1069]. inermis [1248]. Inertia [2526]. Infaunal [372]. Infected [132, 553, 647]. Infection [107, 1666, 1619]. Infections [1507]. Infectious [2476, 1009, 675, 926, 1812, 1943, 1944, 736, 553]. Infective [1564, 941]. Inference [1183]. Inferences [755]. Inferred [1730, 2159, 1360, 2289, 2058]. Inferring [2300, 1173]. infertile [2334]. Infestation [1597]. Infestations [1783]. influencant [1621]. Influence [1271, 1779, 1404, 177, 996, 1436, 583, 32, 2122, 1340, 477, 1661, 650, 2262, 2572, 542, 964, 2068, 2078, 324, 719, 2125, 2116, 1784, 1971, 76, 1698, 2551, 2244, 1541, 1850, 1027, 2131, 2405, 1184, 1734, 1485, 1322, 1297, 2260, 1821, 797, 1688, 1505, 579, 5, 1435, 971, 2617, 2548, 1890, 2420, 2117, 125, 1848, 437, 1140, 1679, 1491, 594, 2324, 694, 2011, 1643, 2363, 1899, 1386, 1556, 1237, 1957, 1995, 1982, 1176, 1644, 2413, 2414, 2547]. Influenced [1708, 853, 1679]. Influences [896, 1029, 2170, 982, 330, 1565, 1144, 1536, 1515, 2030, 19, 545, 2032]. Influencing [2510, 2522, 1895, 127, 780, 2515, 888, 224, 703, 2188, 962, 1896, 1066, 587, 1626, 1621, 1527]. Influenza [1031]. Influx [2348]. Information [1350, 891, 2588, 184, 2003, 1982]. Infrared [2433]. Ingested [151]. Ingestion [737, 2484, 2311, 2176, 2294]. Ingestion-Limited [737]. Inhabiting [2506, 318]. Inhibition [1240, 186, 1544, 103, 2233]. Initial [733, 470, 1259, 2448, 2111, 1282]. Injection [45]. Inland [503]. Inlet [1043, 179, 159, 1822, 155, 1044]. Inorganic [2357, 1964, 1140, 1193, 1176, 1649, 1211, 1525]. inorganique [1525]. inositol [2381]. Input [1116, 1748, 143]. Inputs [1197, 1371, 704, 1330]. Inseason [2194, 2604]. Insect [477, 2326, 2396, 332, 2018, 81]. Insecticide [1404, 692]. Insects [98]. Inseminated [1488]. Inshore [626, 1366, 1519, 868, 1388, 2028, 1368]. Insight [2008]. Insights [1153]. Insoluble [1186]. Instar [1710]. Instream [469, 1938]. Instrumentation [435]. Insular [839, 1895]. Intake [1040]. Intakes [1202, 1607]. Integral [453]. Integrated [252, 888, 2605, 1892, 249]. Integration [1886, 1937, 1045]. integrifolia [311]. Integrity [2150]. Intensity [202, 2078, 653, 315, 2483, 431, 1490]. Inter [1281, 1073, 2530, 2619, 1465, 2427, 1135]. Inter- [1281, 1073, 1465, 1135]. Inter-Annual [2530, 2427, 2619]. Interaction [2370, 2110, 1609, 2596, 1125, 1746, 1869, 1604, 1370, 2033, 1125, 1013]. Interactions [725, 1257, 677, 529, 52, 2355, 145, 566, 261, 2214, 2087, 1122, 451, 2011, 1386, 2051]. Interactive [165, 1481, 2591]. Interannual [1559, 1446, 2093, 1995, 1490]. interannuelle [1490]. Intercalibration [1886]. Intercept [440]. Interface [2087]. Interfaced [351]. Interflow [800]. intermedia [1504]. Intermittently [1067]. Internal [2217, 639, 1570, 1364, 938, 1117, 612]. International [264, 479, 480, 206, 520, 205, 263, 1756]. Interpretation [1214]. Interpretations [2075]. Interpreting [2001]. Interrelation [2000].

Interrelationship [2270]. Interrenal [1020, 1589, 2233]. interruptus [1765, 548]. Interspecies [1733]. Interspecific [14, 2328, 455]. Interstitial [1263, 1283]. Intertidal [1204, 1389, 911, 899]. intertidales [899]. Intervals [37]. Intervention [1603, 1578, 595]. Intoxication [2258, 688, 306]. Intra [394, 1073, 1465, 1135]. Intra-Population [1073, 394, 1135]. intra-populations [1135]. Intra-specific [1465]. Intracellular [307]. Intragravel [1728]. Intrapopulation [2146]. Intraspecific [2013, 1281, 2230, 1231, 498]. Introduced [2351, 2245, 2195, 503, 420]. Introducing [2481]. Introduction [480, 1756, 956, 2423, 1913, 744, 1629, 1200, 205, 1540, 2298, 1313]. Introductions [1660]. Introductory [1757]. Introgressive [2328, 413]. Invariant [1514]. Invasion [222]. Inverse [1973]. Invertebrate [2436, 1443, 407, 2542, 944, 372, 2298]. Invertebrates [1661, 1224, 417, 231, 643, 903, 1484, 2297, 642, 552, 858, 125, 1121, 883]. Investigated [980]. Investigating [1202]. Investigation [823, 1749, 1749]. Investigations [2444, 2445]. Investment [2340, 1955, 993, 994, 2187]. Ion [2253, 2277, 423, 2013, 2356, 1883, 1752, 2087, 1803, 2120, 1414, 2042]. Ionic [1987, 445, 1111, 1679, 1344]. Ionization [1111, 1904]. Ions [1445, 2015, 1628, 2278]. Ireland [494]. Irish [2184, 2185]. Iron [166, 1049, 937, 1541, 1956, 917]. Irradiance [2016]. Irrational [2040]. Irritants [1329]. irroratus [613, 591, 2561, 1244]. Ischnura [2019]. Island [2507, 618, 2471, 2195, 2624, 248, 2491, 2511, 1403, 1929, 1930]. Islands [1767, 1862, 870, 2435]. Isolate [464]. Isolated [2473, 2015, 135, 2348, 194]. Isolates [551, 736, 641]. Isolating [487]. Isolation [807, 1009, 132, 1195, 268, 553, 1109]. Isonychia [1381]. Isopleths [1467]. Isopod [476]. Isotope [2575, 1461, 945, 2252]. Isozyme [147, 105, 210]. Isozymes [1632]. Israel [877]. israelensis [1709, 2577]. Issue [42]. issues [1135]. Isurus [977, 581]. Italy [2274]. Iteroparity [416]. IV [852, 1013]. Ivoirian [1867]. Ivory [2565].

James [1395, 654, 143, 1745]. Japan [2175, 2595]. Japanese [996]. Japanese-Style [996]. japonica [744]. japonicus [2154, 1324]. Jasus [1773, 1761]. Jellyfish [2231]. Jenkins [1940]. Jevenile [1005]. Jolly [434]. Jonah [1778]. Jordanella [101, 1676, 1580]. jordani [1933]. Juvenile [1459, 2162, 1337, 814, 219, 2401, 1422, 583, 333, 1764, 1141, 1521, 1990, 476, 952, 845, 719, 2375, 1155, 2232, 796, 2218, 2345, 2521, 1850, 2131, 288, 1331, 1571, 1852, 1662, 653, 1746, 2515, 2047, 1592, 1715, 1485, 2550, 776, 2200, 1293, 566, 2574, 1994, 2601, 1688, 1959, 1900, 1674, 561, 1909, 1963, 2163, 391, 2548, 1832, 2564, 468, 1360, 1494, 2, 696, 695, 2305, 648, 1208, 328, 843, 1072, 1870, 1617, 1696, 2427, 2531, 2383, 2528, 2533, 817, 1028, 430, 936, 1903, 193, 2275, 1929, 1930, 389, 1723, 2439, 1692]. Juveniles [1842, 2207].

Kalama [1675]. Katsuwonus [1952]. Keeper [821, 996]. Kelp [613, 1047]. Kelts [2441, 1902]. Kenai [1335]. Keogh [2491, 2226, 2584]. Kepone [412].

Keta [2530, 821, 996, 814, 1305, 1466, 1286, 1534, 1851, 2026, 2175, 2500, 1842, 1850, 1331, 1571, 653, 593, 1102, 781, 107, 2163, 938, 2438, 1194, 1028, 2275]. Ketone [2473]. Key [208]. Keys [723, 1777]. kHz [2066, 2605, 87]. Kidney [363, 1841]. Killer [1894]. Kills [271]. Kinds [858]. Kinematics [548]. Kinetic [361]. Kinetics [72, 165, 860, 737, 2074, 315, 467, 321, 1543]. King [2471]. Kinneret [877]. kisutch [1308, 1422, 872, 1985, 583, 342, 1396, 2524, 823, 2375, 1155, 2232, 2218, 1977, 2345, 1331, 1571, 606, 1742, 2151, 2530, 2245, 2491, 2265, 2616, 176, 2508, 1838, 391, 2548, 904, 1565, 1805, 2266, 1483, 891, 2505, 696, 695, 1208, 2181, 1870, 352, 1145, 2438, 1804, 2384, 611, 102, 2531, 2383, 1495, 1496, 2528, 817, 2054, 1917, 335, 1038]. Knowns [263]. Kodiak [2195]. Kokanee [2371, 2266]. Kootenay [764, 800]. Kraft [2272, 32, 386]. Krill [875].

L. [1390, 2094, 2523, 276]. Labeled [151, 97, 1425]. Labeling [303, 456]. Labelling [1296]. Laboratory [1681, 14, 1976, 2303, 2403, 2055, 1439, 2543, 692, 642, 2425, 1494, 2197, 2597, 1617, 2016, 1769, 659, 2207]. Laboratory-Reared [2055]. Labour [2520]. Labrador [603, 2146, 1726, 1971, 1566, 925, 496, 949]. Labyrinth [1457]. Lac [2204, 187, 1119, 2479]. Lack [1680, 1998, 2381, 1878, 567, 778]. lacs [43]. Lactate [1632]. Lactating [1912]. Lacustrine [1059, 2142, 360, 2043, 2266]. lacustris [1618, 325, 1270, 2167]. Lagoon [1001]. lagoons [870]. lagunes [870]. Lahontan [1461]. Lake [2482, 2229, 910, 89, 2529, 2473, 929, 1205, 2401, 1601, 2276, 11, 2255, 1436, 1093, 987, 626, 1096, 1098, 1319, 60, 861, 1033, 509, 968, 807, 1249, 1233, 511, 860, 2509, 1610, 2140, 1349, 2552, 1092, 2078, 2077, 2468, 508, 2519, 1574, 2116, 1306, 1460, 973, 1129, 1461, 110, 764, 822, 1048, 693, 1097, 1027, 2319, 2010, 957, 2008, 507, 28, 2337, 1986, 197, 1052, 1987, 2622, 1806, 2544, 1690, 1822, 33, 58, 2264, 2186, 933, 1447, 1311, 1814, 52, 70, 59, 728, 661, 2204, 1341, 2400, 512, 2216, 105, 277, 1734, 800, 622, 1266, 161, 2166, 712]. Lake [604, 1384, 1410, 282, 240, 855, 2171, 2195, 2331, 246, 777, 1549, 2352, 287, 1450, 1624, 1602, 2263, 503, 773, 318, 1667, 1295, 2316, 815, 1023, 2617, 1865, 2408, 1528, 68, 1262, 2464, 245, 948, 654, 2176, 1673, 428, 1881, 2121, 47, 1085, 2446, 2, 2005, 1401, 1570, 2145, 671, 183, 2621, 1045, 1200, 635, 768, 1054, 784, 247, 1651, 1515, 1334, 98, 184, 38, 1989, 80, 2390, 2090, 122, 1094, 1577, 1503, 2483, 1231, 1510, 1695, 51, 1363, 602, 2419, 2478, 1518, 243, 1022, 1539, 844, 1330, 1343, 2531, 248, 1815, 406, 564, 1543, 2350, 608, 2160, 1936]. Lake [2546, 1012, 1215, 1826, 255, 2227, 415, 244, 1903, 1176, 1348, 1649, 449, 1091, 917, 1864, 2393, 257, 185, 2255, 801, 1175, 201, 509, 2123, 511, 65, 7, 242, 2217, 375, 662, 508, 660, 2311, 1306, 2223, 1405, 1519, 1530, 1879, 1880, 1625, 1475, 2496, 2497, 708, 507, 877, 2081, 610, 734, 1086, 1100, 1088, 1087, 933, 23, 622, 1388, 1547, 515, 945, 2261, 246, 2091, 2079, 2070, 2073, 2074, 2083, 1549, 773, 826, 2087, 448, 922, 1039, 1890, 1262, 227, 130, 315, 1647, 1544, 2086, 974, 1084, 467, 2088, 1090, 644, 630, 959, 2076, 2075, 2082, 2385].

Lake [1193, 2080, 247, 1623, 2382, 2090, 111, 1556, 2109, 2242, 48, 229, 602, 2071, 243, 565, 2072, 844, 2042, 248, 1225, 2085, 2084, 267, 513, 409, 358, 2051, 739, 1099, 244, 1034, 1383, 834, 300, 1582, 2089]. Lakes [2486, 1871, 1277, 492, 1948, 1671, 963, 1470, 1445, 168, 1095, 2126, 1021, 2397, 1606, 344, 951, 1075, 2321, 2510, 1811, 1855, 198, 956, 964, 2353, 759, 1707, 1920, 1542, 67, 2394, 2174, 2256, 2563, 1719, 2387, 2386, 1737, 2032, 1905, 1726, 1895, 2147, 2594, 1971, 829, 2396, 66, 206, 64, 2392, 1382, 1951, 502, 1433, 856, 2389, 2410, 1702, 1068, 59, 1071, 2336, 1287, 53, 691, 2297, 878, 1595, 1819, 2313, 1946, 1221, 1148, 1660, 2616, 62, 960, 503, 773, 2025, 448, 24, 2316, 349, 2503, 203, 2408, 889, 1262, 420, 2250, 429, 199, 1487, 2292]. Lakes [451, 1401, 1911, 1931, 1896, 1089, 2542, 2504, 758, 916, 2202, 755, 1515, 1934, 2259, 1117, 1352, 2363, 1106, 48, 44, 2395, 2419, 2460, 1600, 966, 1813, 1018, 958, 2388, 2293, 1434, 1725, 1330, 1975, 1284, 2017, 1935, 42, 2203, 1469, 1314, 1315, 1642, 2179, 2364, 587, 954, 1908, 275, 1599, 961, 56, 1703, 510, 2169, 592, 838, 483, 2593, 1913, 231, 2553, 487, 512, 2216, 153, 241, 240, 485, 705, 578, 579, 2263, 262, 211, 224, 175, 427, 223, 225, 1553, 222, 259, 294, 208, 1991]. Lakeward [1592]. lalandii [1761]. Lamellae [2285]. Laminaria [1250, 2157]. LAMP [2385]. Lampetra [232, 625, 1882, 2407, 2097, 1048, 236, 239, 1874, 238, 237]. Lamprey [264, 232, 233, 625, 2036, 2407, 2255, 505, 242, 2097, 230, 1048, 206, 231, 252, 228, 251, 253, 1191, 236, 604, 240, 258, 210, 515, 260, 261, 262, 239, 1924, 250, 245, 227, 213, 224, 2005, 223, 225, 247, 1874, 249, 229, 238, 205, 222, 256, 243, 248, 1815, 259, 226, 235, 237, 1590, 255, 263, 244, 1345, 257, 216]. Lamprey-Induced [256, 257]. Lampreys [209, 219, 2583, 220, 458, 241, 221, 217, 234, 211, 212, 723, 218, 229, 254, 235]. Lampricide [628, 1224, 1831, 226]. Lampricides [1691]. Lampsilis [1567]. LAN [13]. Lance [2595, 397, 812, 2152]. Land [167, 184]. Land-Locked [167]. Landing [1448]. Landings [2006, 1578]. Landlocked [1187, 762]. Large [2403, 956, 1552, 2190, 2245, 800, 188, 1148, 2571, 1615, 1386, 111, 1192, 2298, 1782, 1261]. Large-Scale [956, 1552, 1615]. Large-Volume [2190, 1192]. Largemouth [1648, 1230, 506, 2610, 1596]. Larger [2365]. Larvae [729, 996, 2428, 1179, 1051, 676, 1076, 866, 2098, 2586, 2141, 734, 1080, 674, 888, 1636, 236, 2573, 2402, 2492, 1218, 1655, 2294, 2577, 214, 1416, 2532, 1288, 1423, 1979, 1482, 2227, 2513, 2207, 677, 1621, 1767]. Larval [2038, 1717, 1030, 1820, 2053, 2430, 2522, 1005, 476, 952, 2002, 278, 1856, 1784, 126, 1768, 1758, 932, 1974, 314, 657, 621, 1478, 672, 2560, 728, 1849, 2196, 1954, 1209, 806, 1575, 1602, 2288, 2420, 213, 1429, 1848, 2064, 1793, 1761, 2305, 212, 1765, 2588, 1872, 2257, 1572, 1769, 2427, 1869, 2557, 1652, 138, 1824, 2231, 851]. larves [677, 1621, 1767]. Larvicides [251]. Late [1229, 2624]. Latent [1987]. Lateral [985]. Latitude [2493]. Latitudinal [2032, 395]. l’Atlantique [6, 677]. latroscan [1904]. Laurent [6, 1267, 677, 1767]. Laurentian [492]. lavaretus [2223, 239]. Law [823]. Lawrence [74, 6, 819, 2360, 1416, 1288, 441, 2184, 1267, 677, 376, 127, 780, 1767, 2196, 2609, 1318, 1441, 2235, 323, 1321, 319]. lawrencianus [438]. Layer [1390, 1904, 1216]. Layers [627, 2511]. Laying [2492, 1064]. LDH [331]. Leachate [295]. Leachates [1834]. Leaching [311]. Lead [110, 764, 822, 22, 688, 1354, 1947, 1956]. Lead-210 [110, 822]. Leaf [295, 1942, 15, 944, 2493]. Learned [2351]. Learning [1347, 2368]. Least [327, 2453, 318, 164]. Least-Squares [327]. Leaved [1813]. Leeches [1492]. Legs [324]. Leidy [838]. Leiostomus [942]. Length [1286, 336, 2037, 2300, 327, 128, 1419, 2000, 1868, 866, 2206, 1250, 1485, 915, 794, 225, 2099, 172, 409, 2130]. Length-at-Maturity [2130]. Length-Based [1868]. Length/Aperture [409]. Lengths [1638, 1663, 2584]. Lentic [1874]. Lepomis [779, 2014, 1648, 745, 180, 2094, 2599, 2523, 276, 954]. Leptophlebia [2338, 1504]. leptopus [204]. Lesions [727]. Leslie [146, 2062, 298]. Lessons [1615]. Lethal [1179, 1365]. Lethality [1340, 2125, 2506]. Letter [2603]. Letters [118, 170, 39, 285, 301, 337, 383, 460, 574, 598, 666, 683, 716, 731, 789, 808, 827, 906, 930, 1015, 1114, 1131, 1169, 1246, 1272, 1299, 1397, 1412, 1640, 1684, 1711, 1755, 1817, 1876, 1901, 1949, 2023, 2041, 2136, 2268, 2309, 2374, 2443, 2490, 136]. Letters/Lettres [39, 285]. lettre [2603]. Lettres [118, 170, 39, 285, 301, 337, 383, 460, 574, 598, 666, 683, 716, 731, 789, 808, 827, 906, 930, 1015, 1114, 1131, 1169, 1246, 1272, 1299, 1397, 1412, 1640, 1684, 1711, 1755, 1817, 1876, 1901, 1949, 2023, 2041, 2136, 2268, 2309, 2374, 2443, 2490, 136]. leucas [2592, 2020, 2539, 1333, 1745, 2450]. Leucine [13]. Leucocytes [1627]. Leuroglossus [1387]. leurs [714]. Level [664, 1027, 2222, 686, 2171, 787, 2195, 2536, 2322]. Levels [1550, 108, 872, 2498, 1095, 1606, 2456, 1129, 1113, 615, 53, 1734, 2183, 445, 1318, 1566, 1921, 2250, 630, 2054, 954]. LHRH [762]. Life [1146, 2036, 1882, 2535, 510, 1141, 1154, 356, 1950, 919, 395, 653, 1073, 2506, 444, 512, 1266, 2273, 705, 2091, 1675, 1000, 2425, 420, 654, 1516, 979, 833, 1239, 2093, 1183, 1130, 1326, 2383, 2226, 1321, 1982]. Life-History [2425]. Life-stage [919]. Life-Table [1183]. Light [2122, 2262, 2078, 1784, 1460, 2213, 1356, 1297, 57, 315, 1544, 431, 2031, 1518, 1304, 2270, 1991, 1176, 56]. Light-Saturation [1356]. Lightly [1719]. Lights [1438]. Like [1411]. Likelihood [2485, 1888, 2130, 1922]. Lilljeborg [1635]. Lily [1249]. Limanda [1861]. limbata [2412, 2186]. Limestone [197, 1735]. Liming [2229, 2396, 2389, 2390, 2395, 1383]. Limitation [1717, 1873, 860, 287, 889, 474, 1036]. Limitations [1968, 1426]. Limited [2517, 737, 1842, 1390, 2178, 1153, 582, 1524]. Limiting [1539, 1696]. Limits [116, 374, 2418]. Limnetic [1723, 1557, 2396, 2391, 463, 2017]. Limnocalanus [1811]. Limnocorral [1986]. Limnocorrals [1256, 1255]. Limnological [1313]. Limnology [1017, 1642]. limosa [2460]. Line [115, 584, 139, 440, 2539, 750, 1083, 292]. Linear [1857, 1514]. Lines [411, 612]. linked [1841]. Lipid [2499, 2340, 1475, 2497, 903, 1505, 1904]. Lipids [25, 359]. Liquid [313, 628, 1709, 1172]. Liquid-Chromatographic [628]. Litter [15]. Littoral [1319, 1408, 2353, 1999, 1384, 638, 2259, 1094, 1215]. Littoral-Dwelling [1384]. Live [26, 955, 1483]. Liver [2105, 2473, 1037, 1847, 2342, 1893, 21]. Livers [203]. lives [913]. Living [2272, 665, 746, 353]. lncius [2148]. lnduced [1521]. Load [1570, 2072]. Loading [1405, 1547, 1528, 1117, 2109, 1421]. Loadings [1595, 1819, 1678]. Lobster [1790, 1290, 1271, 1779, 620, 1781, 1955, 1789, 1773, 1292, 1776, 978, 1597, 1756, 1764, 1786, 1990, 324, 1179, 124, 1758, 1791, 2228, 2069, 1775, 672, 780, 134, 2047, 2196, 2558, 1293, 2442, 1762, 1774, 2188, 1763, 1766, 1761, 2435, 1765, 1792, 1785, 1289, 1291, 466, 1767, 870]. Lobsters [1680, 1776, 946, 1276, 1771, 1770, 2582, 845, 1294, 1784, 1768, 1662, 1777, 591, 1959, 1760, 1501, 2561, 873, 2189, 846, 1772, 648, 1778, 1488, 1782, 2370, 1862]. Local [728]. Localization [1291]. Location [454]. Locations [1202]. Loch [239]. Loci [210]. Locked [167]. Locomotor [1695]. Lodge [2022]. Log [1857, 1463]. Log-Linear [1857]. Logging [2333, 820, 2151, 304, 1692, 140, 1145, 972, 2528, 817]. Logistic [163]. Lohi [834]. Loligo [1966, 605]. Lomond [239]. Long [2253, 872, 2303, 2050, 2484, 1202, 1624, 2283, 2284, 1881, 1802, 149, 1556, 2285, 2225, 815]. Long-Term [2253, 872, 2303, 2050, 2484, 1202, 1624, 2283, 2284, 1881, 1802, 149, 1556, 2285]. Longear [180]. Longevity [941]. longicruris [1250]. Longitudinal [690]. Longline [2465]. lonoregulation [1493]. Look [2066]. lors [265]. Loss [434, 1597, 1691, 2186, 1845, 2307, 1491, 2042]. Losses [268, 46]. Lotic [1456, 2122, 2517, 1217, 1839, 174, 1743, 2455, 2579, 1892, 81]. lourds [658]. Low [900, 2254, 2277, 1606, 1609, 457, 1950, 1987, 1464, 1607, 1060, 2336, 1752, 2506, 2171, 1344, 50, 2281, 2283, 2284, 2536, 103, 755, 2120, 935, 1106, 2286, 890, 1694, 2460, 1893, 2282, 720, 133]. Low-Alkalinity [2336, 1106]. Low-Frequency [1607]. Low-Level [2171, 2536]. Lower [175, 223]. Lowland [1407]. lucius [1098, 2436, 2552, 187, 203, 1885, 1013, 1477]. luetkeana [311]. Lumped [2301]. Lumped-Parameter [2301]. lurus [1507]. lutrensis [573].

m [2302, 1054]. même [714]. mêmes [1135]. machrochirus [745]. Mackerel [109, 399, 2064, 2197, 2154, 1324, 1321]. Macoma [911, 432]. Macro [427]. Macro- [427]. Macroalgal [896]. Macrobenthic [2618]. Macrobenthos [1068, 1881, 1258, 1091]. macrocephalus [1998, 1878, 2296, 1604, 2130, 1448]. macrochirus [779, 2014, 1648, 2094, 2599, 276]. Macrocystis [311]. Macrofaunal [1801, 2426]. Macroinverfebrate [1407]. Macroinvertebrate [741, 174, 934, 1263, 469, 2544, 2343, 1942, 840, 663, 1058]. Macroinvertebrates [1217, 2469, 1475, 2124, 1430, 982, 140]. Macrophyte [1408, 1336, 2211, 1889, 1813]. Macrophytes [951, 1075, 2262, 2321, 557, 2021, 2409, 1480, 1737, 2116]. macrostoma [625]. Macrozoarces [617]. Macrozooplankton [2002, 1432]. macrurus [1811]. maculatum [2013]. Made [143]. Madeleine [2047, 2196, 1862, 1767, 1862, 870]. maderensis [2565]. maenas [2033]. magellanicus [591, 799, 2252, 1979, 1828]. magister [802, 1025, 1788, 1362, 1605, 279, 668, 1578, 2556, 2555]. magna [898, 307, 1464, 697, 757]. Magnetic [938]. Magnitude [2319]. main [1621]. Maine [2428, 1511, 2430, 2228, 2410, 780, 2141, 306, 1358]. Maintenance [166]. Maja [2442]. Major [1519, 1048, 1541, 1752, 1023, 1448]. Making [2187]. Mako [977, 581]. Malacostraca [1555]. Maldanidae [370]. Male [1584, 1716, 839, 2335, 2470, 2317, 1944, 1656, 2556, 2533]. Mallotus [94, 2572, 677, 1621, 1974, 465, 314, 657, 621, 919, 1174, 995, 1209, 1157, 323]. Mallows [2210]. malma [1346, 2246, 1259, 2265]. Mammal [2398]. Mammalian [2054]. Mammals [1116, 317]. Man [143, 522]. Man-Made [143]. Management [483, 2520, 514, 366, 1310, 2481, 469, 117, 2040, 713, 518, 1701, 291, 485, 484, 521, 891, 506, 1298, 871, 1047, 497, 249, 2329, 1133, 378, 308, 523, 824, 1941, 254, 494, 2153, 2308, 2598]. Managerial [1418]. Managing [1981, 2518, 689, 1594, 1418, 522]. Manan [1771]. Mandibular [875]. Manganese [948]. Manipulation [150, 762, 1263, 152]. Manipulations [1939, 145, 1487]. Manitoba [2297, 2103, 1096, 1095, 1098, 2552, 1092, 1097, 1027, 1986, 2186, 203, 1089, 122, 1094, 2487, 1091, 185]. Manitounuk [1371, 471]. marginatus [1762, 2435]. Margins [2243]. Marine [2398, 1116, 1953, 1404, 814, 2554, 1134, 2304, 2303, 2055, 2559, 2459, 317, 2006, 1185, 1113, 9, 1625, 1981, 288, 299, 1356, 145, 825, 2422, 1527, 642, 296, 552, 805, 1565, 1803, 36, 157, 718, 2332, 2239, 2591, 1400, 353, 102, 1158, 452, 1809, 2039, 401, 1066, 152, 835, 311, 359, 2344, 883, 426]. marinus [233, 2255, 505, 242, 230, 231, 252, 458, 228, 251, 253, 1191, 241, 604, 240, 258, 210, 515, 260, 261, 262, 250, 245, 227, 224, 2005, 223, 225, 247, 249, 229, 222, 243, 248, 1815, 259, 254, 226, 1590, 244, 1345]. Maritime [1898, 104]. Maritimes [978, 1199]. Mark [434, 848, 1857, 1858, 341, 920, 2411]. Marked [814, 1961, 73]. Markers [2612]. Markets [1153]. Marking [904, 1180, 442, 1997]. Marks [240]. Marsh [561, 1696]. Marshes [1747, 1748]. Mass [2267, 937, 657, 1883, 2496, 75, 1962, 850]. Massachusetts [1967]. Masses [682]. massive [408]. Matamec [1125]. Match [621, 1139]. Match/Mismatch [621]. Matched [1514]. Material [1942]. Materials [896, 1541, 516, 312, 2089]. Maternal [2340]. Matings [874]. Matrix [146, 398, 781, 2377]. Mats [1646]. Matter [923, 1833, 2467, 55, 1332, 1661, 2325, 1389, 1986, 2170, 277, 155, 2086, 1441, 1704]. Maturation [1680, 1996, 2138, 2139, 1777, 343, 1902, 1181, 1656, 2099, 563, 2576]. Maturational [1694]. Mature [1681, 1584, 2470, 2289]. Matured [1611]. Maturity [583, 1716, 124, 2335, 2405, 1000, 2477, 1232, 1446, 1536, 290, 921, 2354, 192, 308, 1151, 2130]. Mauritania [1490]. Mauritanie [1490]. Maximum [2215, 631, 664, 2032, 2232, 2485, 1721, 686, 1888, 130, 784, 1282, 2160, 2130, 1922]. Maximum-Likelihood [1922]. Mayflies [742]. Mayfly [985, 1918, 2338, 2412, 2286, 1504]. McGregor [947]. McLay [1361]. Meadow [942]. Meal [795, 1645]. Mean [1538, 2267, 1218]. Meaningful [1124]. Means [316]. Measure [597, 387, 454]. Measured [970, 937, 2077, 2257]. Measurement [2038, 11, 2110, 110, 822, 857, 1958, 160, 638, 380, 1663, 1613, 1904, 2063, 596, 2003, 379, 88]. Measurements [1128, 91, 1670, 1458, 2325, 1214, 69, 1019, 1886, 777, 262, 826, 351, 711, 87, 840, 1908, 2411]. Measures [2127, 1126, 246, 1482]. Measuring [459, 1707, 2057, 1323, 400, 1753, 803, 2180]. Meat [134]. Mechanical [2287]. Mechanics [634]. Mechanism [2500, 1230, 2344]. Mechanisms [989, 2459, 295, 1541, 117, 487, 1688, 2288, 1145]. Media [1164, 1698]. Mediated [678, 1033, 2428, 1415, 2495]. Medium [858, 1109]. Medusae [729]. megalotis [180]. Meganyctiphanes [1248, 478]. Megaptera [2589, 284, 1081]. Meigen [1482]. Meio [427]. Meio-benthos [427]. Meiobenthic [1509, 717]. Meiofauna [353]. mélangé [1621]. Melanogrammus [1820, 2428, 1474, 1821, 2540, 1594]. Meltwater [1470]. Membrane [1844]. Memory [2287]. Memphremagog [945, 630, 959]. Menhaden [1418]. Menidia [1563]. mentella [1598, 925, 1455, 404, 747]. Menten [1543]. mer [543]. mercure [1525]. Mercury [2486, 1095, 775, 1734, 2297, 2103, 773, 448, 2408, 1678, 766, 1334, 113, 1010, 1011, 1012, 1013, 1588, 954, 720, 1908, 1525, 1014]. Mercury-Contaminated [113]. Merganser [1929, 1930, 1403, 1929, 1930]. Mergansers [1403]. Mergus [1403, 1929, 1930]. Meristic [2247, 603, 94, 1063, 2373, 747, 2152]. Meristics [1065]. Merluccius [202, 4, 866, 976, 2574, 824, 1941]. Meromictic [1806, 2160]. Meromixis [166]. Meseta [2624]. Mesh [77, 1485, 427, 300]. Mesopelagic [2600]. Mesoscale [2591, 2511, 292]. Mesotrophic [1319, 2202]. mesure [88]. Metabolic [1732, 1468, 112]. Metabolism [1456, 2105, 2499, 2401, 156, 2178, 686, 1082, 605, 2587, 1893, 835, 1906, 1965]. Metabolite [401, 1066]. Metabolites [1201]. Metal [1871, 1671, 423, 1249, 2410, 52, 1580, 53, 1595, 27, 2402, 203, 1127, 738, 1957, 545, 81, 275]. Metal-Contaminated [275]. Metalimnetic [855]. Metallothionein [738]. Metallothioneins [2239]. Metals [592, 72, 165, 1497, 2326, 2386, 643, 2573, 1510, 51, 1018, 1258, 1035, 1196, 2549, 658]. Metamorphosed [2097, 226, 208]. Metamorphosing [212]. Metamorphosis [2055, 892, 594, 225, 1652, 216]. métaux [658]. Metazoan [947]. Meteorological [465, 1157]. Methanesulfonate [615]. Methanol [1251]. Methemoglobin [615]. Method [1833, 418, 562, 11, 459, 106, 603, 1161, 681, 1699, 1707, 2453, 628, 1868, 986, 1323, 2378, 252, 2190, 614, 903, 691, 825, 1112, 2369, 1172, 2172, 1846, 2062, 1962, 158, 45, 1003, 1024, 162, 2350, 804, 1750, 543, 714]. méthode [543, 714]. Methodological [1105]. Methodology [584, 1351, 2130]. Methods [146, 168, 2249, 63, 2192, 117, 1356, 516, 1529, 1040, 1581, 1420, 1921, 1655, 571, 977, 637, 1531, 924, 1938, 280, 61, 426]. Methoxychlor [1676, 1849, 2487]. Methyl [1734, 2103, 1334, 720, 1908]. Methylation [895, 1201, 359]. Methylisoborneol [1926, 185]. Methylmercury [2498, 2621, 462, 862]. Methyltestosterone [2372, 1269, 382]. Methyltin [1835]. Mg [2167]. Michaelis [1543]. Michigan [222, 201, 509, 2123, 2593, 2217, 375, 662, 1306, 1519, 1879, 1880, 450, 1475, 2496, 2497, 734, 728, 105, 23, 622, 1388, 1547, 2261, 826, 1667, 922, 1039, 130, 1881, 247, 1651, 1556, 2109, 565, 844, 564, 358, 2051, 244, 1034, 300]. Micro [1432]. Micro- [1432]. Microalgae [1371, 1803]. Microanalysis [592]. Microanalytic [387]. Microbial [1953, 55, 1839, 1545, 1295, 2121, 671, 1942, 2587]. Microbiological [2592]. Microcomputer [1753]. Microcontaminants [326]. Microcosms [1430]. Microcrustacean [1677]. Microdetritus [1423]. Microdistribution [1918, 934, 1626]. Microfossils [1730, 2519, 2008]. Microgadus [1008]. Microhabitat [2341, 1832, 1655, 1072]. Microhabitat-Preference [1655]. Microlayers [2190]. Micrometer [2038]. Micronekton [627]. Microorganisms [2103, 1543]. microps [1527]. Micropterus [2486, 1609, 1648, 2408, 1230, 506, 2610, 1596]. Microquantitation [903]. Microscope [2287]. Microscopy [470, 2083, 350]. Microstructore [2053]. Microstructure [1373, 600, 1360, 2045, 2557, 1028]. Microzooplankton [463]. Mid [2433, 2493]. Mid-Infrared [2433]. Mid-Latitude [2493]. Middle [1288, 677]. Midlake [2084]. Midwestern [782]. Migrant [861]. Migrating [2162, 2491, 2541, 1718, 2584, 392, 2275]. Migration [1242, 2474, 558, 394, 1858, 1732, 1560, 2291, 2431, 23, 1489, 2159, 1772, 328, 1870, 2224, 1738, 1929, 2516, 389, 1611]. Migrations [1866, 510, 1811, 932, 395, 2488, 1592, 1322, 870, 1177, 2236]. Migratory [1771, 230]. milieu [899, 88]. milieux [2502]. Milk [1912]. Mill [2272, 32]. Miller [2213]. Milling [1865]. Mills [386]. Mimic [868]. Mine [1136]. Mineral [1953, 414]. Mineral-Based [1953]. Minerals [2596]. Minimum [823, 1101, 2590, 401, 1066]. Mining [1865, 1897, 1196]. Minnesota [742, 1022]. Minnow [2100, 849, 1032]. Minnows [2417, 84, 1663]. minutus [901, 1635]. Miramichi [329]. Mirex [1894, 2541, 1307, 1521, 412]. Mismatch [1139, 621]. Mississippi [1926]. Mite [735]. Mitigate [2389]. Mitigation [2385]. Mitochondria [2473, 1037]. Mitochondrial [2535, 1137, 2328, 2169, 1733, 2461, 2337, 2608, 1970]. Mixed [932, 1063, 1338, 1888, 1122, 1512, 2295, 1278, 1352, 804, 2612, 1621]. Mixed-Species [1122, 1512, 2295, 804]. Mixed-Stock [1278, 2612]. Mixing [1470, 960, 2310, 867, 1022]. mixtum [2338]. Mixture [353, 1922]. Mixtures [165, 1580, 2167]. Mobilization [51]. Mode [872, 2512]. Model [1254, 2429, 2300, 2428, 557, 1635, 1918, 144, 2583, 1004, 34, 2220, 438, 1608, 730, 1631, 893, 1868, 2481, 1625, 273, 398, 597, 556, 624, 722, 2543, 70, 1915, 1980, 373, 1715, 1701, 1070, 2472, 258, 975, 1361, 1467, 330, 815, 2477, 99, 1077, 1468, 2577, 1937, 607, 1848, 157, 718, 2048, 920, 1424, 1914, 914, 1418, 1577, 2109, 439, 864, 1714, 1928, 2444, 2445, 706, 1363, 412, 1133, 2555, 1194, 844, 1069, 824, 2028, 326, 1922, 2502, 37, 158, 164]. modèle [2502]. Modeling [128, 760, 1940, 1365, 184, 1941]. Modelling [1277, 1787, 2029]. Models [1550, 1520, 2106, 2585, 2520, 2004, 147, 1798, 2228, 2481, 469, 1857, 117, 1356, 2525, 2355, 1940, 2301, 1378, 279, 1615, 163, 2128, 2294, 1968, 1516, 298, 181, 2475, 2489, 1934, 1989, 2600, 769, 1252, 1935, 2167]. Modes [366]. Modification [872, 903, 1525, 2150, 1893, 1525]. Modifications [145].

Modified [95, 1349, 1849, 837, 655]. Modifiers [1235]. Modifies [862]. Modify [1430]. Modifying [405]. Modulated [1919]. Modulating [2499]. Modulation [527]. Molecular [1033, 490, 2117, 2567, 83]. Mollusc [2553, 2239, 1196]. Mollusca [1106, 911]. Mollusks [2554, 704]. Molt [1716]. Molting [2580, 2189]. Molybdate [564]. Molybdenum [89, 61]. Moments [2300]. Monimolimnion [166]. Monitor [1083]. Monitored [2519]. Monitoring [1972, 2124, 750]. Mono [2439]. Mono- [2439]. monoceros [1333]. Monoclonal [1967]. Monocyclic [391]. Monodon [1333]. Monomictic [671]. Monooxygenase [865]. Monotypic [2602]. Monoxide [2362]. Montana [1826, 2044]. Montane [1719]. Monte [2111, 164]. Monthly [1625]. mordax [677, 1621, 847, 1602, 1416, 2093, 1288]. morhua [1820, 2428, 2208, 2200, 190, 1474, 1821, 639, 2540, 617, 1921, 1444, 727, 274, 1619, 2236, 2591, 2427, 2199]. Morone [1004, 388, 1184, 1172, 1312, 2366, 19, 369, 629]. Morphoedaphic [2418]. Morphological [2568, 1310, 289, 1598, 2043, 2214, 875, 294, 1482, 2227, 2513]. Morphology [1851, 1141, 512, 328, 2531, 1495, 1289, 216]. Morphometric [2161, 1063, 2373, 2447, 2233, 2285, 1152]. Morphometrics [2274, 1065]. Morphometries [35]. Morphometry [198, 2116, 309]. Mortality [1790, 1243, 814, 1079, 1161, 783, 952, 760, 2002, 664, 1029, 465, 1174, 1625, 272, 1379, 643, 653, 2614, 2263, 2009, 2501, 2380, 1280, 245, 926, 1512, 833, 2, 1147, 2045, 2597, 247, 1872, 2556, 256, 29, 936, 1903, 257]. Mortostilifera [1289]. Mosquitoes [2560]. Moss [2184, 2185]. Motion [16]. Motor [2287]. Mountain [310, 2343, 1924, 2395, 1671, 2315, 1401, 15]. Mountains [2010, 304, 2388, 504, 1574]. Mouth [2038]. Movement [868, 606, 2230, 2488, 1177]. Movements [1276, 1771, 2538, 1362, 1775, 1620, 1794, 2245, 776, 503, 904, 853, 2149]. moyen [677]. Mud [1567, 1926]. Muddy [1509, 185]. Mudflat [1204, 178]. Muds [620, 324]. Mulispecies [2475]. Multi [472, 2520, 1435]. Multi-Objective [2520]. Multi-Sea-Winter [1435]. Multi-species [472]. Multiarea [2475]. Multiattribute [1188]. Multicohort [1069]. Multidisciplinary [1809]. Multifactorial [980]. Multifunctionalism [1633]. Multilocus [730]. Multiplate [741]. Multiple [1417, 2451, 93, 1782, 830]. Multiple-Impoundment [1417]. Multiplicative [273]. Multispecies [851, 2102, 2422, 345, 323]. Multispecific [677]. multispécifiques [677]. Multistage [1212, 279, 1516]. Multitube [752]. Multivariate [316, 603, 2161, 94, 2274, 1715, 925, 1714, 2293, 1152]. Mummichogs [2606]. Muscle [1988, 913]. Musculature [623, 1666]. Musculium [854]. Mussel [1029, 1389, 104, 816, 2009, 2501, 1899]. Mussels [2529, 1216]. Mya [336]. Mycoplasma [1031]. mykiss [2596, 2551, 2607, 2584, 2570]. Mylocheilus [2241]. Myo [2381]. Myo-inositol [2381]. Myofibrogranuloma [1988]. Myopathy [1988]. Myriophyllum [1408]. Mysid [979, 1173]. Mysidacea [1555]. Mysis [120, 547, 764, 1660, 420, 571, 974, 1555, 1718]. mysticetus [1074]. Mytilus [1029, 1390, 2009, 2501, 1525, 1899]. Myxosoma [132, 422, 266, 647]. Myxosporea [422]. Myxosporean [1056]. Myxozoa [1409, 422].

N [2459, 1833, 2123, 2378, 1203]. N-Cycling [1833]. N-Dimethylformamide [1203]. N.B [329]. N.S. [1620]. N.W.T [2578]. N.W.T. [2529, 643, 1315, 1642, 2179, 2364]. NaCl [767]. namaycush [2473, 2401, 2255, 346, 987, 509, 508, 1574, 1306, 507, 2337, 1052, 1987, 1341, 2216, 282, 240, 2171, 246, 2263, 503, 773, 2617, 2408, 1528, 245, 247, 844, 248, 1815, 2546, 255, 244, 257]. Namsen [2404]. Nanaimo [1141]. Nannoplankton [415]. Nanoflagellate [2593]. Nanoplankton [2082, 293, 2203]. Narcotic [849]. Narrow [2478]. Narwhals [1333]. Natal [1679]. National [2044, 1826]. Native [2351, 507, 503, 71, 329, 229]. Natural [2162, 2258, 1161, 1260, 344, 1894, 664, 1726, 1461, 2098, 1625, 272, 569, 2139, 2264, 1112, 1547, 1221, 1576, 2614, 221, 2132, 2540, 2009, 2501, 1840, 2367, 833, 85, 468, 1587, 2197, 1740, 2489, 655, 1414, 152, 1642, 2179, 2364, 83, 1664, 743, 1582, 56, 1862, 478]. Naturally [910, 2008, 981, 755]. Nature [486]. naturel [1862, 478]. NBS [88]. Neanthes [36, 157]. Near [433, 2527, 2556, 565, 275, 1447, 203, 1018, 2578, 319, 834]. Near-Substrate [2527]. Nearshore [2021, 2081, 1086, 2515, 1220, 1651, 2235, 2236, 1034]. nebulosus [79, 566]. Necessary [114, 485]. Nécrologie [1639]. Necrosis [2476, 1009, 107, 675, 926, 1812, 1943, 1944, 274, 1619, 736, 553]. Need [1206]. Needle [1092]. Neglected [1168]. neiges [1267]. Nematoda [987, 1666]. Nematode [894, 838, 149]. Nemertean [1292, 1597, 1289, 1291, 1783]. Neocalanus [2221]. Nephropidae [873]. Nereocystis [311]. nerka [2162, 151, 26, 1825, 2249, 861, 342, 1723, 2004, 609, 1458, 2371, 2194, 1625, 560, 159, 1560, 2431, 1287, 2195, 1705, 600, 1023, 675, 926, 1812, 1944, 2266, 644, 1238, 1239, 1446, 1536, 2058, 2349, 2039, 1903, 193, 702, 2612, 2516, 1960]. Nest [2400, 1038]. Nesting [2523]. Net [2277, 77, 307, 1620, 1266, 2621, 1057, 93, 2203, 2279, 1908]. Net- [2203]. Nets [1167]. Neuroendocrine [530]. Neuronal [1453]. Neurons [324]. Neurophysiology [218]. Neurotoxic [2332]. Neutral [2410]. Neutralization [2391, 2393]. Nevada [2346, 1461]. Newfoundland [603, 724, 839, 1895, 124, 1317, 393, 190, 2159, 585, 1859, 496, 1026, 1368, 812, 2152]. Newly [1620, 444, 399, 1569, 2119]. Newport [787]. Niagara [1405, 1972, 1457]. Niche [1415, 2094]. Nickel [1803, 1339]. Nighttime [2307]. nigromaculatus [79]. Niño [1559, 1324, 2024]. Nitinat [1466]. Nitrate [89, 175]. Nitrification [82, 1907]. Nitrite [615, 130, 1230, 340, 436, 2569]. Nitrogen [150, 968, 1461, 64, 2319, 792, 63, 2178, 1112, 670, 2261, 2331, 2079, 2352, 1890, 1803, 1193, 1391, 1906, 1176, 1348, 1649, 743, 917]. Nitrogen-15 [63]. Nitrogen-Limited [2178]. Nitzschia [2507, 2332]. NO [1019]. Nocturnal [868]. Nomenclature [209]. Non [2184, 1316]. Non-anadromous [1316]. Non-Dragraked [2184]. Nonanadromous [1584, 2371, 2404]. Nonessential [2573]. Nonevasive [548]. Nonfeeding [1997]. Nongame [640]. Nonhomogeneous [1800]. Nonlinear [2050, 644]. Nonnesting [2523]. Nonparametric [2604]. Nonparasitic [2036]. Nonrandom [2156]. Nonrandomly [2518]. Nonresident [5]. Nonsalmonid [2214, 1473]. Nontarget [2487]. Nonviable [4]. Normal [661, 712, 318, 813, 2600]. Normally [1269]. Normally-Growing [1269]. North [283, 1319, 1367, 618, 2494, 744, 1729, 2616, 1053, 1931, 475, 1473, 2149, 1215, 1956, 1345, 233, 364, 2500, 128, 2174, 1248, 2313, 975, 539, 1660, 1440, 1220, 2361, 2600, 1885, 966, 2224, 1637, 319]. North-Central [2313]. North-Temperate [1319, 1473, 1215]. Northeast [3, 975, 1517, 1559]. Northeastern [2159, 404, 292, 1956, 1345]. Northern [2474, 2155, 1095, 1098, 333, 2436, 2552, 1092, 187, 2148, 1986, 180, 203, 2093, 2611, 1885, 2533, 1013, 1477, 1313, 812, 297, 774, 802, 1025, 1788, 2146, 1362, 2594, 105, 2491, 2297, 2103, 1946, 234, 279, 2292, 2312, 1089, 1515, 2235, 122, 238]. Northshore [2084]. Northwest [2426, 848, 1854, 2161, 2050, 1399, 94, 376, 1598, 1228, 1821, 747, 1232, 871, 1568, 1359, 581, 1970, 1923, 2152]. Northwestern [67, 1382, 59, 1549, 448, 887, 699, 2310, 1489, 969, 2435]. norvegica [1248, 478]. Norway [2474, 1248, 2248, 1729, 661, 712, 912, 2404, 1383]. Norwegian [698, 2193, 320, 1622, 1735, 1551, 2113]. Notching [2069]. Note [2186, 2039]. Notemigonus [2512, 2452]. Notes [1761]. Notropis [573, 868]. nouveau [265]. nouveautés [2225]. Nova [1279, 558, 1276, 1770, 2582, 1229, 1526, 672, 780, 2515, 563, 2427, 2252, 466, 1043, 613, 1883, 816, 433, 1994, 2601, 1044, 2296, 1432, 268, 627, 1244, 2157, 818]. novaeangliae [2589, 284, 1081]. Nucleic [1030]. Null [2543]. Number [900, 73, 1162, 2241]. Numbers [330, 1435]. Numerical [404, 111]. numéro [43]. Nuphar [1249]. Nurture [486]. Nutrient [1550, 74, 2065, 2262, 1873, 1408, 2409, 682, 1552, 1129, 1417, 2357, 2497, 1690, 569, 58, 75, 1019, 1748, 2070, 287, 1039, 889, 1968, 400, 1083, 2076, 778, 2504, 1166, 1892, 1551, 1827, 2109, 2242, 45, 1518, 1330, 949, 2034, 841, 720, 1599]. Nutrient-Enriched [1166]. Nutrient-Poor [1599]. Nutrients [1445, 348, 1747, 2503, 2086, 2446, 1035, 2364]. Nutrition [1786, 160]. Nutritional [1785, 545, 2496]. Nymph [1698]. Nymphs [985, 1727].

O [646, 1201]. O-de-ethylase [646]. O-Methylation [1201]. O. [1286, 342, 1267, 1331, 1571, 2530, 2163, 2266, 2438, 1194]. Obituary [1639]. Obituary/Nécrologie [1639]. Objective [2520]. Objectives [830]. Objects [2052]. Obligate [745]. obliquas [2031]. obscurus [601]. Observational [2326]. Observations [124, 321, 706]. Observed [558]. Obtained [2202]. occidentale [854]. Occurrence [2500, 1292, 2021, 2494, 130, 1544, 1555, 2028]. occurrences [2377]. Occurring [1900]. Ocean [665, 617, 2508, 1565, 1238, 1239, 2235, 2236, 1569, 2119, 1843, 3, 2500, 1559, 2101, 292]. Ocean- [1569, 2119]. Oceanic [2143, 1123, 2354, 2502]. océanique [2502]. Oceanographic [435, 1358]. Oceanography [882, 1954, 1017, 879]. Oceans [1019]. ocellata [835].

Odobenus [334]. Odonata [2019]. Odor [185]. Odorant [2367]. Off [2582, 787, 2159, 2296, 758, 2383, 2108, 1276, 1771, 1770, 2218, 1229, 1526, 1560, 2047, 2196, 1966, 1605, 1000, 2565, 1368, 1050]. Off-Channel [2383]. offs [833]. Offshore [626, 1366, 2081, 1429, 358, 1448]. Ohio [741]. Oikopleura [2028]. Oil [2345, 927, 101, 160, 1481, 928, 1304]. Oils [757]. Oklahoma [2264]. Old [1985, 332, 1474]. Old-growth [1985]. Oldman [690]. Olfaction [1453]. Olfactory [2367, 2368]. Oligochaeta [792]. Oligochaete [1120]. Oligochaetes [1573, 2007, 943]. Oligotrophic [860, 1707, 1382, 59, 1570, 1193]. Oligotrophy [1215, 1935]. Ommastrephid [35]. Omnivorous [1741, 2173]. Onchorhynchus [1699]. Onchorynchus [2564]. Oncorhynchus [1308, 2162, 821, 996, 814, 984, 1305, 1431, 1466, 1286, 1534, 1851, 2026, 2258, 1422, 872, 151, 26, 2175, 2500, 1985, 583, 1825, 2249, 861, 342, 1723, 1335, 1141, 1396, 2524, 823, 2004, 609, 1458, 2596, 1842, 2375, 1155, 2232, 1844, 796, 76, 2218, 1977, 2371, 2345, 1850, 2194, 1625, 560, 1913, 1910, 288, 159, 1560, 1721, 2494, 1331, 1571, 606, 653, 1073, 2431, 733, 544, 1742, 2151, 2530, 1287, 2245, 2491, 593, 1102, 2265, 781, 1961, 2607, 705, 2195, 2616, 2615, 1900, 1674, 107, 2251, 2163, 1705, 600, 365, 176, 2508, 1838, 391, 1023, 2548, 2347, 675]. Oncorhynchus [926, 1812, 1944, 904, 703, 1360, 1565, 1805, 2005, 2266, 1483, 644, 891, 1238, 1239, 1446, 1536, 2505, 696, 695, 1208, 938, 1142, 1177, 2058, 2181, 1870, 1282, 352, 1145, 2438, 1938, 1804, 2384, 1194, 611, 102, 2349, 2531, 2383, 1495, 1496, 1569, 2119, 2528, 2039, 817, 2054, 1028, 2584, 1917, 1903, 1860, 193, 702, 1970, 2275, 2516, 1960, 2570, 335, 1611, 1038]. Oncorhynehus [2612]. Oncothynchus [2551, 2488]. One [2036, 2410, 1676, 1474, 2172, 416]. One-Day-Old [1474]. Oneida [2464, 2]. Onset [1863, 2090]. Onshore [1535]. Ontario [1582, 2486, 2255, 801, 1175, 1811, 511, 242, 964, 67, 2311, 2256, 1405, 829, 1382, 502, 708, 2081, 2018, 59, 2336, 2506, 691, 161, 1595, 1819, 2313, 1105, 1221, 2091, 2079, 2070, 2073, 2074, 2083, 1549, 503, 2025, 1394, 2087, 448, 2316, 763, 2408, 1890, 2342, 2562, 2086, 2446, 467, 2088, 2075, 2082, 1727, 2080, 2090, 1106, 1577, 111, 2242, 1363, 2460, 2071, 1018, 2150, 2293, 2072, 641, 1343, 1225, 2085, 2084, 275, 834, 961, 2089]. onto [949]. Ontogenesis [670]. Ontogenetic [2451, 2094, 2305, 2257]. Ontogeny [2559]. Oocysts [1983]. Oocyte [1680, 1996, 1151]. Oocytes [4, 1499]. opalescens [1966, 605]. Open [1046, 2248, 1498]. Open-Work [1046]. Operated [2469]. Opercula [948]. opilio [1716, 1267, 1310, 2335, 2062, 308]. Opossum [120, 1718, 420]. Opportunistic [919, 1080]. Optical [1087, 266]. Optimal [549, 993, 994, 1366, 2127, 2437, 1612, 1915, 1513, 2301, 37, 1516, 1512, 1282, 586, 1069, 19, 21]. Optimization [1437]. Optimizing [831, 1472]. Optimum [1188, 1476, 374, 1024, 141, 377]. Oral [382]. Orconectes [2468, 1382, 1752, 1889, 50, 601, 1628]. Order [332, 112]. Ordination [174]. Oregon [1966, 196, 333, 823, 2218, 1605, 787, 304, 112, 1565, 1805, 891, 1823, 2048, 2505, 941, 2011, 2181, 352, 830, 2240]. oregonensis [2155, 333]. Organic [923, 1833, 2467, 168, 55, 1332, 1661, 1396, 2325, 295, 898, 2357, 1389, 1323, 2170, 1546, 2007, 1126, 1846, 749, 2086, 1964, 1441, 2181, 1035, 1391, 1704, 1348, 1525]. Organics [1120]. organique [1525]. Organismal [490]. Organisms [2229, 1976, 2353, 2487, 1750]. Organization [2043, 2455, 1524]. Organochlorine [1912, 2456, 1457, 299, 604]. Organochlorines [2398]. Orientation [2488, 938, 1177, 87]. Oriented [1133]. Origin [492, 894, 2430, 367, 1278, 281, 658]. Origins [2456, 736]. Orthophosphate [1033, 1542, 1687, 62, 1840, 916, 61, 2350, 449, 587]. Orthophosphorus [564]. Oscillations [1478]. osculus [725]. Osmerus [2046, 677, 1621, 847, 1415, 1602, 1416, 1288]. Osmolality [29]. Osmoregulation [219, 214, 2284]. Osmoregulatory [1223, 2265]. Ostrea [1898]. Other [646, 2415, 2055, 2469, 1306, 674, 2103, 2442, 773, 1329, 2316, 2446, 644, 1632, 312, 2383, 1933, 2207, 954]. Otic [1590]. Otolith [2053, 2271, 1360, 1429, 2045, 2366, 2557, 1028]. Otoliths [1654, 650, 1373, 600, 703, 2588, 2365, 1682, 193, 1426]. Otter [2465]. ouest [1267]. Outdoor [944]. Outlet [2176, 1826]. Output [1748, 1837]. Ovarian [1973, 1996, 2339, 1499]. Ovary [115]. Overestimation [564]. Overexploited [1915]. Overfishing [1927]. Overlap [803]. Overlying [1204, 1215]. Overview [2070, 2385]. Overwinter [2597, 817]. Overwintering [2144, 695]. Ovigerous [1771]. Ovulation [611, 1537, 2054]. Ovulatory [1977]. Oxaloacetate [21]. Oxic [1104]. Oxidation [2473, 160, 709]. Oxidized [2345]. Oxine [1529]. Oxygen [1277, 1948, 929, 342, 156, 1670, 198, 7, 1920, 660, 1433, 691, 20, 670, 1907, 1978, 2359, 1450, 2307, 24, 2292, 183, 1381, 462, 2252, 419, 2160]. oxyrinchus [977, 581]. Oxytetracycline [1925]. Oyster [874, 669]. Oysters [1898, 126, 2244, 2306]. Ozone [388].

P [2123, 865, 1547, 944, 292, 850]. P-450-Dependent [865]. P-Cresol [944]. P. [2471]. P/B [850]. P450 [1967]. Pacific [3, 2500, 1559, 292, 729, 984, 232, 1492, 1409, 1051, 1048, 4, 893, 976, 618, 1118, 1143, 1878, 2461, 339, 2198, 744, 1958, 1996, 2138, 2139, 2040, 1952, 1961, 1440, 1000, 341, 1065, 2564, 1177, 1424, 2611, 499, 2257, 935, 1874, 1156, 706, 831, 2101, 1938, 1324, 590, 2205, 824, 1941, 1995, 1604, 2576, 2130, 1448, 669, 1970, 1152, 1929, 1930]. Packhorse [1773]. Pagophilus [357]. Paired [207]. Paiute [413]. Paleolimnological [610]. Pallas [2553]. pallasi [729, 1118, 339, 2198, 1958, 1996, 2138, 2139, 1065, 2611, 2257, 1156, 706, 831, 590, 2205, 1995, 1604, 2576]. Panache [761]. Panagrellus [149]. Pancreatic [1009, 736, 553]. Pandalopsis [1933]. Pandalus [1933, 1644]. Panope [612]. Panulirus [1776, 1764, 1775, 1777, 1760, 1762, 1774, 1763, 2435, 1765]. Paradox [2098]. Paragnetina [1698]. Paralithodes [2471]. Paralytic [799]. Parameter [1520, 34, 2301, 380, 2369, 794, 298]. Parameterizations [1593]. Parameters [146, 2156, 2300, 86, 2206, 1266, 915, 341, 112, 1130, 1510, 1326, 1695, 439, 2293, 812, 2280]. Parasite [947, 1056, 2058, 1933]. Parasites [947, 623, 2162, 1489, 1783]. Parasitic [625, 2036, 2407, 242, 2097, 2583, 476]. Parasitism [2142, 2005]. Parental [177, 1536]. Park [2044, 1826]. Parophrys [1226, 2339, 786, 1823, 2048, 1481]. Parr [1584, 1037, 2538, 839, 1107, 1317, 2405, 1945, 446, 2470, 1838, 1992, 1656, 2099, 1125]. Parr-to-Smolt [1838]. Parsnip [947]. Partial [1195, 424]. Partially [932, 2160, 1621]. Participation [2040]. Particle [619, 983, 801, 934, 2244, 358, 2051]. Particle-Size [358, 2051]. Particle-Size-Conversion [619, 983, 801]. Particles [1635, 2496, 1411, 958]. Particular [535, 207]. Particulate [11, 1197, 2325, 1389, 2170, 155, 630, 2115, 1054, 1843, 1704]. partiellement [1621]. partir [2502]. Partition [1678]. Partitioning [201, 1249, 1542, 1460, 1702, 1430, 1415, 180, 1837, 1808, 1231, 1196]. Passage [1845]. Passamaquoddy [1693, 1581]. Passing [1473]. Passive [1620, 1167]. Past [2572, 1478, 1006]. Pasteurella [752]. Patches [2605]. Patchiness [371]. Pathogenesis [126, 2060, 675, 1619, 1507]. Pathogenic [752]. Pathogenicity [430]. Pathogens [350]. Pathways [295, 1071]. Patinopecten [1837]. Pattern [609, 187, 2550, 1474, 600]. Patterns [1459, 2321, 1217, 2509, 1737, 1107, 1317, 1974, 792, 1475, 288, 859, 1662, 1677, 433, 306, 2491, 670, 1228, 763, 407, 585, 2094, 1446, 2579, 697, 2478, 1434, 2224, 2523, 2203, 2240, 397]. Pb [1349, 1401, 1136]. PCB [2105, 622, 913]. PCB-Induced [2105]. pCO [1946, 1234]. PDF [1916]. Peak [1916]. Peamouth [2241]. Peatlands [1168]. Pelagic [619, 983, 2437, 1389, 2622, 882, 1210, 1697, 1867, 2565, 1848, 1147, 1766, 735, 958, 2427]. pélagiques [2225]. pelamis [1952]. Pelecypoda [911]. Pellet [173]. peltastes [180]. Penetrating [320]. penicillatus [1242]. Peninsula [6, 1236, 2624]. péninsule [6]. Pennate [2507]. Pennsylvania [123]. People [1247]. Perca [2376, 2068, 1449, 1506, 2317, 1077, 1884, 2464, 2292, 1962, 2, 1200, 2045, 2305, 2597, 570, 1392, 1231, 1696, 2479, 1524]. Perch [2376, 2068, 1449, 1506, 2317, 1312, 1077, 1884, 2464, 2292, 1962, 2, 1200, 2045, 2305, 2597, 570, 1392, 1231, 2101, 1696, 2479, 369, 629, 548]. Performance [2401, 1825, 628, 621, 2012, 1172, 1312, 605, 2384, 1496, 1099, 504]. Perfusate [1234]. Perfused [410, 1234, 2348]. Period [1563, 2594, 2367, 1432, 111]. Periodicity [1064]. Periods [1385, 2484]. Periphytic [2122, 2517]. Periphyton [2304, 2303, 1855, 2067, 1739, 348, 1743, 982, 1731, 49, 1964, 2455, 2022, 1892, 972, 2566]. Perivitelline [2607, 1140, 1679]. perlan [677, 1621]. Perlid [2551]. Perlidae [1698]. Permeability [2015, 2411]. Permethrin [1256, 312, 1255]. Permits [1153]. Peroxide [2509]. Persistence [1831]. Persistent [321]. Perspective [1779, 153, 484, 1927, 1809]. Perspectives [1787, 1760, 1774, 493, 523]. Persson [2106]. Persulfate [1846]. Perturbation [1520, 2622, 2385, 1724]. Perturbations [1484, 2242]. Peruvian [2024]. Pest [252, 249]. Pesticide [356, 624]. Pesticides [312]. Peter’s [877]. Petroleum [1318, 1566, 642, 846, 757, 392]. Petroleum- [757]. Petromyzon [233, 2255, 505, 242, 230, 231, 252, 458, 228, 251, 253, 1191, 241, 604, 240, 258, 210, 515, 260, 261, 262, 250, 245, 227, 224, 2005, 223, 225, 247, 249, 229, 222, 243, 248, 1815, 259, 254, 226, 1590, 244, 1345]. Petromyzonidae [208]. Petromyzoniformes [207]. pH [2254, 2277, 1340, 1021, 1661, 1439, 1609, 1689, 133, 1730, 2256, 2125, 2594, 1950, 1987, 2543, 1585, 643, 1464, 1752, 2506, 53, 1946, 2607, 1344, 1529, 1688, 50, 2316, 2281, 1487, 2121, 2292, 88, 1381, 103, 853, 1140, 1127, 755, 2168, 1334, 2120, 1235, 935, 2286, 340, 890, 1694, 1957, 1537, 1893, 2282, 954, 834]. pH-Altered [2125]. Pharyngeal [1663]. Phase [2498, 242, 241, 1766, 2003]. Phenolic [896]. Phenomenon [1127, 2024]. Phenotype [1807]. Phenotypes [331]. Phenotypic [1113, 911, 2035]. Pheromone [254]. philippinarum [2502]. Phoca [848, 918, 1399, 1309, 1031, 270, 871, 1568, 2212, 2549]. Phosphatase [1542, 1297, 2075, 474]. Phosphate [1939, 1268, 1546, 2087, 315, 1487, 1544, 2121, 2031, 1686, 1543, 1215]. Phosphorus [1550, 1257, 11, 2122, 2517, 1548, 968, 940, 344, 557, 860, 2217, 937, 1542, 1687, 1545, 110, 1541, 417, 722, 2432, 1546, 2264, 782, 2327, 57, 777, 2073, 2074, 1549, 2352, 2360, 2571, 1890, 1808, 85, 199, 858, 315, 1544, 1104, 1570, 2145, 1911, 1931, 2406, 1144, 2088, 962, 630, 1089, 2114, 2115, 635, 758, 1054, 1651, 184, 38, 1643, 1117, 1352, 697, 1421, 1553, 2478, 474, 1258, 1022, 2072, 463, 1225, 2224, 1975, 1284, 1348, 83, 1864, 56, 1540, 1554]. Photo [1402]. Photo-Electric [1402]. Photocycle [939]. Photoinhibition [2080]. Photomicroscopy [148]. Photoperiod [1271, 1680, 2370, 76, 1902, 1992, 389, 1611]. Photoperiodic [873]. Photosynthesis [89, 1487, 2031]. Photosynthetic [60, 2178, 1743, 131]. Phototrophic [2593, 2082]. Phthalate [1226, 399]. physalus [1517]. Physeter [283, 2296]. Physical [910, 807, 2572, 887, 2336, 253, 132, 2573, 471, 2408, 2259, 2587, 1869]. Physicochemical [961]. Physiological [2272, 1337, 1254, 2429, 2144, 892, 58, 981, 107, 1316, 1647, 1494, 2191, 1870, 1735, 1628, 2279, 2280]. Physiology [220, 221, 1829, 216]. Phytomacrofaunal [2211]. Phytopiankton [1545]. phytoplancton [408]. Phytoplankter [1906]. Phytoplankton [1550, 459, 1445, 168, 2403, 1385, 2397, 95, 1873, 2459, 1043, 1687, 1552, 656, 67, 1726, 1895, 1417, 1879, 1880, 66, 2081, 2178, 1690, 117, 569, 58, 1395, 2620, 1547, 57, 360, 945, 2261, 579, 471, 191, 349, 1840, 1968, 1647, 1544, 2446, 962, 1083, 778, 2080, 131, 1551, 1386, 1827, 1486, 1518, 1203, 1330, 1686, 2034, 1225, 371, 1391, 2270, 739, 2364, 1649, 359, 56]. Phytoplanktonic [1839]. Picogram [2498]. Picoplankton [1530, 2083, 2082, 1809]. Pigment [11, 630]. Pigmentation [845]. Pike [1098, 2436, 2552, 187, 2148, 203, 1885, 1013, 1477]. pileus [1526]. Pimephales [2100, 2417, 84, 1663, 849, 1032]. Pink [1431, 1286, 128, 2358, 1913, 288, 1560, 1331, 2488, 705, 2163, 365, 2005, 1194, 1860]. Pinnipeds [1357]. Piona [735]. Pisces [852, 1136]. Piscicola [1492]. Piscivority [640, 167]. Piscivorous [1909]. Piscivory [2363]. Pituitary [1668, 533, 215, 2233]. Placer [1897]. Placopecten [591, 799, 2252, 1979, 1828]. Plaice [617]. Plain [1552, 1827, 1219]. Plains [742]. planctonique [677]. Planktivore [745]. Planktivores [2484]. Planktivorous [1811, 2112, 454, 2571, 1736, 588]. Planktivory [2622, 1384].

Plankton [1204, 2078, 2077, 1741, 2466, 2297, 777, 2073, 2074, 266, 54, 448, 922, 1896, 1090, 867, 959, 2076, 2109, 1724, 1138, 1192, 1869, 1540, 1554, 1657, 2051, 2173, 1935, 743, 677]. Planktonic [14, 1021, 2027, 957, 299, 1546, 1342, 2196, 578, 2025, 1487, 671, 381, 1064, 1979, 152, 1469]. Plans [2124]. Plant [2211, 1636, 1603, 570, 1245]. Planted [502]. Plants [1905, 231, 806]. Plasma [2182, 108, 2258, 1977, 1113, 2234, 1222, 1172, 1992, 1695, 2606, 2450, 436, 2054]. Plasmid [2495]. Plasmid-Mediated [2495]. Plastic [312]. Plasticity [488]. Plate [2190]. platessoides [617]. Platichthys [650, 783, 594]. platyceros [1644]. platypus [2471]. Plecoptera [1698, 1626]. Pleurobrachia [1526]. plicatilis [97]. Plots [750]. Pluie [2204]. plumbeus [1367]. Plume [347]. plumosus [161]. Pneumonia [1031]. Pocket [351]. Poecilia [290, 2365, 936]. Point [815, 1678, 1801]. Points [1927]. Poison [799]. Poisoned [2589]. poisson [1307]. poissons [714, 2225]. Poland [1394]. Polar [174]. Policies [1369, 1915, 153]. Policy [582, 523, 830, 263]. Pollachius [1682]. Pollock [623, 2059, 142, 188, 2416, 1682]. Polluted [2272, 1734]. Pollution [738, 885, 1302, 81]. Polychaeta [370, 3, 36, 157]. Polychaete [545]. Polychaetes [825]. Polychlorinated [913, 1481]. Polycultural [538]. Polycyclic [2322]. Polyethylene [2476]. Polygenic [329]. polymorpha [2553, 1810]. Polymorphic [2273]. Polymorphism [2535, 2146, 2222]. polynyma [444]. Polyorchis [1242]. Pomatoschistus [1527]. Pomoxis [79, 1736]. Pond [2008, 1648, 1677, 123, 1423, 2366, 2566]. Pond-reared [2366]. Ponding [1282]. Ponds [1873, 2307, 1926, 696, 695, 1245, 2383]. Pontopareia [2311]. Pontoporeia [2568, 2497, 2166, 2567]. Pool [2333, 1054]. Pooled [1308]. Pools [1573, 1135]. Poor [1599]. Poorly [1434]. Pope [1800]. Population [146, 2376, 678, 2415, 489, 491, 1308, 2529, 2036, 2407, 2535, 1357, 1098, 2413, 2414, 2547, 1854, 1074, 1796, 1004, 970, 2453, 2468, 2220, 613, 1608, 376, 2092, 1113, 1791, 1382, 1950, 976, 1119, 986, 142, 1118, 1143, 28, 911, 1073, 933, 2102, 2400, 105, 2452, 1952, 167, 1961, 2470, 2614, 484, 1800, 345, 447, 874, 2617, 1455, 1673, 2048, 695, 2154, 2029, 1623, 424, 871, 1568, 1418, 1928, 2101, 281, 1804, 1745, 2177, 844, 1993, 248, 254, 1933, 2546, 1477, 369, 629, 942, 1891, 2130, 2024, 257, 394, 1135]. Populations [1790, 495, 2104, 510, 2169, 1788, 505, 1723, 413, 1887, 1918, 230, 1224, 719, 1719, 1443, 1306, 124, 1461, 347, 1202, 1519, 2466, 2396, 1999, 1027, 104, 816, 1910, 332, 569, 2139, 33, 2246, 2525, 1341, 1259, 1759, 1622, 740, 515, 1807, 2043, 1467, 1195, 1800, 1603, 2009, 2501, 318, 689, 420, 654, 1812, 2536, 1656, 1962, 1805, 298, 2595, 1190, 644, 2542, 2058, 2611, 329, 1392, 1003, 1145, 1885, 1139, 2243, 1495, 1496, 1569, 2119, 1936, 372, 498, 237, 1637, 2290, 2608, 1514, 244, 1970, 1007, 1345, 6, 1125, 1135, 1810]. Pore [69]. Porewater [2478]. Positional [2287]. Positions [450]. Possible [947, 1882, 1560, 868, 284, 2442, 700, 1036, 1801, 2344]. Postemergent [1344, 390]. Postglacial [894, 2568, 1236, 1885]. Postimpoundment [1097, 1099]. Postlarval [2047]. Postmolt [2580]. Postmortem [623].

Postsettlement [1343]. Postsmolt [2159, 2449]. Poststocking [2245]. Potent [2375]. Potential [1175, 357, 1492, 336, 978, 1511, 726, 1452, 2466, 1974, 1478, 2299, 252, 772, 2204, 2132, 287, 1489, 1983, 250, 1201, 1432, 1140, 2489, 1364, 2623, 2030, 1872, 2478, 2028, 2546, 2513, 939]. Potentials [1679, 2015]. Potentiated [770]. Pothole [1948, 183]. Potomac [1004]. pour [2225, 408]. Pout [617]. Power [91, 1511, 2618, 1253, 1636, 806, 1603, 2049, 2505, 570, 525, 1448, 1801]. Power-Cost [2618]. Practical [1787, 686, 1645, 864]. Practical-Type [1645]. Practice [1660]. Prairie [1948, 1445, 968, 1606, 1233, 2174, 2170, 1734, 428, 429, 183, 1696]. Prawn [1285, 1644]. Preamplexus [1875]. Precambrian [64, 2506, 773, 2147, 1487, 47, 1570, 1200, 758, 954]. Precipitation [2476, 1391, 1642]. Precision [434, 2038, 418, 2165, 73, 1753]. Precocious [839, 1181]. Predation [14, 729, 2340, 2255, 793, 477, 333, 2353, 120, 547, 2002, 1478, 1850, 1331, 734, 2560, 1814, 2400, 2142, 591, 1415, 1602, 2163, 979, 2, 1736, 1722, 1848, 2064, 2532, 247, 1872, 2314, 1556, 2109, 1244, 1162, 248, 1815, 1524, 276, 1929, 1930, 2231, 2610, 1080]. Predator [678, 796, 2014, 657, 868, 2591, 276]. Predator-Induced [2014]. Predators [1462, 1698, 1478, 2102, 2452, 304, 435, 645, 2305, 2298, 1032]. Predatory [1548, 964, 1055, 241, 577, 735, 325, 1270]. Predatory-Phase [241]. Predict [1635, 272, 692]. Predictability [2312, 2579]. Predicted [2546]. Predicting [2394, 2466, 556, 691, 2434, 776, 2307, 2172, 2294, 833, 1392, 706, 1258, 2205, 1368, 1403, 2516]. Prediction [344, 1075, 1631, 465, 2432, 559, 1068, 1100, 27, 360, 1494, 2145, 1763, 184, 1816, 2418, 1813, 2578, 1599]. Predictions [2006]. Predictive [1918, 144, 1558, 469]. Predictor [768, 1975]. Predictors [630, 966]. Predicts [1977]. Preexposure [1849, 1899, 1947]. Preference [2484, 96, 601, 1655, 645, 648]. Preferences [2302, 1531]. Preferenda [573]. Prehistoric [660]. Preliminary [562, 1898, 259]. Preparation [2348, 550]. Preseason [2249, 668, 718]. Presence [2412, 2281, 1103, 1525, 1307, 1525]. Present [2572, 1478, 387, 883]. Presenting [38]. Preservation [2421, 552]. Preservatives [2186]. Preserved [1638, 2063]. Preserving [2057]. Presmolts [1805, 2528]. Prespawning [2138]. Prespring [1432]. Pressions [2225]. Pressure [313, 979]. Prevalance [2058]. Prevalence [190, 727]. Previous [1579]. Previously [1110, 775]. Prey [678, 2019, 2414, 1079, 1820, 2436, 2068, 2001, 2484, 1462, 1155, 1306, 796, 1698, 1076, 2521, 657, 2098, 1672, 228, 2560, 1346, 2102, 955, 1384, 454, 1909, 617, 836, 1884, 1483, 645, 1848, 2305, 735, 2314, 2591, 276, 548, 1028, 2480, 1032, 1596, 138, 325, 1426]. Preying [1571]. Pribilof [2471, 2534]. Primary [2507, 2397, 324, 1879, 857, 1986, 569, 1356, 1088, 277, 800, 1322, 1940, 2619, 2620, 855, 293, 1014, 596, 2364]. Prince [2507]. principaux [1621]. Prior [1337, 507]. Pristine [749]. Probabilities [2580]. Probability [2194, 398, 1167]. Probable [1519]. Probe [1733, 2463, 351]. Probes [1967]. Probing [828]. Problem [2103, 912, 2505, 141]. Problems [708, 1191, 1760, 96, 601, 250, 1173, 1359, 1769, 1922]. Procedure [1844, 447, 184]. Procedures [2415, 312, 2205]. Proceedings [480, 205]. Process [2020, 2336, 1522]. Processes [989, 1787, 1758, 2010, 435, 1651, 1151]. Processing [1332, 15, 2493]. Procladius [1814, 2227, 2513]. Produced [821, 2036, 1201, 1483, 2332]. Producers [2397]. Producing [1048, 1457]. Product [1668]. Production [923, 472, 1955, 1985, 72, 165, 619, 1175, 2429, 848, 650, 1511, 1770, 1521, 823, 1309, 760, 67, 2174, 1417, 1879, 66, 717, 1659, 4, 986, 1250, 1986, 876, 2378, 795, 569, 772, 2186, 2103, 2166, 1940, 1410, 2619, 2620, 855, 2601, 471, 287, 742, 1394, 2501, 1564, 971, 37, 296, 2425, 163, 2128, 703, 953, 2379, 1565, 1896, 2621, 1964, 437, 2542, 2475, 1740, 1334, 293, 2063, 1285, 2395, 1486, 1407, 1463, 1993, 1330, 1036, 1537, 850, 401, 1066, 164, 1843, 739, 204, 2203, 2364, 1801, 2426, 1908, 297, 426]. Productive [916, 1117]. Productivity [198, 1382, 857, 1356, 1088, 277, 800, 309, 2132, 1581, 1731, 768, 1014, 596, 2157, 1724]. Products [464, 726, 999, 752]. productus [4, 976, 824, 1941]. produit [1307]. Profiles [631, 691, 1865]. Profundal [1068, 1091]. Progeny [2283, 1611]. Program [2423, 252, 1023]. Programs [519, 1437, 608]. Progress [2457]. Prohibited [1024]. Project [2276, 2385, 44]. Projection [148]. Projections [1932]. Projects [1148]. Prolactin [2234]. Prologue [526]. Prolonged [593, 2281, 1897, 1496]. promelas [2100, 2417, 84, 1663, 849, 1032]. Promotion [2107, 1449]. Propagation [501, 163]. Properties [2573, 361, 2539, 1593]. propos [43]. Proposal [645]. Proposed [157]. Prorocentrum [401, 1066]. Prosimulium [2338]. Prosopium [654]. Prospects [539, 249]. Prostaglandins [534]. Protease [1056, 1182]. Proteases [361]. Protected [1942]. Protection [153]. Protective [551, 751]. Protein [1030, 13, 2061, 2077, 1113, 2133, 1631, 1662, 2012, 2208, 1505, 639, 2076, 1291]. Protein-Based [1662]. Proteins [1630]. Protozoa [699, 1409]. Protozoan [947, 422]. Provide [2008]. Provides [2359]. Provinces [1898, 104, 497]. Proximate [2433, 2138]. Psammophifous [1407]. Pseudocarcinonemertes [1290, 1292, 1597, 1289, 1291]. pseudoharengus [1607, 1447, 728, 23, 1485, 2550, 167, 362, 1667]. Pseudophyllidea [861]. Pseudopleuronectes [865, 1174, 572, 2434, 617, 2322, 1425, 1967, 430]. Pseudopleutonectes [2055]. Pseudoterranova [2492]. Ptychocheilus [2155, 333]. Pu [2140]. publications [41, 171, 322, 576, 685, 791, 908, 1062, 1171, 1301, 1428, 1562, 1713, 1877, 2137, 2269, 2399, 2537]. publications/Corrections [1062, 1171, 1301, 1428, 1562, 2269, 41, 171, 322, 576, 685, 791, 908, 1713, 1877, 2137, 2399, 2537]. Pueruli [1774]. Puerulus [1762]. Puget [814, 1431, 1260, 1887, 2618, 2339, 176]. pulex [636]. Pulp [2272, 386]. Pulse [1987, 1676, 1133]. Pulses [1422, 2034]. Pumpkinseed [180, 2094, 276]. Pumps [1192]. punctatus [1110, 2105, 2499, 2381, 652, 615, 997, 1926, 788, 411, 421, 1230, 1507, 436]. pungens [2507, 2332]. Pup [1357, 848, 1309, 2177, 297]. Pure [874]. Purified [845]. purpuratus [2024]. Purse [1253, 2355]. Purse-Seine [2355]. Pushnet [1485]. Put [416]. pygmaea [1727]. pyrene [278]. Pyruvate [21].

Quadrat [1198]. Quadratic [1455]. Qualitative [2479]. Quality [821, 996, 2078, 2485, 862, 1556, 46, 2150, 1504, 1826, 545]. Quantification [1274, 903, 2457, 2239, 1219]. Quantify [837, 2150]. Quantifying [1844, 1578, 635]. Quantitative [623, 2238, 1443, 87, 2479, 554]. Quantity [2413]. Quebec [2373, 2047, 2196, 1862, 1862, 1703, 1538, 1119, 1125, 2299, 945, 2176, 496, 390, 1582, 1125, 1236, 319]. Question [1653]. Questions [1375]. Quid [540]. Quinoline [2440]. Quinte [161, 2446, 1343, 1421]. Quo [540].

r [272]. r-K [272]. Ra [2442]. Race [560]. Races [1675]. radiata [1567]. Radioactive [1510]. Radiobioassay [2350]. Radiocarbon [60, 59]. Radiochemical [62]. Radionuclides [1865, 51]. Radiophosphorus [860, 406]. Radiotracer [777, 1549, 1295, 1510, 463]. Radiotracers [52]. Radium [1071]. Raiding [2400]. Rain [1168, 1852, 1615]. Rain-Forest [1852]. Rainbow [1668, 2107, 1186, 2254, 2253, 108, 1337, 2104, 510, 1165, 13, 1340, 1020, 1589, 1521, 1894, 2372, 2356, 2596, 2440, 133, 405, 1372, 2125, 1719, 457, 1614, 2551, 847, 1160, 1222, 1223, 2219, 197, 897, 795, 1268, 1847, 1822, 1746, 686, 981, 765, 1645, 22, 30, 1354, 2012, 160, 955, 289, 1493, 1919, 186, 715, 710, 2607, 2359, 1688, 1829, 1830, 1505, 1602, 2060, 2362, 1708, 1312, 2342, 99, 813, 1494, 467, 913, 451, 331, 1532, 1416, 410, 181, 1234, 2334, 2324, 1288, 2168, 2120, 738, 462, 862, 1523, 340, 1213, 2447, 1414, 2096, 2348, 419, 1588, 751, 2570, 21]. Rainbow [265, 1749]. Rainforest [972]. Raised [1857]. Rakers [488]. Random [2488, 689, 437, 1177]. Randomness [1522]. Range [1444]. Ranges [813]. Rank [2484, 1302]. Rapid [2107, 2440, 628, 1493, 149]. Rapidly [166, 1284, 835]. Rare [5]. raschii [1248]. Rate [32, 2122, 2517, 1033, 1161, 2000, 2232, 76, 116, 730, 621, 273, 1150, 688, 572, 1945, 1266, 2614, 874, 2368, 1516, 1173, 1147, 1183, 2063, 655, 2533]. Rates [1277, 459, 1319, 1025, 115, 1320, 1635, 2267, 77, 547, 1349, 952, 2027, 2002, 2311, 465, 1174, 307, 792, 2497, 1164, 556, 63, 1721, 2378, 1662, 1311, 2208, 1940, 1450, 617, 1039, 1865, 2128, 2294, 2271, 833, 1647, 953, 2145, 2406, 1083, 2489, 143, 2382, 2257, 2236, 1008, 431, 474, 2555, 1414, 2349, 1686, 2160, 204, 1176, 2207, 1908]. Ratio [409]. Ration [151, 26, 2106, 1842, 2059, 581]. Rational [2514]. Rationalizing [2040]. Rations [288, 1028, 335]. Ratios [604, 945, 1890, 2257, 850]. Raven [1113]. Ray [68, 592]. Rays [1699, 442]. Reacidification [2394, 2387, 1383]. React [1304]. Reaction [1155, 928]. Reactions [391, 392]. Reactive [779, 2327, 1104, 83]. Reality [2053]. Reanalysis [2418]. Reappraisal [1501]. Reared [177, 2055, 686, 981, 1900, 1674, 268, 195, 1213, 2366]. Rearing [593, 446, 365, 905, 1859, 2531]. Reassessment [2061]. Rebuild [1805]. Rebuilding [1594]. Recapture [434, 848, 1614, 1857, 1858, 341, 920]. Receiving [46]. récentes [41, 171, 322, 576, 685, 791, 908, 1713, 1877, 2137, 2399, 2537]. Recently [2097, 759, 2103]. récents [1062, 1171, 1301, 1428, 1562, 2269]. Reciprocal [174, 1478, 1567]. Recognition [609, 1474, 700]. Recombinant [2107, 2375]. Recommendations [1799, 523, 263]. Reconditioned [2454]. Reconditioning [1902]. Reconstructed [1238]. Reconstruction [2349]. Reconstructions [1730]. Record [660, 2242]. Records [5]. Recovered [1426]. Recoveries [1362, 713, 2159]. Recovery [2255, 1589, 2519, 1863, 2352, 2172, 1121, 248, 1754, 1664]. Recreational [169, 366, 1006]. Recruit [2156, 1101]. Recruitment [1779, 1789, 1773, 1776, 1787, 1764, 2110, 1786, 724, 2004, 1795, 1558, 2437, 1758, 2466, 2201, 1791, 1798, 560, 398, 33, 1447, 1369, 1338, 1476, 1759, 2260, 1228, 1821, 380, 1760, 1762, 279, 2288, 1122, 2, 1594, 2093, 1761, 1765, 1792, 424, 1583, 2445, 1927, 378, 2205, 1151, 824, 2581, 379, 377, 2598, 129, 1514, 1982, 1923, 1756]. Recurrence [271]. Recurrent [1001]. Recycling [74, 2217, 1545]. Red [573, 2494, 445, 835, 1001, 679]. Red- [2494]. Redd [859]. Redds [1283, 2181]. Redeposition [166]. Redfish [1146, 925, 1429]. Redfishes [1598, 925, 1455, 404, 1232, 747]. redivivus [149]. Redmouth [770]. Redside [2011]. Reduce [1024, 1012]. Reduced [2095, 2029, 419, 2290, 1652]. Reduces [1735]. Reducing [45, 300]. Reductant [2145]. Reductant-Soluble [2145]. Reduction [698, 63, 1690, 721, 1178, 1500, 2088, 255, 359]. Reductions [1889, 2072, 369, 629]. Redundancy [836]. Reef [1652, 248]. Reevaluation [2112]. Reference [1671, 1557, 2552, 535, 2059, 1726, 487, 1622, 763, 207, 1094, 1927, 767, 2270]. Reflectance [11]. Reflection [1478]. Reflections [989, 2052]. Regard [221]. Regarding [1760]. Regeneration [1548, 1019, 1039]. Régime [1621, 899, 899, 265]. Regimes [2527, 292, 834]. Region [1519, 393, 512, 2216, 387, 47, 2595, 208, 43]. Regional [1954, 2579, 2419, 308, 1917]. Regionally [2297]. Regions [1566, 637]. Registration [250]. Regression [2050, 823, 1720, 1802, 164]. Regressions [1419]. Regulated [469, 15]. Regulating [1320, 2400, 1738]. Regulation [1680, 2370, 573, 144, 1835, 314, 621, 687, 2402, 1157, 199, 1245, 29]. Regulations [2590]. Regulatory [1531]. Rehabilitate [246]. Rehabilitation [1915, 519]. Reinhardtius [2161, 376, 699]. Related [1205, 1606, 97, 124, 228, 1449, 1302]. Relation [2486, 1955, 665, 1601, 151, 26, 1597, 1811, 1336, 2021, 1839, 1856, 693, 1390, 887, 2246, 348, 782, 1954, 190, 2573, 2607, 1283, 806, 262, 2316, 245, 2295, 1555, 644, 2076, 496, 247, 462, 2090, 431, 1358, 1724, 2224, 1936, 138]. Relations [1266, 437, 2502]. Relationship [1790, 1116, 1789, 1866, 1445, 2467, 1764, 1154, 724, 2127, 464, 823, 2552, 664, 2563, 764, 1229, 1382, 215, 2261, 1517, 794, 2087, 2408, 2406, 671, 1144, 183, 758, 2259, 1632, 965, 2366, 1022, 1225, 1028, 2226, 1514, 466, 702, 589]. Relationships [963, 11, 1134, 2568, 2126, 1249, 2110, 128, 1197, 682, 2211, 67, 796, 124, 1798, 2357, 1910, 1338, 2111, 787, 349, 2503, 1697, 112, 1911, 962, 1, 2504, 1352, 2382, 1583, 1421, 1728, 1330, 1196, 2567, 849, 379, 1264, 415, 2502]. Relative [633, 1770, 557, 896, 970, 1451, 934, 2213, 653, 1221, 773, 2446, 2082, 1778, 2483, 1518, 2096, 1975, 204, 1448, 1454, 954]. Release [583, 307, 417, 792, 2497, 2172, 858, 2145, 2406, 1651, 697, 463, 720]. Release/Recovery [2172]. Released [814, 898]. Releases [823, 2248, 2508, 971]. Releasing [1977]. Releaslng [2054]. Relevance [1073, 2417, 523]. Reliability [803]. Reliable [459]. Relict [281, 2568]. relicta [120, 547, 764, 1660, 420, 571, 974, 1555, 1718]. Remaining [2172]. remontées [1490]. Remote [2319, 435]. Removal [295, 2497, 2412, 1499, 1003]. rendu [575, 931]. Rendus [18, 40, 100, 119, 137, 154, 189, 286, 338, 354, 384, 402, 425, 443, 461, 555, 616, 632, 649, 667, 684, 701, 753, 771, 790, 809, 842, 863, 886, 907, 967, 990, 1016, 1042, 1061, 1115, 1132, 1149, 1170, 1189, 1207, 1227, 1273, 1300, 1328, 1377, 1398, 1413, 1427, 1442, 1479, 1502, 1533, 1561, 1591, 1616, 1641, 1665, 1685, 1712]. Renewal [47]. Renibacterium [1844, 1841]. Repair [2020]. répartition [677, 1767]. Repeat [1983]. Repeat-Spawning [1983]. Repetition [73]. Replacement [2001, 657, 634]. Replication [115]. Report [261, 262, 260]. Reproducibility [681]. Reproduction [2398, 1797, 1508, 983, 527, 636, 2389, 101, 221, 1517, 833, 1201, 2189, 530, 290, 1623, 914, 1213, 534, 774]. Reproductive [346, 357, 978, 2184, 762, 529, 1732, 897, 1676, 1266, 2471, 712, 1210, 1195, 1837, 2623, 2030, 416, 748, 697, 655, 2460, 1936, 939]. Require [1293]. Required [1972]. Requirement [2031, 1469]. Requirements [2061, 342, 2448, 85, 1803, 608]. rerio [1978]. Research [541, 1799, 1168, 1212, 920, 1753, 1010, 2308, 1313, 1058]. Reserves [1871, 1850, 1644]. Reservoir [923, 1092, 2014, 2178, 277, 634, 2312, 122, 1237, 2224, 1085]. Reservoirs [2129, 722, 2264, 782, 2103, 2360, 1089, 1094]. Residence [2140, 2538, 1102, 1173]. Residency [561]. Resident [2108, 1744, 230, 5, 1930]. Residual [788, 2393]. Residue [604]. Residues [1912, 1318, 1566]. Resin [2432]. Resistance [775, 652, 1950, 2611, 1947, 2495]. Resistant [2200]. Resolute [2578]. Resolution [2605]. Resorption [2500]. Resource [444, 518, 1415, 180, 1517, 1693, 1432, 1298, 1355, 1047, 1231]. Resources [665, 1952, 485, 671, 522]. Respiration [1707, 217, 757, 1315, 2449]. Respiratory [1037, 854, 605]. Respirometer [1058]. Response [1165, 1098, 1079, 2397, 2381, 2123, 1406, 2097, 2468, 2387, 2116, 2396, 1977, 1055, 2543, 981, 2102, 1395, 1715, 1902, 2558, 2313, 57, 997, 1449, 2271, 1494, 331, 435, 2334, 2455, 2022, 2305, 1465, 1989, 1438, 1510, 1124, 1714, 229, 243, 2072, 1569, 595, 1627, 1596, 1403]. Responses [472, 1337, 2122, 2517, 1020, 1589, 1263, 1453, 2323, 1223, 1371, 687, 1287, 2195, 1994, 330, 2540, 2317, 601, 1897, 1262, 2367, 950, 468, 1494, 1381, 2242, 1156, 1121, 102, 1936, 255, 2153, 1628, 2344, 1582]. Rest [108]. Restabilization [1085]. Resting [902, 204, 679]. Restoration [1036, 608]. Restricted [1887, 327, 1951]. Restriction [1137]. Result [1079]. Resulting [626, 419]. Results [1681, 1243, 2326, 2251, 2365, 1386, 259, 164].

Resuspension [626, 1233]. Retardation [1754]. Retention [1396, 77, 937, 110, 1705, 1765, 1288, 1488]. reticulata [290, 2365, 936]. Retinal [2246]. Retirement [16]. Return [2232, 2516]. Returned [1776]. Returning [365]. Returns [583, 2204, 2604]. Reuse [1528]. Reversal [382]. Review [1781, 2055, 1758, 882, 884, 234, 642, 1329, 1761, 1765, 1780, 885, 2152, 575, 931]. review/Compte [575, 931]. Reviews [18, 40, 100, 119, 137, 154, 189, 286, 338, 354, 384, 402, 425, 443, 461, 555, 616, 632, 649, 667, 684, 701, 753, 771, 790, 809, 842, 863, 886, 907, 967, 990, 1016, 1042, 1061, 1115, 1132, 1149, 1170, 1189, 1207, 1227, 1273, 1300, 1328, 1377, 1398, 1413, 1427, 1442, 1479, 1502, 1533, 1561, 1591, 1616, 1641, 1665, 1685, 1712]. Reviews/Comptes [18, 40, 100, 119, 137, 154, 189, 286, 338, 354, 384, 402, 425, 443, 461, 555, 616, 632, 649, 667, 684, 701, 753, 771, 790, 809, 842, 863, 886, 907, 967, 990, 1016, 1042, 1061, 1115, 1132, 1149, 1170, 1189, 1207, 1227, 1273, 1300, 1328, 1377, 1398, 1413, 1427, 1442, 1479, 1502, 1533, 1561, 1591, 1616, 1641, 1665, 1685, 1712]. Revised [2568, 960]. Revisited [1999]. Reynolds [900]. Rheotactic [2119]. Rhinichthys [725, 1055]. rhotheus [477]. Rhythms [527]. Richards [86]. Richardsonius [2011]. Richness [1889, 2316]. Ricker [1476]. Riffle [725, 2487]. Right [1498]. Rigler [2350]. Ring [1050]. Ringed [1399, 270, 2212, 2549]. Rings [815, 193]. Riparian [2170, 887, 2056]. Risk [2142, 2163, 689, 1615]. Risks [1418, 1285]. River [1466, 1534, 232, 364, 2474, 1538, 1096, 1095, 333, 2575, 1335, 1141, 690, 2146, 1317, 1125, 2454, 2131, 1746, 1088, 1087, 2291, 1395, 1734, 800, 236, 1675, 1674, 561, 1963, 239, 1579, 2564, 1084, 2005, 1536, 1, 1727, 1142, 329, 843, 1072, 1735, 2487, 281, 238, 1407, 1375, 1993, 2383, 237, 17, 1261, 2584, 1091, 390, 2162, 1882, 741, 1004, 2004, 1458, 952, 1159, 1452, 1451, 690, 1405, 1972, 560, 1729, 2491, 1435, 1678, 2621, 1288, 738, 1010, 1011, 1014, 1012, 1013, 369, 629, 2404, 2226, 1860, 2516]. Riverine [896, 2297, 696, 695, 1237]. Rivers [947, 724, 1659, 567, 1622, 236, 1994, 2601, 1163, 237, 818, 1035]. rivière [1125]. RNA [1668, 2257]. Roberts [1]. Robust [2369, 1669]. Robustness [2062]. robustus [1752]. rocheux [899]. Rock [1773, 1776, 1764, 613, 591, 2561, 1763, 1761, 1244]. Rocky [1671, 1574, 1401, 15]. Role [1786, 2356, 1453, 9, 28, 252, 253, 2400, 199, 1117, 1245, 534, 1696, 2383, 248, 1960]. Roles [1454]. Roots [557, 2409]. rosea [468]. rosmarus [334]. rostrata [2456, 1269, 1307, 1700, 658]. Rot [195]. Rotifer [97, 2025, 10]. Rotifers [381]. Round [654]. Routes [2372, 2516]. rubens [10]. ruckeri [641, 1109]. Ruditapes [2502]. Run [1485, 2195]. Running [236, 237, 2545]. Runoff [1395, 1491]. Runs [1335, 176, 227]. Rural [1126]. rusticus [1752, 1889]. Rutilus [2207]. Ryder [2418].

S [739]. S. [346, 133, 1614, 1822, 1746, 1598, 925, 1455, 404, 747, 451, 2334, 329, 1241]. sa [1749]. Sablefish [562, 2108, 1000, 1617, 659]. Sac [1052, 1379]. Sacramento [548]. Safe [2322]. Saffron [1051]. Sagittae [2358, 1440, 702]. Saguenay [761]. saida [580]. Saint [6, 1267, 677, 877, 819, 1767]. Saint-Laurent [6, 1267, 677, 1767]. saisonnières [870]. Salamanders [2174, 887]. salar [177, 364, 1187, 1538, 1584, 1037, 2247, 558, 724, 1973, 762, 2341, 2538, 839, 133, 719, 1453, 1107, 1317, 2133, 2441, 1125, 2454, 2131, 2248, 859, 1379, 2405, 1746, 1060, 1729, 2193, 2291, 2439, 1592, 343, 1945, 567, 1622, 2448, 446, 1902, 1322, 1845, 2470, 1344, 1283, 1994, 2601, 1327, 1435, 1316, 1992, 2159, 1832, 2367, 2368, 1181, 1656, 1078, 268, 1859, 103, 853, 1140, 2015, 496, 2623, 1278, 328, 329, 843, 1072, 1735, 1057, 1503, 1816, 497, 890, 1326, 2099, 921, 1123, 1163, 2354, 195, 1694, 1151, 494, 1637, 2290, 1241, 2404, 2113, 2449]. salar [390, 2569]. Salicylate [106]. Saline [1445, 1606, 2319, 1806]. Salinity [674, 2431, 1493, 1992, 1679]. salminicola [1056, 1409]. Salmo [1459, 1668, 2107, 1186, 2254, 2253, 177, 108, 1337, 364, 1187, 1538, 510, 1584, 1037, 1165, 2247, 13, 1340, 558, 1020, 1589, 413, 1260, 1887, 724, 1973, 1521, 1894, 2372, 762, 1744, 2341, 2538, 2356, 1689, 839, 2440, 133, 719, 405, 1372, 2125, 1719, 1453, 457, 1107, 1317, 450, 1614, 2133, 2441, 1125, 1160, 640, 2454, 2131, 1222, 1223, 2219, 197, 2248, 859, 897, 1379, 2405, 795, 1268, 1847, 1822, 1746, 1060, 1729, 2193, 2291, 2230, 2302, 1346, 2246, 686, 981, 765, 1645, 22, 30, 1354, 2439, 2012, 1259, 1853, 160, 1592, 343, 955, 289, 1945, 567, 1622, 2448]. Salmo [446, 1493, 1902, 1322, 1919, 186, 1845, 715, 710, 2470, 1344, 1283, 1994, 2601, 2359, 1688, 1829, 1830, 1505, 1327, 1195, 1675, 2060, 2362, 2044, 1435, 1316, 1992, 1708, 1312, 2342, 99, 813, 2159, 2315, 1832, 2367, 2368, 1944, 1181, 1656, 1494, 2379, 1078, 467, 913, 451, 331, 1532, 1190, 410, 268, 181, 1859, 1234, 103, 853, 1140, 2015, 2334, 2324, 496, 2489, 2168, 2623, 1749, 1278, 2011, 2120, 2361, 328, 329, 843, 1072, 738, 462, 862, 1523, 1735, 340, 1057, 1503, 1816, 8, 497, 890, 1326, 2099, 921, 1123, 1163, 2354, 195, 1694, 1067, 1213]. Salmo [2035, 2447, 1414, 2096, 2348, 1151, 2243, 494, 419, 1637, 2290, 1241, 2404, 1588, 2113, 2226, 818, 319, 2449, 390, 2569, 1454, 751, 389, 21, 265]. salmoides [1648, 1230, 506, 2610, 1596]. Salmon [1308, 177, 2162, 821, 996, 814, 984, 1305, 1431, 1466, 1286, 1534, 1851, 2026, 1882, 2258, 1187, 1422, 872, 1538, 151, 26, 2175, 2500, 1985, 1056, 583, 1584, 1825, 1037, 2247, 2249, 1492, 861, 1409, 342, 1723, 1335, 1141, 1396, 2524, 724, 1973, 1699, 823, 2004, 609, 1458, 762, 2341, 2538, 839, 1842, 133, 719, 1453, 2375, 1155, 2232, 1844, 796, 1107, 1317, 76, 1048, 2218, 1977, 2057, 2133, 2371, 2345, 2441, 1850, 2194, 1625, 560, 2454, 1913, 1910, 2131, 288, 159, 1857, 1858, 2248, 859, 1560, 1721, 1379, 2405, 2494, 1331, 1571, 606, 653, 1073, 1746, 1060]. Salmon [1729, 2193, 2291, 2431, 733, 544, 1253, 2040, 2488, 2439, 1742, 2151, 2530, 1592, 1287, 343, 2245, 2491, 593, 1102, 1945, 567, 1622, 2448, 20, 2265, 446, 1902, 1322, 781, 1961, 582, 1845, 705, 2195, 2470, 1344, 1283, 1994, 2601, 387, 484, 2616, 1900, 1674, 561, 107, 2251, 2163, 520, 1705, 600, 365, 1435, 176, 2508, 1838, 391, 971, 1316, 1992, 500, 1023, 2548, 2347, 2159, 1832, 2367, 2368, 675, 926, 1812, 1944, 785, 1673, 904, 2564, 1181, 1656, 703, 1360, 1078, 1565, 1805, 2604, 2005, 2266, 1483, 268, 1859, 71, 437, 718, 644, 891, 1238]. Salmon [1239, 1446, 1536, 2505, 103, 696, 695, 853, 1140, 2015, 496, 1208, 938, 1142, 1177, 2058, 2623, 1278, 499, 328, 329, 1199, 843, 1072, 2181, 1870, 935, 1282, 1735, 1057, 1503, 1816, 497, 890, 1326, 2099, 352, 921, 1123, 1163, 2329, 2354, 195, 864, 1694, 1145, 2438, 1938, 1804, 2384, 1194, 611, 102, 2349, 1151, 2531, 2383, 1495, 1496, 1569, 2119, 2528, 2039, 494, 817, 2054, 1637, 2290, 1028, 1241, 2404, 830, 2113, 1374, 818, 392, 1917, 1903, 2449, 390, 1860, 2569, 193, 702, 1152, 2275, 1403, 1929, 1930, 2612, 2516, 1960, 335, 1611, 1038, 558]. salmon [1327]. Salmon-Producing [1048]. salmonicida [464, 551, 726, 1406, 1182, 2463, 268]. Salmonid [1137, 1608, 2457, 1715, 520, 1564, 1943, 1841, 988, 1997]. Salmonidae [1135]. Salmonidés [1135]. Salmonids [2182, 1128, 333, 535, 2351, 1852, 501, 373, 1963, 422, 905, 1692, 467, 2541, 2383, 498, 194]. salmoninarum [1844, 1841]. Salmonine [1556]. Salmonines [1306]. salmoninus [92]. salmositica [1492]. Salt [441]. salvelini [861]. Salvelinus [646, 310, 2473, 2401, 2474, 2276, 2255, 346, 1703, 987, 2277, 509, 551, 726, 1744, 2356, 508, 2146, 1574, 1306, 2396, 450, 1709, 502, 1119, 2234, 2451, 2389, 2299, 507, 1984, 2337, 1052, 1987, 1346, 2246, 661, 2439, 1341, 1259, 2216, 320, 2448, 2265, 1863, 2183, 712, 2273, 282, 240, 291, 2171, 246, 1499, 2263, 503, 773, 2214, 1316, 2281, 2617, 2408, 1528, 245, 948, 2271, 2283, 2284, 912, 181, 247, 1623, 1235, 1523, 1057, 2391, 2395, 1751, 281, 844, 248, 1815, 1537, 1754, 1893, 2233, 1947, 2285, 2546, 2282, 255, 504, 244, 390, 2278, 2279, 2280, 257]. same [714, 1135]. Sample [283]. Sampled [2238]. Sampler [2190]. Samplers [741, 2213, 93]. Samples [2476, 1309, 2421, 1191, 709, 2063]. Sampling [633, 2469, 1972, 1202, 2618, 2098, 1212, 1437, 571, 1198, 122, 1472, 2158, 554]. Sand [1046, 1567, 2595, 1166, 1407, 397, 812, 2152]. Sandbar [1367]. Sandy [1034]. sapidissima [2535, 952, 1159, 1452, 1451, 1558, 395, 394, 1579]. Saqvaqjuac [1313, 1315, 1642, 1991, 2179, 2364]. Sardina [2602]. sardinae [1983]. Sardinella [2565, 318]. Sardines [2602]. Sardinops [2602]. Sargasso [2600]. Sarotherodon [877]. Sars [2167]. Saskatchewan [690, 2227]. Satisfaction [1279]. Satisfy [1546]. Saturation [1356]. Sauger [2328, 711, 1261]. saumons [1327]. sauvages [1327]. Saving [2065]. Savings [2124]. saxatilis [1004, 388, 1184, 1172, 2366, 19]. Saxitoxin [2344]. Scale [2175, 2500, 956, 609, 1552, 2000, 1076, 1845, 2310, 68, 1615, 707, 1843, 88]. Scaled [2256]. Scales [117, 567, 387, 948, 707, 2584, 714]. Scalings [2429]. Scallop [1979, 1828, 2024]. Scallops [591, 799, 1837, 2252]. Scandinavian [1660, 493]. Scanning [470]. Scarring [239, 442]. Scats [1426]. Scattering [627, 2511]. Scenedesmus [2031, 1339]. Schaefer [37, 158, 1252, 164]. Schedules [1161]. Schemes [2618]. schmidti [1387]. Schnute [1500]. Schools [2164, 2003]. Sciaenops [835]. Science [486, 2222]. Sciences [665]. Scomber [109, 399, 2064, 2197, 2154, 1324, 1321]. scombri [109]. scombrus [109, 399, 2064, 2197, 1321]. Scotia [1279, 558, 1276, 1770, 2582, 1043, 613, 1229, 1526, 1883, 672, 780, 433, 2515, 1994, 2601, 1044, 2296, 1244, 563, 2157, 2427, 2252, 466]. Scotian [1432, 1650, 816, 2574, 5, 2425, 2289, 268, 1772, 1572, 627, 651, 2377, 1358, 832, 850, 818, 1050]. Scotland [239]. Screen [427]. Screens [300]. Sculpin [725, 477, 1311]. Sea [177, 233, 2474, 2255, 980, 1538, 1776, 505, 1973, 242, 230, 1113, 231, 1371, 2248, 252, 458, 653, 2431, 228, 819, 251, 253, 343, 1191, 591, 799, 241, 604, 240, 258, 210, 515, 787, 260, 261, 471, 262, 1435, 250, 1501, 245, 227, 224, 2005, 223, 225, 247, 2623, 1199, 2235, 2236, 2591, 249, 1163, 1244, 596, 229, 205, 222, 243, 248, 1815, 25, 259, 2252, 1995, 254, 226, 1979, 1828, 1590, 244, 466, 669, 1345, 543, 881, 694, 264, 580, 2059, 376, 206, 142, 444, 1954, 975, 2600, 308, 1158, 263]. Sea-Age [1538, 2623]. Sea-Ice [1371]. Sea-Type [2431]. Seabird [2159]. Seabirds [74, 1158]. Seagrass [744, 942]. Seal [1912, 848, 918, 793, 1309, 871, 1568, 2177, 2534, 774]. Seals [357, 1854, 1399, 1031, 1393, 270, 2212, 2533, 2549, 297, 1426]. Sealworm [2492]. Search [2112, 1265, 2329]. Searching [745]. Season [786, 1667]. Seasonal [1292, 993, 762, 2311, 1113, 1883, 1475, 2323, 2131, 859, 1734, 2619, 62, 287, 84, 1808, 979, 784, 1008, 2085, 2084, 358, 1935, 2578, 1649, 669, 870]. Seasonality [1384]. Seasonally [2297, 1678]. Seasons [1816]. Seaward [606, 1322, 1870, 1738, 1929, 1611]. Seawater [106, 861, 1842, 97, 2378, 299, 593, 1102, 611, 2606]. Seawater-Adapted [2606]. Seaweeds [1501]. Sebastes [1146, 1598, 969, 925, 1455, 404, 747, 1232, 1429, 2101]. Seber [434]. Secchi [344, 1336]. Second [1471, 1324]. Secondarily [32]. Secondary [2174, 772, 2501, 2285, 401, 426]. secteur [1490]. Sectional [1308]. Sections [333, 1729, 2379, 1364]. sector [1490]. securis [854]. Sedentary [2230]. Sediment [2229, 1671, 1422, 477, 1670, 1408, 1509, 82, 1349, 2116, 110, 764, 822, 1986, 999, 1086, 52, 70, 2327, 1430, 945, 797, 2087, 1295, 1865, 1897, 2342, 749, 1401, 2406, 1678, 2181, 2382, 2242, 412, 1258, 2089]. Sediment-Degraded [999]. Sediment-Derived [70]. Sediment-Forming [2089]. Sediment-Lead [764]. Sediment/Water [1258]. Sedimentary [2126, 1730, 145, 2406]. Sedimentation [626, 1097, 1086, 2571, 1865, 372]. Sedimentological [178]. Sediments [2467, 313, 1249, 1233, 2217, 1610, 2140, 937, 1691, 2147, 829, 2323, 1323, 2412, 2103, 1595, 1819, 2007, 1683, 1566, 2145, 1120, 1931, 648, 1334, 80, 1011, 1421, 1400, 51, 353, 2478, 1196, 1215, 679, 720, 1908]. Seed [1978]. Seepage [2132]. Segregated [1259]. Segregation [661, 1623, 1267, 1267]. sein [677]. Seine [1253, 2355, 1937]. Selected [1347, 438, 2594, 829, 1759, 1629, 227, 421, 305, 570, 2487]. Selection [1781, 542, 1861, 1462, 1634, 272, 1742, 955, 1384, 1527, 1925, 943, 1884, 1832, 1180, 1072, 651, 2241, 358, 2393]. Selective [730, 1174, 1331, 653, 1814, 2045, 1162, 1903, 1270]. sélectives [2225]. Selectivities [2213]. Selectivity [2001, 1571, 735, 1167, 276]. Selenastrum [1297]. seleniris [413]. Selenium [775, 30, 113, 1011, 1014, 1386, 1012, 1013, 1469, 1525, 1525]. Self [1083]. Self-Cleaning [1083]. Semicontinuous [1686]. Seminatural [1859]. semireductus [161]. Semivoltine [1154]. Semotilus [1055]. Senecella [1154, 1811]. sensible [543]. Sensing [2319]. Sensitive [1563, 2367, 543]. Sensitivities [2154]. Sensitivity [1608, 2577, 1353, 1922]. Separable [1961]. Separate [982]. Separating [699]. Separation [807, 132, 1065, 1943, 1904, 6, 6]. Sept [319]. Sept-Îles [319]. Sequence [2461]. Sequential [1366, 2124, 1812]. Series [1308, 1654, 2110, 1417, 2228, 2434, 1802, 1006, 1869, 595, 2581, 1264]. Serological [641]. serovar [1709]. Serum [2146, 21]. Serville [2186]. Serving [1803]. Session [1797, 1799, 1796, 1795, 1798, 1794, 1793, 1792]. Seston [269, 1707, 1836, 78, 62, 2176, 468, 709, 2051, 11, 2076, 409]. Set [418, 2180]. SETCOL [459]. Settlement [1774]. Settling [173, 888, 2496]. Seulpins [708]. Seventy [2563, 1041]. Seventy-Fifth [1041]. Seventy-Two [2563]. Several [446, 121, 1706, 1803]. Sex [1563, 535, 1721, 1506, 35, 531, 382]. Sexing [284]. Sexual [1781, 984, 343, 1506, 1000, 290, 2099, 921, 192]. Sexually [1584]. Seymour [2624]. Shad [2535, 952, 1159, 1452, 1451, 395, 394, 1579, 1558]. Shadowing [148]. Shagawa [1022]. Shallow [781, 1416, 916, 1117, 2478]. Shape [1654]. Shark [1367]. Sharks [793, 2134]. Sharp [404]. Sharp-beaked [404]. shasta [941]. Sheet [558]. Shelf [433, 787, 2511, 840, 1448, 1650, 618, 2574, 1566, 2310, 5, 2425, 2289, 1432, 1772, 1572, 651, 2377, 1358, 949, 1657, 832, 850, 1050]. Shell [336, 2336, 612, 2529]. Shellfish [2507, 2006, 306, 799, 902]. Shelter [1778]. Shield [1610, 1033, 2140, 2147, 64, 2323, 1071, 2506, 1410, 62, 1549, 773, 1487, 47, 1570, 758, 954, 1599]. Shift [2214, 843]. Shifting [1407]. Shifts [662, 332, 1415, 2094]. Shiner [2512, 868, 2452, 2011]. Shiners [573]. Ship [1438]. Shoal [2572]. Shock [2121]. Shoots [2409]. Shore [319]. Shoreline [1085, 2312]. Short [1456, 1422, 2303, 1217, 2484, 2386, 2392, 1939, 1296, 2251, 365, 452, 2581, 2598, 2397]. Short-Finned [1296]. Short-Term [2303, 1217, 2484, 2386, 2392, 1939, 2251, 2598, 1456, 1422, 365, 452, 2397]. Shortfin [977, 581]. Shrimp [128, 120, 1933, 1718, 420]. Shrinkage [674, 2588]. Si [2123, 1547]. sicilis [358]. Side [905]. Side-Channel [905]. Sierra [2346]. Sieve [875]. Sigmoid [2503]. Signal [2359, 2164]. Significance [1254, 2568, 1526, 501, 2102, 2470, 1575, 2209, 2212, 2243, 1225]. Significant [1680, 1907]. Silica [2217, 2261, 1421, 1553, 353]. Silicon [841]. Silt [648]. Silt-Clay [648]. Siltation [178]. Silver [202, 1990, 866, 2574, 1738, 2306]. Silverside [1563, 634]. SIMCA [1474]. Similarities [1002]. Similarity [2461, 291, 2613]. Simple [1833, 11, 459, 1707, 373, 149, 1543]. simplex [1666]. Simplified [1476, 1846, 663, 550]. simulans [2186]. Simulated [1835, 2481, 1356, 955, 29]. Simulation [998, 1980, 776, 440, 2472, 786, 2062, 2048, 2534, 1353, 164]. Simulations [2415, 10]. Simuliid [2128]. Simuliidae [2522, 1655]. Simultaneous [2475]. Simultaneously [628]. Single [128, 816, 878, 2296, 1916]. Single-Transducer [1916]. Sinking [459]. Site [1029, 1052, 487, 888, 1734, 1463, 1006]. Site-Specific [1734]. Sites [1160, 2227]. Sitka [1560]. Situ [2106, 1356, 69, 1126, 448, 2597, 1166, 2123, 1320, 1635, 2077, 2027, 796, 2202, 131, 87, 596, 1824, 2231, 56, 1862]. Six [2274, 2506, 1968, 1106]. Size [1790, 283, 1955, 1466, 1851, 26, 583, 619, 983, 801, 2429, 779, 2593, 901, 1154, 2067, 2055, 970, 77, 1197, 1458, 2453, 934, 2002, 2220, 820, 2232, 124, 76, 116, 2521, 1850, 986, 1987, 1702, 2544, 780, 1571, 653, 2431, 1476, 819, 688, 1742, 2491, 374, 566, 1821, 1101, 1449, 2617, 2288, 2464, 407, 2580, 1516, 427, 2590, 1232, 959, 594, 2045, 2597, 2185, 694, 2030, 1778, 499, 1167, 1916, 2600, 431, 1928, 2444, 2445, 453, 1145, 2366, 1804, 707, 192, 308, 958, 1724, 432, 2034, 1995, 1225, 2528, 1524, 1477, 152, 358, 2051, 2226, 2584, 2180]. Size [1903, 1270, 1403, 1864, 1038]. Size-Biased [2584]. Size-Class [2185]. Size-Dependent [2597]. Size-Limited [1524]. Size-Related [1449]. Size-Selective [653, 2045, 1903, 1270]. Size-Specific [2580]. Size-Structure [819]. Size-Structured [1928, 2444, 2445]. Sizes [955, 1493, 1198, 1138]. Skin [783]. Skipjack [1952]. Skuse [2513]. Slimy [1311]. SLIS [264, 205, 260, 262]. Slope [2108, 627, 840]. Slot [1473]. Slow [1269, 2020, 1493, 200]. Slow- [1269]. Slower [2365]. Slub [2028]. Slurry [1735]. Small [1470, 60, 1396, 1635, 1610, 2335, 1076, 1382, 1850, 1433, 28, 2622, 1212, 2192, 1852, 1071, 2056, 2457, 2245, 1747, 1748, 188, 155, 2470, 304, 1200, 1423, 843, 1072, 2181, 2363, 1363, 2587, 1192, 2298, 1843, 1314]. Small-Scale [1076, 1843]. Smallmouth [2486, 1609, 2408]. Smeaton [1035]. Smelt [2046, 847, 1415, 1602, 1416, 1288]. Smelter [203]. Smoking [647, 643]. Smolt [177, 1584, 1037, 446, 1838, 971, 1316, 1992, 1565, 437, 2226, 2449, 1611]. Smolt-to-Adult [437, 2226]. Smoltification [2368]. Smolting [177, 1107, 1317, 2405, 890]. Smolts [1973, 2248, 2530, 2491, 2265, 1322, 1845, 1483, 644, 102, 2584]. Snail [2566]. Snapper [2358]. Snow [1953, 929, 1716, 1310, 2335, 2062, 1991, 1267]. Snowcover [1491]. Snowmelt [2010, 1052, 2338, 2562]. Snowpack [2042]. Soak [2555]. Socio [2520]. Socioeconomic [1199]. Sockeye [2162, 151, 26, 1825, 2249, 861, 342, 1723, 2004, 609, 1458, 2371, 2194, 1625, 560, 159, 1560, 2431, 1287, 2195, 1705, 600, 1023, 675, 926, 1812, 1944, 644, 1238, 1239, 1446, 1536, 2058, 2349, 2039, 1903, 193, 702, 2612, 2516, 1960]. Sodium [2234, 2281, 2279]. Soft [2253, 336, 822, 1464, 145, 1580, 53, 1414, 2348, 442]. Soft-Bottom [145]. Soft-Shell [336]. Soft-Water [53]. Softwater [1021, 1469]. Sol [353]. solar [818]. Sole [1226, 2339, 786, 1823, 2048, 1481]. Solids [1806, 768]. Soluble [13, 2327, 1104, 2145, 1841, 478]. solubles [478]. Solution [458, 721, 912]. Solutions [2514]. Solvents [1385]. Solving [1161, 546]. Somatic [983, 2271, 2366, 2479]. Somatotropin [2375]. Some [146, 2415, 370, 1787, 650, 2125, 613, 1739, 1247, 1529, 2301, 107, 1218, 1705, 182, 386, 2539, 298, 2611, 1156, 1145, 1728, 259, 817, 164, 818, 1622, 1283, 1023, 1140, 237]. Sonar [2066, 1916]. Sorbitol [353, 21]. Sound [1953, 814, 2618, 1371, 1607, 2339, 471, 2511, 1431, 1260, 1887, 176]. Sounder [2238, 87, 554]. Sounds [1156]. Source [2507, 845, 945, 1613, 1803, 1678, 98, 1704, 335]. Sources [923, 1221, 182, 531, 85, 1872, 1013]. Souris [2487]. sous [2306]. South [690, 433, 1775, 1447, 2166, 1760, 1774, 2460, 961]. South-Central [2460, 961]. Southeast [2572, 2564, 1035, 1560, 537, 2528, 2612]. Southeastern [2129, 142, 444, 1954, 1667]. Southern [1098, 840, 1305, 1431, 1286, 1093, 1096, 2582, 1811, 2184, 1092, 829, 2318, 1097, 1986, 78, 332, 339, 1086, 1100, 1088, 1087, 2373, 2343, 287, 763, 1881, 1085, 1084, 1090, 1727, 1651, 122, 1094, 2150, 564, 1657, 1099, 1321, 1917, 1091]. Southshore [2084]. Southward [320]. Southwest [1279, 2515, 563]. Southwestern [1276, 1770, 2430, 1267, 1317, 1229, 1526, 2427]. Soya [2595]. Soybean [950]. sp [892]. Space [900, 1822, 1259, 2310, 1420]. Space- [2310]. Spacing [1717, 2000]. Spatial [2529, 2509, 1730, 2147, 117, 2246, 2550, 1595, 1636, 191, 2565, 2086, 2478, 1138, 1637, 1374, 2240]. Spatially [1980]. spatio [677]. spatio-temporal [677]. spatio-temporelle [677]. Spawn [1958, 706, 1472]. Spawning [1271, 1680, 2370, 510, 1335, 2572, 2524, 1974, 2454, 394, 1052, 772, 1822, 2193, 2291, 2373, 1476, 20, 740, 786, 705, 1210, 2609, 221, 1195, 503, 211, 1000, 1489, 1983, 2477, 1579, 227, 1416, 2532, 1874, 2438, 1139, 590, 2039, 19, 1782, 2576, 1923]. Spawnings [339]. Special [1671, 1557, 2552, 2059, 1726, 487, 1622, 221, 1094, 42, 43]. Specialist [1633]. Specialized [1906]. Speciation [1529, 2621, 766, 2168]. Species [2486, 1871, 625, 1093, 156, 2055, 2127, 985, 1263, 708, 1573, 1331, 70, 2506, 622, 878, 752, 2615, 1624, 1889, 836, 1693, 2214, 2316, 763, 634, 1867, 2565, 1122, 1512, 2295, 2209, 1555, 1931, 2406, 2455, 207, 755, 1515, 2164, 340, 2600, 1827, 965, 972, 736, 308, 1024, 1933, 513, 498, 2298, 1325, 939, 1596, 804, 2207, 1626, 2152, 1582, 472, 714]. Specific [1734, 2580, 1008, 431, 850, 1908, 1465]. Specificity [2570]. Specimens [208]. Speckled [725]. Spectra [1724]. Spectral [2215]. Spectroscopy [2433]. Spectrum [2051]. Speed [1436, 77, 1698, 2213, 1860]. Speeds [187, 141]. Sperm [283, 2296]. Spermatogenesis [532, 1834]. Sphaerium [854]. Spherical [2052]. spicatum [1408]. Spider [2442]. Spill [2022, 928]. Spine [1471]. Spiny [1471, 1775, 1777, 1760, 1762, 1774, 1766, 2435, 1765]. Spiralling [400]. Spisula [444]. Spore [266, 1564]. Spores [132, 647]. Sport [2299]. Sports [1531]. Spot [942]. spotted [2035]. spp [1146, 1954, 1429, 1938, 2152]. Spring [1271, 1470, 1043, 2010, 2323, 1052, 2338, 1700, 2373, 740, 1946, 1900, 1280, 1044, 2477, 1491, 111, 1145, 1696, 1923, 1611]. Spring- [2373, 2477]. Spring-Spawning [1923]. Springwater [2258, 2056]. Spurious [2467, 2614]. Squalus [1471]. Squamish [1463]. Squares [327, 2453, 164]. Squawfish [2155, 333]. Squid [1296, 1966, 35, 605]. squinado [2442]. St [441, 2184, 1267, 376, 127, 780, 1767, 2196, 2609, 1318, 1890, 1441, 2235, 323, 1321, 319]. St. [74, 672, 1620, 2360, 1416, 1288]. Stability [1229, 2318, 183]. Stable [2403, 2575, 1461, 2299, 945, 279, 439, 769]. Stackhouse [2184, 2185]. Stage [1950, 2287, 2492, 1564, 386, 941, 697, 919]. Stages [476, 762, 388, 2506, 343, 2616, 2615, 2477, 2093]. Stain [904]. Staining [1844]. Stamina [872]. Standard [692]. Standardization [1806]. Standardized [99]. Standardizing [2555]. Standing [1445, 1480, 1410, 1952, 840]. Standpipe [2411]. Starry [650, 783, 594]. Starvation [221, 1872, 1644, 1609]. State [963, 2126, 951, 853, 478]. States [2129, 887, 506]. Stations [1041, 2084]. Statistic [2210]. Statistical [681, 2000, 1329, 96, 601, 121, 2209, 2049, 2505, 1424, 1183, 1593, 1653, 608]. Statistically [439]. Statolith [1296]. Statoliths [1590]. Status [2482, 1306, 2070, 1544, 1551, 1600, 1414, 1434, 1284, 415, 1628, 2278, 883]. Ste [187, 2479]. Steel [1944]. Steelhead [1459, 1260, 1689, 1853, 1195, 1675, 2143, 1190, 2011, 2361, 1067, 2226, 2584, 319, 389]. Steeppass [1825]. stellatus [650, 783, 594]. stenolepis [1143, 1424]. Sterility [252]. Sterilization [458]. Sterilized [2266]. Steroid [529]. Steroids [531]. Stickleback [2142, 1807, 2043]. Sticklebacks [2400, 1162]. stigmatura [894, 838]. Stikine [1458]. Still [84]. Stimulation [89, 1544]. Stizostedion [2328, 2169, 514, 2552, 2521, 139, 1988, 1575, 1909, 711, 2420, 2, 1261, 2608]. Stochastic [1004, 2220, 1513, 1418, 824, 1941]. Stochasticity [2301]. Stock [495, 491, 2156, 2162, 1789, 1305, 1431, 2026, 480, 2247, 486, 336, 1161, 1787, 2110, 724, 490, 609, 998, 1029, 2358, 1758, 393, 1798, 1974, 1063, 1868, 1878, 1721, 501, 1338, 1476, 1980, 487, 444, 516, 105, 1191, 1410, 721, 1178, 1500, 1952, 740, 1821, 485, 387, 2614, 261, 2616, 2526, 380, 521, 2165, 1265, 500, 1065, 1888, 1455, 585, 1516, 785, 2143, 2289, 1594, 2602, 496, 1278, 1583, 517, 1804, 1359, 378, 523, 2349, 379, 129, 1514, 1982, 1370, 1922, 2613, 1960, 1307, 1862, 2612, 479]. Stock-Isolating [487]. Stock-Recruit [2156]. Stock-Recruitment [1787, 724]. Stocked [1985, 640, 2389, 1280, 363, 2334]. Stocking [108, 877, 1853, 2380, 255]. Stocks [1882, 872, 483, 603, 509, 978, 511, 514, 2274, 1458, 507, 159, 1721, 780, 1369, 1915, 518, 512, 2216, 878, 190, 699, 1228, 1821, 2616, 488, 828, 521, 262, 345, 1705, 506, 1932, 8, 493, 497, 921, 1123, 2354, 831, 452, 494, 513, 1374, 2612]. STOCS [480]. Stomach [1173, 1722]. Stone [1750]. Stone-dwelling [1750]. Stonefly [1698, 2551, 1727]. Stony [837]. Storage [2170, 188, 1463, 2160]. Stored [1129, 995]. Stores [2250]. Strain [268]. Strains [502, 2463, 1708, 1815, 504]. Strait [893, 1566, 879, 867, 1604, 14, 1882, 1399, 882, 772, 883, 881, 884, 880, 1387, 2257, 1472, 840, 949, 885]. Strait/Foxe [1566]. Strategies [1462, 1898, 34, 2198, 1622, 1210, 2301, 828, 1633, 748, 1551, 1982]. Strategy [712, 520]. Stratification [65, 1453, 2090, 2071]. Stratified [1976, 60, 179, 1702, 1352, 1231]. Stratigraphy [1671]. Straying [1142]. Stream [310, 1274, 2333, 725, 1332, 1661, 573, 82, 1197, 934, 1744, 1224, 719, 450, 2318, 1850, 986, 2323, 2170, 332, 1150, 2018, 692, 606, 687, 2230, 2302, 2056, 2151, 2530, 348, 955, 1484, 2343, 1126, 982, 1361, 742, 1376, 1437, 1731, 904, 950, 112, 2379, 125, 1587, 1964, 1658, 455, 655, 1363, 1728, 15, 972, 1121, 2150, 2587, 2531, 1569, 2119, 1826, 1704, 2240, 1454, 1930, 2245, 1351, 2066]. Stream-Dwelling [1454]. Stream-Resident [1930]. Stream-Type [1569, 2119]. Streambeds [837]. Streamflow [2579, 352]. Streams [196, 1985, 1260, 2262, 1396, 2524, 2326, 2325, 2144, 295, 2351, 1883, 78, 1212, 2192, 887, 2457, 1297, 1221, 1283, 1148, 1053, 1394, 227, 2315, 1808, 304, 1692, 140, 400, 1805, 1942, 2334, 1740, 2114, 2115, 2489, 1646, 2181, 46, 2016, 944, 2493, 2528, 1219, 2113, 1403]. Streamside [926]. Streamwater [2562, 1577]. Strength [952, 1159, 2223, 1157, 1111, 1679, 369, 629]. Strengths [2037]. Stress [1205, 1337, 2304, 1605, 1863, 997, 448, 2536, 2095, 1465, 1735, 102, 1627, 599, 1482, 1628, 2279, 2280]. Stress-Related [1205]. Stressed [2234, 289, 54, 1018, 2233, 2227]. Stressors [2380]. Strike [1032]. Striped [1004, 388, 1184, 1172, 2366, 19]. Strips [752, 1692, 140]. Strongly [179]. Strongylocentrotus [980, 1244, 25, 466]. Strontium [1296, 948, 1997, 714]. Structural [1329, 2455, 1121]. Structure [2376, 489, 1205, 2535, 505, 1887, 2593, 1796, 1839, 1741, 393, 2092, 1631, 1119, 142, 911, 2622, 1648, 1721, 473, 2246, 819, 470, 105, 975, 568, 1053, 345, 642, 2617, 2289, 2135, 1144, 1896, 2602, 718, 959, 2579, 2185, 1623, 475, 1583, 655, 172, 831, 2101, 2016, 2587, 2034, 2566, 292, 2546, 2298, 849, 81, 589, 899]. Structured [1788, 1004, 34, 1467, 1513, 828, 1378, 1928, 2444, 2445, 824, 1514, 1891, 893]. Structuring [678, 2363]. Students [541]. Studied [470]. Studies [1681, 1833, 55, 956, 580, 1552, 469, 2553, 1191, 1863, 2007, 2070, 852, 96, 1925, 711, 1613, 718, 305, 1515, 73, 274, 131, 1827, 1769, 2233, 2252, 1824]. Study [947, 623, 1439, 2274, 1542, 393, 816, 1648, 2139, 2040, 888, 53, 440, 1549, 109, 448, 875, 1455, 2143, 2322, 2197, 111, 2534, 2173, 1626]. Studying [1976]. Stunted [1574]. Sturgeon [1411, 2204, 1]. Style [996]. Subadult [1294]. Subalpine [1719, 420, 98]. Subarctic [1703, 1248, 2221, 1911, 965, 1600, 899]. subarctique [899]. Subareas [308]. Subcellular [2074, 642]. Subchronic [2322]. Subice [2158]. Subitaneous [204]. Subject [1512, 1133]. Subjected [108, 1528, 2527]. Sublethal [2253, 1179, 2125, 2183, 710, 846, 2599, 1537]. Sublethally [1030]. Submarine [2310]. Submerged [2032, 1905, 2116, 122]. Submersed [1408, 1336, 2021, 1889]. Submersible [852]. Subsampling [663]. Subsequent [1503, 1947, 1644]. Substocks [1133]. substrat [899]. Substrata [1166]. Substrate [620, 477, 1509, 934, 1698, 2302, 470, 781, 943, 858, 648, 1727, 2527]. Substrates [880, 950, 1587, 1034]. Substratum [1166, 1892]. Subtidal [980]. Subtropical [1993, 1219]. Subyearling [1523]. Success [2437, 1856, 897, 2286, 1816, 655, 2460, 2039, 19]. Successes [1006]. Succession [332, 1209, 1343, 2084, 899]. Successive [1484]. Sucker [693, 1849, 203, 2250, 2536, 1936]. Suckers [1601, 775, 2410, 1136]. Suction [2469, 2158]. sud [1267]. sud-ouest [1267]. Sudbury [2256, 2563, 2025, 275, 834]. Sufficient [1687]. Suggestion [1664]. Suitability [819, 1473]. Suitable [1192, 2017]. suite [1862]. Sulfate [2078, 2077, 2596, 162]. Sulfide [1049]. Sulfite [1354]. Sulfonamide [770]. Sulfoxide [95, 1203]. Sulfur [856]. Sulphate [1363]. Summary [1797, 1799, 1796, 1795, 1798, 1794, 1793, 1792]. Summer [65, 1546, 75, 1675, 1450, 2315, 1832, 183, 867, 1072, 1145, 2071, 1696]. Sunfish [779, 180, 2094, 2523, 954]. superba [875]. Superior [515, 248, 508, 1530, 507, 246, 227, 315, 1647, 1544, 2005, 247, 229, 602, 222, 2042, 267, 513]. Supersaturation [1715, 2171, 1280]. Supplementation [1268]. Supplies [1892]. Supply [1687, 586, 2224]. Support [1306, 289, 2091, 1324]. Suppression [1011]. Surf [444]. Surface [269, 2147, 1880, 2357, 2190, 289, 1978, 1124, 966, 564, 936, 1908]. Surfaces [875]. Surficial [1566, 1931]. Surplus [114]. Survey [2026, 1723, 508, 422, 2289, 831]. Surveys [2483, 1472, 554]. Survival [1776, 1990, 1451, 2521, 314, 919, 502, 1097, 2131, 2389, 197, 1052, 2338, 101, 1853, 1845, 1344, 1283, 503, 2251, 874, 50, 1667, 365, 2508, 68, 2420, 2283, 2465, 1565, 2266, 718, 1381, 1735, 1586, 988, 2445, 1694, 2438, 1067, 1535, 1324, 102, 1815, 2039, 2533, 817, 757, 2226, 2584, 138]. SUS [261]. Susceptibility [973, 1709, 1031, 1184, 736]. Suspended [1422, 558, 477, 746, 1986, 1747, 1011]. Suspension [2221]. Sustain [1687]. Sustainable [664]. Sustained [2182, 2232, 1208]. SW [776]. Sweden [2312, 238]. Swedish [2229, 1813]. Swimbladder [894, 838, 2100, 2600]. Swimbladdered [2600]. Swimming [2401, 1436, 872, 187, 1768, 621, 2100, 200, 1312, 1078, 605, 1208, 2384, 1496, 141, 1860]. Switches [2512]. Symmetric [1914]. Sympatric [719, 1732, 1346, 2246, 1195, 318, 2595, 1626]. Symposium [264, 480, 205, 479, 206, 879, 1540, 1554, 263]. Synchrony [2581]. Syndrome [1900]. Synopsis [1554]. Synoptic [2066]. synthèse [1307]. Synthesis [2077, 1294, 780, 1765, 748, 1010, 263]. Synthetic [1529]. System [1912, 2303, 2065, 2575, 1839, 690, 820, 1461, 1382, 538, 1734, 1410, 2091, 1838, 1528, 1468, 2621, 2475, 738, 1753, 1010, 1011, 1014, 1012, 1013, 2017]. Systematics [1289]. Systemic [350]. Systems [2303, 1886, 1529, 378, 750].

Table [356, 1183]. tacons [1125]. Tactics [1032]. Tadenac [773]. Tadpoles [1439]. Tag [434, 1362, 2525, 713, 2172, 2159]. Tagged [946, 1276, 1771, 2582]. Tagging [2206, 825, 1952, 12, 1026]. Tags [2162, 1180, 1933]. Tailings [1136]. Taking [2066]. Taku [2564]. Tallow [335]. tamarensis [902]. Tangential [2476]. Tanks [1535]. Tanner [1954, 308]. Tannins [2493]. Tantaré [1119]. Tanypodinae [1093]. Target [2037]. Tasersuaq [1623]. Task [260, 261, 262]. Taupo [1193]. Taurine [2033]. Taurocholate [1919]. Taxocenes [1053]. Taxonomic [1302]. Taxonomy [838, 2602]. Technical [1152]. Technique [1349, 284, 132, 387, 709, 1166, 837, 192, 2237]. Techniques [1723, 807, 633, 174, 2057, 1858, 776, 1598, 38]. Technological [1122]. Technologically [459]. Technology [2463]. Teknar(R) [1709]. Telemetry [1312]. Teleost [532, 527, 1733, 531, 1679, 1032, 1596]. Teleosts [642, 530, 1130, 525]. Temiskaming [1582]. Temperate [1277, 1319, 896, 1185, 800, 1450, 1053, 1220, 965, 1473, 2224, 1215]. Temperature [678, 2370, 2019, 2473, 1466, 1436, 151, 2122, 573, 2428, 542, 1563, 2078, 1372, 2228, 1806, 733, 544, 1447, 289, 2448, 1902, 1322, 1365, 786, 1435, 96, 601, 296, 1708, 1890, 2271, 468, 1565, 853, 594, 225, 2011, 2030, 965, 966, 707, 1504, 1995, 292, 939, 1176, 2207, 720, 389]. Temperature-Dependent [296]. Temperature-Induced [289]. Temperature-Mediated [678, 2428]. Temperature-Sensitive [1563]. Temperatures [26, 115, 1060, 2151, 2530, 446, 445, 715, 710, 787, 1002, 2090, 2235, 2236, 2591, 570, 1728, 757, 2113]. Temporal [2529, 2247, 357, 2509, 394, 117, 473, 2550, 1595, 2565, 2478, 677]. Temporary [1852]. temporelle [677]. Temps [1327]. teneur [714]. Tenny [1801]. Tension [156]. Teratogenesis [282]. Terebellidae [3]. Term [2253, 872, 2303, 2397, 2050, 1217, 2484, 2386, 1202, 2392, 1939, 1624, 2251, 2283, 2284, 1881, 1802, 149, 1556, 2285, 2598, 1456, 1422, 365, 452]. terme [2225]. Terminal [1716, 864]. Terrestrial [2419]. Territorial [1422]. Territories [1399]. Territory [318, 1600, 281]. Tertiary [1631]. tertiolecta [359]. Test [1717, 2303, 92, 2321, 822, 1875, 752, 975, 2251, 2365]. Testing [2006, 147, 1023, 2062, 1944]. Testosterone [1977]. Tests [2351, 692, 1812]. Tetra [1201]. Tetra-Chloroguaiacol [1201]. Tetra-Chloroveratrole [1201]. Tetrahymena [751]. Tetrodotoxin [2344]. TFM [628]. Th [2442, 1865]. Th/ [2442]. Thalassoma [1652]. Their [623, 1882, 533, 650, 1154, 2456, 2127, 682, 1687, 1076, 1974, 127, 2098, 1913, 395, 2412, 819, 1598, 291, 2214, 763, 456, 68, 2128, 304, 703, 1201, 798, 671, 1432, 2602, 1364, 416, 87, 1151, 1869, 1196, 1391, 1241, 2513, 2568, 714, 2596, 2246, 261, 2314]. Theoretical [1787, 876, 2493]. Theory [680, 2194, 272, 1073, 1500, 260, 1660, 1265, 1376, 1303, 1139]. Theragra [623, 2059, 142, 188, 2416]. Therapy [770]. There [7]. Thermal [201, 2127, 375, 662, 347, 1087, 982, 123, 183, 2090, 570, 2212, 834, 2089]. Thermally [2014, 1702, 1919]. Thermocline [662]. Thermodynamic [367]. Thermogravimetric [2089]. thermophila [751]. Thermoregulation [2014]. Thermoregulatory [597]. these [1679]. Thickness [702]. Thin [1904]. Thin-Layer [1904]. Thingvallavatn [2273]. Thinning [1612]. Thiourea [457]. Third [2492]. Third-Stage [2492]. Thirteen [2486]. thomasi [1548]. Thompson [1860]. Three [1871, 1254, 156, 978, 1141, 2325, 2174, 1155, 2147, 856, 2410, 236, 2616, 1624, 1040, 211, 2380, 1655, 858, 1727, 2181, 924, 48, 565, 1138, 757, 1325, 2610, 1982]. Three-Dimensional [1155, 1727]. Threespine [2400, 2142, 1807, 2043]. Threshold [1596]. Throughout [1444]. Thunberg [2306]. Thunnus [2461, 1440, 1722, 1364]. thuringiensis [1709, 2577]. Thymallus [2112, 1897, 588]. thynnus [1364]. Thyroid [2616, 2615, 1992, 2368]. Thyroidal [1020, 1589]. Thysanoessa [1248]. Tidal [2050, 1511, 1371, 2619, 1416, 867, 942, 1801]. Tidally [2619, 2620]. Tide [1001, 679]. Tiger [2174]. tiglium [1978]. tigrinum [2174]. Tilapia [1230, 2173]. Time [1308, 2474, 583, 2065, 2110, 128, 2228, 2434, 1485, 2614, 2310, 1420, 1921, 1173, 2095, 916, 328, 2090, 1282, 1802, 546, 1006, 1869, 595, 2581, 1264, 1514, 1327]. Time-Invariant [1514]. Time-Saving [2065]. Time-Scale [2310].

Time-Series [1308, 1869, 595]. Timed [2508]. Times [1335, 2140, 314, 1089, 1282, 2555, 587]. Timing [510, 2559, 653, 2193, 2431, 2530, 2491, 1322, 941, 864, 1139]. Tin [1835, 589, 1211]. Tissue [1990, 2433, 2088, 550, 478]. Tissues [414, 13, 2134, 1196, 1828]. tissulaire [478]. Titrations [1510]. Tobin [2227]. Tofino [2556]. Tolerable [686]. Tolerance [2304, 775, 1950, 2431, 2402, 1992]. tomcod [1008]. Tool [2222]. Top [2135]. Top-down [2135]. Topographical [2212]. Torrent [477]. Total [619, 344, 722, 2192, 1112, 2145, 768, 1054, 1975, 1864]. Towing [77]. Tows [1620]. Toxaphene [999]. Toxic [2320, 386, 1414, 1482, 679]. Toxicant [767]. Toxicants [1030, 1329]. Toxicities [1689]. Toxicity [1497, 2013, 2356, 1179, 405, 1372, 388, 692, 1464, 22, 1849, 1580, 27, 715, 710, 1978, 1688, 1532, 1230, 2322, 1127, 2168, 1899, 340, 1386, 149, 1400, 1957, 902, 419, 849, 2167, 589]. toxicodynamics [1749]. toxicodynamique [1749]. Toxicologtcal [798]. Toxicology [157]. Toxin [2507, 2589]. Toxins [271, 2344]. Trace [2386, 2358, 2412, 1580, 53, 1819, 1510, 1035, 1196]. Tracer [1019, 2007]. Tracers [1461]. Trachylina [200]. Tracings [1905]. Track [57]. Tracking [558, 2417, 1962]. Tract [2155, 2596, 1291]. Trade [833]. Trade-offs [833]. Training [2012]. Traits [2247]. Trajectories [957]. Transaminase [21]. Transducer [1916]. Transects [2605]. Transfaunations [947]. Transfer [1912, 2328, 983, 278, 2133, 9, 299, 1940, 2495]. Transferred [271]. Transferrin [1637]. Transformation [1584, 1037, 1999, 446, 1838, 1992, 1611]. Transformations [2056, 2079, 1045]. Transient [2153]. Transit [1237]. Transition [156, 919, 398, 1072, 111]. Translation [1668]. Transmission [1703, 2433, 2083, 363, 2270]. Transmitters [1312]. transmontanus [1411]. Transparency [2180, 834]. Transplant [1567]. Transport [108, 869, 985, 1610, 1920, 676, 932, 1323, 20, 2281, 2176, 1045, 1510, 2541, 2478, 1237, 1035, 1718, 1339, 1176, 1649]. Transportable [2017]. Transportation [102]. Trap [1620, 2330, 2180]. Trapability [964]. Trapnet [1380]. Traps [584, 2213, 2558, 2561, 122, 2555]. Trawl [2465, 1024, 804]. Trawlers [1448]. Trays [669]. Treated [32, 2394, 2387, 2317]. Treatment [2397, 1224, 2386, 2392, 2393]. Treatments [628, 226]. Trend [604]. Trends [1202, 1595, 2614, 1921, 2049, 1568, 368, 1556]. très [543]. Tri [2459, 1082, 1201, 1965]. Tri- [1082, 1201, 1965]. Tri-N-Butyltin [2459]. Triactinomyxon [1564]. Triacylglycerol [2586]. Trials [1694]. Tributaries [1852, 1746, 630, 329, 229]. Tributary [1822, 1416, 1577, 2531]. Tricaine [615]. Trichoptera [1218]. tridentata [2407, 2097, 1048, 1874]. tridentate [232]. Triphosphatase [389, 550]. Triphosphate [826]. triplicata [1047]. Triploid [1187, 1128]. Triploidy [988, 2384]. Tritiated [69]. Trophic [923, 2482, 963, 2126, 951, 278, 1461, 976, 722, 1734, 2195, 1807, 2043, 1697, 959, 2259, 655, 1600, 1434, 1284, 415, 2502]. Trophic-Dynamics [923]. trophiques [2502]. Trophy [1855]. Tropical [1185, 2493, 2557]. Troughs [32]. Trout [1459, 646, 1668, 2107, 310, 1186, 2254, 2253, 108, 1337, 2401, 2104, 2276, 2255, 510, 987, 1165, 2277, 13, 1340, 509, 1020, 1589, 413, 1260, 1887, 1521, 551, 726, 1894, 2372, 1744, 2356, 1689, 2596, 2440, 133, 508, 405, 1372, 2125, 1719, 1574, 457, 1306, 2396, 450, 116, 730, 1614, 2551, 1709, 502, 1119, 1160, 2234, 640, 1222, 1223, 2219, 2389, 2299, 507, 1984, 2337, 197, 1052, 1987, 897, 795, 1268, 1847, 1822, 1746, 2230, 2302, 1346, 2246, 686, 981, 765, 1645, 22, 30, 1354, 1341, 2012, 1259, 1853, 160, 2216, 955, 289, 374, 567, 2448, 1493, 1919, 282, 240, 186]. Trout [715, 710, 2171, 2607, 2359, 1688, 1829, 1830, 246, 1505, 1195, 1675, 1499, 2263, 503, 773, 2060, 2362, 520, 2044, 1280, 1316, 2281, 2617, 1708, 2408, 1528, 1312, 2342, 99, 813, 363, 245, 2315, 948, 2283, 2284, 1944, 2143, 1494, 2379, 467, 913, 451, 331, 1532, 1190, 410, 181, 1234, 2334, 2324, 2489, 247, 2168, 2011, 2120, 738, 462, 862, 1235, 1523, 340, 1057, 8, 2391, 2395, 1067, 1213, 1751, 2035, 2447, 1414, 2096, 2348, 844, 248, 553, 1815, 1537, 1754, 1893, 2233, 1947, 2243, 419, 2285, 2546, 1241, 1588, 2282, 255, 2226, 2584, 504, 244, 319, 390, 1454]. Trout [751, 647, 2278, 2279, 2280, 2570, 257, 21, 265, 1749, 1624, 318]. Trouts [736]. True [2421]. Truite [265, 1749]. truncatus [1275, 927, 90, 928, 1304, 1627]. trutta [364, 1744, 450, 1614, 640, 1822, 2230, 2302, 567, 2448, 2379, 2489, 8, 2035, 2243, 1241]. Trypanosome [190]. Trypsin [1708]. Trypsinization [266]. tshawytscha [2258, 1335, 1141, 1699, 1844, 796, 76, 1910, 1721, 2494, 1073, 733, 544, 1961, 1900, 1674, 2251, 2347, 703, 1360, 1142, 1282, 1569, 2119, 1970, 1611]. Tubercles [1325]. Tubificid [82, 943, 2382]. Tumors [783, 1969]. Tumours [2342]. Tuna [2461, 2355, 1952, 1722, 1364]. Tundra [2575, 1677, 1658]. Tuning [2472]. Turbidity [2264, 1582]. Turbidostat [1342]. Turbine [1845]. turf [1379]. Turkey [1595]. Turnover [968, 2012, 2208, 1952, 1089, 916, 587]. Tursiops [1275, 927, 90, 928, 1304, 1627]. Twilight [473]. Two [2129, 1871, 1948, 2182, 1703, 872, 1093, 168, 1037, 1335, 633, 1260, 985, 1918, 2563, 2396, 1659, 2392, 288, 1732, 1281, 1672, 2264, 704, 1945, 1598, 1886, 2601, 180, 62, 503, 1774, 1708, 889, 2296, 1808, 1487, 1041, 2210, 571, 331, 637, 329, 1117, 1816, 1163, 2391, 2395, 280, 308, 2388, 1725, 1151, 1815, 1933, 2173, 2549, 2227, 1626, 2427]. Two-Coloured [1945]. Two-Sea-Winter [1163]. Type [2414, 2068, 339, 2431, 1645, 121, 1706, 1632, 651, 1569, 2119]. Types [2036, 1141, 2427]. Typologies [961]. tyrannus [1418].

U [1865]. U.S. [495, 498]. Ultrasonic [558]. Ultrasound [1409]. Ultrastructure [1953, 1381]. Umbra [1524]. Unavailable [2525]. Unbiased [1243]. Uncertain [1271]. Uncertainties [794]. Uncertainty [994, 380, 1039, 2143, 298, 416]. Unconventional [116]. Uncoupling [490, 2271]. undatum [2330]. Undersize [1776]. Underwater [1399, 2215, 591, 1156]. Underyearling [151, 26]. Undisturbed [1197, 1463]. Unexploited [336, 1719]. Ungava [1745, 840]. Unidentified [1110]. Unifying [2504]. Unionid [28, 815]. Unionidae [1567]. Unit [1538, 637, 1424, 280]. United [2129, 887, 506]. Units [713]. Univoltine [1154]. Unknowns [263]. Unproductive [2353]. Unregulated [1069]. Unstable [606]. Unsupported [1401]. Unusually [1606]. Update [12]. Upon [1168, 2172, 19, 603, 542, 794, 1101, 1435, 2189, 125]. Upper [2360, 211, 2263]. Upstream [2407, 2541]. Upsurge [722]. Uptake [89, 1508, 2554, 1226, 60, 1033, 557, 2409, 860, 2078, 1542, 1687, 1160, 63, 1939, 2378, 299, 1585, 1071, 622, 797, 777, 773, 50, 175, 429, 315, 1487, 1544, 2121, 1120, 1083, 410, 1234, 1193, 1103, 462, 862, 707, 1686, 406, 1543, 463, 1211]. Upwelling [2437, 2218, 728, 787, 786, 1565, 352, 2071, 2028, 1490]. Uranium [1865]. Urchin [980, 25, 466]. Urchins [819, 1501, 1244]. Urea [2378, 2134, 1803, 543]. urée [543]. uricotéliques [1810]. Urine [1222]. ursinus [2534, 2533, 774]. USA [2618, 2218, 497, 1826]. Use [2482, 489, 1128, 1357, 375, 1905, 847, 560, 986, 273, 395, 394, 469, 159, 1984, 1852, 1259, 2142, 1674, 852, 1506, 1489, 2548, 2128, 2536, 709, 1805, 1083, 718, 920, 1364, 184, 1503, 1400, 707, 308, 750, 767, 1477, 599, 1590, 335, 2415, 132]. Used [1409, 1709, 1191, 836, 1198, 312]. Useful [2308]. Usefulness [1492, 1733]. User [2040]. Uses [645]. Using [1308, 2162, 1654, 423, 918, 2161, 1699, 860, 1349, 760, 2077, 2052, 110, 893, 1063, 2194, 1857, 272, 2432, 1620, 614, 1607, 69, 1915, 1945, 1715, 776, 1940, 915, 57, 2260, 387, 1172, 35, 2143, 2209, 2605, 1578, 977, 184, 1802, 1916, 87, 149, 824, 2003, 1933, 1353, 1514, 751, 1864, 2411, 1426, 146, 2502, 2519, 2378, 2463, 752, 2526, 2477, 2581, 2612]. Usque [540]. Utility [1356, 2613]. Utilization [549, 1043, 1744, 1746, 544, 1645, 1693, 1963, 2315, 2564, 1432, 1553, 586]. Utilizing [823, 1868]. UV [896].

V [2069, 1014]. V-Notching [2069]. Vaccine [411]. Vadimus [540]. Vaired [304]. Validation [918, 1925, 1924]. Validity [1471, 2303, 1699]. Valley [1961]. Valuation [169]. Value [919, 2590, 2598]. Values [2329]. Vancomycin [134]. Vancouver [2491, 2511, 1929, 1930, 618, 1403]. Vanda [739]. Vangsvatnet [661, 712]. Vapour [2498]. var [1862, 2577, 902]. var. [1862]. Varden [1346, 2246, 1259, 2265]. Vardnes [2474]. variabilis [869, 1257, 1585]. Variabilité [1490, 1135]. Variability [2156, 2413, 2414, 2547, 1730, 1758, 147, 1229, 1380, 2530, 90, 440, 2472, 155, 210, 270, 182, 971, 1437, 1559, 1594, 2093, 2579, 1465, 1792, 31, 1285, 1324, 2039, 513, 2290, 1843, 1152, 1490, 1135]. Variable [398, 1981, 1369, 2330, 1173, 2483, 586, 548]. Variables [2126, 650, 2563, 2006, 560, 836, 2295, 1647, 496, 954]. Variance [440, 280, 554]. Variances [953, 2154]. Variants [159]. Variation [646, 2529, 869, 1257, 1286, 1851, 2535, 2169, 2247, 202, 901, 2055, 514, 1842, 508, 2146, 2148, 1950, 847, 2323, 2451, 142, 2494, 1073, 1729, 1247, 1752, 2273, 287, 1218, 1444, 979, 1656, 2094, 1190, 437, 1446, 98, 2361, 328, 329, 1008, 1326, 1885, 2101, 2035, 308, 1495, 1496, 2119, 2598, 2608, 1917, 669, 2411, 2013, 747]. Variations [1113, 1734, 2009, 654, 2235, 2236, 1123, 2354, 432, 1995]. variegatum [1249]. Various [1030, 762, 645, 1198, 736, 1754, 1588]. Vary [1542]. Varying [1021, 1055, 2264, 2297, 2313, 2518, 2025, 1678, 1140]. Vasa [1681]. Vegetation [1971, 339, 2312]. Velocity [2522, 348, 1485, 806, 2548, 1613, 1646]. VEN [274, 1619]. Vendace [1702]. Ventilation [2282]. Ventilatory [750, 595]. Verification [1550]. Verifying [1296]. Vermont [1746]. verreauxi [1773]. Versatile [439]. Version [1846]. versus [2409, 2436, 2325, 1221, 2317, 1633, 1177, 1069, 406]. Vertebrae [785, 1364, 1960]. Vertebral [330, 1162, 2241]. Vertebrate [1632]. Vertical [1242, 1811, 1920, 631, 932, 1702, 1620, 1108, 433, 23, 2620, 2181, 1392, 1473, 2085, 2224, 1767]. verticale [1767]. verticalis [2019]. Vertically [1976, 1718]. very [543]. vetuls [1823]. vetulus [1226, 2339, 786, 2048, 1481]. VI [2411]. via [278]. Vibrio [1184, 752, 430]. Vibriosis [126]. Vicinity [1435]. victoria [2532]. View [209, 2066, 2135, 1736, 1010]. Viewpoint [491]. viliosus [465]. villosus [94, 2572, 677, 1621, 1974, 314, 657, 621, 919, 1174, 995, 1209, 1157, 323]. Viral [1900, 107, 274, 1619]. virens [1682]. Virginia [640]. virginica [126, 2244, 874]. viridans [1862]. virilis [2468, 1382, 50]. Virtual [2102, 1961, 2614, 1800, 2154, 424]. Virulence [464, 1182]. Virulent [551]. Virus [1110, 2476, 115, 1009, 1031, 675, 1812, 1943, 1944, 411, 736, 553]. Visual [2192, 1346, 1477]. Visually [558]. Vital [904]. Vitamin [652, 1847, 160]. Vitelline [1679]. Vitellogenesis [873]. Vitellogenesss [2284]. Vitellogenic [1893]. Vitellogenin [1521]. vitreum [514, 2521, 139, 1988, 2, 2169, 514, 2552, 2521, 139, 1988, 1575, 1909, 2420, 2, 2608]. Vitro [2492, 274, 1668]. vitulina [1031]. vivo [2270]. Vocalizations [1399, 2296]. voie [2306]. Volatilization [428]. volucellus [868]. Volume [2403, 2238, 2190, 454, 2117, 1192]. volutator [614]. VPA [2514, 2614, 1800]. Vulnerability [454, 1872, 2314, 276, 2610, 1596]. Vulnerable [2560].

W [329]. Wabigoon [1010, 1011, 1014, 1012, 1013, 1678, 2621]. Wabigoon/English [2621]. Walker [2186]. Walking [324, 411]. Wall [1499]. Walleye [623, 2328, 2169, 514, 2552, 2059, 2521, 142, 139, 1988, 188, 1575, 2420, 2, 2416, 2608]. Walleyes [1909]. Walrus [334]. Warm [1050]. Warmwater [542, 1394, 135]. Washington [2618, 370, 814, 2218, 1605, 2339, 1675, 1823, 2048, 1142, 2361, 2349]. Waste [101]. Wastewater [1839]. Watch [520]. Water [2486, 2229, 2272, 620, 2253, 996, 2476, 1466, 11, 1340, 807, 1249, 1204, 542, 2097, 682, 1610, 2552, 628, 1372, 1129, 2485, 1768, 2357, 1229, 657, 1027, 618, 1222, 1052, 1464, 1607, 1447, 52, 69, 1580, 53, 2100, 75, 1322, 1283, 1688, 806, 84, 2087, 1926, 763, 1890, 1528, 1613, 858, 47, 331, 421, 1678, 2621, 635, 1054, 1892, 766, 2011, 735, 862, 1556, 966, 611, 1414, 2348, 1258, 1036, 1754, 406, 1543, 2350, 1215, 1314, 2545, 1908, 1864, 389, 543, 396, 135]. Waterborne [1372, 30, 1354, 462, 862, 2096, 1013]. Waterfoorne [688]. Waters [269, 623, 2108, 1134, 1976, 1033, 896, 511, 2409, 2430, 2509, 580, 1051, 2358, 2201, 507, 2506, 306, 277, 1112, 1566, 191, 852, 1081, 1823, 2048, 247, 1551, 1751, 564, 1956, 504, 1026, 83, 449, 2152]. Watershed [1985, 1197, 2491, 47, 1491]. Watersheds [2357, 749, 46]. Wave [2021]. Waves [2576]. Wawa [1018]. Weather [1556]. Weathervane [1837]. Webs [983, 1158]. Wedge [300]. Wedge-Wire [300]. Weight [646, 1538, 1033, 128, 1419, 1609, 1185, 2421, 2186, 552, 2117, 1282, 1008, 83, 702, 1864]. Weight-Specific [1008]. Weighted [2238]. Weights [1218]. Weir [1673]. Well [1164]. Well-Buffered [1164]. West [2354, 625, 2108, 2010, 1852, 713, 799, 1081, 2354]. West-Central [2010]. Western [1776, 1764, 661, 712, 1763, 283, 1074, 1367, 1719, 2358, 2494, 742, 1911, 758, 2011, 323, 1448]. Wet [709, 1282]. Wetland [1491]. Whale [2592, 2539, 1745, 1368]. Whales [283, 1074, 2020, 2589, 284, 1517, 2296, 1081, 2450, 1368]. Whatman [2202]. Whelk [2330]. Where [260, 482]. Which [393, 2555]. Whirling [132, 422, 1564]. White [1601, 2592, 775, 693, 2410, 2494, 1411, 1849, 203, 1312, 2250, 2536, 1736, 1745, 2243, 1936, 369, 629, 1746]. White-Fleshed [2494]. Whitefish [1205, 1436, 1096, 511, 1097, 33, 933, 512, 105, 1266, 604, 1602, 239, 68, 1262, 654, 2095, 1695, 256, 243]. Whiting [824]. Who [481]. Whole [1409, 2468, 110, 822, 2234, 1268, 52, 70, 59, 2331, 913, 1510, 1330, 2393]. Whole-Body [2234]. Whole-Lake [2468, 70, 59, 2331, 1510, 1330, 2393]. Whose [1133]. Wild [1680, 502, 2491, 2265, 1675, 1674, 520, 363, 1805, 2334, 2058, 1870, 1503, 352, 1327]. Willamette [333]. Wilson [1035]. Wind [1233, 2201, 314, 1605, 2235, 2236, 1869, 2017]. Wind-Aided [2201]. Wind-Exposed [2017]. Wind-Forced [2235, 2236, 1869]. Window [2437]. Windsor [178]. Windstorm [46]. Winter [1277, 1948, 1399, 2055, 1744, 2341, 2346, 865, 1174, 1097, 572, 691, 2434, 1907, 1946, 1675, 1450, 617, 1435, 2548, 2322, 1441, 1727, 2090, 570, 1163, 1425, 1535, 1967, 1995, 1657, 29, 817, 430, 1314, 1315]. Winterkill [929, 2622, 1524]. Wire [300]. Wisconsin [2594, 1946, 1624, 2292, 755, 1515, 602]. Witch [603, 393]. Within [2568, 1509, 719, 347, 2291, 1656, 1386, 2247, 2323, 2056, 1679, 1727, 1463]. Within-River [2291]. Without [147, 140, 1692]. Wolf [2022]. Wood [1219]. Woodland [2056, 455]. Woods [2390]. Woody [2333, 1704]. Work [1279, 1046]. Workers [1728]. Workshop [1756]. Worm [36, 1993]. Worms [82, 1120]. Wound [1275, 2020]. Wupaw [1098].

X [592, 727]. X-Cell [727]. X-ray [592]. xanthurus [942]. Xenobiotic [2105, 2499].

Year [128, 952, 1159, 1609, 2223, 2594, 1984, 1612, 1157, 2045, 2597, 29, 19, 369, 629]. Year-Class [128, 952, 1159, 2223, 1612, 1157, 19, 629, 369]. Yearling [446, 1947]. Yearlings [861, 567]. Years [2333, 2218, 73, 48]. Yellow [2376, 1249, 2068, 1449, 1506, 2317, 1077, 1884, 2464, 2292, 1962, 2, 1200, 2045, 2305, 2597, 570, 1392, 1231, 1696, 2479]. Yellowfïn [1722]. Yellowtail [1861]. Yersinia [641, 1109]. Yield [664, 398, 1721, 559, 1188, 1467, 1101, 1122, 2590, 891, 352, 1123, 769, 2270, 1669]. Yield-per-Recruitment [1122]. Yields [2127, 1467, 1163, 965]. Yolk [1850, 1379, 544, 1291]. York [656, 2010, 622, 2327, 2464, 2, 2390, 2395, 2388, 504]. Young [1882, 342, 2068, 2484, 1451, 2341, 1609, 1984, 22, 955, 1638, 1077, 1884, 2367, 331, 2045, 2597, 2366, 1162, 29]. Young-of-the-Year [1984, 2045, 2597, 29]. Young-of-Year [1609]. Yukon [318, 281, 1897, 1600].

Zealand [1773, 1822, 1193, 83, 587]. zenithicus [267]. Zinc [869, 1257, 1340, 1249, 1990, 1689, 775, 829, 1268, 1585, 2417, 1430, 922, 948, 2250, 290, 1414, 2096, 2348, 1956, 1718]. Zn [1103, 1899, 1136, 2167]. Zoarcidae [396]. Zonation [690]. Zonations [1364]. Zone [1277, 2056, 277, 638, 1094, 965]. Zones [1389]. Zoobenthic [1934, 2259]. Zoobenthos [950]. Zoogeography [2602]. Zooplankton [1871, 963, 26, 1175, 1557, 759, 1043, 120, 1419, 2027, 796, 347, 1202, 1519, 657, 556, 2421, 1939, 1068, 1620, 1690, 745, 1256, 975, 1683, 191, 852, 2214, 123, 182, 361, 1044, 349, 1840, 2464, 979, 1555, 435, 1144, 2064, 2197, 1465, 1643, 2314, 475, 1386, 93, 627, 1586, 2391, 1036, 1225, 2084, 832, 2298, 271, 275, 1599, 1582]. Zostera [744].

References

Pomeroy:1980:AFR

[1] W. M. Pomeroy and C. D. Levings. Association and feeding relationships between Eogammarus confervicolus (Amphipoda, Gammaridae) and benthic algae on sturgeon and Roberts Banks, Fraser River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):1–10, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

      //cdnsciencepub.com/doi/full/10.1139/f80-001.

Nielsen:1980:EWS

[2] Larry A. Nielsen. Effect of walleye (Stizostedion vitreum vitreum) predation on juvenile mortality and recruitment of yellow perch (Perca flavescens) in Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):11–19, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

      com/doi/full/10.1139/f80-002.

Banse:1980:TPN

[3] Karl Banse. Terebellidae (Polychaeta) from the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):20–40, January 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-003.

Foucher:1980:PNO

[4] R. P. Foucher and R. J. Beamish. Production of nonviable oocytes by Pacific hake (Merluccius productus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):41–48, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-004.

Markle:1980:NRR

[5] Douglas F. Markle, W. B. Scott, and A. C. Kohler. New and rare records of Canadian fishes and the influence of hydrography on resident and nonresident Scotian Shelf ichthyofauna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):49–65, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-005.

Cote:1980:SPH

[6] G. Côté, P. Lamoureux, J. Boulva, and G. Lacroix. Séparation des populations de hareng de l’Atlantique (Clupea harengus harengus) de l’estuaire du Saint-Laurent et de la péninsule gaspésienne. (French) [Separation of the Atlantic herring (Clupea harengus harengus) populations of the Saint Lawrence Estuary and the Gaspé Peninsula]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):66–71, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-006.

Charlton:1980:ODL

[7] M. N. Charlton. Oxygen depletion in Lake Erie: Has there been any change? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):72–80, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-007.

Ryman:1980:GCH

[8] Nils Ryman and Gunnar Ståhl. Genetic changes in hatchery stocks of brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (1):82–87, January 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-008.

Fong:1980:RGF

[9] W. Fong and K. H. Mann. Role of gut flora in the transfer of amino acids through a marine food chain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (1):88–96, January 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-009.

Majkowski:1980:SEF

[10] J. Majkowski, J. Pilarska, and R. Z. Klekowski. Simulations of energy flow through the amictic female rotifer Brachionus rubens Ehrbg. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):97–110, January 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-010.

Bergmann:1980:SRM

[11] Martin Bergmann and Robert Henry Peters. A simple reflectance method for the measurement of particulate pigment in lake water and its application to phosphorus–chlorophyll–Seston relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):111–114, January 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-011.

Russell:1980:ADT

[12] Howard J. Russell, Jr. Analysis of double-tagging experiments: an update. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):114–116, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-012.

Bouck:1980:CLA

[13] Gerald R. Bouck. Concentrations of leucine aminonaphthylamidase (LAN) and soluble protein in tissues of rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):116–120, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-013.

Arai:1980:IPC

[14] Mary Needler Arai and J. Roger Jacobs. Interspecific predation of common Strait of Georgia planktonic coelenterates: Laboratory evidence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):120–123, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-014.

Short:1980:LLP

[15] Robert A. Short and James V. Ward. Leaf litter processing in a regulated Rocky Mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (1):123–127, January 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-015.

Roff:1980:MRB

[16] Derek A. Roff. A motion for the retirement of the von Bertalanffy function. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):127–129, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-016.

Vannote:1980:RCC

[17] Robin L. Vannote, G. Wayne Minshall, Kenneth W. Cummins, James R. Sedell, and Colbert E. Cushing. The river continuum concept. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):130–137, January 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-017.

Anonymous:1980:BRCa

[18] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(1):138–140, January 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f80-018.

Ulanowicz:1980:IAS

[19] R. E. Ulanowicz and T. T. Polgar. Influences of anadromous spawning behavior and optimal environmental conditions upon striped bass (Morone saxatilis) year-class success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (2):143–154, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-019.

Johnson:1980:OTS

[20] R. A. Johnson. Oxygen transport in salmon spawning gravels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):155–162, February 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-020.

dApollonia:1980:OAC

[21] S. d’Apollonia and P. D. Anderson. Optimal assay conditions for serum and liver glutamate oxaloacetate transaminase, glutamate pyruvate transaminase, and sorbitol dehydrogenase from the rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):163–169, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-021.

Hodson:1980:EDA

[22] P. V. Hodson, J. W. Hilton, B. R. Blunt, and S. J. Slinger. Effects of dietary ascorbic acid on chronic lead toxicity to young rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2): 170–176, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-022.

Janssen:1980:FEV

[23] John Janssen and Stephen B. Brandt. Feeding ecology and vertical migration of adult alewives (Alosa pseudoharengus) in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):177–184, February 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-023.

Mathias:1980:FCO

[24] Jack A. Mathias and Jan Barica. Factors controlling oxygen depletion in ice-covered lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):

185–194, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-024.

Takagi:1980:DFA

[25] T. Takagi, C. A. Eaton, and R. G. Ackman. Distribution of fatty acids in lipids of the common Atlantic sea urchin Strongylocentrotus droebachiensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):195–202, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-025.

Biette:1980:GUS

[26] R. M. Biette and G. H. Geen. Growth of underyearling sockeye salmon (Oncorhynchus nerka) under constant and cyclic temperatures in relation to live zooplankton ration size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):203–210, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-026.

Kaiser:1980:CPM

[27] Klaus L. E. Kaiser. Correlation and prediction of metal toxicity to aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):211–218, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-027.

Green:1980:RUC

[28] Roger H. Green. Role of a unionid clam population in the calcium budget of a small Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (2):219–224, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-028.

Toneys:1980:MHO

[29] Michael L. Toneys and Daniel W. Coble. Mortality, hematocrit, osmolality, electrolyte regulation, and fat depletion of young-of-the-year freshwater fishes under simulated winter conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):225–232, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-029.

Hodson:1980:ERT

[30] P. V. Hodson, D. J. Spry, and B. R. Blunt. Effects on rainbow trout (Salmo gairdneri) of a chronic exposure to waterborne selenium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):233–240, February 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-030.

Sainsbury:1980:EIV

[31] K. J. Sainsbury. Effect of individual variability on the von Bertalanffy growth equation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):241–247, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-031.

Bothwell:1980:IST

[32] Max L. Bothwell and John G. Stockner. Influence of secondarily treated kraft mill effluent on the accumulation rate of attached algae in experimental continuous-flow troughs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):248–254, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-032.

Healey:1980:GRE

[33] M. C. Healey. Growth and recruitment in experimentally exploited lake whitefish (Coregonus clupeaformis) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):255–267, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-033.

Deriso:1980:HSP

[34] R. B. Deriso. Harvesting strategies and parameter estimation for an age-structured model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2): 268–282, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-034.

Mercer:1980:SDO

[35] M. C. Mercer, R. K. Misra, and G. V. Hurley. Sex determination of the ommastrephid squid Illex illecebrosus using beak morphometries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):283–286, February 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-035.

Pesch:1980:CCM

[36] Gerald G. Pesch and Carol E. Pesch. Chromosome complement of the marine worm Neanthes arenaceodentata (Polychaeta: Annelida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):286–288, February 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-036.

McGaw:1980:CIO

[37] Richard L. McGaw. Confidence intervals for optimal effort estimates from the Schaefer Production Model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):288–289, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-037.

Reckhow:1980:TEP

[38] K. H. Reckhow. Techniques for exploring and presenting data applied to lake phosphorus concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):290–294, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-038.

Anonymous:1980:LLa

[39] Anonymous. Letters/lettres. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (2):295–297, February 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-039.

Anonymous:1980:BRCb

[40] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):298–299, February 1980. CODEN CJFSDX.

ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f80-040.

Anonymous:1980:CRPa

[41] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(2):301, February 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-041.

Watson:1980:FSI

[42] J. Watson. Foreword to the special issue on the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):311, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-042.

Watson:1980:APA

[43] J. Watson. Avant-propos au numéro spécial sur la région des lacs expérimentaux. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):312, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-043.

Schindler:1980:EEL

[44] D. W. Schindler. Evolution of the Experimental Lakes Project. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):313–319, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-044.

Schindler:1980:HIN

[45] D. W. Schindler, T. Ruszczynski, and E. J. Fee. Hypolimnion injection of nutrient effluents as a method for reducing eutrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):320–327, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-045.

Schindler:1980:EWF

[46] D. W. Schindler, R. W. Newbury, K. G. Beaty, J. Prokopowich, T. Ruszczynski, and J. A. Dalton. Effects of a windstorm and forest fire on chemical losses from forested watersheds and on the quality of receiving streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):328–334, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-046.

Newbury:1980:WRE

[47] R. W. Newbury and K. G. Beaty. Water renewal efficiency of watershed and lake combinations in the ELA region of the Precambrian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):335–341, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-047.

Schindler:1980:EAL

[48] D. W. Schindler, R. Wagemann, R. B. Cook, T. Ruszczynski, and J. Prokopowich. Experimental acidification of Lake 223, Experimental Lakes Area: Background data and the first three years of acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):342–354, March 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-048.

Muller:1980:EAA

[49] Peter Müller. Effects of artificial acidification on the growth of periphyton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):355–363, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-049.

Malley:1980:DSC

[50] D. F. Malley. Decreased survival and calcium uptake by the crayfish Orconectes virilis in low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):364–372, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-050.

Schindler:1980:EAM

[51] D. W. Schindler, R. H. Hesslein, R. Wagemann, and W. S. Broecker. Effects of acidification on mobilization of heavy metals and radionuclides from the sediments of a freshwater lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):373–377, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-051.

Hesslein:1980:FMR

[52] R. H. Hesslein, W. S. Broecker, and D. W. Schindler. Fates of metal radiotracers added to a whole lake: Sediment–water interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):378–386, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-052.

Jackson:1980:EST

[53] Togwell A. Jackson, George Kipphut, Raymond H. Hesslein, and David W. Schindler. Experimental study of trace metal chemistry in soft-water lakes at different pH levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):387–402, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-053.

Marshall:1980:DCS

[54] J. S. Marshall and D. L. Mellinger. Dynamics of cadmium-stressed plankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):403–414, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-054.

Brunskill:1980:ESE

[55] G. J. Brunskill, B. W. Graham, and J. W. M. Rudd. Experimental studies on the effect of arsenic on microbial degradation of organic matter and algal growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):415–423, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-055.

deNoyelles:1980:CCN

[56] F. deNoyelles, Jr., R. Knoechel, D. Reinke, D. Treanor, and C. Altenhofen. Continuous culturing of natural phytoplankton communities in the Experimental Lakes Area: Effects of enclosure, in situ incubation, light, phosphorus, and cadmium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):424–433, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-056.

Knoechel:1980:ARH

[57] R. Knoechel and F. deNoyelles, Jr. Analysis of the response of hypolimnetic phytoplankton in continuous culture to increased light or phosphorus using track autoradiography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):434–441, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-057.

Healey:1980:PIN

[58] F. P. Healey and L. L. Hendzel. Physiological indicators of nutrient deficiency in lake phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):442–453, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-058.

Hesslein:1980:WLR

[59] R. H. Hesslein, W. S. Broecker, P. D. Quay, and D. W. Schindler. Wholelake radiocarbon experiment in an oligotrophic lake at the Experimental Lakes Area, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):454–463, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-059.

Bower:1980:GEP

[60] Peter Bower and Daniel McCorkle. Gas exchange, photosynthetic uptake, and carbon budget for a radiocarbon addition to a small enclosure in a stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):464–471, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-060.

Stainton:1980:EMB

[61] M. P. Stainton. Errors in molybdenum blue methods for determining orthophosphate in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):472–478, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-061.

Levine:1980:RAO

[62] S. N. Levine and D. W. Schindler. Radiochemical analysis of orthophosphate concentrations and seasonal changes in the flux of orthophosphate to seston in two Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):479–487, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-062.

Graham:1980:CNU

[63] B. M. Graham, R. D. Hamilton, and N. E. R. Campbell. Comparison of the nitrogen-15 uptake and acetylene reduction methods for estimating the rates of nitrogen fixation by freshwater blue-green algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):488–493, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-063.

Flett:1980:NFC

[64] Robert J. Flett, David W. Schindler, Robert D. Hamilton, and Norman E. R. Campbell. Nitrogen fixation in Canadian Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):494–505, March 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-064.

Chan:1980:DLD

[65] Y. K. Chan and N. E. R. Campbell. Denitrification in Lake 227 during summer stratification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):506–512, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-065.

Fee:1980:IFE

[66] Everett J. Fee. Important factors for estimating annual phytoplankton production in the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):513–522, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-066.

Davies:1980:RBD

[67] I. J. Davies. Relationships between dipteran emergence and phytoplankton production in the Experimental Lakes Area, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):523–533, March 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-067.

Mills:1980:CFR

[68] K. H. Mills and R. J. Beamish. Comparison of fin-ray and scale age determinations for lake whitefish (Coregonus clupeaformis) and their implications for estimates of growth and annual survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):534–544, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-068.

Hesslein:1980:SMP

[69] R. H. Hesslein. In situ measurements of pore water diffusion coefficients using tritiated water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (3):545–551, March 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-069.

Hesslein:1980:WLM

[70] R. H. Hesslein. Whole-lake model for the distribution of sediment-derived chemical species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(3):552–558, March 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-070.

Peterman:1980:DNI

[71] Randall M. Peterman. Dynamics of native Indian food fisheries on salmon in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):561–566, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-071.

Borgmann:1980:EMB

[72] U. Borgmann, R. Cove, and C. Loveridge. Effect of metals on the biomass production kinetics of freshwater copepods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):567–575, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-072.

Reisenbichler:1980:ENM

[73] R. R. Reisenbichler and N. A. Hartmann, Jr. Effect of number of marked fish and years of repetition on precision in studies of contribution to a fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):576–582, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-073.

Bedard:1980:AIN

[74] J. Bédard, J. C. Therriault, and J. Bérubé. Assessment of the importance of nutrient recycling by seabirds in the St. Lawrence Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):583–588, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-074.

Jones:1980:NDC

[75] E. P. Jones and A. R. Coote. Nutrient distributions in the Canadian Archipelago: Indicators of summer water mass and flow characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):589–599, April 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-075.

Ewing:1980:ISG

[76] R. D. Ewing, H. J. Pribble, S. L. Johnson, C. A. Fustish, J. Diamond, and J. A. Lichatowich. Influence of size, growth rate, and photoperiod on cyclic changes in gill (Na + K)–ATPase activity in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):600–605, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-076.

Colton:1980:BNR

[77] J. B. Colton, Jr., J. R. Green, R. R. Byron, and J. L. Frisella. Bongo net retention rates as effected by towing speed and mesh size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):606–623, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-077.

Gurtz:1980:SDS

[78] Martin E. Gurtz, Jackson R. Webster, and J. Bruce Wallace. Seston dynamics in Southern Appalachian streams: Effects of clear-cutting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):624–631, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-078.

Colwell:1980:DIN

[79] A. E. Colwell and C. H. Schaefer. Diets of Ictalurus nebulosus and Pomoxis nigromaculatus altered by diflubenzuron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):632–639, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-079.

Reid:1980:DDA

[80] R. P. Reid, C. H. Pharo, and W. C. Barnes. Direct determination of apatite in lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):640–646, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-080.

Winner:1980:ICS

[81] R. W. Winner, M. W. Boesel, and M. P. Farrell. Insect community structure as an index of heavy-metal pollution in lotic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):647–655, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-081.

Chatarpaul:1980:ETW

[82] Lakeram Chatarpaul, John B. Robinson, and Narinder K. Kaushik. Effects of tubificid worms on denitrification and nitrification in stream sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):656–663, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-082.

White:1980:DBA

[83] E. White and G. Payne. Distribution and biological availability of reactive high molecular weight phosphorus in natural waters in New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):664–669, April 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-083.

Luoma:1980:SFA

[84] Maire E. Luoma and John H. Gee. Seasonal factors affecting buoyancy attained in still water and current by fathead minnows, Pimephales promelas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):670–678, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-084.

Nakashima:1980:NSR

[85] Brian S. Nakashima and William C. Leggett. Natural sources and requirements of phosphorus for fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):679–686, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-085.

Ebert:1980:EPF

[86] Thomas A. Ebert. Estimating parameters in a flexible growth equation, the Richards fonction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):687–692, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-086.

Sameoto:1980:QME

[87] D. D. Sameoto. Quantitative measurements of euphausiids using a 120kHz sounder and their in situ orientation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):693–702, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-087.

Pelletier:1980:MDP

[88] E. Pelletier and J. Lebel. La mesure du pH en milieu estuarien sur l’échelle NBS. (French) [The measurement of pH in an estuarine environment on the NBS scale]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):703–706, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-088.

Axler:1980:SNU

[89] R. P. Axler, R. M. Gersberg, and C. R. Goldman. Stimulation of nitrate uptake and photosynthesis by molybdenum in Castle Lake, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):707–712, April 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-089.

Hui:1980:VDD

[90] Clifford A. Hui. Variability of dentin deposits in Tursiops truncatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):712–716, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-090.

Cammen:1980:EBP

[91] Leon M. Cammen. Estimation of biological power functions from group measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):716–719, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
091.

Bullock:1980:DCS

[92] G. L. Bullock, B. R. Griffin, and H. M. Stuckey. Detection of Corynebacterium salmoninus by direct fluorescent antibody test. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):719–721, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-092.

Sameoto:1980:BND

[93] D. D. Sameoto, L. O. Jaroszynski, and W. B. Fraser. BIONESS, a new design in multiple net zooplankton samplers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):722–724, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-093.

Carscadden:1980:MAM

[94] J. E. Carscadden and R. K. Misra. Multivariate analysis of meristic characters of capelin (Mallotus villosus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):725–729, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-094.

Burnison:1980:MDS

[95] B. K. Burnison. Modified dimethyl sulfoxide (DMSO) extraction for chlorophyll analysis of phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):729–733, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-095.

Mathur:1980:SPS

[96] Dilip Mathur and Carl A. Silver. Statistical problems in studies of temperature preference of fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):733–737, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-096.

Echeverria:1980:ALB

[97] Tina Echeverria. Accumulation of ^{14}C labeled benzene and related compounds in the rotifer Brachionus plicatilis from seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):738–741, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-097.

Rau:1980:CCV

[98] Greg H. Rau. Carbon-13/carbon-12 variation in subalpine lake aquatic insects: Food source implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):742–746, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-098.

Michel:1980:SME

[99] C. Michel. A standardized model of experimental furunculosis in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (4):746–750, April 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-099.

Anonymous:1980:BRCc

[100] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(4):751, April 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-100.

Hedtke:1980:EWO

[101] Steven F. Hedtke and Frank A. Puglisi. Effects of waste oil on the survival and reproduction of the American flagfish, Jordanella floridae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):757–764, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-101.

Specker:1980:SRT

[102] J. L. Specker and C. B. Schreck. Stress responses to transportation and fitness for marine survival in Coho salmon (Oncorhynchus kisutch) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):765–769, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-102.

Peterson:1980:IAS

[103] R. H. Peterson, P. G. Daye, and J. L. Metcalfe. Inhibition of Atlantic salmon (Salmo salar) hatching at low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):770–774, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-103.

Gartner-Kepkay:1980:GDE

[104] K. E. Gartner-Kepkay, L. M. Dickie, K. R. Freeman, and E. Zouros. Genetic differences and environments of mussel populations in the Maritime provinces. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):775–782, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-104.

Imhof:1980:PSS

[105] Mark Imhof, Robb Leary, and Henry E. Booke. Population or stock structure of lake whitefish, Coregonus clupeaformis, in Northern Lake Michigan as assessed by isozyme electrophoresis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):783–793, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-105.

Bower:1980:SHM

[106] Carol E. Bower and Thomas Holm-Hansen. A salicylate–hypochlorite method for determining ammonia in seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):794–798, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-106.

MacMillan:1980:VEN

[107] John R. MacMillan, Dan Mulcahy, and Marsha Landolt. Viral erythrocytic necrosis: Some physiological consequences of infection in chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (5):799–804, May 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-107.

Barton:1980:PCL

[108] Bruce A. Barton, Richard E. Peter, and Christine R. Paulencu. Plasma cortisol levels of fingerling rainbow trout (Salmo gairdneri) at rest, and subjected to handling, confinement, transport, and stocking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):805–811, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-108.

MacLean:1980:SHS

[109] Sharon A. MacLean. Study of Haematractidium scombri in Atlantic mackerel, Scomber scombrus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (5):812–816, May 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-109.

Evans:1980:MWL

[110] R. Douglas Evans and Frank H. Rigler. Measurement of whole lake sediment accumulation and phosphorus retention using lead-210 dating. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):817–822, May 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-110.

Scavia:1980:STP

[111] Donald Scavia and John R. Bennett. Spring transition period in Lake Ontario — a numerical study of the causes of the large biological and chemical gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):823–833, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
111.

Naiman:1980:RBM

[112] Robert J. Naiman and James R. Sedell. Relationships between metabolic parameters and stream order in Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):834–847, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-112.

Rudd:1980:DSM

[113] John W. M. Rudd, Michael A. Turner, Bruce E. Townsend, Alison Swick, and Akira Furutani. Dynamics of selenium in mercury-contaminated experimental freshwater ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):848–857, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-113.

Anderson:1980:NCE

[114] Lee G. Anderson. Necessary components of economic surplus in fisheries economics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):858–870, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-114.

Bowser:1980:GRN

[115] P. R. Bowser and J. A. Plumb. Growth rates of a new cell line from channel catfish ovary and channel catfish virus replication at different temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):871–873, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
115.

Favro:1980:EUS

[116] L. D. Favro, P. K. Kuo, and J. F. McDonald. Effects of unconventional size limits on the growth rate of trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):873–876, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-116.

Harris:1980:TSS

[117] Graham P. Harris. Temporal and spatial scales in phytoplankton ecology. mechanisms, methods, models, and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):877–900, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-117.

Anonymous:1980:LCLa

[118] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):901–902, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-118.

Anonymous:1980:BRCd

[119] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(5):905–906, May 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-119.

Cooper:1980:OSM

[120] Scott D. Cooper and Charles R. Goldman. Opossum shrimp (Mysis relicta) predation on zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):909–919, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-120.

Misra:1980:SCS

[121] R. K. Misra. Statistical comparisons of several growth curves of the von Bertalanffy type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):920–926, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
121.

Rosenberg:1980:SEC

[122] David M. Rosenberg, Allen P. Wiens, and Bohdan Bilyj. Sampling emerging Chironomidae (Diptera) with submerged funnel traps in a new Northern Canadian reservoir, Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):927–936, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-122.

Mathur:1980:ATD

[123] Dilip Mathur, Timothy W. Robbins, and Edmund J. Purdy, Jr. Assessment of thermal discharges on zooplankton in Conowingo Pond, Pennsylvania. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):937–944, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-123.

Ennis:1980:SMR

[124] G. P. Ennis. Size–maturity relationships and related observations in Newfoundland populations of the lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):945–956, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-124.

Peckarsky:1980:IDU

[125] Barbara L. Peckarsky. Influence of detritus upon colonization of stream invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):957–963, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
125.

Elston:1980:PEV

[126] Ralph Elston and Louis Leibovitz. Pathogenesis of experimental vibriosis in larval American oysters, Crassostrea virginica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):964–978, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-126.

Fradette:1980:EBE

[127] Pierre Fradette and Edwin Bourget. Ecology of benthic epifauna of the estuary and Gulf of St. Lawrence: Factors influencing their distribution and abundance on buoys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (6):979–999, June 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-127.

Cohen:1980:MGT

[128] Marc-David Cohen and George S. Fishman. Modeling growth–time and weight–length relationships in a single year-class fishery with examples for North Carolina pink and brown shrimp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1000–1011, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-128.

Ware:1980:BSR

[129] D. M. Ware. Bioenergetics of stock and recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1012–1024, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-129.

Mortonson:1980:ODN

[130] J. A. Mortonson and A. S. Brooks. Occurrence of a deep nitrite maximum in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (6):1025–1027, June 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-130.

Roy:1980:FSD

[131] Suzanne Roy and Louis Legendre. Field studies of DCMU-enhanced fluorescence as an index of in situ phytoplankton photosynthetic activity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1028–1031, June 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-131.

Kozel:1980:IMC

[132] Thomas R. Kozel, Marley Lott, and Robert Taylor. Isolation of Myxosoma cerebralis (whirling disease) spores from infected fish by use of a physical separation technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1032–1035, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-132.

Daye:1980:AAE

[133] P. G. Daye. Attempts to acclimate embryos and alevins of Atlantic salmon, Salmo salar, and rainbow trout, S. gairdneri, to low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1035–1038, June 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-133.

Herbst:1980:DVL

[134] Doris V. Herbst. Determination of vancomycin in lobster meat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1038–1040, June 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-134.

Pyle:1980:NAD

[135] Stephen W. Pyle and Emmett B. Shotts, Jr. A new approach for differentiating flexibacteria isolated from cold-water and warmwater fish.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1040–1042, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-135.

U:1980:LCL

[136] F. H. U. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1043–1055, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-136.

Anonymous:1980:BRCe

[137] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(6):1056–1058, June 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f80-137.

Werner:1980:GSL

[138] Robert G. Werner and J. H. S. Blaxter. Growth and survival of larval herring (Clupea harengus) in relation to prey density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1063–1069, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-138.

Kelly:1980:FCC

[139] R. K. Kelly, H. R. Miller, O. Nielsen, and J. W. Clayton. Fish cell culture: Characteristics of a continuous fibroblastic cell line from walleye (Stizostedion vitreum vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1070–1075, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-139.

Newbold:1980:ELM

[140] J. D. Newbold, D. C. Erman, and K. B. Roby. Effects of logging on macroinvertebrates in streams with and without buffer strips. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1076–1085, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-140.

Trump:1980:OSS

[141] Clifford L. Trump and William C. Leggett. Optimum swimming speeds in fish: The problem of currents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1086–1092, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-141.

Grant:1980:BGV

[142] W. S. Grant and F. M. Utter. Biochemical genetic variation in walleye pollock, Theragra chalcogramma: Population structure in the Southeastern Bering Sea and the Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1093–1100, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-142.

Prinsenberg:1980:MMC

[143] S. J. Prinsenberg. Man-made changes in the freshwater input rates of Hudson and James Bays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1101–1110, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-143.

Clark:1980:TPM

[144] Colin W. Clark. Towards a predictive model for the economic regulation of commercial fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1111–1129, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-144.

Hulberg:1980:CMM

[145] Larry W. Hulberg and John S. Oliver. Caging manipulations in marine soft-bottom communities: Importance of animal interactions or sedimentary habitat modifications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1130–1139, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-145.

Aalto:1980:SME

[146] Sergei Aalto and G. E. (Buck) Newsome. Some methods of estimating the parameters of the Leslie matrix using incomplete population data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1140–1148, July 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-146.

Fairbairn:1980:TGM

[147] Daphne J. Fairbairn and Derek A. Roff. Testing genetic models of isozyme variability without breeding data: Can we depend on the χ^{2} ? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1149–1159, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-147.

Quackenbush:1980:ASD

[148] Dale W. Quackenbush, Max L. Bothwell, and Peter A. Bisson. Aluminum shadowing–direct projection photomicroscopy of diatom frustules. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1160–1166, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-148.

Samoiloff:1980:RSL

[149] M. R. Samoiloff, S. Schulz, Y. Jordan, K. Denich, and E. Arnott. A rapid simple long-term toxicity assay for aquatic contaminants using the nematode Panagrellus redivivus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1167–1174, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-149.

Barica:1980:EMA

[150] J. Barica, H. Kling, and J. Gibson. Experimental manipulation of algal bloom composition by nitrogen addition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1175–1183, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-150.

Biette:1980:FLF

[151] R. M. Biette and G. H. Geen. The fate of ^{14}C labeled food ingested by underyearling sockeye salmon, Oncorhynchus nerka, in relation to temperature and ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1184–1192, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-151.

Turpin:1980:CSM

[152] David H. Turpin and Paul J. Harrison. Cell size manipulation in natural marine, planktonic, diatom communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1193–1195, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-152.

Johnson:1980:GLE

[153] M. G. Johnson. Great Lakes environmental protection policies from a fisheries perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (7):1196–1204, July 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-153.

Anonymous:1980:BRCf

[154] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(7):1205, July 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-154.

Kranck:1980:VPM

[155] Kate Kranck. Variability of particulate matter in a small coastal inlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1209–1215, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-155.

Burton:1980:AOT

[156] Dennis T. Burton and Alan G. Heath. Ambient oxygen tension and transition to anaerobic metabolism in three species of freshwater fish.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1216–1224, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-156.

Pesch:1980:NAP

[157] Gerald G. Pesch and Carol E. Pesch. Neanthes arenaceodentata (Polychaeta: Annelida), a proposed cytogenetic model for marine genetic toxicology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1225–1228, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
157.

Roff:1980:EGM

[158] D. A. Roff and D. J. Fairbairn. An evaluation of Gulland’s method for fitting the Schaefer Model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (8):1229–1235, August 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-158.

Grant:1980:UBG

[159] W. S. Grant, G. B. Milner, P. Krasnowski, and F. M. Utter. Use of biochemical genetic variants for identification of sockeye salmon (Oncorhynchus nerka) stocks in Cook Inlet, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1236–1247, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-159.

Hung:1980:MOF

[160] S. S. O. Hung, C. Y. Cho, and S. J. Slinger. Measurement of oxidation in fish oil and its effect on vitamin E nutrition of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1248–1253, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-160.

Johannsson:1980:EDE

[161] Ora E. Johannsson. Energy dynamics of the eutrophic chironomid Chironomus plumosus f. semireductus from the Bay of Quinte, Lake Ontario.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1254–1265, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-161.

Stauffer:1980:BEC

[162] Robert E. Stauffer. Bias evaluation of the cation-exchange method for sulfate and chloride. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8): 1266–1275, August 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-162.

Mohn:1980:BEP

[163] Robert K. Mohn. Bias and error propagation in logistic production models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1276–1283, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-163.

Uhler:1980:LSR

[164] Russell S. Uhler. Least squares regression estimates of the Schaefer Production Model: Some Monte Carlo simulation results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1284–1294, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-164.

Borgmann:1980:IEM

[165] Uwe Borgmann. Interactive effects of metals in mixtures on biomass production kinetics of freshwater copepods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1295–1302, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f80-165.

Campbell:1980:MIM

[166] P. Campbell and T. Torgersen. Maintenance of iron meromixis by iron redeposition in a rapidly flushed monimolimnion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1303–1313, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-166.

Kohler:1980:PLL

[167] Christopher C. Kohler and John J. Ney. Piscivority in a landlocked alewife (Alosa pseudoharengus) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1314–1317, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-167.

Bitton:1980:OFH

[168] Gabriel Bitton, John R. Beaver, and Samuel R. Farrah. Organic flocculation and hydroextraction: Two new methods for phytoplankton concentration from lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8): 1317–1320, August 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-168.

Cauvin:1980:VRF

[169] D. Cauvin. The valuation of recreational fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1321–1327, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-169.

Anonymous:1980:LCLb

[170] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1328–1332, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-170.

Anonymous:1980:CRPb

[171] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(8):1333, August 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-
171.

Schnute:1980:NAL

[172] Jon Schnute and David Fournier. A new approach to length–frequency analysis: Growth structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (9):1337–1351, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-172.

Bienfang:1980:HDA

[173] P. K. Bienfang. Herbivore diet affects fecal pellet settling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1352–1357, September 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-173.

Culp:1980:RAP

[174] Joseph M. Culp and Ronald W. Davies. Reciprocal averaging and polar ordination as techniques for analyzing lotic macroinvertebrate communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1358–1364, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-174.

Murphy:1980:ANU

[175] T. P. Murphy. Ammonia and nitrate uptake in the Lower Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1365–1372, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-175.

Mathews:1980:FAP

[176] S. B. Mathews and F. W. Olson. Factors affecting Puget Sound Coho salmon (Oncorhynchus kisutch) runs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1373–1378, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-176.

Bailey:1980:IPS

[177] J. K. Bailey, R. L. Saunders, and M. I. Buzeta. Influence of parental smolt age and sea age on growth and smolting of hatchery-reared Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9): 1379–1386, September 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-177.

Turk:1980:SBC

[178] Ted R. Turk, Michael J. Risk, Roy W. M. Hirtle, and Ross K. Yeo. Sedimentological and biological changes in the Windsor Mudflat, an area of induced siltation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1387– 1397, September 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-178.

Freeland:1980:CED

[179] Howard J. Freeland and David M. Farmer. Circulation and energetics of a deep, strongly stratified inlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1398–1410, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-179.

Laughlin:1980:RPT

[180] Dennis R. Laughlin and Earl E. Werner. Resource partitioning in two coexisting sunfish: Pumpkinseed (Lepomis gibbosus) and northern longear sunfish (Lepomis megalotis peltastes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1411–1420, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-180.

Paulson:1980:MAE

[181] Larry J. Paulson. Models of ammonia excretion for brook trout (Salvelinus fontinalis) and rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1421–1425, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-181.

Mayzaud:1980:SSV

[182] P. Mayzaud. Some sources of variability in determination of digestive enzyme activity in zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1426–1432, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-182.

Papst:1980:RBT

[183] M. H. Papst, J. A. Mathias, and J. Barica. Relationship between thermal stability and summer oxygen depletion in a prairie pothole lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1433–1438, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-183.

Reckhow:1980:PUM

[184] K. H. Reckhow and J. T. Simpson. A procedure using modeling and error analysis for the prediction of lake phosphorus concentration from land use information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1439– 1448, September 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-184.

Yurkowski:1980:GMI

[185] M. Yurkowski and Jo-Anne L. Tabachek. Geosmin and 2-methylisoborneol implicated as a cause of muddy odor and flavor in commercial fish from Cedar Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9): 1449–1450, September 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-185.

Klaverkamp:1980:BAI

[186] J. F. Klaverkamp and B. R. Hobden. Brain acetylcholinesterase inhibition and hepatic activation of acephate and fenitrothion in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1450– 1453, September 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-186.

Diana:1980:DAP

[187] James S. Diana. Diel activity pattern and swimming speeds of northern pike (Esox lucius) in Lac Ste. Anne, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1454–1458, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-187.

Kramer:1980:SCS

[188] D. E. Kramer and D. M. A. Nordin. Storage characteristics of small and large walleye pollock (Theragra chalcogramma) held in ice. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1458–1462, September 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-188.

Anonymous:1980:BRCg

[189] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(9):1463–1464, September 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-189.

Khan:1980:PTA

[190] R. A. Khan, J. Murphy, and D. Taylor. Prevalence of a trypanosome in Atlantic cod (Gadus morhua) especially in relation to stocks in the Newfoundland area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10): 1467–1475, October 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-190.

Mackas:1980:SDZ

[191] David L. Mackas, Gregory C. Louttit, and Merelene J. Austin. Spatial distribution of zooplankton and phytoplankton in British Columbian coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1476– 1487, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f80-191.

Somerton:1980:CTE

[192] D. A. Somerton. A computer technique for estimating the size of sexual maturity in crabs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10): 1488–1494, October 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-192.

Wilson:1980:DGR

[193] K. H. Wilson and P. A. Larkin. Daily growth rings in the otoliths of juvenile sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1495–1498, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-193.

Wakabayashi:1980:CFB

[194] H. Wakabayashi, S. Egusa, and J. L. Fryer. Characteristics of filamentous bacteria isolated from a gill disease of salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1499–1504, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-194.

Schneider:1980:BAF

[195] Robert Schneider and Bruce L. Nicholson. Bacteria associated with fin rot disease in hatchery-reared Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1505–1513, October 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-195.

Adams:1980:GBC

[196] J. N. Adams and R. L. Beschta. Gravel bed composition in Oregon coastal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1514– 1521, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f80-196.

Gunn:1980:ESR

[197] J. M. Gunn and W. Keller. Enhancement of the survival of rainbow trout (Salmo gairdneri) eggs and fry in an acid lake through incubation in limestone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1522–1530, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f80-

Charlton:1980:HOC

[198] M. N. Charlton. Hypolimnion oxygen consumption in lakes: Discussion of productivity and morphometry effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1531–1539, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-198.

Nakashima:1980:RFR

[199] Brian S. Nakashima and William C. Leggett. The role of fishes in the regulation of phosphorus availability in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1540–1549, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-199.

Mackie:1980:SSC

[200] G. O. Mackie. Slow swimming and cyclical “Fishing” behavior in Aglantha digitale (Hydromedusae: Trachylina). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1550–1556, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-200.

Brandt:1980:THP

[201] Stephen B. Brandt, John J. Magnuson, and Larry B. Crowder. Thermal habitat partitioning by fishes in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1557–1564, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-201.

Bowman:1980:DVF

[202] Ray E. Bowman and Edgar W. Bowman. Diurnal variation in the feeding intensity and catchability of silver hake (Merluccius bilinearis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1565–1572, October 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-202.

McFarlane:1980:ECC

[203] G. A. McFarlane and W. G. Franzin. An examination of Cd, Cu, and Hg concentrations in livers of northern pike, Esox lucius, and white sucker, Catostomus commersoni, from five lakes near a base metal smelter at Flin Flon, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10): 1573–1578, October 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-203.

Watras:1980:SRE

[204] C. J. Watras. Subitaneous and resting eggs of copepods: Relative rates of clutch production by Diaptomus leptopus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(10):1579–1581, October 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-204.

Smith:1980:IPS

[205] Bernard R. Smith. Introduction to the Proceedings of the 1979 Sea Lamprey International Symposium (SLIS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1585–1587, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-205.

Fetterolf:1980:WGL

[206] Carlos M. Fetterolf, Jr. Why a Great Lakes Fishery Commission and why a Sea Lamprey International Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1588–1593, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-206.

Potter:1980:PPR

[207] I. C. Potter. The petromyzoniformes with particular reference to paired species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1595–1615, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-207.

Vladykov:1980:DKM

[208] Vadim D. Vladykov and Edward Kott. Description and key to metamorphosed specimens and ammocoetes of Petromyzonidae found in the Great Lakes region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 1616–1625, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-208.

Bailey:1980:CCN

[209] Reeve M. Bailey. Comments on the classification and nomenclature of lampreys — an alternative view. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1626–1629, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-209.

Krueger:1980:DVI

[210] Charles C. Krueger. Detection of variability at isozyme loci in sea lamprey, Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (11):1630–1634, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-210.

Manion:1980:SBF

[211] Patrick J. Manion and Lee H. Hanson. Spawning behavior and fecundity of lampreys from the upper three Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1635–1640, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-211.

Potter:1980:ELM

[212] I. C. Potter. Ecology of larval and metamorphosing lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1641–1657, November 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-212.

Moore:1980:FLL

[213] James W. Moore and Jon M. Mallatt. Feeding of larval lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1658–1664, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-213.

Morris:1980:BCO

[214] Ron Morris. Blood composition and osmoregulation in ammocoete larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1665–1679, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-214.

Gorbman:1980:EBP

[215] Aubrey Gorbman. Evolution of the brain–pituitary relationship: Evidence from the agnatha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 1680–1686, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-215.

Youson:1980:MPL

[216] J. H. Youson. Morphology and physiology of lamprey metamorphosis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1687–1710, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-216.

Lewis:1980:RL

[217] S. V. Lewis. Respiration of lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1711–1722, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-217.

Rovainen:1980:NL

[218] Carl M. Rovainen. Neurophysiology of lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1723–1738, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-218.

Beamish:1980:OJA

[219] F. W. H. Beamish. Osmoregulation in juvenile and adult lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1739–1750, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-219.

Farmer:1980:BPF

[220] G. J. Farmer. Biology and physiology of feeding in adult lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1751–1761, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-220.

Larsen:1980:PAL

[221] Lis Olesen Larsen. Physiology of adult lampreys, with special regard to natural starvation, reproduction, and death after spawning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1762–1779, November 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-221.

Smith:1980:SLP

[222] B. R. Smith and J. J. Tibbles. Sea lamprey (Petromyzon marinus) in Lakes Huron, Michigan, and Superior: History of invasion and control, 1936–78. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1780–1801, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-222.

Pearce:1980:SLP

[223] W. A. Pearce, R. A. Braem, S. M. Dustin, and J. J. Tibbles. Sea lamprey (Petromyzon marinus) in the Lower Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1802–1810, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-223.

Morman:1980:FID

[224] R. H. Morman, D. W. Cuddy, and P. C. Rugen. Factors influencing the distribution of sea lamprey (Petromyzon marinus) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1811–1826, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-224.

Purvis:1980:ETM

[225] H. A. Purvis. Effects of temperature on metamorphosis and the age and length at metamorphosis in sea lamprey (Petromyzon marinus) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1827– 1834, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-225.

Torblaa:1980:EIL

[226] R. L. Torblaa and R. W. Westman. Ecological impacts of lampricide treatments on sea lamprey (Petromyzon marinus) ammocoetes and metamorphosed individuals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (11):1835–1850, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-226.

Moore:1980:CSR

[227] H. H. Moore and L. P. Schleen. Changes in spawning runs of sea lamprey (Petromyzon marinus) in selected streams of Lake Superior after chemical control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1851– 1860, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-227.

Heinrich:1980:CBC

[228] John W. Heinrich, Jerry G. Weise, and Bernard R. Smith. Changes in biological characteristics of the sea lamprey (Petromyzon marinus) as related to lamprey abundance, prey abundance, and sea lamprey control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1861–1871, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-228.

Schuldt:1980:CDN

[229] R. J. Schuldt and R. Goold. Changes in the distribution of native lampreys in Lake Superior tributaries in response to sea lamprey (Petromyzon marinus) control, 1953–77. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1872–1885, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-229.

Dahl:1980:ECS

[230] F. H. Dahl and R. B. McDonald. Effects of control of the sea lamprey (Petromyzon marinus) on migratory and resident fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1886–1894, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-230.

Gllderhus:1980:ESL

[231] P. A. Gllderhus and B. G. H. Johnson. Effects of sea lamprey (Petromyzon marinus) control in the Great Lakes on aquatic plants, invertebrates, and amphibians. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1895– 1905, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-231.

Beamish:1980:ABR

[232] Richard J. Beamish. Adult biology of the river lamprey (Lampetra ayresi) and the Pacific lamprey (Lampetra tridentate) from the Pacific Coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1906–1923, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-232.

Beamish:1980:BNA

[233] F. W. H. Beamish. Biology of the North American anadromous sea lamprey, Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (11):1924–1943, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-233.

Maitland:1980:REL

[234] P. S. Maitland. Review of the ecology of lampreys in Northern Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1944–1952, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-234.

Tuunainen:1980:LLF

[235] P. Tuunainen, E. Ikonen, and H. Auvinen. Lampreys and lamprey fisheries in Finland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1953– 1959, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-235.

Kainua:1980:DAE

[236] K. Kainua and T. Valtonen. Distribution and abundance of European river lamprey (Lampetra fluviatilis) larvae in three rivers running into Bothnian Bay, Finland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 1960–1966, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-236.

Valtonen:1980:ERL

[237] Tapani Valtonen. European river lamprey (Lampetra fluviatilis) fishing and lamprey populations in some rivers running into Bothnian Bay, Finland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1967–1973, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-237.

Sjoberg:1980:EER

[238] K. Sjöberg. Ecology of the European river lamprey (Lampetra fluviatilis) in Northern Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 1974–1980, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-238.

Maitland:1980:SWC

[239] P. S. Maitland. Scarring of whitefish (Coregonus lavaretus) by European river lamprey (Lampetra fluviatilis) in Loch Lomond, Scotland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1981–1988, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-239.

King:1980:CSL

[240] Everett Louis King, Jr. Classification of sea lamprey (Petromyzon marinus) attack marks on Great Lakes lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):1989–2006, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-240.

Johnson:1980:PPS

[241] B. G. H. Johnson and William C. Anderson. Predatory-phase sea lampreys (Petromyzon marinus) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2007–2020, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-241.

Christie:1980:PPS

[242] W. J. Christie and D. P. Kolenosky. Parasitic phase of the sea lamprey (Petromyzon marinus) in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2021–2038, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-242.

Spangler:1980:RLW

[243] George R. Spangler and John J. Collins. Response of lake whitefish (Coregonus clupeaformis) to the control of sea lamprey (Petromyzon marinus) in Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 2039–2046, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-243.

Wells:1980:LTS

[244] LaRue Wells. Lake trout (Salvelinus namaycush) and sea lamprey (Petromyzon marinus) populations in Lake Michigan, 1971–78. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2047–2051, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-244.

Moore:1980:CML

[245] J. D. Moore and T. J. Lychwick. Changes in mortality of lake trout (Salvelinus namaycush) in relation to increased sea lamprey (Petromyzon marinus) abundance in Green Bay, 1974–78. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2052–2056, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-245.

Lawrie:1980:EMR

[246] A. H. Lawrie and W. MacCallum. On evaluating measures to rehabilitate lake trout (Salvelinus namaycush) of Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2057–2062, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-246.

Pycha:1980:CML

[247] Richard L. Pycha. Changes in mortality of lake trout (Salvelinus namaycush) in Michigan waters of Lake Superior in relation to sea lamprey (Petromyzon marinus) predation, 1968–78. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2063–2073, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-247.

Swanson:1980:DRL

[248] Bruce L. Swanson and Donald V. Swedberg. Decline and recovery of the Lake Superior Gull Island Reef lake trout (Salvelinus namaycush) population and the role of sea lamprey (Petromyzon marinus) predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2074–2080, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-248.

Sawyer:1980:PIP

[249] Alan J. Sawyer. Prospects for integrated pest management of the sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2081–2092, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-249.

Meyer:1980:PPR

[250] Fred P. Meyer and Rosalie A. Schnick. Potential problems in the registration of sea lamprey (Petromyzon marinus) control agents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2093–2102, November 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-250.

Howell:1980:DSL

[251] John H. Howell, John J. Lech, and John L. Allen. Development of sea lamprey (Petromyzon marinus) larvicides. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2103–2107, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-251.

Hanson:1980:SMP

[252] Lee H. Hanson and Patrick J. Manion. Sterility method of pest control and its potential role in an integrated sea lamprey (Petromyzon marinus) control program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11): 2108–2117, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-252.

Hunn:1980:RPB

[253] J. B. Hunn and W. D. Youngs. Role of physical barriers in the control of sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2118–2122, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-253.

Teeter:1980:PCS

[254] John Teeter. Pheromone communication in sea lampreys (Petromyzon marinus): Implications for population management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2123–2132, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-254.

Walters:1980:RLT

[255] Carl J. Walters, Greg Steer, and George Spangler. Responses of lake trout (Salvelinus namaycush) to harvesting, stocking, and lamprey reduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2133–2145, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-255.

Spangler:1980:ELI

[256] George R. Spangler, Douglas S. Robson, and Henry A. Regier. Estimates of lamprey-induced mortality in whitefish, Coregonus clupeaformis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2146–2150, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-256.

Youngs:1980:ELI

[257] William D. Youngs. Estimate of lamprey-induced mortality in a lake trout (Salvelinus namaycush) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2151–2158, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-257.

Kitchell:1980:BMF

[258] James F. Kitchell and James E. Breck. Bioenergetics model and foraging hypothesis for sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2159–2168, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-258.

Talhelm:1980:BCS

[259] Daniel R. Talhelm and Richard C. Bishop. Benefits and costs of sea lamprey (Petromyzon marinus) control in the Great Lakes: Some preliminary results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2169– 2174, November 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-259.

Lamsa:1980:SLP

[260] A. K. Lamsa, C. M. Rovainen, D. P. Kolenosky, and L. H. Hanson. Sea lamprey (Petromyzon marinus) control — where to from here? Report of the SLIS Control Theory Task Force. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2175–2192, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-260.

Lawrie:1980:IBS

[261] A. H. Lawrie. Interactions between sea lamprey (Petromyzon marinus) and their forage base. a report from the SUS stock effects task force. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2193–2196, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-261.

Maitland:1980:ALF

[262] P. S. Maitland. Assessment of lamprey and fish stocks in the Great Lakes in relation to control of the sea lamprey (Petromyzon marinus): Report from the SLIS Assessment Measurements Task Force. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2197–2201, November 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-262.

Walters:1980:SKU

[263] Carl J. Walters, George Spangler, W. J. Christie, Patrick J. Manion, and James F. Kitchell. A synthesis of knowns, unknowns, and policy recommendations from the Sea Lamprey International Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2202–2208, November 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-263.

Anonymous:1980:SLI

[264] Anonymous. Sea Lamprey International Symposium (SLIS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(11):2209–2214, November 1980. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f80-264.

delaNoue:1980:DCT

[265] J. de la Noüe, G. Choubert, B. Pagniez, J.-M. Blanc, and P. Luquet. Digestibilité chez la truite arc-en-ciel (Salmo gairdneri) lors de l’adaptation à un nouveau régime alimentaire. (French) [Digestibility in rainbow trout (Salmo gairdneri) when adjusting to a new diet]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2218–2224, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-265.

Markiw:1980:MCT

[266] Maria E. Markiw and Ken Wolf. Myxosoma cerebralis: Trypsinization of plankton centrifuge harvests increases optical clarity and spore concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2225–2227, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-266.

Todd:1980:DCZ

[267] Thomas N. Todd and Gerald R. Smith. Differentiation in Coregonus zenithicus in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (12):2228–2235, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f80-267.

Paterson:1980:IIA

[268] W. D. Paterson, D. Douey, and D. Desautels. Isolation and identification of an atypical Aeromonas salmonicida strain causing epizootic losses among Atlantic salmon (Salmo salar) reared in a Nova Scotian hatchery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2236–2241, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f80-268.

Anderson:1980:SES

[269] J. T. Anderson and J. C. Roff. Seston ecology of the surface waters of Hudson Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2242– 2253, December 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-269.

Lowry:1980:VDR

[270] Lloyd F. Lowry, Kathryn J. Frost, and John J. Burns. Variability in the diet of ringed seals, Phoca hispida, in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2254–2261, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-270.

White:1980:RKA

[271] Alan W. White. Recurrence of kills of Atlantic herring (Clupea harengus harengus) caused by dinoflagellate toxins transferred through herbivorous zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12): 2262–2265, December 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-271.

Gunderson:1980:URK

[272] Donald R. Gunderson. Using r-K selection theory to predict natural mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2266–2271, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-272.

Gavaris:1980:UMM

[273] Stratis Gavaris. Use of a multiplicative model to estimate catch rate and effort from commercial data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2272–2275, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-273.

Reno:1980:VEN

[274] Paul W. Reno and Bruce L. Nicholson. Viral erythrocytic necrosis (VEN) in Atlantic cod (Gadus morhua): In vitro studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2276–2281, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-274.

Yan:1980:CZC

[275] N. D. Yan and R. Strus. Crustacean zooplankton communities of acidic, metal-contaminated lakes near Sudbury, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2282–2293, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-275.

Vinyard:1980:DPV

[276] Gary L. Vinyard. Differential prey vulnerability and predator selectivity: Effects of evasive prey on bluegill (Lepomis macrochirus) and pumpkinseed (L. gibhosus) predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2294–2299, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-276.

Jackson:1980:DPP

[277] Togwell A. Jackson and Robert E. Hecky. Depression of primary productivity by humic matter in lake and reservoir waters of the boreal forest zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2300–2317, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f80-277.

Dobroski:1980:ABP

[278] C. J. Dobroski, Jr. and C. E. Epifanio. Accumulation of benzo[a]pyrene in a larval bivalve via trophic transfer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2318–2322, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-278.

McKelvey:1980:SCM

[279] R. McKelvey, D. Hankin, K. Yanosko, and C. Snygg. Stable cycles in multistage recruitment models: an application to the Northern California Dungeness crab (Cancer magister) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2323–2345, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-279.

Smith:1980:CTM

[280] S. J. Smith. Comparison of two methods of estimating the variance of the estimate of catch per unit effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2346–2351, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-280.

Sikstrom:1980:ORP

[281] C. B. Sikstrom. Origin of a relict population of Salvelinus in Bear Creek of the Alsek River Drainage, Yukon Territory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2352–2354, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-281.

Kennedy:1980:TLT

[282] L. A. Kennedy. Teratogenesis in lake trout (Salvelinus namaycush) in an experimentally acidified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2355–2358, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f80-282.

Adler-Fenchel:1980:ADE

[283] Hilary S. Adler-Fenchel. Acoustically derived estimate of the size distribution for a sample of sperm whales (Physeter catodon) in the Western North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2358– 2361, December 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-283.

Jurasz:1980:PST

[284] V. P. Jurasz, D. McSweeney, and C. M. Jurasz. Possible sexing technique for humpback whales (Megaptera novaeangliae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2362–2364, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f80-284.

Anonymous:1980:LLb

[285] Anonymous. Letters/lettres. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37 (12):2365, December 1980. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f80-285.

Anonymous:1980:BRCh

[286] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 37(12):2366, December 1980. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f80-286.

Lin:1981:SVP

[287] C. Kwei Lin and Claire L. Schelske. Seasonal variation of potential nutrient limitation to chlorophyll production in Southern Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):1–9, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-001.

Godin:1981:DPF

[288] Jean-Guy J. Godin. Daily patterns of feeding behavior, daily rations, and diets of juvenile pink salmon (Oncorhynchus gorbuscha) in two marine bays of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (1):10–15, January 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-002.

Jacobs:1981:MCG

[289] Dan Jacobs, Edward F. Esmond, Edward L. Melisky, and Charles H. Hocutt. Morphological changes in gill epithelia of heat-stressed rainbow trout, Salmo gairdneri: Evidence in support of a temperatureinduced surface area change hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):16–22, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-003.

Pierson:1981:ECZ

[290] Keith B. Pierson. Effects of chronic zinc exposure on the growth, sexual maturity, reproduction, and bioaccumulation of the guppy, Poecilia reticulata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):23–31, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-004.

Kornfield:1981:GSA

[291] Irv Kornfield, Kenneth F. Beland, John R. Moring, and Frederick W. Kircheis. Genetic similarity among endemic Arctic char (Salvelinus alpinus) and implications for their management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):32–39, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-005.

Thomson:1981:DTF

[292] Richard E. Thomson and Liliane S. C. Kuwahara. Detecting temperature fine structure and mesoscale frontal regimes along Line P in the Northeastern Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (1):40–51, January 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-006.

Rao:1981:EBP

[293] D. V. Subba Rao. Effect of boron on primary production of nanoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):52–58, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-007.

Todd:1981:EGC

[294] Thomas N. Todd, Gerald R. Smith, and Louella E. Cable. Environmental and genetic contributions to morphological differentiation in ciscoes (Coregoninae) of the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):59–67, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-008.

Dahm:1981:PMR

[295] Clifford N. Dahm. Pathways and mechanisms for removal of dissolved organic carbon from leaf leachate in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):68–76, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-009.

McLaren:1981:TDG

[296] I. A. McLaren and C. J. Corkett. Temperature-dependent growth and production by a marine copepod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):77–83, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-010.

York:1981:PPF

[297] Anne E. York and James R. Hartley. Pup production following harvest of female northern fur seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (1):84–90, January 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-011.

ONeill:1981:SEP

[298] R. V. O’Neill, R. H. Gardner, S. W. Christensen, W. Van Winkle, J. H. Carney, and J. B. Mankin. Some effects of parameter uncertainty in density-independent and density-dependent Leslie models for fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):91–100, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-012.

Harding:1981:IFD

[299] Gareth C. Harding, W. Peter Vass, and Ken F. Drinkwater. Importance of feeding, direct uptake from seawater, and transfer from generation to generation in the accumulation of an organochlorine (p, p′- DDT) by the marine planktonic copepod Calanus finmarchicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):101–119, January 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-013.

Zeitoun:1981:EFM

[300] Ibrahim H. Zeitoun, John A. Gulvas, and Doyle B. Roarabaugh. Effectiveness of fine mesh cylindrical wedge-wire screens in reducing entrainment of Lake Michigan ichthyoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):120–125, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-014.

Anonymous:1981:LCLa

[301] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):126–129, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-015.

Anonymous:1981:AAa

[302] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(1):130, January 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-016.

Michibata:1981:LFAa

[303] Hitoshi Michibata and Reiji Hori. Labeling fish with an activable element. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):133–136, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-017.

Murphy:1981:VEC

[304] Michael L. Murphy and James D. Hall. Vaired effects of clear-cut logging on predators and their habitat in small streams of the Cascade Mountains, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2): 137–145, February 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-018.

Pyle:1981:DHS

[305] Stephen W. Pyle and Emmett B. Shotts, Jr. DNA homology studies of selected flexibacteria associated with fish disease. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):146–151, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-019.

Hurst:1981:PIS

[306] John W. Hurst and Clarice M. Yentsch. Patterns of intoxication of shellfish in the Gulf of Maine coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):152–156, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-020.

Gardner:1981:ICN

[307] Wayne S. Gardner and Warren H. Miller III. Intracellular composition and net release rates of free amino acids in Daphnia magna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):157–162, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-021.

Somerton:1981:RVS

[308] David A. Somerton. Regional variation in the size of maturity of two species of Tanner crab (Chionoecetes bairdi and C. opilio) in the Eastern Bering Sea, and its use in defining management subareas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):163–174, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-022.

Jones:1981:CPA

[309] P. D. Jones and W. T. Momot. Crayfish productivity, allochthony, and basin morphometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2): 175–183, February 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-023.

Allan:1981:DDB

[310] J. David Allan. Determinants of diet of brook trout (Salvelinus fontinalis) in a mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (2):184–192, February 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-024.

Whyte:1981:CCF

[311] J. N. C. Whyte, J. R. Englar, and P. E. Borgmann. Compositional changes on freshwater leaching of the marine algae Nereocystis luetkeana and Macrocystis integrifolia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (2):193–198, February 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-025.

Sharom:1981:ADP

[312] M. S. Sharom and K. R. Solomon. Adsorption and desorption of permethrin and other pesticides on glass and plastic materials used in bioassay procedures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):199–204, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-026.

Brown:1981:ACS

[313] Lewis M. Brown, Barry T. Hargrave, and Michael D. MacKinnon. Analysis of chlorophyll a in sediments by high-pressure liquid chromatography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):205–214, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-027.

Frank:1981:WRE

[314] Kenneth T. Frank and William C. Leggett. Wind regulation of emergence times and early larval survival in capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):215–223, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-028.

Nalewajko:1981:PKL

[315] C. Nalewajko, K. Lee, and H. Shear. Phosphorus kinetics in Lake Superior: Light intensity and phosphate uptake in algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):224–232, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-029.

Bernard:1981:MAM

[316] David R. Bernard. Multivariate analysis as a means of comparing growth in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):233–236, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-030.

DeMaster:1981:EAA

[317] Douglas P. DeMaster. Estimating the average age of first birth in marine mammals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):237–239, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-031.

Mann:1981:CSD

[318] G. J. Mann and P. J. McCart. Comparison of sympatric dwarf and normal populations of least cisco (Coregonus sardinella) inhabiting Trout Lake, Yukon Territory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2): 240–244, February 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-032.

Whoriskey:1981:STS

[319] Frederick G. Whoriskey, Jr., Robert J. Naiman, and Peter H. Heinermann. Steelhead trout (Salmo gairdneri) on the North Shore of the Gulf of St. Lawrence, near Sept-Îles, Québec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):245–246, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-033.

Jensen:1981:AAC

[320] John Willas Jensen. Anadromous Arctic char, Salvelinus alpinus, penetrating southward on the Norwegian Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):247–249, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-034.

Niimi:1981:GGE

[321] A. J. Niimi. Gross growth efficiency of fish (K_{1}) based on field observations of annual growth and kinetics of persistent environmental contaminants. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):250–253, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-035.

Anonymous:1981:CRP

[322] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(2):254, February 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-036.

Vesin:1981:FEC

[323] J-P. Vesin, W. C. Leggett, and K. W. Able. Feeding ecology of capelin (Mallotus villosus) in the estuary and Western Gulf of St. Lawrence and its multispecies implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (3):257–267, March 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-037.

Derby:1981:IDM

[324] Charles D. Derby and Jelle Atema. Influence of drilling muds on the primary chemosensory neurons in walking legs of the lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):268–274, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-038.

Wong:1981:PFB

[325] C. Kim Wong. Predatory feeding behavior of Epischura lacustris (Copepoda, Calanoida) and prey defense. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):275–279, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-039.

Thomann:1981:EMF

[326] Robert V. Thomann. Equilibrium model of fate of microcontaminants in diverse aquatic food chains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (3):280–296, March 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-040.

Clark:1981:RLS

[327] W. G. Clark. Restricted least-squares estimates of age composition from length composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (3):297–307, March 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-041.

Riddell:1981:EAB

[328] Brian E. Riddell and William C. Leggett. Evidence of an adaptive basis for geographic variation in body morphology and time of downstream migration of juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):308–320, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-042.

Riddell:1981:EAP

[329] Brian E. Riddell, William C. Leggett, and Richard L. Saunders. Evidence of adaptive polygenic variation between two populations of Atlantic salmon (Salmo salar) native to tributaries of the S. W. Miramichi River, N.B. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):321–333, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-043.

Lindsey:1981:MRV

[330] C. C. Lindsey and A. N. Arnason. A model for responses of vertebral numbers in fish to environmental influences during development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):334–347, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-044.

Northcote:1981:DRW

[331] T. G. Northcote and B. W. Kelso. Differential response to water current by two homozygous LDH phenotypes of young rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):348–352, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-045.

Haefner:1981:SAI

[332] John D. Haefner and J. Bruce Wallace. Shifts in aquatic insect populations in a first-order Southern Appalachian stream following a decade of old field succession. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (3):353–359, March 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-046.

Buchanan:1981:NSP

[333] David V. Buchanan, Robert M. Hooton, and John R. Moring. Northern squawfish (Ptychocheilus oregonensis) predation on juvenile salmonids in sections of the Willamette River Basin, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):360–364, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-047.

DeMaster:1981:DIW

[334] Douglas P. DeMaster, James B. Faro, James A. Estes, James Taggart, and Cindy Zabel. Drug immobilization of walrus (Odobenus rosmarus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):365–367, March 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-048.

Yu:1981:UBT

[335] T. C. Yu and R. O. Sinnhuber. Use of beef tallow as an energy source in Coho salmon (Oncorhynchus kisutch) rations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):367–370, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-049.

Brethes:1981:EPC

[336] Jean-Claude F. Brêthes and Gaston Desrosiers. Estimation of potential catches of an unexploited stock of soft-shell clam (Mya arenaria) from length composition data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (3):371–374, March 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-050.

Anonymous:1981:LCLb

[337] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):375–376, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-051.

Anonymous:1981:BRCa

[338] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(3):377–378, March 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-052.

Haegele:1981:DED

[339] C. W. Haegele, R. D. Humphreys, and A. S. Hourston. Distribution of eggs by depth and vegetation type in Pacific herring (Clupea harengus pallasi) spawnings in Southern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):381–386, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-053.

Russo:1981:ATN

[340] Rosemarie C. Russo, Robert V. Thurston, and Kenneth Emerson. Acute toxicity of nitrite to rainbow trout (Salmo gairdneri): Effects of pH, nitrite species, and anion species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):387–393, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-054.

McCaughran:1981:EGP

[341] Donald A. McCaughran. Estimating growth parameters for Pacific halibut from mark–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):394–398, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-055.

Brett:1981:ORG

[342] J. R. Brett and J. M. Blackburn. Oxygen requirements for growth of young Coho (Oncorhynchus kisutch) and sockeye (O. nerka) salmon at 15 ^{o}C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):399–404, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-056.

Idler:1981:DSM

[343] D. R. Idler, S. J. Hwang, L. W. Crim, and D. Reddin. Determination of sexual maturation stages of Atlantic salmon (Salmo salar) captured at sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):405–413, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-057.

Canfield:1981:PTP

[344] Daniel E. Canfield, Jr. and Roger W. Bachmann. Prediction of total phosphorus concentrations, chlorophyll a, and secchi depths in natural and artificial lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (4):414–423, April 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-058.

Majkowski:1981:AMA

[345] Jacek Majkowski. Application of a multispecies approach for assessing the population abundance and the age–structure of fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):424–431, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-059.

Berst:1981:RBH

[346] Alfred H. Berst, Alan R. Emery, and George R. Spangler. Reproductive behavior of hybrid charr (Salvelinus fontinalis × S. namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):432–440, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-060.

Evans:1981:DZP

[347] Marlene S. Evans. Distribution of zooplankton populations within and adjacent to a thermal plume. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (4):441–448, April 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-061.

Horner:1981:SPD

[348] R. R. Horner and E. B. Welch. Stream periphyton development in relation to current velocity and nutrients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):449–457, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-062.

McCauley:1981:ERB

[349] Edward McCauley and Jaap Kalff. Empirical relationships between phytoplankton and zooplankton biomass in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):458–463, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-063.

Lewis:1981:IMD

[350] D. H. Lewis. Immunoenzyme microscopy for differentiating among systemic bacterial pathogens of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):463–466, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-064.

McAllister:1981:CFM

[351] Don E. McAllister and R. Jon Planck. Capturing fish measurements and counts with calipers and probe interfaced with a computer or pocket calculator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):466–470, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-065.

Scarnecchia:1981:ESU

[352] Dennis L. Scarnecchia. Effects of streamflow and upwelling on yield of wild Coho salmon (Oncorhynchus kisutch) in Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):471–475, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-066.

Schwinghamer:1981:ELM

[353] Peter Schwinghamer. Extraction of living meiofauna from marine sediments by centrifugation in a silica sol–sorbitol mixture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):476–478, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-067.

Anonymous:1981:BRCb

[354] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):479–480, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-068.

Anonymous:1981:AAb

[355] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(4):483, April 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-069.

Daniels:1981:LTE

[356] Richard E. Daniels and J. David Allan. Life table evaluation of chronic exposure to a pesticide. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (5):485–494, May 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-070.

Bowen:1981:TCR

[357] W. Don Bowen, Charles K. Capstick, and David E. Sergeant. Temporal changes in the reproductive potential of female harp seals (Pagophilus groenlandicus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):495–503, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
071.

Vanderploeg:1981:SPS

[358] H. A. Vanderploeg. Seasonal particle-size selection by Diaptomus sicilis in offshore Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (5):504–517, May 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-072.

Wrench:1981:RMI

[359] J. J. Wrench and R. F. Addison. Reduction, methylation, and incorporation of arsenic into lipids by the marine phytoplankton Dunaliella tertiolecta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):518–523, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-073.

LaZerte:1981:PLP

[360] Bruce D. LaZerte and Susan Watson. The prediction of lacustrine phytoplankton diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (5):524–534, May 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-074.

Mayzaud:1981:KPD

[361] P. Mayzaud and O. Mayzaud. Kinetic properties of digestive carbohydrases and proteases of zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):535–543, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-075.

MacLellan:1981:DEF

[362] Patricia MacLellan, G. E. (Buck) Newsome, and Peter A. Dill. Discrimination by external features between alewife (Alosa pseudoharengus) and blueback herring (A. aestivalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):544–546, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-076.

Mitchum:1981:TBK

[363] Douglas L. Mitchum and Loris E. Sherman. Transmission of bacterial kidney disease from wild to stocked hatchery trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):547–551, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-077.

Beland:1981:ESS

[364] Kenneth F. Beland, Franklin L. Roberts, and Richard L. Saunders. Evidence of Salmo salar × Salmo trutta hybridization in a North American river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):552–554, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-078.

Martin:1981:STR

[365] Roy M. Martin, William R. Heard, and Alex C. Wertheimer. Shortterm rearing of pink salmon (Oncorhynchus gorbuscha) fry: Effect on survival and biomass of returning adults. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):554–558, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-079.

Copes:1981:RFA

[366] Parzival Copes and Jack L. Knetsch. Recreational fisheries analysis: Management modes and benefit implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):559–570, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-080.

Johnson:1981:TOE

[367] Lionel Johnson. The thermodynamic origin of ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):571–590, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-081.

Rosenberg:1981:RTE

[368] David M. Rosenberg, Vincent H. Resh, Steven S. Balling, Mark A. Barnby, Joshua N. Collins, David V. Durbin, Thomas S. Flynn, David D. Hart, Gary A. Lamberti, Eric P. McElravy, John R. Wood, Ted E. Blank, David M. Schultz, Donn L. Marrin, and Donald G. Price. Recent trends in environmental impact assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(5):591–624, May 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-082.

VanWinkle:1981:AAD

[369] W. Van Winkle, D. S. Vaughan, L. W. Barnthouse, and B. L. Kirk. An analysis of the ability to detect reductions in year-class strength of the Hudson River white perch (Morone americana) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):627–632, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-083.

Banse:1981:SCM

[370] Karl Banse. On some cossuridae and maldanidae (polychaeta) from Washington and British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):633–637, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-084.

Therriault:1981:ECP

[371] Jean-Claude Therriault and Trevor Platt. Environmental control of phytoplankton patchiness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (6):638–641, June 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-085.

Turk:1981:ESI

[372] Ted R. Turk and Michael J. Risk. Effect of sedimentation on infaunal invertebrate populations of Cobequid Bay, Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):642–648, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-086.

Iwama:1981:SGM

[373] George K. Iwama and Arthur F. Tautz. A simple growth model for salmonids in hatcheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (6):649–656, June 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-087.

Jensen:1981:OSL

[374] A. L. Jensen. Optimum size limits for trout fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):657–661, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-088.

Crowder:1981:CUF

[375] Larry B. Crowder, John J. Magnuson, and Stephen B. Brandt. Complementarity in the use of food and thermal habitat by Lake Michigan fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):662–668, June 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-089.

Fairbairn:1981:BGA

[376] Daphne J. Fairbairn. Biochemical genetic analysis of population differentiation in Greenland halibut (Reinhardtius hippoglossoides) from the Northwest Atlantic, Gulf of St. Lawrence, and Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):669–677, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-090.

Walters:1981:OEF

[377] Carl J. Walters. Optimum escapements in the face of alternative recruitment hypotheses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):678–689, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
091.

Smith:1981:AMS

[378] A. D. M. Smith and Carl J. Walters. Adaptive management of stock– recruitment systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (6):690–703, June 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-092.

Walters:1981:EME

[379] Carl J. Walters and Donald Ludwig. Effects of measurement errors on the assessment of stock–recruitment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):704–710, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-093.

Ludwig:1981:MEU

[380] Donald Ludwig and Carl J. Walters. Measurement errors and uncertainty in parameter estimates for stock and recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):711–720, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-094.

Stemberger:1981:GAC

[381] Richard S. Stemberger. A general approach to the culture of planktonic rotifers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):721–724, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-095.

Nagy:1981:SRC

[382] A. Nagy, M. Bercsényi, and V. Csányi. Sex reversal in carp (Cyprinus carpio) by oral administration of methyltestosterone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):725–728, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-096.

Anonymous:1981:LCLc

[383] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):729–730, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-097.

Anonymous:1981:BRCc

[384] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):731–733, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-098.

Anonymous:1981:AAc

[385] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(6):735–736, June 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-099.

McKague:1981:STC

[386] A. B. McKague. Some toxic constituents of chlorination-stage effluents from bleached kraft pulp mills. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):739–743, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-100.

Lapi:1981:SSI

[387] Louis A. Lapi and T. J. Mulligan. Salmon stock identification using a microanalytic technique to measure elements present in the freshwater growth region of scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (7):744–751, July 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-101.

Hall:1981:COC

[388] Lenwood W. Hall, Jr., Dennis T. Burton, and Leonard B. Richardson. Comparison of ozone and chlorine toxicity to the developmental stages of striped bass, Morone saxatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):752–757, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-102.

Zaugg:1981:APW

[389] W. S. Zaugg. Advanced photoperiod and water temperature effects on gill Na^{+}–K^{+} adenosine triphosphatase activity and migration of juvenile steelhead (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (7):758–764, July 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-103.

Williams:1981:FDP

[390] D. Dudley Williams. The first diets of postemergent brook trout (Salvelinus fontinalis) and Atlantic salmon (Salmo salar) alevins in a Quebec river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):765–771, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-104.

Maynard:1981:ARJ

[391] Desmond J. Maynard and Douglas D. Weber. Avoidance reactions of juvenile Coho salmon (Oncorhynchus kisutch) to monocyclic aromatics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):772–778, July 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-105.

Weber:1981:ARM

[392] Douglas D. Weber, Desmond J. Maynard, William D. Gronlund, and Vyacheslav Konchin. Avoidance reactions of migrating adult salmon to petroleum hydrocarbons. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (7):779–781, July 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-106.

Fairbairn:1981:WWW

[393] D. J. Fairbairn. Which witch is which? A study of the stock structure of witch flounder (Glyptocephalus cynoglossus) in the Newfoundland region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):782–794, July 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-107.

Glebe:1981:TIP

[394] B. D. Glebe and W. C. Leggett. Temporal, intra-population differences in energy allocation and use by American shad (Alosa sapidissima) during the spawning migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (7):795–805, July 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-108.

Glebe:1981:LDE

[395] B. D. Glebe and W. C. Leggett. Latitudinal differences in energy allocation and use during the freshwater migrations of American shad (Alosa sapidissima) and their life history consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):806–820, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-109.

McAllister:1981:DWE

[396] Don E. McAllister, M. Eric Anderson, and J. G. Hunter. Deep-water eelpouts, Zoarcidae, from Arctic Canada and Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):821–839, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-110.

Winters:1981:GPS

[397] G. H. Winters. Growth patterns in sand lance, Ammodytes dubius, from the Grand Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):841–846, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
111.

Getz:1981:PTM

[398] Wayne M. Getz and Gordon L. Swartzman. A probability transition matrix model for yield estimation in fisheries with highly variable recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):847–855, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-112.

Musial:1981:HPD

[399] Charles J. Musial, John F. Uthe, Gary R. Sirota, B. Garth Burns, Michael W. Gilgan, Vladimir Zitko, and R. Alexander Matheson. Di-n-hexyl phthalate (DHP), a newly identified contaminant in Atlantic herring (Clupea harengus harengus) and Atlantic mackerel (Scomber scombrus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):856–859, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-113.

Newbold:1981:MNS

[400] J. Denis Newbold, Jerry W. Elwood, Robert V. O’Neill, and Webster Van Winkle. Measuring nutrient spiralling in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):860–863, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-114.

Trick:1981:ESM

[401] Charles G. Trick, Paul J. Harrison, and Raymond J. Andersen. Extracellular secondary metabolite production by the marine dinoflagellate Prorocentrum minimum in culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):864–867, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-115.

Anonymous:1981:BRCd

[402] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):868–870, July 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-116.

Anonymous:1981:AAd

[403] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(7):871, July 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-117.

Ni:1981:NCS

[404] I-H. Ni. Numerical classification of sharp-beaked redfishes, Sebastes mentella and S. fasciatus, from Northeastern Grand Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):873–879, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-118.

Dixon:1981:ACR

[405] D. George Dixon and John B. Sprague. Acclimation to copper by rainbow trout (Salmo gairdneri) — a modifying factor in toxicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):880–888, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-119.

Tarapchak:1981:AVB

[406] Stephen J. Tarapchak, David R. Slavens, and Lisa M. Maloney. Abiotic versus biotic uptake of radiophosphorus in lake water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):889–895, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-120.

Mittelbach:1981:PIS

[407] Gary G. Mittelbach. Patterns of invertebrate size and abundance in aquatic habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):896–904, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-121.

Marsot:1981:CDE

[408] Pierre Marsot, Réal Fournier, and Claude Blais. Culture à dialyse: emploi de fibres creuses dialysantes pour la culture massive de phytoplancton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):905–911, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-122.

Vanderploeg:1981:EAL

[409] H. A. Vanderploeg. Effect of the algal length/aperture length ratio on Coulter analyses of Lake Seston. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):912–916, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-123.

Part:1981:UCP

[410] P. Pärt and O. Svanberg. Uptake of cadmium in perfused rainbow trout (Salmo gairdneri) gills. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8): 917–924, August 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-124.

Noga:1981:EWC

[411] Edward J. Noga and James X. Hartmann. Establishment of walking catfish (Clarias batrachus) cell lines and development of a channel catfish (Ictalurus punctatus) virus vaccine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):925–930, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-125.

Skaar:1981:FKM

[412] Donald R. Skaar, B. Thomas Johnson, John R. Jones, and James N. Huckins. Fate of Kepone and Mirex in a model aquatic environment: Sediment, fish, and diet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (8):931–938, August 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-126.

Busack:1981:IHP

[413] Craig A. Busack and G. A. E. Gall. Introgressive hybridization in populations of Paiute cutthroat trout (Salmo clarki seleniris). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):939–951, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-127.

Batterman:1981:DMF

[414] Allan R. Batterman and Philip M. Cook. Determination of mineral fiber concentrations in fish tissues. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (8):952–959, August 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-128.

Watson:1981:RBN

[415] Susan Watson and Jaap Kalff. Relationships between nannoplankton and lake trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (8):960–967, August 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-129.

Roff:1981:RUE

[416] Derek A. Roff. Reproductive uncertainty and the evolution of iteroparity: Why don’t flatfish put all their eggs in one basket? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):968–977, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-130.

Gardner:1981:RPC

[417] Wayne S. Gardner, Thomas F. Nalepa, Michael A. Quigley, and John M. Malczyk. Release of phosphorus by certain benthic invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):978–981, August 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-131.

Beamish:1981:MCP

[418] R. J. Beamish and D. A. Fournier. A method for comparing the precision of a set of age determinations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (8):982–983, August 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-132.

Thurston:1981:ITA

[419] Robert V. Thurston, Glenn R. Phillips, Rosemarie C. Russo, and Susan M. Hinkins. Increased toxicity of ammonia to rainbow trout (Salmo gairdneri) resulting from reduced concentrations of dissolved oxygen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):983–988, August 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-133.

Morgan:1981:ALH

[420] Mark D. Morgan. Abundance, life history, and growth of introduced populations of the opossum shrimp (Mysis relicta) in subalpine California lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):989–993, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-134.

Nusbaum:1981:ASA

[421] Kenneth E. Nusbaum and Emmett B. Shotts, Jr. Absorption of selected antimicrobic drugs from water by channel catfish, Ictalurus punctatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):993–996, August 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-135.

Margolis:1981:APM

[422] L. Margolis, T. E. McDonald, and G. E. Hoskins. Absence of the protozoan Myxosoma cerebralis (Myxozoa: Myxosporea), the cause of whirling disease, in a survey of salmonids from British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):996–998, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-136.

Borgmann:1981:DFM

[423] Uwe Borgmann. Determination of free metal ion concentrations using bioassays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):999–1002, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
137.

Roff:1981:EPR

[424] Derek A. Roff. On estimating partial recruitment in virtual population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):1003–1005, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-

Anonymous:1981:BRCe

[425] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(8):1006–1015, August 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f81-139.

Wildish:1981:MES

[426] D. J. Wildish and D. Peer. Methods for estimating secondary production in marine amphipoda. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9): 1019–1026, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-140.

Nalepa:1981:SMS

[427] Thomas F. Nalepa and Andrew Robertson. Screen mesh size affects estimates of macro- and meio-benthos abundance and biomass in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1027– 1034, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-141.

Murphy:1981:AVH

[428] T. P. Murphy and B. G. Brownlee. Ammonia volatilization in a hypertrophic prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1035– 1039, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-142.

Murphy:1981:BGA

[429] T. P. Murphy and B. G. Brownlee. Blue-green algal ammonia uptake in hypertrophic prairie lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1040–1044, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-143.

Watkins:1981:PVA

[430] W. D. Watkins, R. E. Wolke, and V. J. Cabelli. Pathogenicity of Vibrio anguillarum for juvenile winter flounder, Pseudopleuronectes americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1045–1051, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-144.

Schlesinger:1981:SGR

[431] D. A. Schlesinger, L. A. Molot, and B. J. Shuter. Specific growth rates of freshwater algae in relation to cell size and light intensity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1052–1058, September 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-145.

Strong:1981:VCB

[432] Carol R. Strong and Samuel N. Luoma. Variations in the correlation of body size with concentrations of Cu and Ag in the bivalve Macoma balthica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1059–1064, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-146.

Herman:1981:VHD

[433] Alex W. Herman, Douglas D. Sameoto, and Alan R. Longhurst. Vertical and horizontal distribution patterns of copepods near the shelf break south of Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9): 1065–1076, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-147.

Arnason:1981:BLP

[434] A. N. Arnason and K. H. Mills. Bias and loss of precision due to tag loss in Jolly–Seber estimates for mark–recapture experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1077–1095, September 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-148.

Orr:1981:RAD

[435] Marshall H. Orr. Remote acoustic detection of zooplankton response to field processes, oceanographic instrumentation, and predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1096–1105, September 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-149.

Tomasso:1981:PCD

[436] J. R. Tomasso, Kenneth B. Davis, and Bill A. Simco. Plasma corticosteroid dynamics in channel catfish (Ictalurus punctatus) exposed to ammonia and nitrite. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9): 1106–1112, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-150.

Peterman:1981:FRV

[437] Randall M. Peterman. Form of random variation in salmon smolt-to-adult relations and its influence on production estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1113–1119, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-151.

Doyle:1981:DEA

[438] R. W. Doyle and W. Hunte. Demography of an estuarine amphipod (Gammarus lawrencianus) experimentally selected for high “r”: a model of the genetic effects of environmental change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1120–1127, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-152.

Schnute:1981:VGM

[439] Jon Schnute. A versatile growth model with statistically stable parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1128–1140, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-153.

Kimura:1981:VLI

[440] Daniel K. Kimura and Norman A. Lemberg. Variability of line intercept density estimates (a simulation study of the variance of hydroacoustic biomass estimates). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1141– 1152, September 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-154.

Bugden:1981:SHB

[441] G. L. Bugden. Salt and heat budgets for the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1153–1167, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-155.

Welch:1981:MFS

[442] Harold E. Welch and Kenneth H. Mills. Marking fish by scarring soft fin rays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1168–1170, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-156.

Anonymous:1981:BRCf

[443] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(9):1171, September 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f81-157.

Hughes:1981:SAL

[444] Steven E. Hughes and Neil Bourne. Stock assessment and life history of a newly discovered Alaska surf clam (Spisula polynyma) resource in the Southeastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10): 1173–1181, October 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-158.

Koss:1981:HLR

[445] Ted F. Koss and Arthur H. Houston. Hemoglobin levels and red cell ionic composition in goldfish (Carassius auratus) exposed to constant and diurnally cycling temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1182–1188, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-159.

Johnston:1981:PST

[446] C. E. Johnston and R. L. Saunders. Parr–smolt transformation of yearling Atlantic salmon (Salmo salar) at several rearing temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1189–1198, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-160.

Majkowski:1981:PEF

[447] Jacek Majkowski and Kenneth G. Waiwood. A procedure for evaluating the food biomass consumed by a fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1199–1208, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-161.

Marshall:1981:SSC

[448] J. S. Marshall, J. I. Parker, D. L. Mellinger, and S. G. Lawrence. An in situ study of cadmium and mercury stress in the plankton community of Lake 382, Experimental Lakes Area, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1209–1214, October 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-162.

White:1981:OFL

[449] E. White, G. Payne, S. Pickmere, and F. R. Pick. Orthophosphate and its flux in lake waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10): 1215–1219, October 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-163.

Fausch:1981:CBB

[450] Kurt D. Fausch and Ray J. White. Competition between brook trout (Salvelinus fontinalis) and brown trout (Salmo trutta) for positions in a Michigan stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10): 1220–1227, October 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-164.

Nilsson:1981:RTS

[451] Nils-Arvid Nilsson and Thomas G. Northcote. Rainbow trout (Salmo gairdneri) and cutthroat trout (S. clarki) interactions in coastal British Columbia lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1228– 1246, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f81-165.

Stocker:1981:STF

[452] Max Stocker and Ray Hilborn. Short-term forecasting in marine fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1247–1254, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
166.

Schwinghamer:1981:CSD

[453] Peter Schwinghamer. Characteristic size distributions of integral benthic communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1255– 1263, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f81-167.

Luecke:1981:PLV

[454] Chris Luecke and W. John O’Brien. Prey location volume of a planktivorous fish: a new measure of prey vulnerability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1264–1270, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-168.

Reice:1981:IAW

[455] Seth R. Reice. Interspecific associations in a woodland stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1271–1280, October 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-169.

Michibata:1981:LFAb

[456] Hitoshi Michibata. Labeling fish with an activable element through their diet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1281–1282, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f81-
170.

Eales:1981:EEL

[457] J. G. Eales. Extrathyroidal effects of low concentrations of thiourea on rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1283–1285, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-171.

Hanson:1981:SSL

[458] Lee H. Hanson. Sterilization of sea lampreys (Petromyzon marinus) by immersion in an aqueous solution of bisazir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1285–1289, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-172.

Bienfang:1981:STS

[459] Paul K. Bienfang. SETCOL — a technologically simple and reliable method for measuring phytoplankton sinking rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1289–1294, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-173.

Anonymous:1981:LCLd

[460] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1295–1300, October 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-174.

Anonymous:1981:BRCg

[461] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(10):1301–1304, October 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f81-175.

Rodgers:1981:UWM

[462] D. W. Rodgers and F. W. H. Beamish. Uptake of waterborne methylmercury by rainbow trout (Salmo gairdneri) in relation to oxygen consumption and methylmercury concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1309–1315, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-176.

Taylor:1981:REP

[463] W. D. Taylor and D. R. S. Lean. Radiotracer experiments on phosphorus uptake and release by limnetic microzooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1316–1321, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-177.

Cipriano:1981:ASR

[464] Rocco C. Cipriano, Billy R. Griffin, and Burton C. Lidgerding. Aeromonas salmonicida: Relationship between extracellular growth products and isolate virulence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1322–1326, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-178.

Frank:1981:PED

[465] Kenneth T. Frank and William C. Leggett. Prediction of egg development and mortality rates in capelin (Mallotus viliosus) from meteorological, hydrographic, and biological factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1327–1338, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-179.

Wharton:1981:RBD

[466] W. G. Wharton and K. H. Mann. Relationship between destructive grazing by the sea urchin, Strongylocentrotus droebachiensis, and the abundance of American lobster, Homarus americanus, on the Atlantic Coast of Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11): 1339–1349, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-180.

Niimi:1981:EHH

[467] A. J. Niimi and C. Y. Cho. Elimination of hexachlorobenzene (HCB) by rainbow trout (Salmo gairdneri), and an examination of its kinetics in Lake Ontario salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11): 1350–1356, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-181.

Neill:1981:DRJ

[468] William E. Neill. Developmental responses of juvenile Daphnia rosea to experimental alteration of temperature and natural seston concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1357–1362, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-182.

Gore:1981:PMB

[469] James A. Gore and Robert D. Judy, Jr. Predictive models of benthic macroinvertebrate density for use in instream flow studies and regulated flow management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11): 1363–1370, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-183.

Hudon:1981:ICA

[470] Christiane Hudon and Edwin Bourget. Initial colonization of artificial substrate: Community development and structure studied by scanning electron microscopy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11): 1371–1384, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-184.

Legendre:1981:PCP

[471] L. Legendre, R. G. Ingram, and M. Poulin. Physical control of phytoplankton production under sea ice (Manitounuk Sound, Hudson Bay). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1385–1392, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-185.

Agnello:1981:PRM

[472] R. J. Agnello and L. G. Anderson. Production responses for multi-species fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1393–1404, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-186.

Helfman:1981:TAT

[473] Gene S. Helfman. Twilight activities and temporal structure in a freshwater fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11): 1405–1420, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-187.

Smith:1981:EPL

[474] Ralph E. H. Smith and Jaap Kalff. The effect of phosphorus limitation on algal growth rates: Evidence from alkaline phosphatase. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1421–1427, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-188.

Roff:1981:SCZ

[475] J. C. Roff, W. G. Sprules, J. C. H. Carter, and M. J. Dadswell. The structure of crustacean zooplankton communities in glaciated Eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1428– 1437, November 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-189.

Coyle:1981:LJS

[476] Kenneth O. Coyle and George J. Mueller. Larval and juvenile stages of the isopod Holophryxus alaskensis (Epicarida, Dajidae) parasitic on decapods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1438–1443, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-190.

Brusven:1981:ISC

[477] M. A. Brusven and S. T. Rose. Influence of substrate composition and suspended sediment on insect predation by the torrent sculpin, Cottus rhotheus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1444–1448, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-191.

Poirier:1981:DTD

[478] Louis Poirier and Daniel Cossa. Distribution tissulaire du cadmium chez Meganyctiphanes norvegica (Euphausiacée): état naturel et accumulation expérimentale de formes solubles. (French) [Tissue distribution of cadmium in Meganyctiphanes norvegica (Euphausia): natural state and experimental accumulation of soluble forms]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(11):1449–1453, November 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-192.

Anonymous:1981:SCI

[479] Anonymous. Stock Concept International Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12), December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-193.

Berst:1981:IPS

[480] Alfred H. Berst and Raymond C. Simon. Introduction to the Proceedings of the 1980 Stock Concept International Symposium (STOCS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1457–1458, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-194.

Harlan:1981:WC

[481] Jack R. Harlan. Who’s in charge? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1459–1463, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-195.

Loftus:1981:WWN

[482] K. H. Loftus. Where are we now? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1464–1465, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-196.

Borgeson:1981:CMG

[483] David P. Borgeson. Changing management of Great Lakes fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1466–1468, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-197.

Larkin:1981:PPG

[484] P. A. Larkin. A perspective on population genetics and salmon management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1469–1475, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-198.

Kutkuhn:1981:SDN

[485] Joseph H. Kutkuhn. Stock definition as a necessary basis for cooperative management of Great Lakes fish resources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1476–1478, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-199.

Booke:1981:CSC

[486] Henry E. Booke. The conundrum of the stock concept — are nature and nurture definable in fishery science? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1479–1480, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-200.

Horrall:1981:BSI

[487] Ross M. Horrall. Behavioral stock-isolating mechanisms in Great Lakes fishes with special reference to homing and site imprinting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1481–1496, December 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-201.

Lindsey:1981:SCP

[488] C. C. Lindsey. Stocks are chameleons: Plasticity in gill rakers of coregonid fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1497– 1506, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-202.

Allendorf:1981:UAF

[489] Fred W. Allendorf and Stevan R. Phelps. Use of allelic frequencies to describe population structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1507–1514, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-203.

Clayton:1981:SCU

[490] James W. Clayton. The stock concept and the uncoupling of organismal and molecular evolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12): 1515–1522, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-204.

Altukhov:1981:SCV

[491] Yu. P. Altukhov. The stock concept from the viewpoint of population genetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1523–1538, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-205.

Bailey:1981:OGF

[492] Reeve M. Bailey and Gerald R. Smith. Origin and geography of the fish fauna of the Laurentian Great Lakes Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1539–1561, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-206.

Ryman:1981:GPI

[493] Nils Ryman and Gunnar Ståhl. Genetic perspectives of the identification and conservation of Scandinavian stocks of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1562–1575, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-207.

Thorpe:1981:SAS

[494] J. E. Thorpe and K. A. Mitchell. Stocks of Atlantic salmon (Salmo salar) in Britain and Ireland: Discreteness, and current management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1576–1590, December 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-208.

Altukhov:1981:ASC

[495] Yu. P. Altukhov and E. A. Salmenkova. Applications of the stock concept to fish populations in the U.S. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1591–1600, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-209.

Power:1981:SCC

[496] G. Power. Stock characteristics and catches of Atlantic salmon (Salmo salar) in Quebec, and Newfoundland and Labrador in relation to environmental variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12): 1601–1611, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-210.

Saunders:1981:ASS

[497] Richard L. Saunders. Atlantic salmon (Salmo salar) stocks and management implications in the Canadian Atlantic provinces and New England, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1612–1625, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-211.

Utter:1981:BCD

[498] Fred M. Utter. Biological criteria for definition of species and distinct intraspecific populations of anadromous salmonids under the U.S. Endangered Species Act of 1973. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38 (12):1626–1635, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f81-212.

Ricker:1981:CAS

[499] W. E. Ricker. Changes in the average size and average age of Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1636–1656, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-213.

McDonald:1981:SCA

[500] J. McDonald. The stock concept and its application to British Columbia salmon fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1657– 1664, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-214.

Helle:1981:SSC

[501] John H. Helle. Significance of the stock concept in artificial propagation of salmonids in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12): 1665–1671, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-215.

Fraser:1981:CSG

[502] J. M. Fraser. Comparative survival and growth of planted wild, hybrid, and domestic strains of brook trout (Salvelinus fontinalis) in Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1672–1684, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-216.

MaClean:1981:SGS

[503] J. A. MaClean, D. O. Evans, N. V. Martin, and R. L. DesJardine. Survival, growth, spawning distribution, and movements of introduced and native lake trout (Salvelinus namaycush) in two inland Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1685–1700, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-217.

Webster:1981:PIE

[504] Dwight A. Webster and William A. Flick. Performance of indigenous, exotic, and hybrid strains of brook trout (Salvelinus fontinalis) in waters of the Adirondack Mountains, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1701–1707, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-218.

Brussard:1981:SAF

[505] Peter F. Brussard, Marjorie Collings Hall, and Janet Wright. Structure and affinities of freshwater sea lamprey (Petromyzon marinus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1708–1714, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-219.

Philipp:1981:MID

[506] David P. Philipp, William F. Childers, and Gregory S. Whitt. Management implications for different genetic stocks of largemouth bass (Micropterus salmoides) in the United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1715–1723, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-220.

Goodier:1981:NLT

[507] John L. Goodier. Native lake trout (Salvelinus namaycush) stocks in the Canadian waters of Lake Superior prior to 1955. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1724–1737, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-221.

Dehring:1981:SGV

[508] Terrence R. Dehring, Anne F. Brown, Charles H. Daugherty, and Stevan R. Phelps. Survey of the genetic variation among Eastern Lake Superior lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1738–1746, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-222.

Brown:1981:HED

[509] E. H. Brown, Jr., G. W. Eck, N. R. Foster, R. M. Horrall, and C. E. Coberly. Historical evidence for discrete stocks of lake trout (Salvelinus namaycush) in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1747–1758, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-223.

Biette:1981:LHT

[510] R. M. Biette, D. P. Dodge, R. L. Hassinger, and T. M. Stauffer. Life history and timing of migrations and spawning behavior of rainbow trout (Salmo gairdneri) populations of the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1759–1771, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-224.

Casselman:1981:LWC

[511] J. M. Casselman, J. J. Collins, E. J. Grossman, P. E. Ihssen, and G. R. Spangler. Lake whitefish (Coregonus clupeaformis) stocks of the Ontario waters of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12): 1772–1789, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-225.

Ihssen:1981:LHM

[512] P. E. Ihssen, D. O. Evans, W. J. Christie, J. A. Reckahn, and R. L. DesJardine. Life history, morphology, and electrophoretic characteristics of five allopatric stocks of lake whitefish (Coregonus clupeaformis) in the Great Lakes region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):

1790–1807, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-226.

Todd:1981:AVS

[513] Thomas N. Todd. Allelic variability in species and stocks of Lake Superior ciscoes (Coregoninae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12): 1808–1813, December 1981. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f81-227.

Colby:1981:VAS

[514] Peter J. Colby and Stephen J. Nepszy. Variation among stocks of walleye (Stizostedion vitreum vitreum): Management implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1814–1831, December 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-228.

Krueger:1981:GIS

[515] Charles C. Krueger and George R. Spangler. Genetic identification of sea lamprey (Petromyzon marinus) populations from the Lake Superior Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1832–1837, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-229.

Ihssen:1981:SIM

[516] P. E. Ihssen, H. E. Booke, J. M. Casselman, J. M. McGlade, N. R. Payne, and F. M. Utter. Stock identification: Materials and methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1838–1855, December 1981. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-230.

Ryder:1981:CCF

[517] R. A. Ryder, S. R. Kerr, W. W. Taylor, and P. A. Larkin. Community consequences of fish stock diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1856–1866, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-231.

Hynes:1981:GCF

[518] J. D. Hynes, E. H. Brown, Jr., J. H. Helle, N. Ryman, and D. A. Webster. Guidelines for the culture of fish stocks for resource management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1867–1876, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f81-232.

Krueger:1981:GAF

[519] Charles C. Krueger, Anthony J. Gharrett, Terrence R. Dehring, and Fred W. Allendorf. Genetic aspects of fisheries rehabilitation programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1877–1881, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f81-233.

Maitland:1981:WST

[520] P. S. Maitland, H. A. Regier, G. Power, and N. A. Nilsson. A wild salmon, trout, and char watch: an international strategy for salmonid conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1882–1888, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-234.

MacLean:1981:SCD

[521] J. A. MacLean and D. O. Evans. The stock concept, discreteness of fish stocks, and fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1889–1898, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f81-235.

Thorpe:1981:AMM

[522] J. E. Thorpe, J. F. Koonce, D. Borgeson, B. Henderson, A. Lamsa, P. S. Maitland, M. A. Ross, R. C. Simon, and C. Walters. Assessing and managing man’s impact on fish genetic resources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1899–1907, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f81-236.

Spangler:1981:PPR

[523] G. R. Spangler, A. H. Berst, and J. F. Koonce. Perspectives and policy recommendations on the relevance of the stock concept to fishery management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 38(12):1908–1914, December 1981. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f81-237.

Anonymous:1982:EE

[524] Anonymous. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):1–2, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-001.

Ware:1982:PEF

[525] D. M. Ware. Power and evolutionary fitness of teleosts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):3–13, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-002.

Anonymous:1982:P

[526] Anonymous. Prologue. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1): 16, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-003.

Crim:1982:EMA

[527] L. W. Crim. Environmental modulation of annual and daily rhythms associated with reproduction in teleost fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):17–21, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-004.

Liley:1982:CCF

[528] N. R. Liley. Chemical communication in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):22–35, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-005.

Demski:1982:HCF

[529] L. S. Demski and P. J. Hornby. Hormonal control of fish reproductive behavior: Brain–gonadal steroid interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):36–47, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-006.

Peter:1982:NCR

[530] R. E. Peter. Neuroendocrine control of reproduction in teleosts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):48–55, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-007.

Nagahama:1982:CSS

[531] Y. Nagahama, H. Kagawa, and G. Young. Cellular sources of sex steroids in teleost gonads. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):56–64, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-008.

Billard:1982:ECS

[532] R. Billard, A. Fostier, C. Weil, and B. Breton. Endocrine control of spermatogenesis in teleost fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (1):65–79, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-009.

Burzawa-Gerard:1982:CDP

[533] E. Burzawa-Gerard. Chemical data on pituitary gonadotropins and their implication to evolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (1):80–91, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-010.

Stacey:1982:RPF

[534] N. E. Stacey and F. W. Goetz. Role of prostaglandins in fish reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):92–98, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-011.

Donaldson:1982:SCF

[535] Edward M. Donaldson and George A. Hunter. Sex control in fish with particular reference to salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):99–110, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-012.

Lam:1982:AEF

[536] T. J. Lam. Applications of endocrinology to fish culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):111–137, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-013.

Lam:1982:FCS

[537] T. J. Lam. Fish culture in Southeast Asia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):138–142, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-014.

Hao-Ren:1982:PSF

[538] Lin Hao-Ren. Polycultural system of freshwater fish in China. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):143–150, January 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-015.

Larkin:1982:ANA

[539] P. A. Larkin. Aquaculture in North America: an assessment of future prospects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):151–154, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f82-

Hoar:1982:QUA

[540] William S. Hoar. Quid usque adhuc? Quo vadimus? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):155–159, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-017.

Anonymous:1982:GSR

[541] Anonymous. Graduate students and research associates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):160–161, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-018.

Cherry:1982:ITS

[542] Donald S. Cherry, Stephan R. Larrick, James D. Giattina, John Cairns, Jr., and John Van Hassel. Influence of temperature selection upon the chlorine avoidance of cold-water and warmwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):162–173, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-019.

Aminot:1982:DAU

[543] A. Aminot and R. Kerouel. Dosage automatique de l’urée dans l’eau de mer: une méthode très sensible à la diacétylmonoxime. (French) [Automatic determination of urea in sea water: a very sensitive method for diacetylmonoxime]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1): 174–183, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-020.

Heming:1982:ETU

[544] T. A. Heming. Effects of temperature on utilization of yolk by Chinook salmon (Oncorhynchus tshawytscha) eggs and alevins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):184–190, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-021.

Windom:1982:MAP

[545] H. L. Windom, K. T. Tenore, and D. L. Rice. Metal accumulation by the polychaete Capitella capitata: Influences of metal content and nutritional quality of detritus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (1):191–196, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-022.

Sims:1982:ASC

[546] S. E. Sims. Algorithms for solving the catch equation forward and backward in time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):197–202, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-023.

Cooper:1982:EFA

[547] Scott D. Cooper and Charles R. Goldman. Environmental factors affecting predation rates of Mysis relicta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):203–208, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-024.

Vinyard:1982:VKS

[548] Gary L. Vinyard. Variable kinematics of Sacramento perch (Archoplites interruptus) capturing evasive and nonevasive prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):208–211, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-025.

Anderson:1982:OUF

[549] Lee G. Anderson. Optimal utilization of fisheries with increasing costs of effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):211–214, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-026.

Zaugg:1982:SPA

[550] W. S. Zaugg. A simplified preparation for adenosine triphosphatase determination in gill tissue. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (1):215–217, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-027.

Cipriano:1982:IBT

[551] Rocco C. Cipriano. Immunization of brook trout (Salvelinus fontinalis) against Aeromonas salmonicida: Immunogenicity of virulent and avirulent isolates and protective ability of different antigens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):218–221, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-028.

Mills:1982:EPW

[552] Eric L. Mills, Karin Pittman, and Brent Munroe. Effect of preservation on the weight of marine benthic invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):221–224, January 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-029.

Swanson:1982:IIP

[553] Robert N. Swanson and James H. Gillespie. Isolation of infectious pancreatic necrosis virus from the blood and blood components of experimentally infected trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (1):225–228, January 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-030.

Williamson:1982:CSE

[554] Neal J. Williamson. Cluster sampling estimation of the variance of abundance estimates derived from quantitative echo sounder surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):229–231, January 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-031.

Anonymous:1982:BRCa

[555] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(1):232–234, January 1982. CODEN CJFSDX.

ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f82-032.

Giguere:1982:PER

[556] Louis A. Giguère, André Delâge, Lawrence M. Dill, and Jeroen Gerritsen. Predicting encounter rates for zooplankton: a model assuming a cylindrical encounter field. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (2):237–242, February 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-033.

Carignan:1982:EME

[557] R. Carignan. An empirical model to estimate the relative importance of roots in phosphorus uptake by aquatic macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):243–247, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-034.

Brawn:1982:BAS

[558] Vivien M. Brawn. Behavior of Atlantic salmon (Salmo salar) during suspended migration in an estuary, Sheet Harbour, Nova Scotia, observed visually and by ultrasonic tracking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):248–256, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-035.

Hanson:1982:EPF

[559] John Mark Hanson and William C. Leggett. Empirical prediction of fish biomass and yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2): 257–263, February 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-036.

Gardner:1982:UCD

[560] Peter N. Gardner. Use of cycle dummy variables and race escapement in the Fraser River sockeye salmon (Oncorhynchus nerka) recruitment function. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):264–269, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-037.

Levy:1982:JSR

[561] D. A. Levy and T. G. Northcote. Juvenile salmon residency in a marsh area of the Fraser River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):270–276, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-038.

Beamish:1982:PEM

[562] Richard J. Beamish and Doris E. Chilton. Preliminary evaluation of a method to determine the age of sablefish (Anoplopoma fimbria). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):277–287, February 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-039.

Sinclair:1982:GMS

[563] M. Sinclair, A. Sinclair, and T. D. Iles. Growth and maturation of Southwest Nova Scotia Atlantic herring (Clupea harengus harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):288–295, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-040.

Tarapchak:1982:OOC

[564] Stephen J. Tarapchak, Sylvia M. Bigelow, and Cora Rubitschun. Overestimation of orthophosphorus concentrations in surface waters of Southern Lake Michigan: Effects of acid and ammonium molybdate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):296–304, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-041.

Spigarelli:1982:FAD

[565] S. A. Spigarelli, R. M. Goldstein, W. Prepejchal, and M. M. Thommes. Fish abundance and distribution near three heated effluents to Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):305–315, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-042.

Keen:1982:BIB

[566] W. Hubert Keen. Behavioral interactions and body size differences in competition for food among juvenile brown bullhead (Ictalurus nebulosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):316–320, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-043.

Jensen:1982:DAA

[567] Arne Johan Jensen and Bjørn Ove Johnsen. Difficulties in aging Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) from cold rivers due to lack of scales as yearlings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):321–325, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-044.

Kulka:1982:CSB

[568] D. W. Kulka, S. Corey, and T. D. Iles. Community structure and biomass of euphausiids in the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):326–334, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-045.

Harrison:1982:PPN

[569] W. G. Harrison, Trevor Platt, and Brian Irwin. Primary production and nutrient assimilation by natural phytoplankton populations of the Eastern Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (2):335–345, February 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-046.

Ross:1982:TSP

[570] Marvin Jon Ross and Donald B. Siniff. Temperatures selected in a power plant thermal effluent by adult yellow perch (Perca flavescens) in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):346–349, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-047.

Nero:1982:CTS

[571] R. W. Nero and I. J. Davies. Comparison of two sampling methods for estimating the abundance and distribution of Mysis relicta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):349–355, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-048.

Huebner:1982:RGE

[572] Judith D. Huebner and Richard W. Langton. Rate of gastric evacuation for winter flounder, Pseudopleuronectes americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):356–360, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-049.

Calhoun:1982:SRA

[573] Stuart W. Calhoun, Earl G. Zimmerman, and Thomas L. Beitinger. Stream regulation alters acute temperature preferenda of red shiners, Notropis lutrensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2): 360–363, February 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-050.

Anonymous:1982:LCLa

[574] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):364–366, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-051.

Anonymous:1982:BRCb

[575] Anonymous. Book review/compte rendu. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):367, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-052.

Anonymous:1982:CRPa

[576] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(2):368, February 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-053.

Kerr:1982:EEB

[577] S. R. Kerr. Estimating the energy budgets of actively predatory fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):371–379, March 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-054.

Lederman:1982:BHG

[578] Timoth C. Lederman and G-Yull Rhee. Bioconcentration of a hexachlorobiphenyl in Great Lakes planktonic algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):380–387, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-055.

Lederman:1982:IHG

[579] Timoth C. Lederman and G-Yull Rhee. Influence of a hexachlorobiphenyl in Great Lakes phytoplankton in continuous culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):388–394, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-056.

Craig:1982:ESA

[580] P. C. Craig, W. B. Griffiths, L. Haldorson, and H. McElderry. Ecological studies of Arctic cod (Boreogadus saida) in Beaufort Sea coastal waters, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):395–406, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-057.

Stillwell:1982:FFH

[581] C. E. Stillwell and N. E. Kohler. Food, feeding habits, and estimates of daily ration of the shortfin mako (Isurus oxyrinchus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):407–414, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-058.

Koslow:1982:LEP

[582] J. Anthony Koslow. Limited entry policy and the Bristol Bay, Alaska salmon fishermen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):415–425, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-059.

Bilton:1982:ITS

[583] H. T. Bilton, D. F. Alderdice, and J. T. Schnute. Influence of time and size at release of juvenile Coho salmon (Oncorhynchus kisutch) on returns at maturity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (3):426–447, March 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-060.

Eggers:1982:MEA

[584] Douglas M. Eggers, Neil A. Rickard, Douglas G. Chapman, and Richard R. Whitney. A methodology for estimating area fished for baited hooks and traps along a ground line. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):448–453, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-061.

Moores:1982:GPN

[585] J. A. Moores and G. H. Winters. Growth patterns in a Newfoundland Atlantic herring (Clupea harengus harengus) stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):454–461, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-062.

Silvert:1982:OUV

[586] William Silvert. Optimal utilization of a variable fish supply. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):462–468, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-063.

White:1982:FIO

[587] E. White, G. Payne, S. Pickmere, and F. R. Pick. Factors influencing orthophosphate turnover times: a comparison of Canadian and New Zealand lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):469–474, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-064.

Schmidt:1982:PFE

[588] David Schmidt and W. John O’Brien. Planktivorous feeding ecology of Arctic grayling (Thymallus arcticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):475–482, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-065.

Wong:1982:STR

[589] P. T. S. Wong, Y. K. Chau, O. Kramar, and G. A. Bengert. Structure– toxicity relationship of tin compounds on algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):483–488, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-066.

Stacey:1982:SFB

[590] N. E. Stacey and A. S. Hourston. Spawning and feeding behavior of captive Pacific herring, Clupea harengus pallasi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):489–498, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-067.

Jamieson:1982:PSS

[591] G. S. Jamieson, H. Stone, and M. Etter. Predation of sea scallops (Placopecten magellanicus) by lobsters (Homarus americanus) and rock crabs (Cancer irroratus) in underwater cage enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):499–505, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-068.

Bistricki:1982:AXR

[592] T. Bistricki and M. Munawar. Application of X-ray microanalysis for the detection of heavy metals in Great Lakes algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):506–509, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-069.

Iwata:1982:DSA

[593] Munehico Iwata, Sanae Hasegawa, and Tetsuya Hirano. Decreased seawater adaptability of chum salmon (Oncorhynchus keta) fry following prolonged rearing in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (3):509–514, March 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-070.

Policansky:1982:IAS

[594] David Policansky. Influence of age, size, and temperature on metamorphosis in the starry flounder, Platichthys stellatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):514–517, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-071.

Thompson:1982:ATS

[595] Kenneth W. Thompson, Michael L. Deaton, Robert V. Foutz, John Cairns, Jr., and Albert C. Hendricks. Application of time-series intervention analysis to fish ventilatory response data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):518–521, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-072.

Schrader:1982:ICS

[596] G. C. Schrader, R. Horner, and G. F. Smith. An improved chamber for in situ measurement of primary productivity by sea ice algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):522–524, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-073.

Giattina:1982:GMT

[597] James D. Giattina and Ronald R. Garton. Graphical model of thermoregulatory behavior by fishes with a new measure of eurythermality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):524–528, March 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-074.

Anonymous:1982:LCLb

[598] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(3):529–530, March 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-075.

Vetter:1982:UAC

[599] Russell D. Vetter and Robert E. Hodson. Use of adenylate concentrations and adenylate energy charge as indicators of hypoxic stress in estuarine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):535–541, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-076.

Marshall:1982:PIM

[600] Scott L. Marshall and Steven S. Parker. Pattern identification in the microstructure of sockeye salmon (Oncorhynchus nerka) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):542–547, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-077.

Mathur:1982:TPA

[601] Dilip Mathur, Robert M. Schutsky, and Edmund J. Purdy, Jr. Temperature preference and avoidance responses of the crayfish, Orconectes obscurus, and associated statistical problems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):548–553, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-078.

Selgeby:1982:DLH

[602] James H. Selgeby. Decline of lake herring (Coregonus artedii) in Lake Superior: an analysis of the Wisconsin herring fishery, 1936–78. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):554–563, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-079.

Bowering:1982:CWF

[603] W. R. Bowering and R. K. Misra. Comparisons of witch flounder (Glyptocephalus cynoglossus) stocks of the Newfoundland–Labrador area, based upon a new multivariate analysis method for meristic characters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):564–570, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-080.

Kaiser:1982:ETD

[604] Klaus L. E. Kaiser. Early trend determination of organochlorine contamination from residue ratios in the sea lamprey (Petromyzon marinus) and its lake whitefish (Coregonus clupeaformis) host. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):571–579, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-081.

ODor:1982:RMS

[605] R. K. O’Dor. Respiratory metabolism and swimming performance of the squid, Loligo opalescens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (4):580–587, April 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-082.

Hartman:1982:SMC

[606] G. F. Hartman, B. C. Andersen, and J. C. Scrivener. Seaward movement of Coho salmon (Oncorhynchus kisutch) fry in Carnation Creek, an unstable coastal stream in British Colombia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):588–597, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-083.

Paloheimo:1982:GMD

[607] J. E. Paloheimo, S. J. Crabtree, and W. D. Taylor. Growth model of Daphnia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):598–606, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-084.

Trautmann:1982:SDD

[608] Nancy Morton Trautmann, Charles E. McCulloch, and Ray T. Oglesby. Statistical determination of data requirements for assessment of lake restoration programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (4):607–610, April 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-085.

Cook:1982:SIS

[609] Rodney C. Cook. Stock identification of sockeye salmon (Oncorhynchus nerka) with scale pattern recognition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):611–617, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-086.

Harris:1982:PEE

[610] Graham P. Harris and Richard A. Vollenweider. Paleolimnological evidence of early eutrophication in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):618–626, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-087.

Sower:1982:HIO

[611] Stacia A. Sower, Carl B. Schreck, and Edward M. Donaldson. Hormoneinduced ovulation of Coho salmon (Oncorhynchus kisutch) held in seawater and fresh water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (4):627–632, April 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-088.

Shaul:1982:GPG

[612] Warren Shaul and Lynn Goodwin. Geodock (Panope generosa: Bivalvia) age as determined by internal growth lines in the shell. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):632–636, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-089.

Drummond-Davis:1982:SEP

[613] N. C. Drummond-Davis, K. H. Mann, and R. A. Pottle. Some estimates of population density and feeding habits of the rock crab, Cancer irroratus, in a kelp bed in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):636–639, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-090.

Hawkins:1982:AEC

[614] Christopher M. Hawkins and Paul D. Keizer. Ammonia excretion in Corophium volutator: Using an automated method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):640–643, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-091.

Huey:1982:MLC

[615] D. W. Huey and T. L. Beitinger. Methemoglobin levels in channel catfish, Ictalurus punctatus, exposed to nitrite and tricaine methanesulfonate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):643–645, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-092.

Anonymous:1982:BRCc

[616] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(4):646, April 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-093.

Macdonald:1982:RDD

[617] J. Stevenson Macdonald, Kenneth G. Waiwood, and Roger H. Green. Rates of digestion of different prey in Atlantic cod (Gadus morhua), ocean pout (Macrozoarces americanus), winter flounder (Pseudopleuronectes americanus), and American plaice (Hippoglossoides platessoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):651–659, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-094.

Gardner:1982:BHE

[618] Grant A. Gardner. Biological and hydrographical evidence for Pacific equatorial water on the Continental Shelf north of Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):660–667, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f82-
095.

Borgmann:1982:PSC

[619] Uwe Borgmann. Particle-size-conversion efficiency and total animal production in pelagic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):668–674, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-096.

Atema:1982:EDM

[620] Jelle Atema, Dale F. Leavitt, Diana E. Barshaw, and M. Carmela Cuomo. Effects of drilling muds on behavior of the American lobster, Homarus americanus, in water column and substrate exposures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):675–690, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-097.

Frank:1982:ERG

[621] Kenneth T. Frank and William C. Leggett. Environmental regulation of growth rate, efficiency, and swimming performance in larval capelin (Mallotus villosus), and its application to the Match/Mismatch Hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):691–699, May 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-098.

Jensen:1982:PUF

[622] A. L. Jensen, S. A. Spigarelli, and M. M. Thommes. PCB uptake by five species of fish in Lake Michigan, Green Bay of Lake Michigan, and Cayuga Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (5):700–709, May 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-099.

Arthur:1982:QSE

[623] J. R. Arthur, L. Margolis, D. J. Whitaker, and T. E. McDonald. A quantitative study of economically important parasites of walleye pollock (Theragra chalcogramma) from British Columbian waters and effects of postmortem handling on their abundance in the musculature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):710–726, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-100.

Griesbach:1982:AMP

[624] Sara Griesbach, Robert Henry Peters, and Sami Youakim. An allometric model for pesticide bioaccumulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):727–735, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-101.

Beamish:1982:LMN

[625] Richard J. Beamish. Lampetra macrostoma, a new species of freshwater parasitic lamprey from the West Coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):736–747, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-102.

Bloesch:1982:IOS

[626] J. Bloesch. Inshore–offshore sedimentation differences resulting from resuspension in the Eastern Basin of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):748–759, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-103.

Sameoto:1982:ZMA

[627] D. D. Sameoto. Zooplankton and micronekton abundance in acoustic scattering layers on the Nova Scotian slope. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):760–777, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-104.

Dawson:1982:RHP

[628] V. K. Dawson. A rapid high-performance liquid-chromatographic method for simultaneously determining the concentrations of TFM and Bayer 73 in water during lampricide treatments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):778–782, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-105.

Vaughan:1982:CAA

[629] Douglas S. Vaughan and Webster Van Winkle. Corrected analysis of the ability to detect reductions in year-class strength of the Hudson River white perch (Morone americana) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):782–785, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-106.

Peters:1982:CDP

[630] Robert Henry Peters and Martin Bergmann. A comparison of different phosphorus fractions as predictors of particulate pigment levels in Lake Memphremagog and its tributaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(5):785–790, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-107.

Cullen:1982:DCM

[631] John J. Cullen. The deep chlorophyll maximum: Comparing vertical profiles of chlorophyll a. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (5):791–803, May 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-108.

Anonymous:1982:BRCd et aquatiques, 39(5):804–805, May 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-109.

Cada:1982:RET

[633] Glenn F. Cada and James M. Loar. Relative effectiveness of two ichthyoplankton sampling techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):811–814, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-110.

McComas:1982:SRR

[634] Steven R. McComas and Ray W. Drenner. Species replacement in a reservoir fish community: Silverside feeding mechanics and competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):815–821, June 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-111.

Prepas:1982:IQP

[635] E. E. Prepas and F. H. Rigler. Improvements in quantifying the phosphorus concentration in lake water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):822–829, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-112.

Geiger:1982:HDG

[636] J. G. Geiger and A. L. Buikema, Jr. Hydrocarbons depress growth and reproduction of Daphnia pulex (Cladocera). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):830–836, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-113.

Quinn:1982:CTM

[637] Terrance J. Quinn II, Stephen H. Hoag, and G. Morris Southward. Comparison of two methods of combining catch-per-unit-effort data from geographic regions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):837–846, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f82-

Kent:1982:ILZ

[638] Clement Kent and Jonathan Wong. An index of littoral zone complexity and its measurement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (6):847–853, June 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-115.

Lied:1982:DPD

[639] E. Lied, K. Julshamn, and O. R. Braekkan. Determination of protein digestibility in Atlantic cod (Gadus morhua) with internal and external indicators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):854–861, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-116.

Garman:1982:PSB

[640] Greg C. Garman and Larry A. Nielsen. Piscivority by stocked brown trout (Salmo trutta) and its impact on the nongame fish community of Bottom Creek, Virginia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (6):862–869, June 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-117.

Stevenson:1982:BSC

[641] R. M. W. Stevenson and J. G. Daly. Biochemical and serological characteristics of Ontario isolates of Yersinia ruckeri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):870–876, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-118.

Malins:1982:ACS

[642] Donald C. Malins. Alterations in the cellular and subcellular structure of marine teleosts and invertebrates exposed to petroleum in the laboratory and field: a critical review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (6):877–889, June 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-119.

Havas:1982:AIS

[643] Magda Havas and T. C. Hutchinson. Aquatic invertebrates from the Smoking Hills, N.W.T.: Effect of pH and metals on mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):890–903, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-120.

Peterman:1982:NRB

[644] Randall M. Peterman. Nonlinear relation between smolts and adults in Babine Lake sockeye salmon (Oncorhynchus nerka) and implications for other salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (6):904–913, June 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-121.

Pearre:1982:EPP

[645] Sifford Pearre, Jr. Estimating prey preference by predators: Uses of various indices, and a proposal of another based on χ^{2}. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):914–923, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-122.

Addison:1982:VHE

[646] R. F. Addison and D. E. Willis. Variation of hepatic ethoxycoumarin O-de-ethylase activity with body weight and other factors in brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (6):924–926, June 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-123.

Wolf:1982:MCI

[647] Ken Wolf and Maria E. Markiw. Myxosoma cerebralis: Inactivation of spores by hot smoking of infected trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):926–928, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-124.

Pottle:1982:SPB

[648] Robert A. Pottle and Robert W. Elner. Substrate preference behavior of juvenile American lobsters, Homarus americanus, in gravel and silt-clay sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):928–932, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-125.

Anonymous:1982:BRCe

[649] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(6):933–934, June 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-126.

Campana:1982:DGI

[650] Steven E. Campana and John D. Neilson. Daily growth increments in otoliths of starry flounder (Platichthys stellatus) and the influence of some environmental variables in their production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):937–942, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-127.

Scott:1982:SBT

[651] J. S. Scott. Selection of bottom type by groundfishes of the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):943–947, July 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-128.

Durve:1982:VCD

[652] V. S. Durve and R. T. Lovell. Vitamin C and disease resistance in channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (7):948–951, July 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-129.

Healey:1982:TRI

[653] M. C. Healey. Timing and relative intensity of size-selective mortality of juvenile chum salmon (Oncorhynchus keta) during early sea life. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):952–957, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-130.

Morin:1982:LHV

[654] Roderick Morin, Julian J. Dodson, and Geoffrey Power. Life history variations of anadromous cisco (Coregonus artedii), lake whitefish (C. clupeaformis), and round whitefish (Prosopium cylindraceum) populations of Eastern James–Hudson Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):958–967, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-131.

Schlosser:1982:TSR

[655] Isaac J. Schlosser. Trophic structure, reproductive success, and growth rate of fishes in a natural and modified headwater stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):968–978, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-132.

Dagg:1982:ICG

[656] Michael J. Dagg and Jefferson T. Turner. The impact of copepod grazing on the phytoplankton of Georges Bank and the New York Bight. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):979–990, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-133.

Frank:1982:CWM

[657] Kenneth T. Frank and William C. Leggett. Coastal water mass replacement: Its effect on zooplankton dynamics and the predator–prey complex associated with larval capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):991–1003, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-134.

Moreau:1982:MLC

[658] Guy Moreau and Claude Barbeau. Les métaux lourds comme indicateurs d’origine géographique de l’anguille d’Amérique (Anguilla rostrata). (French) [Heavy metals as indicators of the geographical origin of the American eel (Anguilla rostrata)]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1004–1011, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-135.

Sullivan:1982:ESA

[659] Kathleen M. Sullivan and Kenneth L. Smith, Jr. Energetics of sablefish, Anoplopoma fimbria, under laboratory conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1012–1020, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-136.

Delorme:1982:LEO

[660] L. D. Delorme. Lake Erie oxygen; the prehistoric record. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1021–1029, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-137.

Hindar:1982:HFS

[661] Kjetil Hindar and Bror Jonsson. Habitat and food segregation of dwarf and normal Arctic charr (Salvelinus alpinus) from Vangsvatnet Lake, Western Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1030– 1045, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-138.

Crowder:1982:THS

[662] Larry B. Crowder and John J. Magnuson. Thermal habitat shifts by fishes at the thermocline in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1046–1050, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-139.

Wrona:1982:MSS

[663] F. J. Wrona, Joseph M. Culp, and Ronald W. Davies. Macroinvertebrate subsampling: a simplified apparatus and approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1051–1054, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-140.

Deriso:1982:RFM

[664] R. B. Deriso. Relationship of fishing mortality to natural mortality and growth at the level of maximum sustainable yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1054–1058, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-141.

Bakun:1982:OSR

[665] Andrew Bakun, Jan Beyer, Daniel Pauly, John G. Pope, and Gary D. Sharp. Ocean sciences in relation to living resources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1059–1070, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-142.

Anonymous:1982:LCLc

[666] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1071–1072, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-143.

Anonymous:1982:BRCf

[667] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(7):1073–1074, July 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f82-144.

Methot:1982:EPA

[668] Richard D. Methot, Jr. and Louis W. Botsford. Estimated preseason abundance in the California Dungeness crab (Cancer magister) fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1077–1083, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-145.

Whyte:1982:SVC

[669] J. N. C. Whyte and J. R. Englar. Seasonal variation in the chemical composition and condition indices of Pacific oyster, Crassostrea gigas, grown in trays or on the sea bed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (8):1084–1094, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-146.

Kaushik:1982:PNE

[670] Sadasivam J. Kaushik, Konrad Dabrowski, and Pierre Luquet. Patterns of nitrogen excretion and oxygen consumption during ontogenesis of common carp (Cyprinus carpio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (8):1095–1105, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-147.

Pace:1982:PCT

[671] Michael L. Pace. Planktonic ciliates: Their distribution, abundance, and relationship to microbial resources in a monomictic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1106–1116, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-148.

Harding:1982:ALA

[672] Gareth C. Harding, W. Peter Vass, and Ken F. Drinkwter. Aspects of larval American lobster (Homarus americanus) ecology in St. Georges Bay, Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8): 1117–1129, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-149.

Loder:1982:HEC

[673] John W. Loder, Daniel G. Wright, Christopher Garrett, and Barbara-Ann Juszko. Horizontal exchange on Central Georges Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1130–1137, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-150.

Hay:1982:FSH

[674] D. E. Hay. Fixation shrinkage of herring larvae: Effects of salinity, formalin concentration, and other factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1138–1143, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-151.

Mulcahy:1982:PIH

[675] Dan Mulcahy, John Burke, Ron Pascho, and C. K. Jenes. Pathogenesis of infectious hematopoietic necrosis virus in adult sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8): 1144–1149, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-152.

Fortier:1982:FTD

[676] L. Fortier and W. C. Leggett. Fickian transport and the dispersal of fish larvae in estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8): 1150–1163, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-153.

Courtois:1982:IMR

[677] R. Courtois, M. Simoneau, and J. J. Dodson. Interactions multispécifiques: répartition spatio-temporelle des larves de capelan (Mallotus villosus), d’éperlan (Osmerus mordax) et de hareng de l’Atlantique (Clupea harengus harengus) an sein de la communauté planctonique de l’estuaire moyen du Saint-Laurent. (French) [Multispecific interactions: spatio-temporal distribution of capelin larvae (Mallotus villosus), perlan (Osmerus mordax) and Atlantic herring (Clupea harengus harengus) in the plankton community of the middle estuary of the Saint Lawrence]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1164–1174, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-154.

Adams:1982:SPP

[678] S. Marshall Adams, R. B. McLean, and M. M. Huffman. Structuring of a predator population through temperature-mediated effects on prey availability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1175–1184, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f82-
155.

White:1982:RCT

[679] Alan W. White and Carrie M. Lewis. Resting cysts of the toxic, Red Tide dinoflagellate Gonyaulax excavate in Bay of Fundy sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1185–1194, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-156.

Fournier:1982:GTA

[680] David Fournier and Chris P. Archibald. A general theory for analyzing catch at age data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8): 1195–1207, August 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-157.

Chang:1982:SME

[681] William Y. B. Chang. A statistical method for evaluating the reproducibility of age determination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1208–1210, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-158.

Coote:1982:NDT

[682] A. R. Coote and E. P. Jones. Nutrient distributions and their relationships to water masses in Baffin Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1210–1214, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-159.

Anonymous:1982:LCLd

[683] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1215–1221, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-160.

Anonymous:1982:BRCg

[684] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1222–1223, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f82-161.

Anonymous:1982:CRPb

[685] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(8):1225– 1227, August 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f82-162.

Hilton:1982:MTL

[686] J. W. Hilton, J. L. Atkinson, and S. J. Slinger. Maximum tolerable level, digestion, and metabolism of D-glucose (cerelose) in rainbow trout (Salmo gairdneri) reared on a practical trout diet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1229–1234, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-163.

Hauer:1982:ERH

[687] F. Richard Hauer and Jack A. Stanford. Ecological responses of hydropsychid caddisflies to stream regulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1235–1242, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-164.

Hodson:1982:EGR

[688] Peter V. Hodson, D. George Dixon, Douglas J. Spry, D. M. Whittle, and John B. Sprague. Effect of growth rate and size of fish on rate of intoxication by waterfoorne lead. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9): 1243–1251, September 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-165.

Mendelssohn:1982:DFR

[689] Roy Mendelssohn. Discount factors and risk aversion in managing random fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9): 1252–1257, September 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-166.

Culp:1982:ALZ

[690] Joseph M. Culp and Ronald W. Davies. Analysis of longitudinal zonation and the river continuum concept in the Oldman–South Saskatchewan River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1258– 1266, September 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-167.

Jackson:1982:MPW

[691] Michael B. Jackson and David C. Lasenby. A method for predicting winter oxygen profiles in ice-covered Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1267–1272, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-168.

Hansen:1982:AST

[692] Stephen R. Hansen and Ronald R. Garton. Ability of standard toxicity tests to predict the effects of the insecticide diflubenzuron on laboratory stream communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9): 1273–1288, September 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-169.

Fraser:1982:ECB

[693] Grant A. Fraser and Harold H. Harvey. Elemental composition of bone from white sucker (Catostomus commersoni) in relation to lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1289–1296, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-170.

Ralston:1982:IHS

[694] Stephen Ralston. Influence of hook size in the Hawaiian deep-sea handline fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1297– 1302, September 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-171.

Peterson:1982:PCJ

[695] N. P. Peterson. Population characteristics of juvenile Coho salmon (Oncorhynchus kisutch) overwintering in riverine ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1303–1307, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-172.

Peterson:1982:IJC

[696] N. P. Peterson. Immigration of juvenile Coho salmon (Oncorhynchus kisutch) into riverine ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (9):1308–1310, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-173.

Scavia:1982:PRP

[697] Donald Scavia and Margaret J. McFarland. Phosphorus release patterns and the effects of reproductive stage and ecdysis in Daphnia magna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1310–1314, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-174.

Flaam:1982:CRN

[698] S. D. Flåm and S. Storøy. Capacity reduction on Norwegian industrial fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1314–1317, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-175.

Khan:1982:BPA

[699] R. A. Khan, M. Dawe, R. Bowering, and R. K. Misra. Blood protozoa as an aid for separating stocks of Greenland halibut, Reinhardtius hippoglossoides, in the Northwestern Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1317–1322, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-176.

Rigler:1982:RPA

[700] F. H. Rigler. Recognition of the possible: an advantage of empiricism in ecology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1323–1331, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-177.

Anonymous:1982:BRCh

[701] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(9):1332, September 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f82-178.

Wilson:1982:RBT

[702] Kenneth H. Wilson and P. A. Larkin. Relationship between thickness of daily growth increments in sagittae and change in body weight of sockeye salmon (Oncorhynchus nerka) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1335–1339, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-179.

Neilson:1982:OCS

[703] John D. Neilson and Glen H. Geen. Otoliths of Chinook salmon (Oncorhynchus tshawytscha): Daily growth increments and factors influencing their production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10): 1340–1347, October 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-180.

Incze:1982:CIB

[704] Lewis S. Incze, Lawrence M. Mayer, Evelyn B. Sherr, and Stephen A. Macko. Carbon inputs to bivalve mollusks: a comparison of two estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1348–1352, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-181.

Kwain:1982:SBE

[705] Wen hwa Kwain. Spawning behavior and early life history of pink salmon (Oncorhynchus gorbuscha) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1353–1360, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-182.

Schweigert:1982:MPP

[706] J. F. Schweigert and D. Fournier. A model for predicting Pacific herring (Clupea harengus pallasi) spawn density from diver observations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1361–1365, October 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-183.

Smagula:1982:TSS

[707] Carl M. Smagula and Ira R. Adelman. Temperature and scale size errors in the use of [^{14}C]Glycine uptake by scales as a growth index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1366–1372, October 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-184.

Godkin:1982:PSI

[708] C. M. Godkin, W. J. Christie, and D. E. McAllister. Problems of species identity in the Lake Ontario seulpins Cottus bairdi and C. cognatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1373–1382, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-185.

Newell:1982:EWO

[709] Roger I. E. Newell. An evaluation of the wet oxidation technique for use in determining the energy content of seston samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1383–1388, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-186.

Kovacs:1982:STC

[710] Tibor G. Kovacs and Gérard Leduc. Sublethal toxicity of cyanide to rainbow trout (Salmo gairdneri) at different temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1389–1395, October 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-187.

Minton:1982:MGC

[711] Joseph W. Minton and Richard B. McLean. Measurements of growth and consumption of sauger (Stizostedion canadense): Implication for fish energetics studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10): 1396–1403, October 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-188.

Jonsson:1982:RSD

[712] Bror Jonsson and Kjetil Hindar. Reproductive strategy of dwarf and normal Arctic charr (Salvelinus alpinus) from Vangsvatnet Lake, Western Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1404–1413, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f82-
189.

Hourston:1982:HCW

[713] A. S. Hourston. Homing by Canada’s West Coast herring to management units and divisions as indicated by tag recoveries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1414–1422, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-190.

Castonguay:1982:CMD

[714] Martin Castonguay and Gérard J. FitzGerald. Critique de la méthode de distinction entre poissons anadromes et dulcicoles de la même espèce par la teneur en strontium de leurs écailles. (French) [Critique of the method of distinguishing between anadromous and freshwater fish of the same species by the strontium content of their scales]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1423–1425, October 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-191.

Kovacs:1982:ATC

[715] Tibor G. Kovacs and Gérard Leduc. Acute toxicity of cyanide to rainbow trout (Salmo gairdneri) acclimated at different temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1426–1429, October 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-192.

Anonymous:1982:LCLe

[716] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(10):1430–1431, October 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-193.

Feller:1982:EEC

[717] Robert J. Feller. Empirical estimates of carbon production for a meiobenthic harpacticoid copepod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (11):1435–1443, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-194.

Peterman:1982:MSA

[718] Randall M. Peterman. Model of salmon age structure and its use in preseason forecasting and studies of marine survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1444–1452, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-195.

Dickson:1982:IHE

[719] T. A. Dickson and H. R. MacCrimmon. Influence of hatchery experience on growth and behavior of juvenile Atlantic salmon (Salmo salar) within allopatric and sympatric stream populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1453–1458, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-196.

Wright:1982:RMM

[720] David R. Wright and R. D. Hamilton. Release of methyl mercury from sediments: Effects of mercury concentration, low temperature, and nutrient addition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1459– 1466, November 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-197.

Kimura:1982:SRA

[721] Daniel K. Kimura and Jack V. Tagart. Stock reduction analysis, another solution to the catch equations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1467–1472, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-198.

Grimard:1982:TUN

[722] Y. Grimard and H. G. Jones. Trophic upsurge in new reservoirs: a model for total phosphorus concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1473–1483, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-199.

Richards:1982:DKA

[723] J. E. Richards, R. J. Beamish, and F. W. H. Beamish. Descriptions and keys for ammocoetes of lampreys from British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1484–1495, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-200.

Chadwick:1982:SRR

[724] E. M. P. Chadwick. Stock-recruitment relationship for Atlantic salmon (Salmo salar) in Newfoundland rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1496–1501, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-201.

Baltz:1982:CIB

[725] Donald M. Baltz, Peter B. Moyle, and Ned J. Knight. Competitive interactions between benthic stream fishes, riffle sculpin, Cottus gulosus, and speckled dace, Rhinichthys osculus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1502–1511, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-202.

Cipriano:1982:IPG

[726] Rocco C. Cipriano. Immunogenic potential of growth products extracted from cultures of Aeromonas salmonicida for brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1512–1518, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-203.

Morrison:1982:HPX

[727] C. M. Morrison, G. Shum, R. G. Appy, P. Odense, and C. Annand. Histology and prevalence of X-cell lesions in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1519–1530, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f82-204.

Heufelder:1982:EUL

[728] George R. Heufelder, David J. Jude, and Frank J. Tesar. Effects of upwelling on local abundance and distribution of larval alewife (Alosa pseudoharengus) in Eastern Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1531–1537, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-205.

Arai:1982:PMP

[729] Mary Needler Arai and Douglas E. Hay. Predation by medusae on Pacific herring (Clupea harengus pallasi) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1537–1540, November 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-206.

Favro:1982:MGM

[730] L. D. Favro, P. K. Kuo, J. F. McDonald, D. D. Favro, and A. D. Kuo. A multilocus genetic model applied to the effects of selective fishing on the growth rate of trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11): 1540–1543, November 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-207.

Anonymous:1982:LCLf

[731] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(11):1544–1548, November 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-208.

Anonymous:1982:AA

[732] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1553, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-209.

Heming:1982:ETI

[733] T. A. Heming, J. E. McInerney, and D. F. Alderdice. Effect of temperature on initial feeding in alevins of Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1554– 1562, December 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-210.

Hartig:1982:CPL

[734] John H. Hartig, David J. Jude, and Marlene S. Evans. Cyclopoid predation on Lake Michigan fish larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1563–1568, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-211.

Riessen:1982:PBP

[735] Howard P. Riessen. Predatory behavior and prey selectivity of the pelagic water mite Piona constricta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (12):1569–1579, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-212.

Silim:1982:STD

[736] A. Silim, M. A. S. Y. Elazhary, and A. Lagacé. Susceptibility of trouts of different species and origins to various isolates of infectious pancreatic necrosis virus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1580– 1584, December 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-213.

Condrey:1982:ILG

[737] Richard E. Condrey. Ingestion-limited growth of aquatic animals: The case for blackman kinetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39 (12):1585–1595, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f82-214.

Roch:1982:HMR

[738] M. Roch, J. A. McCarter, A. T. Matheson, M. J. R. Clark, and R. W. Olafson. Hepatic metallothionein in rainbow trout (Salmo gairdneri) as an indicator of metal pollution in the Campbell River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1596–1601, December 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-215.

Vincent:1982:FCP

[739] Warwick F. Vincent and Connie L. Vincent. Factors controlling phytoplankton production in Lake Vanda (77^{o}S). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1602–1609, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-216.

Kornfield:1982:SDA

[740] Irv Kornfield, Bruce D. Sidell, and P. S. Gagnon. Stock definition in Atlantic herring (Clupea harengus harengus): Genetic evidence for discrete fall and spring spawning populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1610–1621, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-217.

Beckett:1982:MCM

[741] David C. Beckett and Michael C. Miller. Macroinvertebrate colonization of multiplate samplers in the Ohio River: the effect of dams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1622–1627, December 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-218.

MacFarlane:1982:APC

[742] M. B. MacFarlane and T. F. Waters. Annual production by caddisflies and mayflies in a Western Minnesota Plains stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1628–1635, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-219.

Wurtsbaugh:1982:ECN

[743] Wayne A. Wurtsbaugh and Alexander J. Horne. Effects of copper on nitrogen fixation and growth of blue-green algae in natural plankton associations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1636–1641, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-220.

Harrison:1982:RIS

[744] Paul Garth Harrison and Richard E. Bigley. The recent introduction of the seagrass Zostera japonica Aschers. and Graebn. to the Pacific Coast of North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12): 1642–1648, December 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-221.

Janssen:1982:CSB

[745] John Janssen. Comparison of searching behavior for zooplankton in an obligate planktivore, blueback herring (Alosa aestivalis) and a facultative planktivore, bluegill (Lepomis machrochirus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1649–1654, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-222.

Cammen:1982:DAA

[746] Leon M. Cammen and Judith A. Walker. Distribution and activity of attached and free-living suspended bacteria in the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1655–1663, December 1982. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-223.

Ni:1982:MVB

[747] I-H. Ni. Meristic variation in beaked redfishes, Sebastes mentella and S. fasciatus, in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1664–1685, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-224.

Roff:1982:RSF

[748] Derek A. Roff. Reproductive strategies in flatfish: a first synthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1686–1698, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f82-225.

Naiman:1982:CSO

[749] Robert J. Naiman. Characteristics of sediment and organic carbon export from pristine boreal forest watersheds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1699–1718, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f82-226.

Sydor:1982:UBL

[750] W. J. Sydor, W. R. Miller III, J. Cairns, Jr., and D. Gruber. Use of box and line plots to assess fish ventilatory behavior in biological monitoring systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1719–1722, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f82-227.

Wolf:1982:III

[751] Ken Wolf and Maria E. Markiw. Ichthyophthiriasis: Immersion immunization of rainbow trout (Salmo gairdneri) using Tetrahymena thermophila as a protective immunogen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1722–1725, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f82-228.

Kent:1982:CIP

[752] Michael L. Kent. Characteristics and identification of Pasteurella and Vibrio species pathogenic to fishes using API-20E (Analytab Products) multitube test strips. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12): 1725–1729, December 1982. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f82-229.

Anonymous:1982:BRCi

[753] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 39(12):1730, December 1982. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f82-230.

Anonymous:1983:EEa

[754] Anonymous. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):1–2, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-001.

Rahel:1983:LPA

[755] Frank J. Rahel and John J. Magnuson. Low pH and the absence of fish species in naturally acidic Wisconsin lakes: Inferences for cultural acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):3–9, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-002.

Madenjian:1983:EBE

[756] C. P. Madenjian and D. J. Jude. Error bounds for estimates of entrained ichthyoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (1):10–16, January 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-003.

Ullrich:1983:SRF

[757] Sterling O. Ullrich, Jr. and Raymond E. Millemann. Survival, respiration, and food assimilation of Daphnia magna exposed to petroleum- and coal-derived oils at three temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):17–26, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-004.

Prepas:1983:EPC

[758] E. E. Prepas and D. O. Trew. Evaluation of the phosphorus–chlorophyll relationship for lakes off the Precambrian Shield in Western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):27–35, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-005.

Confer:1983:ZDB

[759] John L. Confer, Thomas Kaaret, and Gene E. Likens. Zooplankton diversity and biomass in recently acidified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):36–42, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-006.

Csirke:1983:PMU

[760] J. Csirke and J. F. Caddy. Production modeling using mortality estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):43–51, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-007.

Lebel:1983:PDS

[761] J. Lebel, E. Pelletier, M. Bergeron, N. Belzile, and G. Marquis. Le panache du Saguenay. (French) [The panache of Saguenay]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):52–60, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-008.

Crim:1983:MSR

[762] L. W. Crim, D. M. Evans, and B. H. Vickery. Manipulation of the seasonal reproductive cycle of the landlocked Atlantic salmon (Salmo salar) by LHRH analogues administered at various stages of gonadal development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):61–67, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-009.

Maude:1983:BCW

[763] Stephen H. Maude and D. Dudley Williams. Behavior of crayfish in water currents: Hydrodynamics of eight species with reference to their distribution patterns in Southern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):68–77, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-010.

Evans:1983:RBB

[764] R. D. Evans and D. C. Lasenby. Relationship between body-lead concentration of Mysis relicta and sediment-lead concentration in Kootenay Lake, B.C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):78–81, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
011.

Hilton:1983:EID

[765] J. W. Hilton, J. L. Atkinson, and S. J. Slinger. Effect of increased dietary fiber on the growth of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):81–85, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-012.

Ramamoorthy:1983:MSW

[766] S. Ramamoorthy, T. C. Cheng, and D. J. Kushner. Mercury speciation in water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):85–89, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-013.

Threader:1983:UNR

[767] R. W. Threader and A. H. Houston. Use of NaCl as a reference toxicant for goldfish, Carassius auratus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):89–92, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-014.

Prepas:1983:TDS

[768] E. E. Prepas. Total dissolved solids as a predictor of lake biomass and productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):92–95, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-015.

Sinclair:1983:CSA

[769] M. Sinclair, R. O’Boyle, and T. D. Iles. Consideration of the stable age distribution assumption in “Analytical” yield models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):95–101, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-016.

Bullock:1983:PST

[770] G. L. Bullock, G. Maestrone, C. Starliper, and B. Schill. Potentiated sulfonamide therapy of enteric redmouth disease. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):101–102, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-017.

Anonymous:1983:BRCa

[771] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(1):103–104, January 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f83-018.

Hay:1983:PSP

[772] D. E. Hay and J. Fulton. Potential secondary production from herring spawning in the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):109–113, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-019.

MacCrimmon:1983:MUL

[773] Hugh R. MacCrimmon, Christopher D. Wren, and Barra L. Gots. Mercury uptake by lake trout, Salvelinus namaycush, relative to age, growth, and diet in Tadenac Lake with comparative data from other PreCambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2): 114–120, February 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-020.

York:1983:AAF

[774] Anne E. York. Average age at first reproduction of the northern fur seal (Callorhinus ursinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (2):121–127, February 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-021.

Duncan:1983:TRC

[775] D. A. Duncan and J. F. Klaverkamp. Tolerance and resistance to cadmium in white suckers (Catostomus commersoni) previously exposed to cadmium, mercury, zinc, or selenium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):128–138, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-022.

Jovellanos:1983:PMJ

[776] Cesario L. Jovellanos and David E. Gaskin. Predicting the movements of juvenile Atlantic herring (Clupea harengus harengus) in the SW Bay of Fundy using computer simulation techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):139–146, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-023.

Lean:1983:CRM

[777] D. R. S. Lean and E. White. Chemical and radiotracer measurements of phosphorus uptake by lake plankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):147–155, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-024.

Planas:1983:LEA

[778] Dolores Planas and Alain Lamarche. Lack of effect of arsenic on phytoplankton communities in different nutrient conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):156–161, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-025.

Breck:1983:EFS

[779] James E. Breck and Michael J. Gitter. Effect of fish size on the reactive distance of bluegill (Lepomis macrochirus) sunfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):162–167, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-026.

Harding:1983:FIS

[780] Gareth C. Harding, Ken F. Drinkwater, and W. Peter Vass. Factors influencing the size of American lobster (Homarus americanus) stocks along the Atlantic Coast of Nova Scotia, Gulf of St. Lawrence, and Gulf of Maine: a new synthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (2):168–184, February 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-027.

Kapuscinski:1983:DCS

[781] Anne R. D. Kapuscinski and James E. Lannan. On density of chum salmon (Oncorhynchus keta) eggs in shallow matrix substrate incubators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):185–191, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-028.

Hoyer:1983:FAR

[782] Mark V. Hoyer and John R. Jones. Factors affecting the relation between phosphorus and chlorophyll a in midwestern reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):192–199, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-029.

Campana:1983:MSF

[783] Steven E. Campana. Mortality of starry flounders (Platichthys stellatus) with skin tumors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2): 200–207, February 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-030.

Priscu:1983:SDD

[784] John C. Priscu and Charles R. Goldman. Seasonal dynamics of the deep-chlorophyll maximum in Castle Lake, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):208–214, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-031.

Mulligan:1983:SSI

[785] T. J. Mulligan, L. Lapi, R. Kieser, S. B. Yamada, and D. L. Duewer. Salmon stock identification based on elemental composition of vertebrae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):215–229, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-032.

Kruse:1983:STU

[786] Gordon H. Kruse and Albert V. Tyler. Simulation of temperature and upwelling effects on the English sole (Parophrys vetulus) spawning season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):230–237, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-033.

Kruse:1983:RAS

[787] Gordon H. Kruse and Adriana Huyer. Relationships among shelf temperatures, coastal sea level, and the coastal upwelling index off Newport, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):238–242, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-034.

Mitchell:1983:ACG

[788] Stephen J. Mitchell and Joseph J. Cech, Jr. Ammonia-caused gill damage in channel catfish (Ictalurus punctatus): Confounding effects of residual chlorine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):242–247, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-035.

Anonymous:1983:LCLa

[789] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):248–251, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-036.

Anonymous:1983:BRCb

[790] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):253, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-037.

Anonymous:1983:CRPa

[791] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(2):255–256, February 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-038.

Gardner:1983:PRN

[792] Wayne S. Gardner, Thomas F. Nalepa, David R. Slavens, and Gwenyth A. Laird. Patterns and rates of nitrogen release by benthic chironomidae and oligochaeta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (3):259–266, March 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-039.

Brodie:1983:PSG

[793] Paul Brodie and Brian Beck. Predation by sharks on the grey seal (Halichoerus grypus) in Eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):267–271, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-040.

Majkowski:1983:EPU

[794] J. Majkowski and J. Hampton. The effect of parameter uncertainties in an age–length relationship upon estimating the age composition of catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):272–280, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-041.

Hardy:1983:CMR

[795] Ronald W. Hardy and Craig V. Sullivan. Canola meal in rainbow trout (Salmo gairdneri) production diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):281–286, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-042.

English:1983:PPR

[796] Karl K. English. Predator–prey relationships for juvenile Chinook salmon, Oncorhynchus tshawytscha, feeding on zooplankton in “in situ” enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):287–297, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-043.

Landrum:1983:ISA

[797] Peter F. Landrum and Donald Scavia. Influence of sediment on anthracene uptake, depuration, and biotransformation by the amphipod Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):298–305, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-044.

Niimi:1983:BTE

[798] A. J. Niimi. Biological and toxicologtcal effects of environmental contaminants in fish and their eggs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (3):306–312, March 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-045.

Jamieson:1983:PSP

[799] G. S. Jamieson and R. A. Chandler. Paralytic shellfish poison in sea scallops (Placopecten magellanicus) in the West Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):313–318, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-046.

Jasper:1983:PPL

[800] Steve Jasper, Eddy C. Carmack, Ralph J. Daley, Colin B. J. Gray, Chris H. Pharo, and Ron C. Wiegand. Primary productivity in a large, temperate lake with river interflow: Kootenay Lake, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):319–327, March 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-047.

Borgmann:1983:PSC

[801] Uwe Borgmann and D. M. Whittle. Particle-size-conversion efficiency and contaminant concentrations in Lake Ontario biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):328–336, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-048.

Botsford:1983:EDN

[802] Louis W. Botsford, Richard D. Methot, Jr., and Warren E. Johnston. Effort dynamics of the Northern California Dungeness crab (Cancer magister) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):337–346, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-049.

Wallace:1983:RMD

[803] Richard K. Wallace and John S. Ramsey. Reliability in measuring diet overlap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):347–351, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-050.

Westrheim:1983:NMA

[804] S. J. Westrheim. A new method for allotting effort to individual species in a mixed-species trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (3):352–360, March 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-051.

Mork:1983:IMF

[805] Jarle Mork, Per Solemdal, and Gunnar Sundnes. Identification of marine fish eggs: a biochemical genetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):361–369, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-052.

Leslie:1983:CWF

[806] J. K. Leslie. Cooling water flow velocity in relation to collection of entrained larval fish at power plants. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):370–373, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-053.

Burnison:1983:ICF

[807] B. Kent Burnison and Gary G. Leppard. Isolation of colloidal fibrils from lake water by physical separation techniques. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):373–381, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-054.

Anonymous:1983:LCLb

[808] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):382–384, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-055.

Anonymous:1983:BRCc

[809] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(3):385–389, March 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-056.

Anonymous:1983:AAa

[810] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):393–397, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-057.

Dill:1983:AFF

[811] Lawrence M. Dill. Adaptive flexibility in the foraging behavior of fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):398–408, April 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-058.

Winters:1983:ABD

[812] G. H. Winters. Analysis of the biological and demographic parameters of northern sand lance, Ammodytes dubius, from the Newfoundland Grand Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):409–419, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-059.

Miller:1983:NRD

[813] W. Roscoe Miller III, Albert C. Hendricks, and John Cairns, Jr. Normal ranges for diagnostically important hematological and blood chemistry characteristics of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):420–425, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-060.

Bax:1983:EMM

[814] Nicholas J. Bax. Early marine mortality of marked juvenile chum salmon (Oncorhynchus keta) released into Hood Canal, Puget Sound, Washington, in 1980. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):426–435, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
061.

McCuaig:1983:UGC

[815] Janet M. McCuaig and Roger H. Green. Unionid growth curves derived from annual rings: a baseline model for Long Point Bay, Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):436–442, April 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-062.

Gartner-Kepkay:1983:GDF

[816] K. E. Gartner-Kepkay, E. Zouros, L. M. Dickie, and K. R. Freeman. Genetic differentiation in the face of gene flow: a study of mussel populations from a single Nova Scotian embayment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):443–451, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-063.

Tschaplinski:1983:WDJ

[817] P. J. Tschaplinski and G. F. Hartman. Winter distribution of juvenile Coho salmon (Oncorhynchus kisutch) before and after logging in Carnation Creek, British Columbia, and some implications for overwinter survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):452–461, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-064.

Watt:1983:EAS

[818] W. D. Watt, C. D. Scott, and W. J. White. Evidence of acidification of some Nova Scotian rivers and its impact on Atlantic salmon, Salmo solar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):462–473, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-065.

Himmelman:1983:SUS

[819] John H. Himmelman, Yves Lavergne, Fritz Axelsen, André Cardinal, and Edwin Bourget. Sea urchins in the Saint Lawrence Estuary: Their abundance, size-structure, and suitability for commercial exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):474–486, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-066.

Dickie:1983:ECL

[820] L. M. Dickie, R. G. Dowd, and P. R. Boudreau. An echo counting and logging system (ECOLOG) for demersal fish size distributions and densities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):487–498, April 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-067.

Bams:1983:EGQ

[821] R. A. Bams. Early growth and quality of chum salmon (Oncorhynchus keta) produced in keeper channels and gravel incubators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):499–505, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-068.

Evans:1983:TLD

[822] R. Douglas Evans and Frank H. Rigler. A test of lead-210 dating for the measurement of whole lake soft sediment accumulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):506–515, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-069.

Clark:1983:IRB

[823] Jerry Clark and Bruce McCarl. An investigation of the relationship between Oregon Coho salmon (Oncorhynchus kisutch) hatchery releases and adult production utilizing law of the minimum regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):516–523, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-070.

Swartzman:1983:MAP

[824] Gordon L. Swartzman, Wayne M. Getz, Robert C. Francis, Robert T. Haar, and Kenneth Rose. A management analysis of the Pacific whiting (Merluccius productus) fishery using an age-structured stochastic recruitment model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):524–539, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
071.

Joule:1983:EMT

[825] Bruce J. Joule. An effective method for tagging marine polychaetes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):540–541, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-072.

Maki:1983:MDA

[826] James S. Maki, Michael E. Sierszen, and Charles C. Remsen. Measurements of dissolved adenosine triphosphate in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):542–547, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-073.

Anonymous:1983:LCLc

[827] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(4):548, April 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-074.

Ludwig:1983:APS

[828] D. Ludwig and R. Hilborn. Adaptive probing strategies for agestructured fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (5):559–569, May 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-075.

Evans:1983:AZC

[829] Hayla E. Evans, P. Jane Smith, and Peter J. Dillon. Anthropogenic zinc and cadmium burdens in sediments of selected Southern Ontario lakes.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):570–579, May 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-076.

Walker:1983:AMO

[830] K. D. Walker, R. B. Rettig, and R. Hilborn. Analysis of multiple objectives in Oregon Coho salmon policy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):580–587, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-077.

Schweigert:1983:OSD

[831] J. F. Schweigert and J. R. Sibert. Optimizing survey design for determining age structure of fish stocks: an example from British Columbia Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):588–597, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-078.

Tremblay:1983:CGS

[832] M. John Tremblay and John C. Roff. Community gradients in the Scotian Shelf zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (5):598–611, May 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-079.

Myers:1983:PNM

[833] R. A. Myers and R. W. Doyle. Predicting natural mortality rates and reproduction–mortality trade-offs from fish life history data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):612–620, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-080.

Yan:1983:ECP

[834] Norman D. Yan. Effects of changes in pH on transparency and thermal regimes of Lohi Lake, near Sudbury, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):621–626, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-081.

Vetter:1983:EMR

[835] Russell D. Vetter, Robert E. Hodson, and Connie Arnold. Energy metabolism in a rapidly developing marine fish egg, the red drum (Sciaenops ocellata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):627–634, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
082.

Macdonald:1983:RVU

[836] J. Stevenson Macdonald and R. H. Green. Redundancy of variables used to describe importance of prey species in fish diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):635–637, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-083.

Pugsley:1983:MFC

[837] C. W. Pugsley and H. B. N. Hynes. A modified freeze-core technique to quantify the depth distribution of fauna in stony streambeds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):637–643, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-084.

Black:1983:TSN

[838] Geoff A. Black. Taxonomy of a swimbladder nematode, Cystidicola stigmatura (Leidy), and evidence of its decline in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):643–647, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-085.

Dalley:1983:PMA

[839] E. L. Dalley, C. W. Andrews, and J. M. Green. Precocious male Atlantic salmon parr (Salmo salar) in insular Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):647–652, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-086.

Stewart:1983:MBM

[840] Patrick L. Stewart. Measurements of benthic macroinvertebrate standing crop from the Canadian Continental Shelf and Slope of Southern Davis Strait and Ungava Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (5):652–657, May 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-087.

Tarapchak:1983:SCD

[841] Stephen J. Tarapchak, David R. Slavens, Michael A. Quigley, and Judith S. Tarapchak. Silicon contamination in diatom nutrient enrichment experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):657–664, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
088.

Anonymous:1983:BRCd

[842] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(5):665–666, May 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-089.

Rimmer:1983:AHS

[843] D. M. Rimmer, U. Paim, and R. L. Saunders. Autumnal habitat shift of juvenile Atlantic salmon (Salmo salar) in a small river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):671–680, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-090.

Stewart:1983:EML

[844] Donald J. Stewart, David Weininger, Donald V. Rottiers, and Thomas A. Edsall. An energetics model for lake trout, Salvelinus namaycush: Application to the Lake Michigan population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):681–698, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-091.

DAbramo:1983:CSP

[845] Louis R. D’Abramo, Nancy A. Baum, Clark E. Bordner, and Douglas E. Conklin. Carotenoids as a source of pigmentation in juvenile lobsters fed a purified diet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):699–704, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
092.

Payne:1983:SEP

[846] Jerry F. Payne, J. Kiceniuk, R. Misra, Garth Fletcher, and R. Thompson. Sublethal effects of petroleum hydrocarbons on adult American lobsters (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (6):705–717, June 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-093.

Frechet:1983:UVB

[847] A. Fréchet, J. J. Dodson, and H. Powles. Use of variation in biological characters for the classification of anadromous rainbow smelt (Osmerus mordax) groups. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):718–727, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
094.

Bowen:1983:MRE

[848] W. D. Bowen and D. E. Sergeant. Mark–recapture estimates of harp seal pup (Phoca groenlandica) production in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):728–742, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-095.

Veith:1983:STR

[849] Gilman D. Veith, Daniel J. Call, and L. T. Brooke. Structure–toxicity relationships for the fathead minnow, Pimephales promelas: Narcotic industrial chemicals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (6):743–748, June 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-096.

Tremblay:1983:PES

[850] M. John Tremblay and John C. Roff. Production estimates for Scotian Shelf copepods based on mass specific P/B ratios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):749–753, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-097.

Frank:1983:MLF

[851] Kenneth T. Frank and William C. Leggett. Multispecies larval fish associations: Accident or adaptation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):754–762, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-098.

Mackie:1983:UPI

[852] George O. Mackie and Claudia E. Mills. Use of the Pisces IV submersible for zooplankton studies in coastal waters of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):763–776, June 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-099.

Peterson:1983:EMA

[853] R. H. Peterson and D. J. Martin-Robichaud. Embryo movements of Atlantic salmon (Salmo salar) as influenced by pH, temperature, and state of development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):777–782, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
100.

McKee:1983:RAF

[854] Paul M. McKee and Gerald L. Mackie. Respiratory adaptations of the fingernail clams Sphaerium occidentale and Musculium securis to ephemeral habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):783–791, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-101.

Konopka:1983:EMP

[855] Allan Konopka. Epilimnetic and metalimnetic primary production in an Indiana hardwater lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (6):792–798, June 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-102.

Galloway:1983:EAS

[856] James N. Galloway, Carl L. Schofield, Norman E. Peters, George R. Hendrey, and Elmar R. Altwicker. Effect of atmospheric sulfur on the composition of three Adirondack lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):799–806, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-103.

Gentner:1983:MPP

[857] S-R. Gentner. Measurement of primary productivity in dense algal cultures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):807–810, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-104.

Nalepa:1983:PRT

[858] Thomas F. Nalepa, Wayne S. Gardner, and John M. Malczyk. Phosphorus release by three kinds of benthic invertebrates: Effects of substrate and water medium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (6):810–813, June 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-105.

Gustafson-Marjanen:1983:SDP

[859] K. I. Gustafson-Marjanen and H. B. Dowse. Seasonal and diel patterns of emergence from the redd of Atlantic salmon (Salmo salar) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):813–817, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-106.

Chow-Fraser:1983:APL

[860] Patricia Chow-Fraser and Hamish C. Duthie. Assessment of phosphorus limitation in an oligotrophic lake using radiophosphorus uptake kinetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):817–821, June 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-107.

Boyce:1983:ESC

[861] N. P. Boyce and W. Craig Clarke. Eubothrium salvelini (Cestoda: Pseudophyllidea) impairs seawater adaptation of migrant sockeye salmon yearlings (Oncorhynchus nerka) from Babine Lake, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):821–824, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-108.

Rodgers:1983:WQM

[862] D. W. Rodgers and F. W. H. Beamish. Water quality modifies uptake of waterborne methylmercury by rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):824–828, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-109.

Anonymous:1983:BRCe

[863] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(6):829–830, June 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-110.

Schnute:1983:STF

[864] Jon Schnute and John Sibert. The salmon terminal fishery: a practical, comprehensive timing model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (7):835–853, July 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-111.

Foureman:1983:BEW

[865] Gary L. Foureman, Nathaniel B. White, Jr., and John R. Bend. Biochemical evidence that winter flounder (Pseudopleuronectes americanus) have induced hepatic cytochrome P-450-dependent monooxygenase activities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):854–865, July 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-112.

Fowler:1983:LCS

[866] G. M. Fowler and S. J. Smith. Length changes in silver hake (Merluccius bilinearis) larvae: Effects of formalin, ethanol, and freezing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):866–870, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-113.

Perry:1983:TMS

[867] R. Ian Perry, Bruce R. Dilke, and Timothy R. Parsons. Tidal mixing and summer plankton distributions in Hecate Strait, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):871–887, July 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-114.

Hanych:1983:NIM

[868] David A. Hanych, Michael R. Ross, Robert E. Magnien, and Amy L. Suggars. Nocturnal inshore movement of the mimic shiner (Notropis volucellus): a possible predator avoidance behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):888–894, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-115.

Bates:1983:VZA

[869] Stephen S. Bates, Martin Létourneau, André Tessier, and Peter G. C. Campbell. Variation in zinc adsorption and transport during growth of Chlamydomonas variabilis (Chlorophyceae) in batch culture with daily addition of zinc. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):895–904, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
116.

Munro:1983:MSD

[870] Jean Munro and Jean-Claude Therriault. Migrations saisonnières du homard (Homarus americanus) entre la côte et les lagunes des Îles-de-la-Madeleine. (French) [Seasonal migrations of the lobster (Homarus americanus) between the coast and the lagoons of the Madeleine Islands]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):905–918, July 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-117.

Roff:1983:PDM

[871] Derek A. Roff and W. Don Bowen. Population dynamics and management of the Northwest Atlantic harp seal (Phoca groenlandica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):919–932, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-118.

Besner:1983:MSM

[872] M. Besner and L. S. Smith. Modification of swimming mode and stamina in two stocks of Coho salmon (Oncorhynchus kisutch) by differing levels of long-term continuous exercise. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):933–939, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-119.

Nelson:1983:PEC

[873] Keith Nelson, Dennis Hedgecock, and Will Borgeson. Photoperiodic and ecdysial control of vitellogenesis in lobsters (Homarus) (Decapoda, Nephropidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):940–947, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
120.

Mallet:1983:GRS

[874] Andre L. Mallet and L. E. Haley. Growth rate and survival in pure population matings and crosses of the oyster Crassostrea virginica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):948–954, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-121.

McClatchie:1983:MSS

[875] Sam McClatchie and Carl M. Boyd. Morphological study of sieve efficiencies and mandibular surfaces in the Antarctic krill, Euphausia superba.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):955–967, July 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-122.

Hannesson:1983:BPF

[876] Rögnvaldur Hannesson. Bioeconomic production function in fisheries: Theoretical and empirical analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):968–982, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-123.

Gophen:1983:CSD

[877] Moshe Gophen, Ray W. Drenner, and Gary L. Vinyard. Cichlid stocking and the decline of the Galilee Saint Peter’s fish (Sarotherodon galilaeus) in Lake Kinneret, Israel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (7):983–986, July 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-124.

Johnson:1983:HCS

[878] Lionel Johnson. Homeostatic characteristics of single species fish stocks in Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):987–1024, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
125.

Parsons:1983:SFO

[879] T. R. Parsons. Symposium on the Fisheries and Oceanography of the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1025, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-126.

Luternauer:1983:SSG

[880] J. L. Luternauer, J. J. Clague, and C. H. Pharo. Substrates of the Strait of Georgia, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (7):1026–1032, July 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-127.

LeBlond:1983:SGF

[881] Paul H. LeBlond. The Strait of Georgia: functional anatomy of a coastal sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1033–1063, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-128.

Harrison:1983:RBO

[882] P. J. Harrison, J. D. Fulton, F. J. R. Taylor, and T. R. Parsons. Review of the biological oceanography of the Strait of Georgia: Pelagic environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1064–1094, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-129.

Ketchen:1983:HPS

[883] K. S. Ketchen, N. Bourne, and T. H. Butler. History and present status of fisheries for marine fishes and invertebrates in the Strait of Georgia, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1095– 1119, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-130.

Levings:1983:RBS

[884] C. D. Levings, R. E. Foreman, and V. J. Tunnicliffe. Review of the benthos of the Strait of Georgia and contiguous fjords. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1120–1141, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-131.

Waldichuk:1983:PSG

[885] Michael Waldichuk. Pollution in the Strait of Georgia: a review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(7):1142–1167, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-132.

Anonymous:1983:BRCf et aquatiques, 40(7):1168, July 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-133.

Hawkins:1983:DFS

[887] Charles P. Hawkins, Michael L. Murphy, N. H. Anderson, and Margaret A. Wilzbach. Density of fish and salamanders in relation to riparian canopy and physical habitat in streams of the Northwestern United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1173–1185, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
134.

Hudon:1983:ISF

[888] Christiane Hudon, Edwin Bourget, and Pierre Legendre. An integrated study of the factors influencing the choice of the settling site of Balanus crenatus cyprid larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8): 1186–1194, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-135.

Mccoy:1983:NLT

[889] George A. Mccoy. Nutrient limitation in two Arctic lakes, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1195–1202, August 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-136.

Saunders:1983:ELE

[890] Richard L. Saunders, Eugene B. Henderson, Paul R. Harmon, C. Edward Johnston, and J. Geoffrey Eales. Effects of low environmental pH on smolting of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1203–1211, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-137.

Peterman:1983:EMO

[891] Randall M. Peterman and Richard D. Routledge. Experimental management of Oregon Coho salmon (Oncorhynchus kisutch): Designing for yield of information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):

1212–1223, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-138.

Forstner:1983:TBD

[892] H. Forstner, S. Hinterleitner, K. Mähr, and W. Wieser. Towards a better definition of “Metamorphosis” in Coregonus sp.: Biochemical, histological, and physiological data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (8):1224–1232, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-139.

Fournier:1983:AHS

[893] D. A. Fournier. An analysis of the Hecate Strait Pacific cod fishery using an age-structured model incorporating density-dependent effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1233–1243, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-140.

Black:1983:ODP

[894] Geoff A. Black. Origin, distribution, and postglacial dispersal of a swimbladder nematode, Cystidicola stigmatura. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1244–1253, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-141.

Baker:1983:MAF

[895] M. D. Baker, P. T. S. Wong, Y. K. Chau, C. I. Mayfield, and W. E. Inniss. Methylation of arsenic by freshwater green algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1254–1257, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-142.

Carlson:1983:RIR

[896] David J. Carlson and Lawrence M. Mayer. Relative influences of riverine and macroalgal phenolic materials on UV absorbance in temperate coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8): 1258–1263, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-143.

Halliburton:1983:RSA

[897] Richard Halliburton, R. E. Pipkin, and G. A. E. Gall. Reproductive success of artificially hybridized golden trout (Salmo aguabonita) and rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (8):1264–1269, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-144.

Fish:1983:COC

[898] William Fish and François M. M. Morel. Characterization of organic copper-complexing agents released by Daphnia magna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1270–1277, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-145.

Archambault:1983:IDR

[899] Diane Archambault and Edwin Bourget. Importance du régime de dénudation sur la structure et la succession des communautés intertidales de substrat rocheux en milieu subarctique. (French) [Importance of the denudation regime on the structure and succession of intertidal bedrock communities in the subarctic environment]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1278–1292, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-146.

Andrews:1983:DAS

[900] John C. Andrews. Deformation of the active space in the low Reynolds number feeding current of calanoid copepods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1293–1302, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-147.

Carter:1983:FGG

[901] J. C. H. Carter, W. G. Sprules, M. J. Dadswell, and J. C. Roff. Factors governing geographical variation in body size of Diaptomus minutus (Copepoda, Calanoida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8): 1303–1307, August 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-148.

Thayer:1983:DRC

[902] P. E. Thayer, J. W. Hurst, C. M. Lewis, R. Selvin, and C. M. Yentsch. Distribution of resting cysts of Gonyaulax tamarensis var. excavata and shellfish toxicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1308– 1314, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f83-149.

Herbes:1983:LQF

[903] Stephen E. Herbes and Carol P. Allen. Lipid quantification of freshwater invertebrates: Method modification for microquantitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1315–1317, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-150.

Mundie:1983:MCS

[904] J. H. Mundie and R. E. Traber. Movements of Coho salmon (Oncorhynchus kisutch) fingerlings in a stream following marking with a vital stain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1318–1319, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-151.

Mundie:1983:CCE

[905] J. H. Mundie and R. E. Traber. Carrying capacity of an enhanced side-channel for rearing salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1320–1322, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-152.

Anonymous:1983:LCLd

[906] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1323–1324, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-153.

Anonymous:1983:BRCg

[907] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1325, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-154.

Anonymous:1983:CRPb

[908] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(8):1327, August 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
155.

Gordon:1983:DEE

[909] Donald C. Gordon, Jr. and Con Desplanque. Dynamics and environmental effects of ice in the Cumberland Basin of the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1331–1342, September 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-156.

Ashley:1983:HAN

[910] Kenneth Ian Ashley. Hypolimnetic aeration of a naturally eutrophic lake: Physical and chemical effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1343–1359, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-157.

Green:1983:AIP

[911] Roger H. Green, Shiva M. Singh, Beverly Hicks, and Janet M. McCuaig. An Arctic intertidal population of Macoma balthica (Mollusca, Pelecypoda): Genotypic and phenotypic components of population structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1360–1371, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-158.

Nordeng:1983:SCP

[912] Hans Nordeng. Solution to the “Char problem” based on Arctic char (Salvelinus alpinus) in Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1372–1387, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-159.

Niimi:1983:BHL

[913] A. J. Niimi and B. G. Oliver. Biological half-lives of polychlorinated biphenyl (PCB) congeners in whole fish and muscle of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1388– 1394, September 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-160.

Roff:1983:AMG

[914] Derek A. Roff. An allocation model of growth and reproduction in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1395–1404, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-161.

Kirkwood:1983:EBG

[915] G. P. Kirkwood. Estimation of von Bertalanffy growth curve parameters using both length increment and age–length data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1405–1411, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-162.

Prepas:1983:OTT

[916] E. E. Prepas. Orthophosphate turnover time in shallow productive lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1412–1418, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-163.

Wurtsbaugh:1983:IEC

[917] Wayne A. Wurtsbaugh and A. J. Horne. Iron in eutrophic Clear Lake, California: Its importance for algal nitrogen fixation and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1419–1429, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-164.

Bowen:1983:VAE

[918] W. D. Bowen, D. E. Sergeant, and T. Øritsland. Validation of age estimation in the harp seal, Phoca groenlandica, using dentinal annuli. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1430–1441, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-165.

Frank:1983:SVO

[919] Kenneth T. Frank and William C. Leggett. Survival value of an opportunistic life-stage transition in capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1442–1448, September 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-166.

Pollock:1983:UAD

[920] K. H. Pollock and R. H. K. Mann. Use of an age-dependent mark– recapture model in fisheries research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1449–1455, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-167.

Scarnecchia:1983:ASM

[921] Dennis L. Scarnecchia. Age at sexual maturity in Icelandic stocks of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (9):1456–1468, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-168.

Marshall:1983:BEC

[922] J. S. Marshall, J. I. Parker, D. L. Mellinger, and C. Lei. Bioaccumulation and effects of cadmium and zinc in a Lake Michigan plankton community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1469–1479, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-169.

Adams:1983:SOM

[923] S. M. Adams, B. L. Kimmel, and G. R. Ploskey. Sources of organic matter for reservoir fish production: a trophic-dynamics analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1480–1495, September 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-170.

Roff:1983:ACE

[924] Derek A. Roff. Analysis of catch/effort data: a comparison of three methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1496–1506, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-171.

Misra:1983:DBR

[925] R. K. Misra and I-H. Ni. Distinguishing beaked redfishes (deepwater redfish, Sebastes mentella and Labrador redfish, S. fasciatus) by discriminant analysis (with covariance) and multivariate analysis of covariance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1507–1511, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-172.

Mulcahy:1983:MDI

[926] Dan Mulcahy, Ron Pascho, and C. K. Jenes. Mortality due to infectious hematopoietic necrosis of sockeye salmon (Oncorhynchus nerka) fry in streamside egg incubation boxes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1511–1516, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-173.

Geraci:1983:BDT

[927] J. R. Geraci, D. J. St. Aubin, and R. J. Reisman. Bottlenose dolphins, Tursiops truncatus, can detect oil. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1516–1522, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-174.

Smith:1983:RBD

[928] T. G. Smith, J. R. Geraci, and D. J. St. Aubin. Reaction of bottlenose dolphins, Tursiops truncatus, to a controlled oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1522–1525, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-175.

Barica:1983:FSC

[929] J. Barica, J. Gibson, and W. Howard. Feasibility of snow clearing to improve dissolved oxygen conditions in a winterkill lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1526–1531, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-176.

Anonymous:1983:LCLe

[930] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1532–1538, September 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-177.

Anonymous:1983:BRCh

[931] Anonymous. Book review/compte rendu. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(9):1539, September 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-178.

Fortier:1983:VMT

[932] L. Fortier and W. C. Leggett. Vertical migrations and transport of larval fish in a partially mixed estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1543–1555, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-179.

Henderson:1983:DEP

[933] B. A. Henderson, J. J. Collins, and J. A. Reckahn. Dynamics of an exploited population of lake whitefish (Coregonus clupeaformis) in Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1556–1567, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
180.

Culp:1983:RIS

[934] Joseph M. Culp, Sandra J. Walde, and Ronald W. Davies. Relative importance of substrate particle size and detritus to stream benthic macroinvertebrate microdistribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1568–1574, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-181.

Rombough:1983:ELP

[935] Peter J. Rombough. Effects of low pH on eyed embryos and alevins of Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1575–1582, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-182.

Weber:1983:EHS

[936] Jean-Michel Weber and Donald L. Kramer. Effects of hypoxia and surface access on growth, mortality, and behavior of juvenile guppies, Poecilia reticulata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1583–1588, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-183.

Cross:1983:PIR

[937] Patricia M. Cross and Frank H. Rigler. Phosphorus and iron retention in sediments measured by mass budget calculations and directly. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1589–1597, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-184.

Quinn:1983:OCS

[938] T. P. Quinn and C. Groot. Orientation of chum salmon (Oncorhynchus keta) after internal and external magnetic field alteration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1598–1606, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-185.

Watras:1983:RCD

[939] Carl J. Watras. Reproductive cycles in diaptomid copepods: Effects of temperature, photocycle, and species on reproductive potential. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1607–1613, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-186.

Burnison:1983:CPE

[940] B. Kent Burnison. “Colloidal” phosphorus errors in gel chromatography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1614–1621, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-187.

Ratliff:1983:CSL

[941] Donald E. Ratliff. Ceratomyxa shasta: Longevity, distribution, timing, and abundance of the infective stage in Central Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1622–1632, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-188.

Weinstein:1983:PDE

[942] Michael P. Weinstein. Population dynamics of an estuarine-dependent fish, the spot (Leiostomus xanthurus), along a tidal creek–seagrass meadow coenocline. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1633–1638, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-189.

McMurtry:1983:SST

[943] M. J. McMurtry, D. J. Rapport, and K. E. Chua. Substrate selection by tubificid oligochaetes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1639–1646, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-190.

Stout:1983:EPC

[944] R. Jean Stout and William E. Cooper. Effect of P-Cresol on leaf decomposition and invertebrate colonization in experimental outdoor streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1647–1657, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-191.

LaZerte:1983:SCI

[945] Bruce D. LaZerte. Stable carbon isotope ratios: Implications for the source of sediment carbon and for phytoplankton carbon assimilation in Lake Memphremagog, Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1658–1666, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-192.

Campbell:1983:GTA

[946] Alan Campbell. Growth of tagged American lobsters, Homarus americanus, in the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (10):1667–1675, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-193.

Arai:1983:PMP

[947] Hisao P. Arai and Dwight R. Mudry. Protozoan and metazoan parasites of fishes from the headwaters of the Parsnip and McGregor Rivers, British Columbia: a study of possible parasite transfaunations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1676–1684, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-194.

Moreau:1983:ZMS

[948] G. Moreau, C. Barbeau, J. J. Frenette, J. Saint-Onge, and M. Simoneau. Zinc, manganese, and strontium in opercula and scales of brook trout (Salvelinus fontinalis) as indicators of lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1685–1691, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-195.

Sutcliffe:1983:NFL

[949] W. H. Sutcliffe, Jr., R. H. Loucks, K. F. Drinkwater, and A. R. Coote. Nutrient flex onto the Labrador Shelf from Hudson Strait and its biological consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1692–1701, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-196.

Mundie:1983:RSZ

[950] J. H. Mundie, S. M. McKinnell, and R. E. Traber. Responses of stream zoobenthos to enrichment of gravel substrates with cereal grain and soybean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1702–1712, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
197.

Canfield:1983:TSC

[951] Daniel E. Canfield, Jr., Kenneth A. Langeland, Michael J. Maceina, William T. Haller, Jerome V. Shireman, and John R. Jones. Trophic state classification of lakes with aquatic macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1713–1718, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-198.

Crecco:1983:DMR

[952] V. Crecco, T. Savoy, and L. Gunn. Daily mortality rates of larval and juvenile American shad (Alosa sapidissima) in the Connecticut River with changes in year-class strength. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1719–1728, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-199.

Newman:1983:EFP

[953] Raymond M. Newman and Frank B. Martin. Estimation of fish production rates and associated variances. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1729–1736, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-200.

Wren:1983:MLS

[954] Christopher D. Wren and Hugh R. MacCrimmon. Mercury levels in the sunfish, Lepomis gibbosus, relative to pH and other environmental variables of Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1737–1744, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-201.

Irvine:1983:SYR

[955] J. R. Irvine and T. G. Northcote. Selection by young rainbow trout (Salmo gairdneri) in simulated stream environments for live and dead prey of different sizes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1745–1749, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-202.

Collins:1983:ILS

[956] Nicholas C. Collins and W. G. Sprules. Introduction to large-scale comparative studies of lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1750–1751, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-203.

Gates:1983:PBT

[957] Michael A. Gates, Ann P. Zimmerman, W. Gary Sprules, and Roy Knoechel. Planktonic biomass trajectories in lake ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1752–1760, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-204.

Sprules:1983:SDP

[958] W. Gary Sprules, J. M. Casselman, and B. J. Shuter. Size distribution of pelagic particles in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):

1761–1769, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-205.

Peters:1983:SSP

[959] Robert Henry Peters. Size structure of the plankton community along the trophic gradient of Lake Memphremagog. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1770–1778, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-206.

Lewis:1983:RCL

[960] William M. Lewis, Jr. A revised classification of lakes based on mixing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1779–1787, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-207.

Zimmerman:1983:PTS

[961] A. P. Zimmerman, K. M. Noble, M. A. Gates, and J. E. Paloheimo. Physicochemical typologies of South-Central Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1788–1803, October 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-208.

Paloheimo:1983:FIP

[962] J. E. Paloheimo and A. P. Zimmerman. Factors influencing phosphorus– phytoplankton relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (10):1804–1812, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-209.

Bays:1983:ZTS

[963] J. S. Bays and T. L. Crisman. Zooplankton and trophic state relationships in Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1813–1819, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-210.

Collins:1983:IPF

[964] Nicholas C. Collins, Harold H. Harvey, Ann Jane Tierney, and David W. Dunham. Influence of predatory fish density on trapability of crayfish in Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10): 1820–1828, October 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-211.

Schlesinger:1983:RBE

[965] David A. Schlesinger and Henry A. Regier. Relationship between environmental temperature and yields of subarctic and temperate zone fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1829–1837, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f83-
212.

Shuter:1983:EPA

[966] B. J. Shuter, D. A. Schlesinger, and A. P. Zimmerman. Empirical predictors of annual surface water temperature cycles in North American lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1838–1845, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-213.

Anonymous:1983:BRCi

[967] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(10):1846–1849, October 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f83-214.

Brownlee:1983:NFP

[968] B. G. Brownlee and T. P. Murphy. Nitrogen fixation and phosphorus turnover in a hypertrophic prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1853–1860, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-215.

McGlade:1983:EIS

[969] J. M. McGlade, M. C. Annand, and T. J. Kenchington. Electrophoretic identification of Sebastes and Helicolenus in the Northwestern Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1861–1870, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-216.

Collie:1983:EPS

[970] Jeremy S. Collie and Michael P. Sissenwine. Estimating population size from relative abundance data measured with error. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1871–1879, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-217.

McCarl:1983:IHS

[971] B. A. McCarl and R. B. Rettig. Influence of hatchery smolt releases on adult salmon production and its variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1880–1886, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-218.

Shortreed:1983:PBS

[972] Ken S. Shortreed and John G. Stockner. Periphyton biomass and species composition in a coastal rainforest stream in British Columbia: Effects of environmental changes caused by logging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1887–1895, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-219.

Eilers:1983:HCL

[973] J. M. Eilers, G. E. Glass, K. E. Webster, and J. A. Rogalla. Hydrologic control of lake susceptibility to acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1896–1904, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-220.

Nero:1983:DMR

[974] R. W. Nero and D. W. Schindler. Decline of Mysis relicta during the acidification of Lake 223. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11): 1905–1911, November 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-221.

Koslow:1983:ZCS

[975] J. Anthony Koslow. Zooplankton community structure in the North Sea and Northeast Atlantic: Development and test of a biological model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1912–1924, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-222.

Francis:1983:PTD

[976] Robert C. Francis. Population and trophic dynamics of Pacific hake (Merluccius productus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11): 1925–1943, November 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-223.

Pratt:1983:AGS

[977] Harold L. Pratt, Jr. and John G. Casey. Age and growth of the shortfin mako, Isurus oxyrinchus, using four methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1944–1957, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-224.

Campbell:1983:RPT

[978] A. Campbell and D. G. Robinson. Reproductive potential of three American lobster (Homarus americanus) stocks in the Canadian maritimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1958–1967, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-225.

Murtaugh:1983:MLH

[979] Paul A. Murtaugh. Mysid life history and seasonal variation in predation pressure on zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11): 1968–1974, November 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-226.

Bernstein:1983:SUS

[980] B. B. Bernstein, S. C. Schroeter, and K. H. Mann. Sea urchin (Strongylocentrotus droebachiensis) aggregating behavior investigated by a subtidal multifactorial experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (11):1975–1986, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-227.

Hilton:1983:CAP

[981] J. W. Hilton, P. V. Hodson, H. E. Braun, J. L. Leatherland, and S. J. Slinger. Contaminant accumulation and physiological response in rainbow trout (Salmo gairdneri) reared on naturally contaminated diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1987–1994, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-228.

Lamberti:1983:GES

[982] Gary A. Lamberti and Vincent H. Resh. Geothermal effects on stream benthos: Separate influences of thermal and chemical components on periphyton and macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):1995–2009, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-229.

Borgmann:1983:ESG

[983] Uwe Borgmann. Effect of somatic growth and reproduction on biomass transfer up pelagic food webs as calculated from particle-size-conversion efficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2010–2018, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-230.

Beacham:1983:SDA

[984] Terry D. Beacham and Clyde B. Murray. Sexual dimorphism in the adipose fin of Pacific salmon (Oncorhynchus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2019–2024, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-231.

Ciborowski:1983:DLT

[985] Jan J. H. Ciborowski. Downstream and lateral transport of nymphs of two mayfly species (Ephemeroptera). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2025–2029, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-232.

Garman:1983:USF

[986] Greg C. Garman and Thomas F. Waters. Use of the size–frequency (Hynes) method to estimate annual production of a stream fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2030–2034, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-233.

Black:1983:CFN

[987] Geoff A. Black. Cystidicola farionis (Nematoda) as an indicator of lake trout (Salvelinus namaycush) of Bering ancestry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2034–2040, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-234.

Scheerer:1983:ISS

[988] Paul D. Scheerer and Gary H. Thorgaard. Increased survival in salmonid hybrids by induced triploidy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (11):2040–2044, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-235.

Balon:1983:EMR

[989] Eugene K. Balon. Epigenetic mechanisms: Reflections on evolutionary processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2045–2058, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-236.

Anonymous:1983:BRCj

[990] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(11):2059–2060, November 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-237.

Anonymous:1983:AAb

[991] Anonymous. Announcements/avis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2065, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-238.

Anonymous:1983:EEb

[992] Anonymous. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2066–2068, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-239.

Charles:1983:OFIa

[993] Anthony T. Charles. Optimal fisheries investment: Comparative dynamics for a deterministic seasonal fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2069–2079, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-240.

Charles:1983:OFIb

[994] Anthony T. Charles. Optimal fisheries investment under uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2080–2091, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-241.

Kjosbakken:1983:BDA

[995] J. Kjosbakken, I. Storrø, and H. Larsen. Bacteria decomposing amino acids in bulk-stored capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2092–2097, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-242.

Bams:1983:IDW

[996] R. A. Bams and C. N. H. Lam. Influence of deteriorating water quality on growth and development of chum salmon (Oncorhynchus keta) larvae in a Japanese-style keeper channel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2098–2104, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-243.

Limsuwan:1983:SRB

[997] Chalor Limsuwan, Tasanee Limsuwan, John M. Grizzle, and John A. Plumb. Stress response and blood characteristics of channel catfish (Ictalurus punctatus) after anesthesia with etomidate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2105–2112, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-244.

Cook:1983:SAS

[998] Rodney C. Cook. Simulation and application of stock composition estimators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2113–2118, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-245.

Harder:1983:AET

[999] Howard W. Harder, Taryn V. Carter, and Terry F. Bidleman. Acute effects of toxaphene and its sediment-degraded products on estuarine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2119–2125, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-246.

Mason:1983:SMF

[1000] J. C. Mason, R. J. Beamish, and G. A. McFarlane. Sexual maturity, fecundity, spawning, and early life history of sablefish (Anoplopoma fimbria) off the Pacific Coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2126–2134, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-247.

Robinson:1983:RRT

[1001] M. G. Robinson and L. N. Brown. A recurrent Red Tide in a British Columbia coastal lagoon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12): 2135–2143, December 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-248.

Mathur:1983:SAT

[1002] Dilip Mathur, Robert M. Schutsky, Edmund J. Purdy, Jr., and Carl A. Silver. Similarities in avoidance temperatures of freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2144–2152, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-249.

Schnute:1983:NAE

[1003] Jon Schnute. A new approach to estimating populations by the removal method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2153–2169, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-250.

Cohen:1983:SAS

[1004] Joel E. Cohen, Sigurd W. Christensen, and C. Phillip Goodyear. A stochastic age-structured population model of striped bass (Morone saxatilis) in the Potomac River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40 (12):2170–2183, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f83-251.

Cobb:1983:BDL

[1005] J. S. Cobb, T. Gulbransen, B. F. Phillips, D. Wang, and M. Syslo. Behavior and distribution of larval and early jevenile Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2184–2188, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-252.

Steinnes:1983:TSA

[1006] Donald N. Steinnes and Raymond L. Raab. A time series approach to forecasting angling activity at a recreational site based on past successes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2189–2193, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-253.

Worlund:1983:EDI

[1007] Donald D. Worlund and Gib Taylor. Estimation of disease incidence in fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12): 2194–2197, December 1983. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f83-254.

Salinas:1983:SVW

[1008] I. M. Salinas and I. A. McLaren. Seasonal variation in weight-specific growth rates, feeding rates, and growth efficiencies in Microgadus tomcod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2197–2200, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f83-255.

Dea:1983:CII

[1009] S. Dea and M. A. S. Y. Elazhary. Counterimmunoelectrophoresis for identification of infectious pancreatic necrosis virus after isolation in cell culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2200–2203, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f83-256.

Rudd:1983:EWRa

[1010] John W. M. Rudd, Michael A. Turner, Akira Furutani, Alison L. Swick, and B. E. Townsend. The English–Wabigoon River system: I. A synthesis of recent research with a view towards mercury amelioration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2206–2217, December 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-257.

Rudd:1983:EWRb

[1011] John W. M. Rudd and Michael A. Turner. The English–Wabigoon River system: II. Suppression of mercury and selenium bioaccumulation by suspended and bottom sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2218–2227, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-258.

Turner:1983:EWRa

[1012] Michael A. Turner and John W. M. Rudd. The English–Wabigoon River system: III. selenium in lake enclosures: Its geochemistry, bioaccumulation, and ability to reduce mercury bioaccumulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2228–2240, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-259.

Turner:1983:EWRb

[1013] Michael A. Turner and Alison L. Swick. The English–Wabigoon River system: IV. Interaction between mercury and selenium accumulated from waterborne and dietary sources by northern pike (Esox lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2241–2250, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-260.

Rudd:1983:EWRc

[1014] John W. M. Rudd and Michael A. Turner. The English–Wabigoon River system: V. Mercury and selenium bioaccumulation as a function of aquatic primary productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2251–2259, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f83-261.

Anonymous:1983:LCLf

[1015] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2260–2262, December 1983. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f83-262.

Anonymous:1983:BRCk

[1016] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 40(12):2263–2267, December 1983. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f83-263.

Legendre:1984:TDB

[1017] L. Legendre and S. Demers. Towards dynamic biological oceanography and limnology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (1):2–19, January 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-001.

Somers:1984:AFC

[1018] Keith M. Somers and Harold H. Harvey. Alteration of fish communities in lakes stressed by acid deposition and heavy metals near Wawa, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):20–29, January 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-002.

Jones:1984:NRD

[1019] E. P. Jones, D. Dyrssen, and A. R. Coote. Nutrient regeneration in deep Baffin Bay with consequences for measurements of the conservative tracer NO and fossil fuel CO_{2} in the oceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):30–35, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-003.

Brown:1984:ITC

[1020] S. B. Brown, J. G. Eales, R. E. Evans, and Toshiaki J. Hara. Interrenal, thyroidal, and carbohydrate responses of rainbow trout (Salmo gairdneri) to environmental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):36–45, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-004.

Brezonik:1984:PCF

[1021] Patrick L. Brezonik, Thomas L. Crisman, and Randy L. Schulze. Planktonic communities in Florida softwater lakes of varying pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):46–56, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-005.

Stauffer:1984:LMR

[1022] Robert E. Stauffer and David E. Armstrong. Lake mixing and its relationship to epilimnetic phosphorus in Shagawa Lake, Minnesota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):57–69, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-006.

McDonald:1984:BLS

[1023] J. McDonald and J. M. Hume. Babine lake sockeye salmon (Oncorhynchus nerka) enhancement program: Testing some major assumptions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):70–92, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-007.

Somerton:1984:CBM

[1024] David A. Somerton and Jeffrey June. A cost-benefit method for determining optimum closed fishing areas to reduce the trawl catch of prohibited species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):93–98, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-008.

Botsford:1984:EIG

[1025] Louis W. Botsford. Effect of individual growth rates on expected behavior of the Northern California Dungeness crab (Cancer magister) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):99–107, January 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-009.

Wheeler:1984:HAH

[1026] J. P. Wheeler and G. H. Winters. Homing of Atlantic herring (Clupea harengus harengus) in Newfoundland waters as indicated by tagging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):108–117, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-010.

Gaboury:1984:IWL

[1027] M. N. Gaboury and J. W. Patalas. Influence of water level drawdown on the fish populations of Cross Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):118–125, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-011.

Volk:1984:RBO

[1028] Eric C. Volk, Robert C. Wissmar, Charles A. Simenstad, and Douglas M. Eggers. Relationship between otolith microstructure and the growth of juvenile chum salmon (Oncorhynchus keta) under different prey rations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):126–133, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-012.

Dickie:1984:ISS

[1029] L. M. Dickie, P. R. Boudreau, and K. R. Freeman. Influences of stock and site on growth and mortality in the blue mussel (Mytilus edulis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):134–140, January 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-013.

Barron:1984:NAP

[1030] Mace G. Barron and Ira R. Adelman. Nucleic acid, protein content, and growth of larval fish sublethally exposed to various toxicants. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):141–150, January 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-014.

Geraci:1984:SGH

[1031] J. R. Geraci, D. j. St. Aubin, I. K. Barker, V. S. Hinshaw, R. G. Webster, and H. L. Ruhnke. Susceptibility of grey (Halichoerus grypus) and harp (Phoca groenlandica) seals to the influenza virus and mycoplasma of epizootic pneumonia of harbor seals (Phoca vitulina). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):151–156, January 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-015.

Webb:1984:BFF

[1032] P. W. Webb. Body and fin form and strike tactics of four teleost predators attacking fathead minnow (Pimephales promelas) prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):157–165, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-016.

Brassard:1984:OUR

[1033] P. Brassard and J. C. Auclair. Orthophosphate uptake rate constants are mediated by the 10^{3}–10^{4} molecular weight fraction in Shield lake waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):166–173, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-017.

Winnell:1984:AAC

[1034] Michael H. Winnell and David J. Jude. Associations among chironomidae and sandy substrates in nearshore Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):174–179, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-018.

Sugai:1984:TDO

[1035] Susan F. Sugai and David C. Burrell. Transport of dissolved organic carbon, nutrients, and trace metals from the Wilson and Blossom Rivers to Smeaton Bay, Southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):180–190, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-019.

Taggart:1984:HAZ

[1036] Christopher T. Taggart. Hypolimnetic aeration and zooplankton distribution: a possible limitation to the restoration of cold-water fish production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):191–198, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-020.

Blake:1984:PST

[1037] Robert L. Blake, Franklin L. Roberts, and Richard L. Saunders. Parr– smolt transformation of Atlantic salmon (Salmo salar): Activities of two respiratory enzymes and concentrations of mitochondria in the liver. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):199–203, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-021.

vandenBerghe:1984:FSN

[1038] Eric P. van den Berghe and Mart R. Gross. Female size and nest depth in Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):204–206, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-022.

McCormick:1984:UAC

[1039] Michael J. McCormick and Stephen J. Tarapchak. Uncertainty analysis of calculated nutrient regeneration rates in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):206–211, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-023.

Majkowski:1984:CTM

[1040] Jacek Majkowski and William S. Hearn. Comparison of three methods for estimating the food intake of a fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):212–215, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-024.

Needler:1984:SFA

[1041] A. W. H. Needler. The seventy-fifth anniversary of two Canadian biological stations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):216–224, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-025.

Anonymous:1984:BRCa

[1042] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(1):225–226, January 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f84-026.

Conover:1984:UPZ

[1043] R. J. Conover and P. Mayzaud. Utilization of phytoplankton by zooplankton during the spring bloom in a Nova Scotia inlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):232–244, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-027.

Mayzaud:1984:DDE

[1044] P. Mayzaud and R. J. Conover. Distribution of digestive enzymes in zooplankton during the spring bloom in a Nova Scotia inlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):245–252, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-028.

Perrin:1984:IFL

[1045] C. J. Perrin, K. S. Shortreed, and J. G. Stockner. An integration of forest and lake fertilization: Transport and transformations of fertilizer elements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):253–262, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-029.

Carling:1984:DFC

[1046] Paul A. Carling. Deposition of fine and coarse sand in an open-work gravel bed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):263–270, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-030.

Roland:1984:RMB

[1047] William G. Roland. Resource management biology for the edible kelp Cymathere triplicata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):

271–277, February 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-031.

Farlinger:1984:RCM

[1048] Susan P. Farlinger and Richard J. Beamish. Recent colonization of a major salmon-producing lake in British Columbia by Pacific lamprey (Lampetra tridentata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2): 278–285, February 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-032.

Cook:1984:DFI

[1049] Robert B. Cook. Distributions of ferrous iron and sulfide in an anoxic hypolimnion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):286–293, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-033.

Wroblewski:1984:IAW

[1050] J. S. Wroblewski and Jerry Cheney. Ichthyoplankton associated with a warm core ring off the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):294–303, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-034.

Dunn:1984:DLS

[1051] Jean R. Dunn and Beverly M. Vinter. Development of larvae of the saffron cod, Eleginus gracilis, with comments on the identification of gadid larvae in Pacific and Arctic waters contiguous to Canada and Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):304–318, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-035.

Gunn:1984:SSW

[1052] J. M. Gunn and W. Keller. Spawning site water chemistry and lake trout (Salvelinus namaycush) sac fry survival during spring snowmelt. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):319–329, February 1984.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-036.

Mahon:1984:DSF

[1053] Robin Mahon. Divergent structure in fish taxocenes of north temperate streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):330–350, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-037.

Prepas:1984:CPP

[1054] E. E. Prepas and J. Vickery. Contribution of particulate phosphorus (> 250 μ m) to the total phosphorus pool in lake water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):351–363, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-038.

Fraser:1984:BRB

[1055] Douglas F. Fraser and Edward E. Emmons. Behavioral response of blacknose dace (Rhinichthys atratulus) to varying densities of predatory creek chub (Semotilus atromaculatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):364–370, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-039.

Bilinski:1984:CPM

[1056] E. Bilinski, N. P. Boyce, R. E. E. Jonas, and M. D. Peters. Characterization of protease from the myxosporean salmon parasite, Henneguya salminicola. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2): 371–376, February 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-040.

Ryan:1984:FNC

[1057] P. M. Ryan. Fyke net catches as indices of the abundance of brook trout, Salvelinus fontinalis, and Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):377–380, February 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-041.

Wrona:1984:IFT

[1058] Frederick J. Wrona and Ronald W. Davies. An improved flow-through respirometer for aquatic macroinvertebrate bioenergetic research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):380–385, February 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-042.

Burnison:1984:EFL

[1059] B. Kent Burnison and Gary G. Leppard. Ethanol fraction of lacustrine colloidal fibrils. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2): 385–388, February 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-043.

Heggberget:1984:IEA

[1060] Tor G. Heggberget and Jeff C. Wallace. Incubation of the eggs of Atlantic salmon, Salmo salar, at low temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):389–391, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-044.

Anonymous:1984:BRCb

[1061] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):392–393, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f84-045.

Anonymous:1984:CRPa

[1062] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(2):396, February 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-046.

Fournier:1984:ESC

[1063] D. A. Fournier, T. D. Beacham, B. E. Riddell, and C. A. Busack. Estimating stock composition in mixed stock fisheries using morphometric, meristic, and electrophoretic characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):400–408, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-047.

Tessier:1984:PEL

[1064] Alan J. Tessier. Periodicity of egg laying and egg age distributions in planktonic Cladocera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):409–413, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-048.

Meng:1984:EMM

[1065] Hans Jürg Meng and Max Stocker. An evaluation of morphometrics and meristics for stock separation of Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):414–422, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-049.

Trick:1984:EFI

[1066] Charles G. Trick, Raymond J. Andersen, and Paul J. Harrison. Environmental factors influencing the production of an antibacterial metabolite from a marine dinoflagellate, Prorocentrum minimum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):423–432, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-050.

Seim:1984:GSD

[1067] W. K. Seim, L. R. Curtis, S. W. Glenn, and G. A. Chapman. Growth and survival of developing steelhead trout (Salmo gairdneri) continuously or intermittently exposed to copper. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):433–438, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-051.

Hanson:1984:EPC

[1068] John Mark Hanson and Robert Henry Peters. Empirical prediction of crustacean zooplankton biomass and profundal macrobenthos biomass in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):439–445, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-052.

Stollery:1984:OVU

[1069] K. Stollery. Optimal versus unregulated industry behavior in a Beverton– Holt multicohort fishery model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):446–450, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-053.

Kieser:1984:AEC

[1070] R. Kieser and T. J. Mulligan. Analysis of echo counting data: a model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):451–458, March 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-054.

Hesslein:1984:GPB

[1071] R. H. Hesslein and E. Slavicek. Geochemical pathways and biological uptake of radium in small Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):459–468, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-055.

Rimmer:1984:CSM

[1072] D. M. Rimmer, U. Paim, and R. L. Saunders. Changes in the selection of microhabitat by juvenile Atlantic salmon (Salmo salar) at the summer–autumn transition in a small river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):469–475, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-056.

Healey:1984:IIP

[1073] M. C. Healey and W. R. Heard. Inter- and intra-population variation in the fecundity of Chinook salmon (Oncorhynchus tshawytscha) and its relevance to life history theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):476–483, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-057.

Breiwsck:1984:PDW

[1074] Jeffrey M. Breiwsck, L. Lee Eberhardt, and Howard W. Braham. Population dynamics of Western Arctic bowhead whales (Balaena mysticetus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):484–496, March 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-058.

Canfield:1984:PCC

[1075] Daniel E. Canfield, Jr., Jerome V. Shireman, Douglas E. Colle, William T. Haller, Curtis E. Watkins II, and Michael J. Maceina. Prediction of chlorophyll a concentrations in Florida lakes: Importance of aquatic macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):497–501, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-059.

Fortier:1984:SSC

[1076] L. Fortier and W. C. Leggett. Small-scale covariability in the abundance of fish larvae and their prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):502–512, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-060.

Mills:1984:GEM

[1077] Edward L. Mills, Richard C. Ready, Michael Jahncke, Clare R. Hanger, and Cynthia Trowbridge. A gastric evacuation model for young yellow perch, Perca flavescens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):513–518, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-061.

Nichols:1984:DCF

[1078] D. J. Nichols and M. Weisbart. Dual cannulation of free-swimming Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (3):519–521, March 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-062.

Boulding:1984:CRP

[1079] E. G. Boulding and T. K. Hay. Crab response to prey density can result in density-dependent mortality of clams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):521–525, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-063.

Hartig:1984:OCP

[1080] John H. Hartig and David J. Jude. Opportunistic cyclopoid Prédation on fish larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):526–532, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-064.

Perkins:1984:ADH

[1081] Judith S. Perkins, Kenneth C. Balcomb III, George Nichols, Jr., and Marisa DeAvilla. Abundance and distribution of humpback whales (Megaptera novaeangliae) in West Greenland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):533–536, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-065.

Maguire:1984:AMT

[1082] R. J. Maguire, P. T. S. Wong, and J. S. Rhamey. Accumulation and metabolism of tri-n-butyltin cation by a green alga, Ankistrodesmus falcatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):537–540, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-066.

Parslow:1984:USC

[1083] J. S. Parslow, P. J. Harrison, and P. A. Thompson. Use of a selfcleaning, in-line filter to continuously monitor phytoplankton nutrient uptake rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(3):540–544, March 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-067.

Newbury:1984:SIL

[1084] R. W. Newbury, G. K. McCullough, and R. E. Hecky. The Southern Indian Lake impoundment and Churchill River Diversion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):548–557, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-068.

Newbury:1984:SER

[1085] R. W. Newbury and G. K. McCullough. Shoreline erosion and restabilization in the Southern Indian Lake Reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):558–566, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-069.

Hecky:1984:EID

[1086] R. E. Hecky and G. K. McCullough. Effect of impoundment and diversion on the sediment budget and nearshore sedimentation of Southern Indian Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):567–578, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-070.

Hecky:1984:TOC

[1087] R. E. Hecky. Thermal and optical characteristics of Southern Indian Lake before, during, and after impoundment and Churchill River Diversion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):579–590, April 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-071.

Hecky:1984:PPS

[1088] R. E. Hecky and S. J. Guildford. Primary productivity of Southern Indian Lake before, during, and after impoundment and Churchill River Diversion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):591–604, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-072.

Planas:1984:CPT

[1089] Dolores Planas and R. E. Hecky. Comparison of phosphorus turnover times in Northern Manitoba reservoirs with lakes of the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):605–612, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-073.

Patalas:1984:EID

[1090] K. Patalas and A. Salki. Effects of impoundment and diversion on the crustacean plankton of Southern Indian Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):613–637, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-074.

Wiens:1984:EIR

[1091] Allen P. Wiens and David M. Rosenberg. Effect of impoundment and river diversion on profundal macrobenthos of Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):638–648, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-075.

Crawford:1984:BCN

[1092] Patrick J. Crawford and David M. Rosenberg. Breakdown of conifer needle debris in a new northern reservoir, Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):649–658, April 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-076.

Bilyj:1984:DTN

[1093] Bohdan Bilyj. Descriptions of two new species of Tanypodinae (Diptera: Chironomidae) from Southern Indian Lake, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):659–671, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-077.

Rosenberg:1984:CDE

[1094] David M. Rosenberg, Bohdan Bilyj, and Allen P. Wiens. Chironomidae (Diptera) emerging from the littoral zone of reservoirs, with special reference to Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):672–681, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-078.

Bodaly:1984:IFM

[1095] R. A. Bodaly, R. E. Hecky, and R. J. P. Fudge. Increases in fish mercury levels in lakes flooded by the Churchill River Diversion, Northern Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):682–691, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-079.

Bodaly:1984:CLW

[1096] R. A. Bodaly, T. W. D. Johnson, R. J. P. Fudge, and J. W. Clayton. Collapse of the lake whitefish (Coregonus clupeaformis) fishery in Southern Indian Lake, Manitoba, following lake impoundment and river diversion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):692–700, April 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-080.

Fudge:1984:PWS

[1097] R. J. P. Fudge and R. A. Bodaly. Postimpoundment winter sedimentation and survival of lake whitefish (Coregonus clupeaformis) eggs in Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (4):701–705, April 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-081.

Bodaly:1984:RBN

[1098] R. A. Bodaly and L. F. W. Lesack. Response of a boreal northern pike (Esox lucius) population to lake impoundment: Wupaw Bay, Southern Indian Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (4):706–714, April 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-082.

Wagner:1984:PCF

[1099] M. W. Wagner. Postimpoundment change in financial performance of the Southern Indian Lake commercial fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(4):715–719, April 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-083.

Hecky:1984:EIP

[1100] R. E. Hecky, R. W. Newbury, R. A. Bodaly, K. Patalas, and D. M. Rosenberg. Environmental impact prediction and assessment: the Southern Indian Lake experience. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (4):720–732, April 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-084.

Majkowski:1984:EDE

[1101] Jacek Majkowski, William S. Hearn, and Ronald L. Sandland. An experimental determination of the effect of increasing the minimum age (or size) of fish at capture upon the yield per recruit. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):736–743, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-085.

Iwata:1984:IER

[1102] Munehico Iwata and Shichiko Komatsu. Importance of estuarine residence for adaptation of chum salmon (Oncorhynchus keta) fry to seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):744–749, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-086.

Ramamoorthy:1984:UZC

[1103] S. Ramamoorthy and Karen Blumhagen. Uptake of Zn, Cd, and Hg by fish in the presence of competing compartments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):750–756, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-087.

Nurnberg:1984:BAS

[1104] Gertrud Nürnberg and Robert H. Peters. Biological availability of soluble reactive phosphorus in anoxic and oxic freshwaters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):757–765, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-088.

LaZerte:1984:FAA

[1105] Bruce D. LaZerte. Forms of aqueous aluminum in acidified catchments of Central Ontario: a methodological analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):766–776, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-089.

Rooke:1984:MSL

[1106] J. Brian Rooke and Gerald L. Mackie. Mollusca of six low-alkalinity lakes in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):777–782, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-090.

Evans:1984:PGS

[1107] Geoffrey T. Evans, Jake C. Rice, and E. Michael P. Chadwick. Patterns in growth and smolting of Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):783–797, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-091.

Haug:1984:VDA

[1108] Tore Haug, Elin Kjørsvik, and Per Solemdal. Vertical distribution of Atlantic halibut (Hippoglossus hippoglossus) eggs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):798–804, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-092.

Waltman:1984:MID

[1109] W. Douglas Waltman and Emmett B. Shotts, Jr. A medium for the isolation and differentiation of Yersinia ruckeri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):804–806, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-093.

Amend:1984:CPU

[1110] Donald F. Amend, Terry McDowell, and Ronald P. Hedrick. Characteristics of a previously unidentified virus from channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):807–811, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-094.

Messer:1984:ISC

[1111] Jay J. Messer, Jimmy Ho, and William J. Grenney. Ionic strength correction for extent of ammonia ionization in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):811–815, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-095.

Kalff:1984:MAT

[1112] J. Kalff and E. Bentzen. A method for the analysis of total nitrogen in natural waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):815–819, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
096.

Fletcher:1984:SPV

[1113] G. L. Fletcher, M. H. Kao, and K. Haya. Seasonal and phenotypic variations in plasma protein antifreeze levels in a population of marine fish, sea raven (Hemitripterus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):819–824, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-097.

Anonymous:1984:LCLa

[1114] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(5):825–834, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-098.

Anonymous:1984:BRCc et aquatiques, 41(5):835–836, May 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-099.

Aguilar:1984:RDD

[1116] Alex Aguilar. Relationship of DDE/Σ DDT in marine mammals to the chronology of DDT input into the ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):840–844, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-100.

Riley:1984:RIP

[1117] E. T. Riley and E. E. Prepas. Role of internal phosphorus loading in two shallow, productive lakes in Alberta, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):845–855, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-101.

Grant:1984:BPGa

[1118] W. Stewart Grant and Fred M. Utter. Biochemical population genetics of Pacific herring (Clupea pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):856–864, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-102.

Frenette:1984:BTS

[1119] Jean-Jacques Frenette and Julian J. Dodson. Brook trout (Salvelinus fontinalis) population structure in acidified Lac Tantaré, Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):865–877, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-103.

Oliver:1984:UCO

[1120] Barry G. Oliver. Uptake of chlorinated organics from anthropogenically contaminated sediments by oligochaete worms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):878–883, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-104.

Specht:1984:SFR

[1121] Winona L. Specht, Donald S. Cherry, Richard A. Lechleitner, and John Cairns, Jr. Structural, functional, and recovery responses of stream invertebrates to fly ash effluent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (6):884–896, June 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-105.

Murawski:1984:MSY

[1122] S. A. Murawski. Mixed-species yield-per-recruitment analyses accounting for technological interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (6):897–916, June 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-106.

Scarnecchia:1984:COV

[1123] Dennis L. Scarnecchia. Climatic and oceanic variations affecting yield of Icelandic stocks of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):917–935, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-107.

Schnute:1984:BMA

[1124] Jon Schnute and Skip McKinnell. A biologically meaningful approach to response surface analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (6):936–953, June 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-108.

Frenette:1984:IDP

[1125] M. Frenette, M. Caron, P. Julien, and R. J. Gibson. Interaction entre le débit et les populations de tacons (Salmo salar) de la rivière Matamec, Québec. (French) [Interaction between the flow and the populations of parr (Salmo salar) of the Matamec River, Quebec]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):954–963, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-109.

Kuserk:1984:SMD

[1126] Frank T. Kuserk, Louis A. Kaplan, and Thomas L. Bott. In situ measures of dissolved organic carbon flux in a rural stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):964–973, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-110.

Peterson:1984:MTA

[1127] Hans G. Peterson, F. Patrick Healey, and Rudolf Wagemann. Metal toxicity to algae: a highly pH dependent phenomenon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):974–979, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-111.

Benfey:1984:UEM

[1128] T. J. Benfey, A. M. Sutterlin, and R. J. Thompson. Use of erythrocyte measurements to identify triploid salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):980–984, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-112.

El-Shaarawi:1984:CNL

[1129] A. H. El-Shaarawi and M. A. Neilson. Changes in nutrient levels of lake water stored at 4 ^{o}C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (6):985–988, June 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-113.

Roff:1984:ELH

[1130] Derek A. Roff. The evolution of life history parameters in teleosts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):989–1000, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-114.

Anonymous:1984:LCLb

[1131] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):1001–1002, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-115.

Anonymous:1984:BRCd

[1132] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(6):1003, June 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-116.

Sluczanowski:1984:MOM

[1133] Philip R. Sluczanowski. A management oriented model of an abalone fishery whose substocks are subject to pulse fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1008–1014, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-117.

Bird:1984:ERB

[1134] D. F. Bird and J. Kalff. Empirical relationships between bacterial abundance and chlorophyll concentration in fresh and marine waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1015–1023, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-118.

Guyomard:1984:EVG

[1135] R. Guyomard, C. Grévisse, F. X. Oury, and P. Davaine. Évolution de la variabilité génétique inter et intra-populations de populations de Salmonidés issues de mêmes pools géniques. (French) [Evolution of the inter- and intra-population genetic variability of populations of Salmonidae derived from the same gene pools]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1024–1029, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-119.

Schmitt:1984:BPZ

[1136] Christopher J. Schmitt, F. James Dwyer, and Susan E. Finger. Bioavailability of Pb and Zn from mine tailings as indicated by erythrocyte δ-aminolevulinic acid dehydratase (ALA-D) activity in suckers (Pisces: Catostomidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1030– 1040, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-120.

Berg:1984:REA

[1137] William J. Berg and Stephen D. Ferris. Restriction endonuclease analysis of salmonid mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1041–1047, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-121.

Stephenson:1984:SDP

[1138] G. L. Stephenson, P. Hamilton, N. K. Kaushik, J. B. Robinson, and K. R. Solomon. Spatial distribution of plankton in enclosures of three sizes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1048–1054, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-122.

Sinclair:1984:TSA

[1139] M. Sinclair and M. J. Tremblay. Timing of spawning of Atlantic herring (Clupea harengus harengus) populations and the Match–Mismatch Theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1055–1065, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-123.

Peterson:1984:IVP

[1140] R. H. Peterson. Influence of varying pH and some inorganic cations on the perivitelline potential of eggs of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1066–1069, July 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-124.

Carl:1984:DEF

[1141] L. M. Carl and M. C. Healey. Differences in enzyme frequency and body morphology among three juvenile life history types of Chinook salmon (Oncorhynchus tshawytscha) in the Nanaimo River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1070–1077, July 1984.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-125.

Quinn:1984:HSC

[1142] Thomas P. Quinn and Kurt Fresh. Homing and straying in Chinook salmon (Oncorhynchus tshawytscha) from Cowlitz River Hatchery, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1078–1082, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-126.

Grant:1984:BPGb

[1143] W. Stewart Grant, David J. Teel, Tokimasa Kobayashi, and Cyreis Schmitt. Biochemical population genetics of Pacific halibut (Hippoglossus stenolepis) and comparison with Atlantic halibut (H. hippoglossus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1083–1088, July 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-127.

Pace:1984:ZCS

[1144] Michael L. Pace. Zooplankton community structure, but not biomass, influences the phosphorus–chlorophyll a relationship. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1089–1096, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-128.

Scrivener:1984:LIS

[1145] J. C. Scrivener and B. C. Andersen. Logging impacts and some mechanisms that determine the size of spring and summer populations of Coho salmon fry (Oncorhynchus kisutch) in Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1097–1105, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
129.

Anderson:1984:ELH

[1146] John T. Anderson. Early life history of redfish (Sebastes spp.) on Flemish Cap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1106–1116, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-130.

Peterson:1984:MRF

[1147] Ingrid Peterson and J. S. Wroblewski. Mortality rate of fishes in the pelagic ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1117– 1120, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-131.

Larkin:1984:CEI

[1148] P. A. Larkin. A commentary on environmental impact assessment for large projects affecting lakes and streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1121–1127, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-132.

Anonymous:1984:BRCe

[1149] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(7):1128, July 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-133.

Hall:1984:EDR

[1150] Ronald J. Hall and Gene E. Likens. Effect of discharge rate on biotic and abiotic chemical flux in an acidified stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1132–1138, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-134.

Sutterlin:1984:AFM

[1151] A. M. Sutterlin and D. MacLean. Age at first maturity and the early expression of oocyte recruitment processes in two forms of Atlantic salmon (Salmo salar) and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1139–1149, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-135.

Winans:1984:MMV

[1152] Gary A. Winans. Multivariate morphometric variability in Pacific salmon: Technical demonstration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1150–1159, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-136.

Karpoff:1984:IML

[1153] Jonathan M. Karpoff. Insights from the markets for limited entry permits in Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1160–1166, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-137.

Carter:1984:USL

[1154] J. C. H. Carter, M. J. Dadswell, and K. Goudie. Univoltine and semivoltine life histories in Senecella calanoides (Copepoda: Calanoida) and their relationship to body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1167–1175, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-138.

Dunbrack:1984:TDP

[1155] Robert L. Dunbrack and Lawrence M. Dill. Three-dimensional prey reaction field of the juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1176–1182, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-139.

Schwarz:1984:RPH

[1156] Abby L. Schwarz and Galen L. Greer. Responses of Pacific herring, Clupea harengus pallasi, to some underwater sounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1183–1192, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-140.

Leggett:1984:MHR

[1157] W. C. Leggett, K. T. Frank, and J. E. Carscadden. Meteorological and hydrographic regulation of year-class strength in capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1193–1201, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
141.

Springer:1984:ECM

[1158] Alan M. Springer, David G. Roseneau, Edward C. Murphy, and Martha I. Springer. Environmental controls of marine food webs: Food habits of seabirds in the Eastern Chukchi Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1202–1215, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-142.

Crecco:1984:EFH

[1159] Victor A. Crecco and Thomas F. Savoy. Effects of fluctuations in hydrographic conditions on year-class strength of American shad (Alosa sapidissima) in the Connecticut River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1216–1223, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-143.

From:1984:SUG

[1160] Jon From and Viggo Hørlyck. Sites of uptake of geosmin, a cause of earthy-flavor, in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1224–1226, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-144.

Caddy:1984:MSN

[1161] J. F. Caddy. Method of solving for natural mortality rate for stock components with different schedules of growth and mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1226–1230, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-145.

Swain:1984:SPV

[1162] D. P. Swain and C. C. Lindsey. Selective predation for vertebral number of young sticklebacks, Gasterosteus aculeatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1231–1233, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-146.

Scarnecchia:1984:FYT

[1163] Dennis L. Scarnecchia. Forecasting yields of two-sea-winter Atlantic salmon (Salmo salar) from Icelandic rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1234–1240, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-147.

Gensemer:1984:GRF

[1164] Robert W. Gensemer and Susan Soltau Kilham. Growth rates of five freshwater algae in well-buffered acidic media. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1240–1243, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-148.

Blazer:1984:EDI

[1165] Vicki S. Blazer and Richard E. Wolke. Effect of diet on the immune response of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1244–1247, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-149.

Pringle:1984:SSF

[1166] Catherine M. Pringle and James A. Bowers. An in situ substratum fertilization technique: Diatom colonization on nutrient-enriched, sand substrata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1247–1251, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
150.

Rudstam:1984:SSP

[1167] Lars G. Rudstam, John J. Magnuson, and William M. Tonn. Size selectivity of passive fishing gear: a correction for encounter probability applied to gill nets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8): 1252–1255, August 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-151.

Gorham:1984:EEA

[1168] Eville Gorham, Suzanne E. Bayley, and David W. Schindler. Ecological effects of acid deposition upon peatlands: a neglected field in “Acid-Rain” research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8): 1256–1268, August 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-152.

Anonymous:1984:LCLc

[1169] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1269–1270, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-153.

Anonymous:1984:BRCf

[1170] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1271–1273, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f84-154.

Anonymous:1984:CRPb

[1171] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(8):1276, August 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
155.

MacFarsane:1984:DCF

[1172] R. Bruce MacFarsane. Determination of corticosteroids in fish plasma by high performance liquid chromatography: Evaluation of the method using striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1280–1286, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-156.

Murtaugh:1984:VGR

[1173] Paul A. Murtaugh. Variable gut residence time: Problems in inferring feeding rate from stomach fullness of a mysid crustacean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1287–1293, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-157.

Frank:1984:SEC

[1174] K. T. Frank and W. C. Leggett. Selective exploitation of capelin (Mallotus villosus) eggs by winter flounder (Pseudopleuronectes americanus): Capelin egg mortality rates, and contribution of egg energy to the annual growth of flounder. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9): 1294–1302, September 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-158.

Borgmann:1984:ZPF

[1175] U. Borgmann, H. Shear, and J. Moore. Zooplankton and potential fish production in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1303–1309, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-159.

Whalen:1984:ITL

[1176] S. C. Whalen and V. Alexander. Influence of temperature and light on rates of inorganic nitrogen transport by algae in an Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1310–1318, September 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-160.

Quinn:1984:PSO

[1177] Thomas P. Quinn and Cornelis Groot. Pacific salmon (Oncorhynchus) migrations: Orientation versus random movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1319–1324, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-161.

Kimura:1984:GSR

[1178] Daniel K. Kimura, James W. Balsiger, and Daniel H. Ito. Generalized stock reduction analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9): 1325–1333, September 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-162.

Derby:1984:LST

[1179] Jennifer G. Smith Derby and Judith M. Capuzzo. Lethal and sublethal toxicity of drilling fluids to larvae of the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1334– 1340, September 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-163.

Nakashima:1984:SET

[1180] Brian S. Nakashima and G. H. Winters. Selection of external tags for marking Atlantic herring (Clupea harengus harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1341–1348, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-164.

Myers:1984:DCP

[1181] R. A. Myers. Demographic consequences of precocious maturation of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (9):1349–1353, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f84-165.

Hackett:1984:EPE

[1182] Jeremy L. Hackett, William H. Lynch, William D. Paterson, and David H. Coombs. Extracellular protease, extracellular haemolysin, and virulence in Aeromonas salmonicida. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1354–1360, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-166.

Rago:1984:SIL

[1183] Paul J. Rago and Robert M. Dorazio. Statistical inference in life-table experiments: The finite rate of increase. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1361–1374, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-167.

Hetrick:1984:ICS

[1184] F. M. Hetrick, L. W. Hall, Jr., S. Wolski, W. C. Graves, B. S. Roberson, and D. T. Burton. Influence of chlorine on the susceptibility of striped bass (Morone saxatilis) to Vibrio anguillarum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1375–1380, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-168.

Edwards:1984:CGW

[1185] R. R. C. Edwards. Comparisons of growth in weight of temperate and tropical marine fish counterparts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1381–1384, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-169.

Atkinson:1984:EAI

[1186] J. L. Atkinson, J. W. Hilton, and S. J. Slinger. Evaluation of acidinsoluble ash as an indicator of feed digestibility in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1384–1386, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-170.

Benfey:1984:GGD

[1187] Tillmann J. Benfey and Arnold M. Sutterlin. Growth and gonadal development in triploid landlocked Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1387–1392, September 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-171.

Healey:1984:MAC

[1188] M. C. Healey. Multiattribute analysis and the concept of optimum yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1393–1406, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-172.

Anonymous:1984:BRCg

[1189] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(9):1407–1408, September 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-173.

Parkinson:1984:GVP

[1190] Eric A. Parkinson. Genetic variation in populations of steelhead trout (Salmo gairdneri) in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1412–1420, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-174.

Jacobson:1984:PSS

[1191] L. D. Jacobson, R. L. Torblaa, and R. H. Morman. Problems with samples of sea lamprey (Petromyzon marinus) ammocoetes used in stock identification studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10): 1421–1427, October 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-175.

Taggart:1984:ELV

[1192] Christopher T. Taggart and William C. Leggett. Efficiency of largevolume plankton pumps, and evaluation of a design suitable for deployment from small boats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10): 1428–1435, October 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-176.

Priscu:1984:INU

[1193] John C. Priscu and Linda R. Priscu. Inorganic nitrogen uptake in oligotrophic Lake Taupo, New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1436–1445, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-177.

Smoker:1984:GED

[1194] William W. Smoker. Genetic effect on the dynamics of a model of pink (Oncorhynchus gorbuscha) and chum (O. keta) salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1446–1453, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-178.

Leider:1984:SCS

[1195] Steven A. Leider, Mark W. Chilcote, and John J. Loch. Spawning characteristics of sympatric populations of steelhead trout (Salmo gairdneri): Evidence for partial reproductive isolation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1454–1462, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-179.

Tessier:1984:RBP

[1196] A. Tessier, P. G. C. Campbell, J. C. Auclair, and M. Bisson. Relationships between the partitioning of trace metals in sediments and their accumulation in the tissues of the freshwater mollusc Elliptio complanata in a mining area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10): 1463–1472, October 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-180.

Conners:1984:PAI

[1197] M. Elizabeth Conners and Robert J. Naiman. Particulate allochthonous inputs: Relationships with stream size in an undisturbed watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1473–1484, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-181.

Pringle:1984:EEV

[1198] J. D. Pringle. Efficiency estimates for various quadrat sizes used in benthic sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1485– 1489, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f84-182.

Ridler:1984:SAS

[1199] Neil B. Ridler. Socioeconomic aspects of sea cage salmon farming in the maritimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1490–1495, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-183.

Post:1984:CBC

[1200] John R. Post and Dan Cucin. Changes in the benthic community of a small Precambrian lake following the introduction of yellow perch, Perca flavescens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1496–1501, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-184.

Neilson:1984:TTC

[1201] Alasdair H. Neilson, Ann-Sofie Allard, Sven Reiland, Mikael Remberger, Anders Tärnholm, Tomas Viktor, and Lars Landner. Tri- and tetra-chloroveratrole, metabolites produced by bacterial O-methylation of tri- and tetra-chloroguaiacol: an assessment of their bioconcentration potential and their effects on fish reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1502–1512, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-185.

Evans:1984:ISL

[1202] Marlene S. Evans and Loren E. Flath. Intakes as sampling locations for investigating long-term trends in zooplankton populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1513–1518, October 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-186.

Speziale:1984:CDD

[1203] Barbara J. Speziale, Stephen P. Schreiner, Paul A. Giammatteo, and James E. Schindler. Comparison of N,N-dimethylformamide, dimethyl sulfoxide, and acetone for extraction of phytoplankton chlorophyll. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1519–1522, October 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-187.

Carlson:1984:EIM

[1204] David J. Carlson, David W. Townsend, Alan L. Hilyard, and Jonathan F. Eaton. Effect of an intertidal mudflat on plankton of the overlying water column. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1523–1528, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f84-
188.

Barnes:1984:SRC

[1205] M. A. Barnes, G. Power, and R. G. H. Downer. Stress-related changes in lake whitefish (Coregonus clupeaformis) associated with a hydroelectric control structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10): 1528–1533, October 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-189.

Gilbertson:1984:NDE

[1206] Michael Gilbertson. Need for development of epidemiology for chemically induced diseases in fish in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1534–1540, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-190.

Anonymous:1984:BRCh

[1207] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(10):1541–1542, October 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f84-191.

Puckett:1984:CSB

[1208] K. J. Puckett and L. M. Dill. Cost of sustained and burst swimming to juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1546–1551, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-192.

Lambert:1984:LCS

[1209] Timothy C. Lambert. Larval cohort succession in herring (Clupea harengus) and capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1552–1564, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-193.

Lambert:1984:RSD

[1210] T. C. Lambert and D. M. Ware. Reproductive strategies of demersal and pelagic spawning fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11): 1565–1569, November 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-194.

Wong:1984:UAI

[1211] P. T. S. Wong, R. J. Maguire, Y. K. Chau, and O. Kramar. Uptake and accumulation of inorganic tin by a freshwater alga, Ankistrodesmus falcatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1570–1574, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-195.

Hankin:1984:MSD

[1212] David G. Hankin. Multistage sampling designs in fisheries research: Applications in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (11):1575–1591, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f84-196.

Shearer:1984:CEC

[1213] Karl D. Shearer. Changes in elemental composition of hatchery-reared rainbow trout, Salmo gairdneri, associated with growth and reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1592–1600, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-197.

Hendzel:1984:EAA

[1214] L. L. Hendzel and F. P. Healey. Extraction of algal ATP and interpretation of measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11): 1601–1608, November 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-198.

Twinch:1984:PEB

[1215] A. J. Twinch and R. H. Peters. Phosphate exchange between littoral sediments and overlying water in an oligotrophy north-temperate lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1609–1617, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-199.

Wildish:1984:IMB

[1216] D. J. Wildish and D. D. Kristmanson. Importance to mussels of the benthic boundary layer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11): 1618–1625, November 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-200.

Ciborowski:1984:STC

[1217] Jan J. H. Ciborowski and Hugh F. Clifford. Short-term colonization patterns of lotic macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1626–1633, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-201.

Mackay:1984:SCV

[1218] Rosemary J. Mackay. Some causes of variation in estimates of mean weights for Hydropsyche larvae (Trichoptera: Hydropsychidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1634–1642, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-202.

Wallace:1984:QWH

[1219] J. Bruce Wallace and Arthur C. Benke. Quantification of wood habitat in subtropical coastal plain streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1643–1652, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-203.

Powles:1984:NIN

[1220] H. Powles, F. Auger, and G. J. FitzGerald. Nearshore ichthyoplankton of a North temperate estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1653–1663, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-204.

LaZerte:1984:RIA

[1221] Bruce D. LaZerte and Peter J. Dillon. Relative importance of anthropogenic versus natural sources of acidity in lakes and streams of Central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1664–1677, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-205.

Giles:1984:EWB

[1222] Michael A. Giles. Electrolyte and water balance in plasma and urine of rainbow trout (Salmo gairdneri) during chronic exposure to cadmium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1678–1685, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-206.

Giles:1984:OHR

[1223] M. A. Giles, H. S. Majewski, and B. Hobden. Osmoregulatory and hematological responses of rainbow trout (Salmo gairdneri) to extended environmental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11): 1686–1694, November 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-207.

Dermott:1984:CPD

[1224] R. M. Dermott and H. J. Spence. Changes in populations and drift of stream invertebrates following lampricide treatment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1695–1701, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-208.

Taylor:1984:PFT

[1225] W. D. Taylor. Phosphorus flux through epilimnetic zooplankton from Lake Ontario: Relationship with body size and significance to phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1702–1712, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-209.

Boese:1984:UEG

[1226] Bruce L. Boese. Uptake efficiency of the gills of English sole (Parophrys vetulus) for four phthalate esters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1713–1718, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-210.

Anonymous:1984:BRCi

[1227] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(11):1719, November 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f84-211.

Koslow:1984:RPN

[1228] J. Anthony Koslow. Recruitment patterns in Northwest Atlantic fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1722–1729, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-212.

Fournier:1984:VCS

[1229] R. O. Fournier, R. Ernst, N. B. Hargreaves, M. Van Det, and D. Douglas. Variability of chlorophyll a off Southwestern Nova Scotia in late fall and its relationship to water column stability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1730–1738, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-213.

Palachek:1984:TNC

[1230] R. M. Palachek and J. R. Tomasso. Toxicity of nitrite to channel catfish (Ictalurus punctatus), Tilapia (Tilapia aurea), and largemouth bass (Micropterus salmoides): Evidence for a nitrite exclusion mechanism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1739–1744, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-214.

Sandheinrich:1984:IRP

[1231] Mark B. Sandheinrich and Wayne A. Hubert. Intraspecific resource partitioning by yellow perch (Perca flavescens) in a stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1745–1752, December 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-215.

Ni:1984:SMN

[1232] I-H. Ni and E. J. Sandeman. Size at maturity for Northwest Atlantic redfishes (Sebastes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12): 1753–1762, December 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-216.

Carper:1984:WRS

[1233] Glenn L. Carper and Roger W. Bachmann. Wind resuspension of sediments in a prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12): 1763–1767, December 1984. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f84-217.

Perry:1984:EPH

[1234] S. F. Perry, P. Payan, and J. P. Girard. Effects of perfusate HCO_{3}^{−} and pCO_{2} on chloride uptake in perfused gills of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1768–1773, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f84-218.

Rodgers:1984:APC

[1235] David W. Rodgers. Ambient pH and calcium concentration as modifiers of growth and calcium dynamics of brook trout, Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1774–1780, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-219.

Legendre:1984:PDF

[1236] Pierre Legendre and Vianney Legendre. Postglacial dispersal of freshwater fishes in the Québec Peninsula. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1781–1802, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f84-220.

Soballe:1984:IRT

[1237] D. M. Søballe and R. W. Bachmann. Influence of reservoir transit on riverine algal transport and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1803–1813, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-221.

Peterman:1984:CCB

[1238] Randall M. Peterman and Fred Y. C. Wong. Cross correlations between reconstructed ocean abundances of Bristol Bay and British Columbia sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1814–1824, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-222.

Peterman:1984:DDG

[1239] Randall M. Peterman. Density-dependent growth in early ocean life of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1825–1829, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-223.

Houston:1984:CIE

[1240] A. H. Houston and J. E. Keen. Cadmium inhibition of erythropoiesis in goldfish, Carassius auratus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41 (12):1829–1834, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f84-224.

Vuorinen:1984:EIA

[1241] Jukka Vuorinen and Jorma Piironen. Electrophoretic identification of Atlantic salmon (Salmo salar), brown trout (S. trutta), and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1834–1837, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-225.

Arkett:1984:DVM

[1242] S. A. Arkett. Diel vertical migration and feeding behavior of a demersal hydromedusan (Polyorchis penicillatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1837–1843, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-226.

Barlow:1984:MEB

[1243] Jay Barlow. Mortality estimation: Biased results from unbiased ages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1843–1847, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-227.

Scheibling:1984:PRC

[1244] Robert E. Scheibling. Predation by rock crabs (Cancer irroratus) on diseased sea urchins (Strongylocentrotus droebachiensis) in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1847–1851, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f84-228.

Spencer:1984:RFR

[1245] Craig N. Spencer and Darren L. King. Role of fish in regulation of plant and animal communities in eutrophic ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1851–1855, December 1984. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f84-229.

Anonymous:1984:LCLd

[1246] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 41(12):1856–1859, December 1984. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f84-230.

Hilborn:1985:FDI

[1247] Ray Hilborn. Fleet dynamics and individual variation: Why some people catch more fish than others. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (1):2–13, January 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-001.

Falk-Petersen:1985:GET

[1248] Stig Falk-Petersen. Growth of the euphausiids Thysanoessa inermis, Thysanoessa raschii, and Meganyctiphanes norvegica in a subarctic fjord, North Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):14–22, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-002.

Campbell:1985:ACZ

[1249] Peter G. C. Campbell, André Tessier, Marc Bisson, and René Bougie. Accumulation of copper and zinc in the yellow water lily, Nuphar variegatum: Relationships to metal partitioning in the adjacent lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):23–32, January 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-003.

Gendron:1985:CBL

[1250] L. Gendron. Conversion of blade length increments into biomass production estimates in Laminaria longicruris. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):33–37, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-004.

Wood:1985:CME

[1251] Lindsay W. Wood. Chloroform–methanol extraction of chlorophyll a. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):38–43, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-005.

Tsoa:1985:GFM

[1252] E. Tsoa, W. E. Schrank, and N. Roy. Generalizing fisheries models: an extension of the Schaefer analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):44–50, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-006.

Hilborn:1985:DCP

[1253] Ray Hilborn and Max Ledbetter. Determinants of catching power in the British Columbia salmon purse seine fleet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):51–56, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-007.

Bevelhimer:1985:ASP

[1254] Mark S. Bevelhimer, Roy A. Stein, and Robert F. Carline. Assessing significance of physiological differences among three esocids with a bioenergefics model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):57–69, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-008.

Solomon:1985:DPL

[1255] K. R. Solomon, J. Y. Yoo, D. Lean, N. K. Kaushik, K. E. Day, and G. L. Stephenson. Dissipation of permethrin in limnocorrals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):70–76, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-009.

Kaushik:1985:IPZ

[1256] N. K. Kaushik, G. L. Stephenson, K. R. Solomon, and K. E. Day. Impact of permethrin on zooplankton communities in limnocorrals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):77–85, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-010.

Bates:1985:ZPI

[1257] Stephen S. Bates, André Tessier, Peter G. C. Campbell, and Martin Létourneau. Zinc–phosphorus interactions and variation in zinc accumulation during growth of Chlamydomonas variabilis (Chlorophyceae) in batch culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):86–94, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-011.

Starkel:1985:PEM

[1258] William M. Starkel. Predicting the effect of macrobenthos on the sediment/water flux of metals and phosphorus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):95–100, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-012.

Hume:1985:ICU

[1259] Jeremy M. B. Hume and Thomas G. Northcote. Initial changes in use of space and food by experimentally segregated populations of Dolly Varden (Salvelinus malma) and cutthroat trout (Salmo clarki). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):101–109, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-013.

Campton:1985:NHB

[1260] Donald E. Campton and Fred M. Utter. Natural hybridization between steelhead trout (Salmo gairdneri) and coastal cutthroat trout (Salmo clarki clarki) in two Puget Sound streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):110–119, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-014.

Wahi:1985:FES

[1261] David H. Wahi and Larry A. Nielsen. Feeding ecology of the sauger (Stizostedion canadense) in a large river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):120–128, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-015.

Mills:1985:RLW

[1262] K. H. Mills. Responses of lake whitefish (Coregonus clupeaformis) to fertilization of Lake 226, the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):129–138, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-016.

Culp:1985:RBM

[1263] Joseph M. Culp and Ronald W. Davies. Responses of benthic macroinvertebrate species to manipulation of interstitial detritus in Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (1):139–146, January 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-017.

Walters:1985:BEF

[1264] Carl J. Walters. Bias in the estimation of functional relationships from time series data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):147–149, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-018.

Mangel:1985:SSD

[1265] Marc Mangel and Jay H. Beder. Search and stock depletion: Theory and applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):150–163, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-019.

Jensen:1985:RAN

[1266] A. L. Jensen. Relations among net reproductive rate and life history parameters for lake whitefish (Coregonus clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):164–168, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-020.

Coulombe:1985:SEB

[1267] Francis Coulombe, Jean-Claude F. Brêthes, Roxane Bouchard, and Gaston Desrosiers. Ségrégation édaphique et bathymétrique chez le crabe des neiges, Chionoecetes opilio (O. Fabr.), dans le sud-ouest du golfe du Saint-Laurent. (French) [Edaphic and bathymetric segregation in snow crabs, Chionoecetes opilio (O. Fabr.), in the Southwestern Gulf of St. Lawrence]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):169–180, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f85-
021.

Hardy:1985:EDC

[1268] Ronald W. Hardy and Karl D. Shearer. Effect of dietary calcium phosphate and zinc supplementation on whole body zinc concentration of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):181–184, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-022.

Degani:1985:EDM

[1269] Gad Degani and Margie Lee Gallagher. Effects of dietary 17αmethyltestosterone and bovine growth hormone on growth and food conversion of slow- and normally-growing American elvers (Anguilla rostrata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):185–189, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-023.

Wong:1985:SSF

[1270] C. K. Wong and W. G. Sprules. Size-selective feeding by the predatory copepod Epischura lacustris Forbes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):189–193, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-024.

Aiken:1985:UIS

[1271] D. E. Aiken and S. L. Waddy. The uncertain influence of spring photoperiod on spawning in the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):194–197, January 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-025.

Anonymous:1985:LCLa

[1272] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):198–200, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-026.

Anonymous:1985:BRCa

[1273] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(1):201–202, January 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-027.

Allan:1985:QSD

[1274] J. David Allan and Estelle Russek. The quantification of stream drift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):210–215, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-028.

Bruce-Allen:1985:WHB

[1275] L. J. Bruce-Allen and J. R. Geraci. Wound healing in the bottlenose dolphin (Tursiops truncatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (2):216–228, February 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-029.

Campbell:1985:MTA

[1276] Alan Campbell and Aivars B. Stasko. Movements of tagged American lobsters, Homarus americanus, off Southwestern Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):229–238, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-030.

Babin:1985:MWO

[1277] J. Babin and E. E. Prepas. Modelling winter oxygen depletion rates in ice-covered temperate zone lakes in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):239–249, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-031.

Reddin:1985:HIO

[1278] D. G. Reddin and R. K. Misra. Hotelling’s T^{2} to identify the origin of Atlantic salmon (Salmo salar) in a mixed-stock fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):250–255, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-032.

Apostle:1985:WSC

[1279] Richard Apostle, Leonard Kasdan, and Arthur Hanson. Work satisfaction and community attachment among fishermen in Southwest Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):256–267, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-033.

Mathias:1985:GSC

[1280] Jack A. Mathias and Jan Barica. Gas supersaturation as a cause of early spring mortality of stocked trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):268–279, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-034.

Hanson:1985:EFE

[1281] John Mark Hanson and William C. Leggett. Experimental and field evidence for inter- and intraspecific competition in two freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):280–286, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-035.

Rombough:1985:IEW

[1282] Peter J. Rombough. Initial egg weight, time to maximum alevin wet weight, and optimal ponding times for Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):287–291, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-036.

Lacroix:1985:SEA

[1283] Gilles L. Lacroix. Survival of eggs and alevins of Atlantic salmon (Salmo salar) in relation to the chemistry of interstitial water in redds in some acidic streams of Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):292–299, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-037.

Uttormark:1985:CPC

[1284] Paul D. Uttormark and Mark L. Hutchins. Cellular phosphorus concentration as a criterion for trophic status in rapidly flushed lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):300–306, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-038.

Samples:1985:EPV

[1285] Karl C. Samples and PingSun Leung. The effect of production variability on financial risks of freshwater prawn farming in Hawaii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):307–311, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-039.

Beacham:1985:VLB

[1286] Terry D. Beacham and Clyde B. Murray. Variation in length and body depth of pink salmon (Oncorhynchus gorbuscha) and chum salmon (O. keta) in Southern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):312–319, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-040.

Hyatt:1985:RSS

[1287] Kim D. Hyatt and John G. Stockner. Responses of sockeye salmon (Oncorhynchus nerka) to fertilization of British Columbia coastal lakes.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):320–331, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-041.

Quellet:1985:DRA

[1288] Patrick Quellet and Julian J. Dodson. Dispersion and retention of anadromous rainbow smelt (Osmerus mordax) larvae in the middle estuary of the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (2):332–341, February 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-042.

Uhazy:1985:MSN

[1289] L. S. Uhazy, D. E. Aiken, and A. Campbell. Morphology and systematics of the Nemertean Pseudocarcinonemertes homari (Hoplonemertea: Mortostilifera) from the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):342–350, February 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-043.

Aiken:1985:ABP

[1290] D. E. Aiken, S. L. Waddy, and L. S. Uhazy. Aspects of the biology of Pseudocarcinonemertes homari and its association with the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):351–356, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-044.

Waddy:1985:ILA

[1291] S. L. Waddy and D. E. Aiken. Immunofluorescent localization of American lobster egg yolk protein in the alimentary tract of the Nemertean Pseudocarcinonemertes homari. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):357–359, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-045.

Brattey:1985:GDS

[1292] J. Brattey, A. Campbell, A. E. Bagnall, and L. S. Uhazy. Geographic distribution and seasonal occurrence of the Nemertean Pseudocarcinonemertes homari on the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):360–367, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-046.

Kean:1985:JLH

[1293] J. C. Kean, J. D. Castell, and D. J. Trider. Juvenile lobster (Homarus americanus) do not require dietary ascorbic acid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):368–370, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-047.

Desjardins:1985:SDA

[1294] L. M. Desjardins, J. D. Castell, and J. C. Kean. Synthesis of dehydroascorbic acid by subadult lobsters (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):370–373, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-048.

Marten:1985:REM

[1295] Gerald G. Marten. Radiotracer estimation of microbial biomass in lake sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):373–379, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-049.

Hurley:1985:SLV

[1296] Geoffrey V. Hurley, Paul H. Odense, Ron K. O’Dor, and Earl G. Dawe. Strontium labelling for verifying daily growth increments in the statolith of the short-finned squid (Illex illecebrosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):380–383, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-050.

Klotz:1985:ILA

[1297] R. L. Klotz. Influence of light on the alkaline phosphatase activity of Selenastrum capricornutum (Chlorophyceae) in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):384–388, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-051.

Pringle:1985:HFF

[1298] J. D. Pringle. The human factor in fishery resource management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):389–392, February 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-052.

Anonymous:1985:LCLb

[1299] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):393–395, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-053.

Anonymous:1985:BRCb

[1300] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):396–397, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-054.

Anonymous:1985:CRPa

[1301] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(2):400, February 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-055.

Waterhouse:1985:IPR

[1302] J. C. Waterhouse and M. P. Farrell. Identifying pollution related changes in chironomid communities as a function of taxonomic rank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):406–413, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-056.

Schnute:1985:GTA

[1303] Jon Schnute. A general theory for analysis of catch and effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):414–429, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-057.

StAubin:1985:HDB

[1304] D. J. St. Aubin, J. R. Geraci, T. G. Smith, and T. G. Friesen. How do bottlenose dolphins, Tursiops truncatus, react to oil films under different light conditions? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):430–436, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-058.

Beacham:1985:BGSa

[1305] Terry D. Beacham, Ruth E. Withler, and Allan P. Gould. Biochemical genetic stock identification of chum salmon (Oncorhynchus keta) in Southern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):437–448, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-059.

Eck:1985:LMC

[1306] G. W. Eck and Edward H. Brown, Jr. Lake Michigan’s capacity to support lake trout (Salvelinus namaycush) and other salmonines: an estimate based on the status of prey populations in the 1970s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):449–454, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-060.

Dutil:1985:DSP

[1307] J.-D. Dutil, B. Légaré, and C. Desjardins. Discrimination d’un stock de poisson, l’anguille (Anguilla rostrata), basée sur la présence d’un produit chimique de synthèse, le mirex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):455–458, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-061.

Anderson:1985:EPD

[1308] James L. Anderson and James E. Wilen. Estimating the population dynamics of Coho salmon (Oncorhynchus kisutch) using pooled time-series and cross-sectional data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):459–467, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-062.

Cooke:1985:EHS

[1309] Justin G. Cooke. Estimation of harp seal (Phoca groenlandica) pup production from age samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):468–473, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-063.

Davidson:1985:MEF

[1310] K. Davidson, J. C. Roff, and R. W. Elner. Morphological, electrophoretic and fecundity characteristics of Atlantic snow crab, Chionoecetes opilio, and implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):474–482, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-064.

Hershey:1985:DDR

[1311] Anne E. Hershey and Michael E. McDonald. Diet and digestion rates of slimy sculpin, Cottus cognatus, in an Alaskan Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):483–487, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-065.

Mellas:1985:SPB

[1312] Ernest J. Mellas and James M. Haynes. Swimming performance and behavior of rainbow trout (Salmo gairdneri) and white perch (Morone americana): Effects of attaching telemetry transmitters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):488–493, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-066.

Welch:1985:ILR

[1313] Harold E. Welch. Introduction to limnological research at Saqvaqjuac, Northern Hudson Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):494–505, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-067.

Welch:1985:WCS

[1314] Harold E. Welch and Martin A. Bergmann. Water circulation in small Arctic lakes in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):506–520, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-068.

Welch:1985:WRL

[1315] Harold E. Welch and Martin A. Bergmann. Winter respiration of lakes at Saqvaqjuac, N.W.T. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):521–528, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-069.

McCormick:1985:PSC

[1316] Stephen D. McCormick, Robert J. Naiman, and Ellyn T. Montgomery. Physiological smolt characteristics of anadromous and non-anadromous brook trout (Salvelinus fontinalis) and Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):529–538, March 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-070.

Evans:1985:PGS

[1317] Geoffrey T. Evans, Jake C. Rice, and E. Michael P. Chadwick. Patterns of growth and smolting of Atlantic salmon (Salmo salar) parr in a Southwestern Newfoundland river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):539–543, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-071.

Levy:1985:BLD

[1318] E. M. Levy. Background levels of dissolved/dispersed petroleum residues in the Gulf of St. Lawrence, 1970–79. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):544–555, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-072.

Boisclair:1985:RFE

[1319] D. Boisclair and W. C. Leggett. Rates of food exploitation by littoral fishes in a mesotrophic north-temperate lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):556–566, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-073.

Chow-Fraser:1985:FRS

[1320] Patricia Chow-Fraser and Roy Knoechel. Factors regulating in situ filtering rates of cladocera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):567–576, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-074.

Ware:1985:ELH

[1321] D. M. Ware and T. C. Lambert. Early life history of Atlantic mackerel (Scomber scombrus) in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):577–592, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-075.

Jonsson:1985:WTP

[1322] Bror Jonsson and Jan Ruud-Hansen. Water temperature as the primary influence on timing of seaward migrations of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):593–595, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-076.

Grant:1985:MMH

[1323] Jonathan Grant. A method for measuring horizontal transport of organic carbon over sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):595–602, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-077.

Sinclair:1985:NNE

[1324] M. Sinclair, M. J. Tremblay, and P. Bernal. El Niño events and variability in a Pacific mackerel (Scomber japonicus) survival index: Support for Hjort’s second hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (3):602–608, March 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-078.

Vladykov:1985:BTT

[1325] Vadim D. Vladykov, Claude B. Renaud, and Sylvie Laframboise. Breeding tubercles in three species of Gadus (cods). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):608–615, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-079.

Saunders:1985:IVL

[1326] Richard L. Saunders and Charles B. Schom. Importance of the variation in life history parameters of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):615–618, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-080.

Legault:1985:TAF

[1327] Michel Legault and Louis-Marie Lalancette. Temps d’ajustement de la flottabilité des saumons atlantiques (Salmo salar) d’élevage et flottabilité comparée de saumons atlantiques d’élevage et sauvages. (French) [Buoyancy adjustment time of farmed Atlantic salmon (Salmo salar) and comparative buoyancy of farmed and wild Atlantic salmon]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):619–623, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-081.

Anonymous:1985:BRCc

[1328] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(3):624–626, March 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-082.

Mallatt:1985:FGS

[1329] Jon Mallatt. Fish gill structural changes induced by toxicants and other irritants: a statistical review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):630–648, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-083.

Stockner:1985:WLF

[1330] J. G. Stockner and K. S. Shortreed. Whole-lake fertilization experiments in coastal British Columbia lakes: Empirical relationships between nutrient inputs and phytoplankton biomass and production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):649–658, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-084.

Hargreaves:1985:SSP

[1331] N. Brent Hargreaves and Robin J. Lebrasseur. Species selective predation on juvenile pink (Oncorhynchus gorbuscha) and chum salmon (O. keta) by Coho salmon (O. kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):659–668, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-085.

Burton:1985:AES

[1332] Thomas M. Burton, Richard M. Stanford, and Jon W. Allan. Acidification effects on stream biota and organic matter processing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):669–675, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-086.

Smith:1985:DAB

[1333] T. G. Smith, M. O. Hammsll, D. J. Burrage, and G. A. Sleno. Distribution and abundance of belugas, Delphinapterus leucas, and narwhals, Monodon monoceros, in the Canadian High Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):676–684, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-087.

Ramial:1985:EPM

[1334] P. S. Ramial, John W. M. Rudd, Akira Furutam, and Luying Xun. The effect of pH on methyl mercury production and decomposition in lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):685–692, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-088.

Burger:1985:CSA

[1335] Carl V. Burger, Richard L. Wilmot, and David B. Wangaard. Comparison of spawning areas and times for two runs of Chinook salmon (Oncorhynchus tshawytscha) in the Kenai River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):693–700, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-089.

Chambers:1985:DDB

[1336] Patricia A. Chambers and Jacob Kaiff. Depth distribution and biomass of submersed aquatic macrophyte communities in relation to secchi depth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):701–709, April 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-090.

Barton:1985:EPA

[1337] Bruce A. Barton, Gary S. Weirter, and Cars B. Schreck. Effect of prior acid exposure on physiological responses of juvenile rainbow trout (Salmo gairdneri) to acute handling stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):710–717, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-091.

Hilborn:1985:ASR

[1338] R. Hilborn. Apparent stock recruitment relationships in mixed stock fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):718–723, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-092.

Watras:1985:NAS

[1339] Carl J. Watras, John MacFarlane, and François M. M. Morel. Nickel accumulation by Scenedesmus and Daphnia: Food-chain transport and geochemical implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):724–730, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-093.

Bradley:1985:IPW

[1340] R. W. Bradley and J. B. Sprague. The influence of pH, water hardness, and alkalinity on the acute lethality of zinc to rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):731–736, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-094.

Horns:1985:DED

[1341] William H. Horns. Differences in early development among lake trout (Salvelinus namaycush) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):737–743, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-095.

Hill:1985:CCT

[1342] S. H. Hill, M. R. Abbott, and K. L. Denman. A computer-controlled turbidostat for the culture of planktonic algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):744–753, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-096.

Stoermer:1985:PDS

[1343] E. F. Stoermer, J. A. Wolin, C. L. Schelske, and D. J. Conley. Postsettlement diatom succession in the Bay of Quinte, Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):754–767, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-097.

Lacroix:1985:ELE

[1344] Gilles L. Lacroix, D. James Gordon, and Dana J. Johnston. Effects of low environmental pH on the survival, growth, and ionic composition of postemergent Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):768–775, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-098.

Wright:1985:SLP

[1345] Janet Wright, Charles C. Krueger, Peter F. Brussard, and Marjorie Collings Hail. Sea lamprey (Petromyzon marinus) populations in Northeastern North America: Genetic differentiation and affinities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):776–784, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-099.

Henderson:1985:VPD

[1346] M. A. Henderson and T. G. Northcote. Visual prey detection and foraging in sympatric cutthroat trout (Salmo clarki clarki) and Dolly Varden (Salvelinus malma). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):785–790, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-100.

Coble:1985:CLA

[1347] Daniel W. Coble, Gordon B. Farabee, and Richard O. Anderson. Comparative learning ability of selected fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):791–796, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-101.

Whalen:1985:NPO

[1348] S. C. Whalen and J. C. Cornwell. Nitrogen, phosphorus, and organic carbon cycling in an Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):797–808, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-102.

Cornwell:1985:SAR

[1349] Jeffrey C. Cornwell. Sediment accumulation rates in an Alaskan Arctic lake using a modified ^{210}Pb technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):809–814, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-103.

Deriso:1985:CAA

[1350] R. B. Deriso, T. J. Quinn II, and P. R. Neal. Catch-age analysis with auxiliary information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (4):815–824, April 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-104.

Mathur:1985:CSF

[1351] Dilip Mathur, William H. Bason, Edmund J. Purdy, Jr., and Carl A. Silver. A critique of the in stream flow incremental methodology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):825–831, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-105.

Riley:1985:CPC

[1352] E. T. Riley and E. E. Prepas. Comparison of the phosphorus–chlorophyll relationships in mixed and stratified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):831–835, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-106.

Tyler:1985:SAD

[1353] Albert V. Tyler, Linda L. Sebring, Margaret C. Murphy, and Lea F. Murphy. A sensitivity analysis of Deriso’s delay-difference equation using simulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):836–841, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-107.

Hodson:1985:ESD

[1354] Peter V. Hodson and Douglas J. Spry. Effect of sulfite dechlorination on the accumulation of waterborne lead by rainbow trout (Salmo gairdneri).

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):841–844, April 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-108.

Regier:1985:FRA

[1355] H. A. Regier and A. P. Grima. Fishery resource allocation: an exploratory essay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(4):845–859, April 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f85-
109.

Harrison:1985:ULS

[1356] W. G. Harrison, T. Platt, and M. R. Lewis. The utility of lightsaturation models for estimating marine primary productivity in the field: a comparison with conventional “Simulated” in situ methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):864–872, May 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-110.

Berkson:1985:UPC

[1357] J. M. Berkson and D. P. DeMaster. Use of pup counts in indexing population changes in pinnipeds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (5):873–879, May 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-111.

Sinclair:1985:AHC

[1358] M. Sinclair and T. D. Iles. Atlantic herring (Clupea harengus) distributions in the Gulf of Maine — Scotian Shelf area in relation to oceanographic features. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):880–887, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f85-
112.

Sinclair:1985:SAP

[1359] M. Sinclair, V. C. Anthony, T. D. Iles, and R. N. O’Boyle. Stock assessment problems in Atlantic herring (Clupea harengus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):888–898, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-113.

Neilson:1985:EGJ

[1360] John D. Neilson, Glen H. Geen, and Daniel Bottom. Estuarine growth of juvenile Chinook salmon (Oncorhynchus tshawytscha) as inferred from otolith microstructure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (5):899–908, May 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-114.

Larkin:1985:EFE

[1361] P. A. Larkin and D. W. McKone. An evaluation by field experiments of the McLay model of stream drift. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):909–918, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-115.

Diamond:1985:MAF

[1362] Nancy Diamond and David G. Hankin. Movements of adult female Dungeness crabs (Cancer magister) in Northern California based on tag recoveries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):919–926, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-116.

Seip:1985:MSC

[1363] Hans M. Seip, Ragnhild Seip, Peter J. Dillon, and Ed de Grosbois. Model of sulphate concentration in a small stream in the Harp Lake Catchment, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):927–937, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-117.

Prince:1985:IZS

[1364] Eric D. Prince, Dennis W. Lee, and Joaquin C. Javech. Internal zonations in sections of vertebrae from Atlantic bluefin tuna, Thunnus thynnus, and their potential use in age determination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):938–946, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-118.

Kilgour:1985:MLE

[1365] D. Marc Kilgour, Robert W. McCauley, and W. Kwain. Modeling the lethal effects of high temperature on fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):947–951, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-119.

Charles:1985:BAS

[1366] Anthony T. Charles and William J. Reed. A bioeconomic analysis of sequential fisheries: Competition, coexistence, and optimal harvest allocation between inshore and offshore fleets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):952–962, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-120.

Casey:1985:AGS

[1367] J. G. Casey, H. L. Pratt, Jr., and C. E. Stillwell. Age and growth of the sandbar shark (Carcharhinus plumbeus) from the Western North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):963–975, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-121.

Whitehead:1985:PIW

[1368] Hal Whitehead and J. E. Carscadden. Predicting inshore whale abundance — whales and capelin off the Newfoundland Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):976–981, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-122.

Hightower:1985:CCE

[1369] Joseph E. Hightower and Gary D. Grossman. Comparison of constant effort harvest policies for fish stocks with variable recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):982–988, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-123.

Winters:1985:IBS

[1370] G. H. Winters and J. P. Wheeler. Interaction between stock area, stock abundance, and catchability coefficient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):989–998, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-124.

Gosselin:1985:RSI

[1371] M. Gosselin, L. Legendre, S. Demers, and R. G. Ingram. Responses of sea-ice microalgae to climatic and fortnightly tidal energy inputs (Manitounuk Sound, Hudson Bay). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (5):999–1006, May 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-125.

Dixon:1985:EAD

[1372] D. G. Dixon and J. W. Hilton. Effects of available dietary carbohydrate and water temperature on the chronic toxicity of waterborne copper to rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):1007–1013, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-126.

Campana:1985:MFO

[1373] Steven E. Campana and John D. Neilson. Microstructure of fish otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):1014–1032, May 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-127.

Walters:1985:EDS

[1374] Carl J. Walters and Peter Cahoon. Evidence of decreasing spatial diversity in British Columbia salmon stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):1033–1037, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-128.

Statzner:1985:QCR

[1375] Bernhard Statzner and Bert Higler. Questions and comments on the river continuum concept. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (5):1038–1044, May 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-129.

Minshall:1985:DSE

[1376] G. Wayne Minshall, Kenneth W. Cummins, Robert C. Petersen, Colbert E. Cushing, Dale A. Bruns, James R. Sedell, and Robin L. Vannote. Developments in stream ecosystem theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):1045–1055, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-130.

Anonymous:1985:BRCd

[1377] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(5):1056–1059, May 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-131.

Ludwig:1985:ASM

[1378] Donald Ludwig and Carl J. Walters. Are age-structured models appropriate for catch–effort data? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (6):1066–1072, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-132.

Hansen:1985:YAY

[1379] T. J. Hansen and D. Møller. Yolk absorption, yolk sac constrictions, mortality, and growth during first feeding of Atlantic salmon (Salmo salar) incubated on astro-turf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6): 1073–1078, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-133.

Hamley:1985:FAV

[1380] John M. Hamley and Thomas P. Howley. Factors affecting variability of trapnet catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6): 1079–1087, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-134.

Peters:1985:RIB

[1381] Gregory T. Peters, Donald S. Cherry, and John Cairns, Jr. Responses of Isonychia bicolor to alkaline pH: an evaluation of survival, oxygen consumption, and chloride cell ultrastructure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1088–1095, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-135.

France:1985:RCO

[1382] R. L. France. Relationship of crayfish (Orconectes virilis) growth to population abundance and system productivity in small oligotrophic lakes in the Experimental Lakes Area, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1096–1102, June 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-136.

Wright:1985:CLH

[1383] Richard F. Wright. Chemistry of Lake Hovvatn, Norway, following liming and reacidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6): 1103–1113, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-137.

Keast:1985:PLD

[1384] Allen Keast. Planktivory in a littoral-dwelling lake fish association: Prey selection and seasonality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (6):1114–1126, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-138.

Bowles:1985:ESE

[1385] N. Dean Bowles, Hans W. Paerl, and Jane Tucker. Effective solvents and extraction periods employed in phytoplankton carotenoid and chlorophyll determinations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6): 1127–1131, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-139.

Salki:1985:IFZ

[1386] A. Salki, M. Turner, K. Patalas, J. Rudd, and D. Findlay. The influence of fish–zooplankton–phytoplankton interactions on the results of selenium toxicity experiments within large enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1132–1143, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-140.

Mason:1985:BBF

[1387] J. C. Mason and A. C. Phillips. Biology of the bathylagid fish, Leuroglossus schmidti, in the Strait of Georgia, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1144–1153, June 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-141.

Jude:1985:RCI

[1388] David J. Jude and Frank J. Tesar. Recent changes in the inshore forage fish of Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (6):1154–1157, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-142.

Frechette:1985:EFB

[1389] Marcel Fréchette and Edwin Bourget. Energy flow between the pelagic and benthic zones: Factors controlling particulate organic matter available to an intertidal mussel bed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1158–1165, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-143.

Frechette:1985:FLG

[1390] Marcel Fréchette and Edwin Bourget. Food-limited growth of Mytilus edulis L. in relation to the benthic boundary layer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1166–1170, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-144.

Timperley:1985:ONC

[1391] M. H. Timperley, R. J. Vigor-Brown, M. Kawashima, and M. Ishigami. Organic nitrogen compounds in atmospheric precipitation: Their chemistry and availability to phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1171–1177, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-145.

Rudstam:1985:PVD

[1392] Lars G. Rudstam and John J. Magnuson. Predicting the vertical distribution of fish populations: Analysis of cisco, Coregonus artedii, and yellow perch, Perca flavescens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (6):1178–1188, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-146.

Kingsley:1985:DAS

[1393] M. C. S. Kingsley, I. Stirling, and W. Calvert. The distribution and abundance of seals in the Canadian High Arctic, 1980–82. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1189–1210, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-147.

Mahon:1985:FPW

[1394] Robin Mahon and Eugene K. Balon. Fish production in warmwater streams in Poland and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1211–1215, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-148.

Ingram:1985:PRF

[1395] R. Grant Ingram, Louis Legendre, Yvan Simard, and Serge Lepage. Phytoplankton response to freshwater runoff: The diversion of the Eastmain River, James Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (6):1216–1221, June 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-149.

Cederholm:1985:RCS

[1396] C. J. Cederholm and N. P. Peterson. The retention of Coho salmon (Oncorhynchus kisutch) carcasses by organic debris in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1222–1225, June 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-150.

Anonymous:1985:LCLc

[1397] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1226–1230, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-151.

Anonymous:1985:BRCe

[1398] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(6):1231–1232, June 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-152.

Calvert:1985:WDR

[1399] Wendy Calvert and Ian Stirling. Winter distribution of ringed seals (Phoca hispida) in the Barrow Strait area, Northwest Territories, determined by underwater vocalizations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1238–1243, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-153.

Schiewe:1985:UBB

[1400] Michael H. Schiewe, Eric G. Hawk, David I. Actor, and Margaret M. Krahn. Use of a bacterial bioluminescence assay to assess toxicity of contaminated marine sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (7):1244–1248, July 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-154.

Norton:1985:EUL

[1401] Stephen A. Norton, Charles T. Hess, Geneva M. Blake, Marilyn L. Morrison, and Jill Baron. Excess unsupported ^{210}Pb in lake sediment from Rocky Mountain lakes: a groundwater effect. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1249–1254, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-155.

Tanasichuk:1985:IPE

[1402] R. W. Tanasichuk, C. Armstrong, and D. M. Ware. An improved photoelectric fish egg counter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (7):1255–1258, July 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-156.

Wood:1985:ARC

[1403] C. C. Wood. Aggregative response of common mergansers (Mergus merganser): Predicting flock size and abundance on Vancouver Island salmon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1259– 1271, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-157.

Antia:1985:IID

[1404] N. J. Antia, P. J. Harrison, D. S. Sullivan, and T. Bisalputra. Influence of the insecticide diflubenzuron (Dimilin) on the growth of marine diatoms and a harpacticoid copepod in culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1272–1277, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-158.

El-Shaarawi:1985:ECL

[1405] A. H. El-Shaarawi, S. R. Esterby, N. D. Warry, and K. W. Kuntz. Evidence of contaminant loading to Lake Ontario from the Niagara River.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1278–1289, July 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-159.

Cipriano:1985:ADI

[1406] Rocco C. Cipriano and Stephen W. Pyle. Adjuvant-dependent immunity and the agglutinin response of fishes against Aeromonas salmonicida, cause of furunculosis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (7):1290–1295, July 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-160.

Soluk:1985:MAP

[1407] Daniel A. Soluk. Macroinverfebrate abundance and production of psammophifous chironomidae in shifting sand areas of a lowland river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1296–1302, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-161.

Carignan:1985:NDL

[1408] R. Carignan. Nutrient dynamics in a littoral sediment colonized by the submersed macrophyte Myriophyllum spicatum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1303–1311, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-162.

Boyce:1985:UIU

[1409] N. P. Boyce. Ultrasound imaging used to detect cysts of Henneguya salminicola (Protozoa: Myxozoa) in the flesh of whole Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1312–1314, July 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-163.

Kelso:1985:SSP

[1410] John R. M. Kelso. Standing stock and production of fish in a cascading lake system on the Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1315–1320, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-164.

Hedrick:1985:ALP

[1411] R. P. Hedrick, J. Speas, M. L. Kent, and T. McDowell. Adenovirus-like particles associated with a disease of cultured white sturgeon, Acipenser transmontanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1321– 1325, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-165.

Anonymous:1985:LCLd

[1412] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1326–1327, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-166.

Anonymous:1985:BRCf

[1413] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(7):1328, July 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-167.

Spry:1985:IFR

[1414] Douglas J. Spry and Chris M. Wood. Ion flux rates, acid–base status, and blood gases in rainbow trout, Salmo gairdneri, exposed to toxic zinc in natural soft water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (8):1332–1341, August 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-168.

Lammens:1985:RPN

[1415] Eddy H. R. R. Lammens, Henrik W. de Nie, Jacobus Vijverberg, and Wim L. T. van Densen. Resource partitioning and niche shifts of bream (Abramis brama) and eel (Anguilla anguilla) mediated by predation of smelt (Osmerus eperlanus) on Daphnia hyalina. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1342–1351, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-169.

Ouellet:1985:TEA

[1416] Patrick Ouellet and Julian J. Dodson. Tidal exchange of anadromous rainbow smelt (Osmerus mordax) larvae between a shallow spawning tributary and the St. Lawrence Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1352–1358, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-170.

Elser:1985:NAP

[1417] James J. Elser and Bruce L. Kimmel. Nutrient availability for phytoplankton production in a multiple-impoundment series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1359–1370, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-171.

Ruppert:1985:SPM

[1418] D. Ruppert, R. L. Reish, R. B. Deriso, and R. J. Carroll. A stochastic population model for managing the Atlantic menhaden (Brevoortia tyrannus) fishery and assessing managerial risks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1371–1379, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-172.

Culver:1985:BFC

[1419] David A. Culver, Mary M. Boucherle, David J. Bean, and James W. Fletcher. Biomass of freshwater crustacean zooplankton from length– weight regressions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8): 1380–1390, August 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-173.

Millard:1985:STC

[1420] Steven P. Millard, John R. Yearsley, and Dennis P. Lettenmaier. Space– time correlation and its effects on methods for detecting aquatic ecological change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1391–1400, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-174.

Schelske:1985:HRB

[1421] Claire L. Schelske, Daniel J. Conley, and William F. Warwick. Historical relationships between phosphorus loading and biogenic silica accumulation in Bay of Quinte sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1401–1409, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-175.

Berg:1985:CTG

[1422] L. Berg and T. G. Northcote. Changes in territorial, gill-flaring, and feeding behavior in juvenile Coho salmon (Oncorhynchus kisutch) following short-term pulses of suspended sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1410–1417, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-176.

Rasmussen:1985:EDM

[1423] Joseph B. Rasmussen. Effects of density and microdetritus enrichment on the growth of chironomid larvae in a small pond. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1418–1422, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-177.

Quinn:1985:CPU

[1424] Terrance J. Quinn II. Catch-per-unit-effort: a statistical model for Pacific halibut (Hippoglossus stenolepis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1423–1429, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-178.

Schauer:1985:BBD

[1425] Paul S. Schauer and Kenneth L. Simpson. Bioaccumulation and bioconversion of dietary labeled fatty acids in Artemia and winter flounder (Pseudopleuronectes americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1430–1438, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-179.

daSilva:1985:LUO

[1426] Jack da Silva and John D. Neilson. Limitations of using otoliths recovered in scats to estimate prey consumption in seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1439–1442, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-180.

Anonymous:1985:BRCg

[1427] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1443–1446, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-181.

Anonymous:1985:CRPb

[1428] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(8):1447, August 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f85-
182.

Penney:1985:GHL

[1429] Randy W. Penney and Geoffrey T. Evans. Growth histories of larval redfish (Sebastes spp.) on an offshore Atlantic fishing bank determined by otolith increment analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1452–1464, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-183.

Krantzberg:1985:BMM

[1430] G. Krantzberg and P. M. Stokes. Benthic macroinvertebrates modify copper and zinc partitioning in freshwater–sediment microcosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1465–1473, September 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-184.

Beacham:1985:BGSb

[1431] Terry D. Beacham, Ruth E. Withler, and Allan P. Gould. Biochemical genetic stock identification of pink salmon (Oncorhynchus gorbuscha) in Southern British Columbia and Puget Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1474–1483, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-185.

Paranjape:1985:MMN

[1432] M. A. Paranjape, R. J. Conover, G. C. Harding, and N. J. Prouse. Microand macrozooplankton on the Nova Scotian Shelf in the prespring bloom period: a comparison of their potential resource utilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1484–1492, September 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-186.

Fulthorpe:1985:HOC

[1433] Roberta R. Fulthorpe and J. E. Paloheimo. Hypolimnetic oxygen consumption in small lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9): 1493–1500, September 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-187.

Stevenson:1985:PEA

[1434] R. J. Stevenson, R. Singer, D. A. Roberts, and C. W. Boylen. Patterns of epipelic algal abundance with depth, trophic status, and acidity in poorly buffered New Hampshire lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1501–1512, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-188.

Martin:1985:IST

[1435] J. H. A. Martin and K. A. Mitchell. Influence of sea temperature upon the numbers of grilse and multi-sea-winter Atlantic salmon (Salmo salar) caught in the vicinity of the River Dee (Aberdeenshire). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1513–1521, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-189.

Bernatchez:1985:ITC

[1436] Louis Bernatchez and Julian J. Dodson. Influence of temperature and current speed on the swimming capacity of lake whitefish (Coregonus clupeaformis) and cisco (C. artedii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1522–1529, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-190.

Morin:1985:VDE

[1437] Antoine Morin. Variability of density estimates and the optimization of sampling programs for stream benthos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1530–1534, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-191.

Sameoto:1985:RBA

[1438] D. Sameoto, N. A. Cochrane, and A. W. Herman. Response of biological acoustic backscattering to ships’ lights. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1535–1543, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-192.

Clark:1985:LSE

[1439] Karen L. Clark and Bruce D. Lazerte. A laboratory study of the effects of aluminum and pH on amphibian eggs and tadpoles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1544–1551, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-193.

Laurs:1985:FIF

[1440] R. M. Laurs, R. Nishimoto, and J. A. Wetherall. Frequency of increment formation on sagittae of North Pacific albacore (Thunnus alalunga). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1552–1555, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-194.

Pocklington:1985:OMG

[1441] Roger Pocklington. Organic matter in the Gulf of St. Lawrence in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1556–1561, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-195.

Anonymous:1985:BRCh

[1442] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(9):1562–1563, September 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-196.

Downing:1985:QEE

[1443] John A. Downing and Hélène Cyr. Quantitative estimation of epiphytic invertebrate populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1570–1579, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-197.

Mork:1985:GVA

[1444] Jarle Mork, Nils Ryman, Gunnar Ståhl, Fred Utter, and Gunnar Sundnes. Genetic variation in Atlantic cod (Gadus morhua) throughout its range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1580–1587, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f85-
198.

Bierhuizen:1985:RBN

[1445] J. F. H. Bierhuizen and E. E. Prepas. Relationship between nutrients, dominant ions, and phytoplankton standing crop in prairie saline lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1588–1594, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-199.

Peterman:1985:PIV

[1446] Randall M. Peterman. Patterns of interannual variation in age at maturity of sockeye salmon (Oncorhynchus nerka) in Alaska and British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1595–1607, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-200.

Henderson:1985:EAW

[1447] Bryan A. Henderson and Edward H. Brown, Jr. Effects of abundance and water temperature on recruitment and growth of alewife (Alosa pseudoharengus) near South Bay, Lake Huron, 1954–82. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1608–1613, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-201.

Westrheim:1985:RFP

[1448] S. J. Westrheim and R. P. Foucher. Relative fishing power for Canadian trawlers landing Pacific cod (Gadus macrocephalus) and important shelf cohabitants from major offshore areas of Western Canada, 1960–81. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1614–1626, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-202.

Malison:1985:HGP

[1449] Jeffrey A. Malison, Cody D. Best, Terrence B. Kayes, Clyde H. Amundson, and Bernard C. Wentworth. Hormonal growth promotion and evidence for a size-related difference in response to Estradiol-17β in yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10): 1627–1633, October 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-203.

Linsey:1985:CSW

[1450] G. A. Linsey and D. C. Lasenby. Comparison of summer and winter oxygen consumption rates in a temperate dimictic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1634–1639, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-204.

Crecco:1985:EBA

[1451] Victor A. Crecco and Thomas F. Savoy. Effects of biotic and abiotic factors on growth and relative survival of young American shad, Alosa sapidissima, in the Connecticut River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1640–1648, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-205.

Crecco:1985:DDC

[1452] Victor A. Crecco and Thomas F. Savoy. Density-dependent catchability and its potential causes and consequences on Connecticut River American shad, Alosa sapidissima. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1649–1657, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-206.

Doving:1985:ROB

[1453] Kjell B. Døving, Håkan Westerberg, and Peter B. Johnsen. Role of olfaction in the behavioral and neuronal responses of Atlantic salmon, Salmo salar, to hydrographic stratification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1658–1667, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-207.

Wilzbach:1985:RRF

[1454] Margaret A. Wilzbach. Relative roles of food abundance and cover in determining the habitat distribution of stream-dwelling cutthroat trout (Salmo clarki). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1668– 1672, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f85-208.

Misra:1985:QDA

[1455] R. K. Misra. Quadratic discriminant analysis with covariance for stock delineation and population differentiation: a study of beaked redfishes (Sebastes mentella and S. fasciatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1672–1676, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-209.

Allard:1985:STE

[1456] Martine Allard and Guy Moreau. Short-term effect on the metabolism of lotic benthic communities following experimental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1676–1680, October 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-210.

Gilbertson:1985:NLH

[1457] Michael Gilbertson. The Niagara labyrinth — the human ecology of producing organochlorine chemicals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(10):1681–1692, October 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-211.

Craig:1985:ISS

[1458] P. C. Craig. Identification of sockeye salmon (Oncorhynchus nerka) stocks in the Stikine River based on egg size measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1696–1701, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-212.

Abbott:1985:PAA

[1459] Jeremy C. Abbott and Lawrence M. Dill. Patterns of aggressive attack in juvenile steelhead trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1702–1706, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-213.

Effler:1985:PLA

[1460] Steven W. Effler, Gary C. Schafran, and Charles T. Driscoll. Partitioning light attenuation in an acidic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1707–1711, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-214.

Estep:1985:SCN

[1461] Marilyn L. F. Estep and Steven Vigg. Stable carbon and nitrogen isotope tracers of trophic dynamics in natural populations and fisheries of the Lahontan Lake System, Nevada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1712–1719, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-215.

Cooper:1985:PSF

[1462] Scott D. Cooper, Daniel W. Smith, and James R. Bence. Prey selection by freshwater predators with different foraging strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1720–1732, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-216.

Stanhope:1985:GPE

[1463] M. J. Stanhope and C. D. Levsngs. Growth and production of Eogammarus confervicolus (Amphipoda: Anisogammaridae) at a log storage site and in areas of undisturbed habitat within the Squamish Estuary, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1733– 1740, November 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-217.

Havas:1985:ABT

[1464] Magda Havas. Aluminum bioaccumulation and toxicity to Daphnia magna in soft water at low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1741–1748, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-218.

Price:1985:IIS

[1465] Edmund E. Price and Michael C. Swift. Inter- and intra-specific variability in the response of zooplankton to acid stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1749–1754, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-219.

Beacham:1985:EFS

[1466] Terry D. Beacham and Clyde B. Murray. Effect of female size, egg size, and water temperature on developmental biology of chum salmon (Oncorhynchus keta) from the Nitinat River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1755–1765, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-220.

Lawson:1985:EYY

[1467] Timothy A. Lawson and Ray Hilborn. Equilibrium yields and yield isopleths from a general age-structured model of harvested populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1766–1771, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-221.

Moon:1985:FHM

[1468] T. W. Moon, P. J. Walsh, and T. P. Mommsen. Fish hepatocytes: a model metabolic system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11): 1772–1782, November 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-222.

Wehr:1985:SRB

[1469] John D. Wehr and Lewis M. Brown. Selenium requirement of a bloomforming planktonic alga from softwater and acidified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1783–1788, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-223.

Bergmann:1985:SMM

[1470] Martin A. Bergmann and Harold E. Welch. Spring meltwater mixing in small Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11): 1789–1798, November 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-224.

Beamish:1985:ADS

[1471] Richard J. Beamish and Gordon A. McFarlane. Annulus development on the second dorsal spine of the spiny dogfish (Squalus acanthias) and its validity for age determination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1799–1805, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-225.

Schweigert:1985:OSD

[1472] J. F. Schweigert, C. W. Haegele, and M. Stocker. Optimizing sampling design for herring spawn surveys in the Strait of Georgia, B.C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1806–1814, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-226.

Schwalme:1985:SVS

[1473] Karl Schwalme, William C. Mackay, and Dieter Lindner. Suitability of vertical slot and Denil fishways for passing north-temperate, nonsalmonid fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1815– 1822, November 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-227.

Knutsen:1985:DBO

[1474] H. Knutsen, E. Moksnes, and N. B. Vogt. Distinguishing between oneday-old cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) eggs by gas chromatography and SIMCA pattern recognition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1823–1826, November 1985. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f85-228.

Gardner:1985:SPL

[1475] Wayne S. Gardner, Thomas F. Nalepa, William A. Frez, Elizabeth A. Cichocki, and Peter F. Landrum. Seasonal patterns in lipid content of Lake Michigan macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1827–1832, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-229.

Hilborn:1985:SCO

[1476] Ray Hilborn. Simplified calculation of optimum spawning stock size from Ricker’s stock recruitment curve. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1833–1834, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-230.

Turner:1985:UVC

[1477] L. J. Turner and W. C. Mackay. Use of visual census for estimating population size in northern pike (Esox lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1835–1840, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-231.

Frank:1985:ROD

[1478] K. T. Frank and W. C. Leggett. Reciprocal oscillations in densities of larval fish and potential predators: a reflection of present or past predation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1841–1849, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-232.

Anonymous:1985:BRCi

[1479] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(11):1852, November 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f85-233.

Downing:1985:ESB

[1480] John A. Downing and M. Robin Anderson. Estimating the standing biomass of aquatic macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1860–1869, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-234.

Rhodes:1985:IEC

[1481] Linda Rhodes, Edmundo Casillas, Barbara McKnight, William Gronlund, Mark Myers, O. Paul Olson, and Bruce McCain. Interactive effects of cadmium, polychlorinated biphenyls, and fuel oil on experimentally exposed English sole (Parophrys vetulus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1870–1880, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-235.

Warwick:1985:MAC

[1482] W. F. Warwick. Morphological abnormalities in Chironomidae (Diptera) larvae as measures of toxic stress in freshwater ecosystems: Indexing antennal deformities in Chironomus Meigen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1881–1914, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-236.

Paszkowski:1985:FBH

[1483] Cynthia A. Paszkowski and Bori L. Olla. Foraging behavior of hatcheryproduced Coho salmon (Oncorhynchus kisutch) smolts on live prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1915–1921, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-237.

Irvine:1985:ESF

[1484] J. R. Irvine. Effects of successive flow perturbations on stream invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1922–1927, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-238.

Jessop:1985:IMC

[1485] B. M. Jessop. Influence of mesh composition, velocity, and run time on the catch and length composition of juvenile alewives (Alosa pseudoharengus) and blueback herring (A. aestivalis) collected by pushnet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1928–1939, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-239.

Smith:1985:BCE

[1486] Daniel Wilkins Smith. Biological control of excessive phytoplankton growth and the enhancement of aquacultural production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1940–1945, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-240.

Nalewajko:1985:EMA

[1487] C. Nalewajko and B. Paul. Effects of manipulations of aluminum concentrations and pH on phosphate uptake and photosynthesis of planktonic communities in two Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1946–1953, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-241.

Waddy:1985:FER

[1488] S. L. Waddy and D. E. Aiken. Fertilization and egg retention in artificially inseminated female American lobsters, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1954–1956, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-242.

McGladdery:1985:PPU

[1489] S. E. McGladdery and M. D. B. Burt. Potential of parasites for use as biological indicators of migration, feeding, and spawning behavior of Northwestern Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1957–1968, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-243.

Arfi:1985:VII

[1490] Robert Arfi. Variabilité interannuelle d’un indice d’intensité des remontées d’eau dans le secteur du cap Blanc (Mauritanie). (French) [Interannual variability of an index of upwelling intensity in the Cape Blanc (Mauritania) sector]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12): 1969–1978, December 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-244.

Pierson:1985:ISD

[1491] D. C. Pierson and C. H. Taylor. Influence of snowcover development and ground freezing on cation loss from a wetland watershed during spring runoff. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1979–1985, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-245.

Bower:1985:PUC

[1492] S. M. Bower, L. Margolis, and R. J. MacKay. Potential usefulness of chlorine for controlling Pacific salmon leeches, Piscicola salmositica, in hatcheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1986–1993, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-246.

Johnston:1985:CAL

[1493] C. E. Johnston and J. C. Cheverie. Comparative analysis of lonoregulation in rainbow trout (Salmo gairdneri) of different sizes following rapid and slow salinity adaptation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):1994–2003, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f85-247.

Neville:1985:PRJ

[1494] C. M. Neville. Physiological response of juvenile rainbow trout, Salmo gairdneri, to acid and aluminum — prediction of field responses from laboratory data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):2004– 2019, December 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-248.

Taylor:1985:VBM

[1495] Eric B. Taylor and J. D. McPhail. Variation in body morphology among British Columbia populations of Coho salmon, Oncorhynchus kisutch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):2020–2028, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-249.

Taylor:1985:VBP

[1496] Eric B. Taylor and J. D. McPhail. Variation in burst and prolonged swimming performance among British Columbia populations of Coho salmon, Oncorhynchus kisutch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12): 2029–2033, December 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-250.

Campbel:1985:ATM

[1497] P. G. C. Campbel and P. M. Stokes. Acidification and toxicity of metals to aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12): 2034–2049, December 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-251.

Townsend:1985:RFE

[1498] Ralph E. Townsend. The right to fish as an external benefit of open access. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):2050–2053, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f85-252.

Lessman:1985:RFW

[1499] Charles A. Lessman, Soosamma Kavumpurath, and Hamid R. Habibi. Removal of follicle wall components from ovarian oocytes of the brook trout, Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42 (12):2053–2058, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f85-253.

Kimura:1985:CSR

[1500] Daniel K. Kimura. Changes to stock reduction analysis indicated by Schnute’s general theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12): 2059–2060, December 1985. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f85-254.

Miller:1985:SSU

[1501] Robert J. Miller. Seaweeds, sea urchins, and lobsters: a reappraisal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):2061–2072, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f85-255.

Anonymous:1985:BRCj

[1502] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 42(12):2073–2077, December 1985. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f85-256.

Ryan:1986:LUW

[1503] P. M. Ryan. Lake use by wild anadromous Atlantic salmon, Salmo salar, as an index of subsequent adult abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):2–11, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-001.

Sweeney:1986:ETF

[1504] Bernard W. Sweeney, Robin L. Vannote, and Peter J. Dodds. Effects of temperature and food quality on growth and development of a mayfly, Leptophlebia intermedia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):12–18, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-002.

LeGrow:1986:IDP

[1505] S. M. LeGrow and F. W. H. Beamish. Influence of dietary protein and lipid on apparent heat increment of rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):19–25, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-003.

Malison:1986:SDU

[1506] Jeffrey A. Malison, Terrence B. Kayes, Cody D. Best, Clyde H. Amundson, and Bernard C. Wentworth. Sexual differentiation and use of hormones to control sex in yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):26–35, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-004.

Shotts:1986:PEE

[1507] E. B. Shotts, V. S. Blazer, and W. D. Waltman. Pathogenesis of experimental Edwardsiella ictaluri infections in channel catfish (Icta lurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):36–42, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
005.

Belanger:1986:UCA

[1508] Scott E. Belanger, Donald S. Cherry, and John Cairns, Jr. Uptake of chrysotile asbestos fibers alters growth and reproduction of Asiatic clams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):43–52, January 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-006.

Chandler:1986:HDC

[1509] G. Thomas Chandler. High-density culture of meiobenthic harpacticoid copepods within a muddy sediment substrate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):53–59, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-007.

Santschi:1986:RRT

[1510] Peter H. Santschi, Urs P. Nyffeler, Robert F. Anderson, Sherry L. Schiff, Patricia O’Hara, and Raymond H. Hesslein. Response of radioactive trace metals to acid–base titrations in controlled experimental ecosystems: Evaluation of transport parameters for application to whole-lake radiotracer experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):60–77, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-008.

Campbell:1986:FTP

[1511] Daniel E. Campbell and J. S. Wroblewski. Fundy tidal power development and potential fish production in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):78–89, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-009.

Murawski:1986:OEA

[1512] S. A. Murawski and J. T. Finn. Optimal effort allocation among competing mixed-species fisheries, subject to fishing mortality constraints. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):90–100, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-010.

Lovejoy:1986:BOA

[1513] William S. Lovejoy. Bounds on the optimal age-at-first-capture for stochastic, age-structured fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):101–107, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-011.

Welch:1986:ISR

[1514] David W. Welch. Identifying the stock–recruitment relationship for agestructured populations using time-invariant matched linear filters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):108–123, January 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-012.

Rahel:1986:BIF

[1515] Frank J. Rahel. Biogeographic influences on fish species composition of Northern Wisconsin lakes with applications for lake acidification studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):124–134, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-013.

Moussalli:1986:OSS

[1516] Elie Moussalli and Ray Hilborn. Optimal stock size and harvest rate in multistage life history models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):135–141, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-014.

Lockyer:1986:BFC

[1517] Christina Lockyer. Body fat condition in Northeast Atlantic fin whales, Balaenoptera physalus, and its relationship with reproduction and food resource. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):142–147, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-015.

Smith:1986:LNE

[1518] Val H. Smith. Light and nutrient effects on the relative biomass of bluegreen algae in lake phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):148–153, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-016.

Evans:1986:RMD

[1519] Marlene S. Evans. Recent major declines in zooplankton populations in the inshore region of Lake Michigan: Probable causes and implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):154–159, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-017.

Bartell:1986:IPP

[1520] S. M. Bartell, J. E. Breck, R. H. Gardner, and A. L. Brenkert. Individual parameter perturbation and error analysis of fish bioenergetics models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):160–168, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-018.

Chen:1986:EAM

[1521] Thomas T. Chen, P. C. Reid, R. Van Beneden, and R. A. Sonstegard. Effect of Aroclor 1254 and Mirex on estradiol-lnduced vitellogenin production in juvenile rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):169–173, January 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-019.

Reed:1986:ACE

[1522] William J. Reed. Analyzing catch–effort data allowing for randomness in the catching process. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):174–186, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-020.

Rose:1986:GDS

[1523] G. A. Rose. Growth decline in subyearling brook trout (Salvelinus fontinalis) after emergence of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):187–193, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-021.

Tonn:1986:SLP

[1524] William M. Tonn and Cynthia A. Paszkowski. Size-limited predation, winterkill, and the organization of Umbra–Perca fish assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):194–202, January 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-022.

Pelletier:1986:MBD

[1525] Emilien Pelletier. Modification de la bioaccumulation du sélénium chez Mytilus edulis en présence du mercure organique et inorganique. (French) [Modification of the bioaccumulation of selenium in Mytilus edulis in the presence of organic and inorganic mercury]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):203–210, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-023.

Frank:1986:ESC

[1526] Kenneth T. Frank. Ecological significance of the ctenophore Pleurobrachia pileus off Southwestern Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):211–222, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-024.

Magnhagen:1986:ADI

[1527] Carin Magnhagen. Activity differences influencing food selection in the marine fish Pomatoschistus microps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):223–227, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-025.

Meade:1986:HCC

[1528] James W. Meade and Roger Lee Herman. Histological changes in cultured lake trout, Salvelinus namaycush, subjected to cumulative loading in a water reuse system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):228–231, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-026.

Lalande:1986:ASS

[1529] Hélène Lalande and William H. Hendershot. Aluminum speciation in some synthetic systems: Comparison of the fast-oxine, pH 5.0 extraction and dialysis methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):231–234, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-027.

Fahnenstiel:1986:IPL

[1530] Gary L. Fahnenstiel, Linda Sicko-Goad, Donald Scavia, and Eugene F. Stoermer. Importance of picoplankton in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):235–240, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-028.

Renyard:1986:SAP

[1531] T. S. Renyard and Ray Hilborn. Sports angler preferences for alternative regulatory methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (1):240–242, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-029.

Orr:1986:AIC

[1532] P. L. Orr, R. W. Bradley, J. B. Sprague, and N. J. Hutchinson. Acclimation-induced change in toxicity of aluminum to rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1): 243–246, January 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-030.

Anonymous:1986:BRCa

[1533] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(1):247, January 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-031.

Beacham:1986:CDB

[1534] Terry D. Beacham and Clyde B. Murray. Comparative developmental biology of chum salmon (Oncorhynchus keta) from the Fraser River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):252–262, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-032.

Shacklock:1986:WSG

[1535] Peter F. Shacklock and James S. Craigie. Winter survival and growth of Chondrus crispus in onshore culture tanks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):263–268, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-033.

Peterman:1986:EPI

[1536] Randall M. Peterman, Michael J. Bradford, and Judith L. Anderson. Environmental and parental influences on age at maturity in sockeye salmon (Oncorhynchus nerka) from the Fraser River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):269–274, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-034.

Tam:1986:ECE

[1537] W. H. Tam and P. D. Payson. Effects of chronic exposure to sublethal pH on growth, egg production, and ovulation in brook trout, Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):275–280, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-035.

Bielak:1986:CMW

[1538] Alex T. Bielak and Geoffrey Power. Changes in mean weight, seaage composition, and catch-per-unit-effort of Atlantic salmon (Salmo salar) angled in the Godbout River, Quebec, 1859–1983. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):281–287, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-036.

Stephenson:1986:LAL

[1539] M. Stephenson and G. L. Mackie. Lake acidification as a limiting factor in the distribution of the freshwater amphipod Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):288–292, February 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-037.

Tarapchak:1986:IPP

[1540] Stephen J. Tarapchak and Czesia Nalewajko. Introduction: Phosphorus– Plankton Dynamics Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):293–301, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-038.

Francko:1986:EPC

[1541] David A. Francko. Epilimnetic phosphorus cycling: Influence of humic materials and iron on coexisting major mechanisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):302–310, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-039.

Currie:1986:DAB

[1542] David J. Currie, Ellen Bentzen, and Jacob Kalff. Does algal–bacterial phosphorus partitioning vary among lakes? A comparative study of orthophosphate uptake and alkaline phosphatase activity in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):311–318, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-040.

Tarapchak:1986:PUM

[1543] Stephen J. Tarapchak and Lynn R. Herche. Phosphate uptake by microorganisms in lake water: Deviations from simple Michaelis–Menten kinetics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):319–328, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-041.

Nalewajko:1986:LHP

[1544] C. Nalewajko, B. Paul, K. Lee, and H. Shear. Light history, phosphorus status, and the occurrence of light stimulation or inhibition of phosphate uptake in Lake Superior phytoplankton and bacteria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):329–335, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-042.

DePinto:1986:MRP

[1545] Joseph V. De Pinto, Thomas C. Young, James S. Bonner, and Paul W. Rodgers. Microbial recycling of phytopiankton phosphorus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):336–342, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-043.

Heath:1986:DOP

[1546] Robert T. Heath. Dissolved organic phosphorus compounds: Do they satisfy planktonic phosphate demand in summer? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):343–350, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-044.

Kilham:1986:DLM

[1547] Susan Soltau Kilham. Dynamics of Lake Michigan natural phytoplankton communities in continuous cultures along a Si:P loading gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):351–360, February 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-045.

Bowers:1986:PRP

[1548] James A. Bowers. Phosphorus regeneration by the predatory copepod Diacyclops thomasi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2): 361–365, February 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-046.

Levine:1986:RSP

[1549] S. N. Levine, M. P. Stainton, and D. W. Schindler. A radiotracer study of phosphorus cycling in a eutrophic Canadian Shield lake, Lake 227, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2): 366–378, February 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-047.

Auer:1986:ICN

[1550] Martin T. Auer, Mark S. Kieser, and Raymond P. Canale. Identification of critical nutrient levels through field verification of models for phosphorus and phytoplankton growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):379–388, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-048.

Sakshaug:1986:NSP

[1551] Egil Sakshaug and Yngvar Olsen. Nutrient status of phytoplankton blooms in Norwegian waters and algal strategies for nutrient competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):389–396, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-049.

DElia:1986:NES

[1552] Christopher F. D’Elia, James G. Sanders, and Walter R. Boynton. Nutrient enrichment studies in a coastal plain estuary: Phytoplankton growth in large-scale, continuous cultures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):397–406, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-050.

Schelske:1986:PES

[1553] Claire L. Schelske, Eugene F. Stoermer, Gary L. Fahnenstiel, and Mark Haibach. Phosphorus enrichment, silica utilization, and biogeochemical silica depletion in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):407–415, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-051.

Tarapchak:1986:SPP

[1554] Stephen J. Tarapchak and Czesia Nalewajko. Synopsis: Phosphorus– Plankton Dynamics Symposium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):416–419, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-052.

Nero:1986:ZSA

[1555] Redwood W. Nero and W. Gary Sprules. Zooplankton species abundance and biomass in relation to occurrence of Mysis relicta (Malacostraca: Mysidacea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2): 420–434, February 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-053.

Scavia:1986:ISP

[1556] Donald Scavia, Gary L. Fahnenstiel, Marlene S. Evans, David J. Jude, and John T. Lehman. Influence of salmonine predation and weather on long-term water quality trends in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):435–443, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-054.

Carter:1986:LZA

[1557] J. C. H. Carter, W. D. Taylor, R. Chengalath, and D. A. Scruton. Limnetic zooplankton assemblages in Atlantic Canada with special reference to acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2): 444–456, February 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-055.

Crecco:1986:EDD

[1558] Victor Crecco, Thomas Savoy, and Walter Whitworth. Effects of densitydependent and climatic factors on American shad, Alosa sapidissima, recruitment: a predictive approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):457–463, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-056.

Mysak:1986:NNI

[1559] Lawrence A. Mysak. El Niño, interannual variability and fisheries in the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (2):464–497, February 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-057.

Hamilton:1986:PES

[1560] Kevin Hamilton and Lawrence A. Mysak. Possible effects of the Sitka Eddy on sockeye (Oncorhynchus nerka) and pink salmon (Oncorhynchus gorbuscha) migration off Southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):498–504, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-058.

Anonymous:1986:BRCb

[1561] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):505–506, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f86-059.

Anonymous:1986:CRPa

[1562] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(2):509, February 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-060.

Conover:1986:TSP

[1563] David O. Conover and Mark H. Fleisher. Temperature-sensitive period of sex determination in the Atlantic silverside, Menidia menidia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):514–520, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-061.

Markiw:1986:SWD

[1564] Maria E. Markiw. Salmonid whirling disease: Dynamics of experimental production of the infective stage — the triactinomyxon spore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):521–526, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-062.

Nickelson:1986:IUO

[1565] Thomas E. Nickelson. Influences of upwelling, ocean temperature, and smolt abundance on marine survival of Coho salmon (Oncorhynchus kisutch) in the Oregon production area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):527–535, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-063.

Levy:1986:BLP

[1566] E. M. Levy. Background levels of petroleum residues in the waters and surficial bottom sediments of the Labrador Shelf and Hudson Strait/ Foxe Basin regions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (3):536–547, March 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-064.

Hinch:1986:GLR

[1567] Scott G. Hinch, Robert C. Bailey, and Roger H. Green. Growth of Lampsilis radiata (Bivalvia: Unionidae) in sand and mud: a reciprocal transplant experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (3):548–552, March 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-065.

Roff:1986:FAP

[1568] Derek Roff and W. Don Bowen. Further analysis of population trends in the Northwest Atlantic harp seal (Phoca groenlandica) from 1967 to 1985. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):553–564, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-066.

Taylor:1986:CRA

[1569] Eric B. Taylor and P. A. Larkin. Current response and agonistic behavior in newly emerged fry of Chinook salmon, Oncorhynchus tshawytscha, from ocean- and stream-type populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):565–573, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-067.

Nurnberg:1986:IPL

[1570] Gertrud K. Nürnberg, Margo Shaw, Peter J. Dillon, and Don J. Mc-Queen. Internal phosphorus load in an oligotrophic Precambrian Shield lake with an anoxic hypolimnion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):574–580, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-068.

Hargreaves:1986:SSC

[1571] N. Brent Hargreaves and Robin J. LeBrasseur. Size selectivity of Coho (Oncorhynchus kisutch) preying on juvenile chum salmon (O. keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):581–586, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-069.

Roff:1986:DAL

[1572] J. C. Roff, L. P. Fanning, and A. B. Stasko. Distribution and association of larval crabs (Decapoda: Brachyura) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):587–599, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-070.

Graney:1986:FAA

[1573] R. L. Graney, T. J. Keilty, and J. P. Giesy. Free amino acid pools of five species of freshwater oligochaetes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):600–607, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-071.

Donald:1986:SLT

[1574] D. B. Donald and D. J. Alger. Stunted lake trout (Salvelinus namaycush) from the Rocky Mountains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (3):608–612, March 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-072.

Loadman:1986:SCL

[1575] N. L. Loadman, G. E. E. Moodie, and J. A. Mathias. Significance of cannibalism in larval walleye (Stizostedion vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):613–618, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-073.

Lamers:1986:NAA

[1576] C. H. J. Lamers and W. B. Van Muiswinkel. Natural and acquired agglutinins to Aeromonas hydrophila in carp (Cyprinus carpio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):619–624, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-074.

Rustad:1986:MSC

[1577] Steinar Rustad, Nils Christophersen, Hans M. Seip, and Peter J. Dillon. Model for streamwater chemistry of a tributary to Harp Lake, Ontario.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):625–633, March 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-075.

Noakes:1986:QCB

[1578] D. Noakes. Quantifying changes in British Columbia Dungeness crab (Cancer magister) landings using intervention analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):634–639, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-076.

Melvin:1986:FAS

[1579] Gary D. Melvin, Michael J. Dadswell, and James D. Martin. Fidelity of American shad, Alosa sapidissima (Gupeidae), to its river of previous spawning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):640–646, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-077.

Hutchinson:1986:TTM

[1580] N. J. Hutchinson and J. B. Sprague. Toxicity of trace metal mixtures to American flagfish (Jordanella floridae) in soft, acidic water and implications for cultural acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):647–655, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-078.

Middlebrook:1986:CME

[1581] Ken Middlebrook and John C. Roff. Comparison of methods for estimating annual productivity of the copepods Acartia hudsonica and Eurytemora herdmani in Passamaquoddy Bay, New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):656–664, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-079.

Zettler:1986:ZCS

[1582] Erik R. Zettler and John C. H. Carter. Zooplankton community and species responses to a natural turbidity gradient in Lake Temiskaming, Ontario–Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):665–673, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-080.

Rosenberg:1986:AEA

[1583] Andrew A. Rosenberg and Roger W. Doyle. Analysing the effect of age structure on stock–recruitment relationships in herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):674–679, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-081.

Birt:1986:PST

[1584] Tim P. Birt and John M. Green. Parr–smolt transformation in female and sexually mature male anadromous and nonanadromous Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):680–686, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-082.

Harrison:1986:EPC

[1585] Gail I. Harrison, Peter G. C. Campbell, and André Tessier. Effects of pH changes on zinc uptake by Chlamydomonas variabilis grown in batch culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):687–693, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-083.

Sanders:1986:DIE

[1586] James G. Sanders. Direct and indirect effects of arsenic on the survival and fecundity of estuarine zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):694–699, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-084.

Peckarsky:1986:CNS

[1587] Barbara L. Peckarsky. Colonization of natural substrates by stream benthos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):700–709, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-085.

Walczak:1986:EMA

[1588] Barbara Z. Walczak, U. Theodore Hammer, and P. Ming Huang. Ecophysiology and mercury accumulation of rainbow trout (Salmo gairdneri) when exposed to mercury in various concentrations of chloride. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):710–714, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-086.

Brown:1986:ITC

[1589] S. B. Brown, R. E. Evans, and Toshiaki J. Hara. Interrenal, thyroidal, carbohydrate, and electrolyte responses in rainbow trout (Salmo gairdneri) during recovery from the effects of acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):714–718, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-087.

Volk:1986:UCO

[1590] Eric C. Volk. Use of calcareous otic elements (statoliths) to determine age of sea lamprey ammocoetes (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):718–722, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-088.

Anonymous:1986:BRCc

[1591] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(3):723–725, March 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-089.

Hutchings:1986:LMJ

[1592] Jeffrey A. Hutchings. Lakeward migrations by juvenile Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):732–741, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-

Ratkowsky:1986:SPA

[1593] D. A. Ratkowsky. Statistical properties of alternative parameterizations of the von Bertalanffy growth curve. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):742–747, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-091.

Overholtz:1986:RVI

[1594] William J. Overholtz, Michael P. Sissenwine, and Stephen H. Clark. Recruitment variability and its implication for managing and rebuilding the Georges Bank haddock (Melanogrammus aeglefinus) stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):748–753, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-092.

Johnson:1986:TST

[1595] M. G. Johnson, L. R. Culp, and S. E. George. Temporal and spatial trends in metal loadings to sediments of the Turkey Lakes, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):754–762, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-093.

Webb:1986:EBF

[1596] P. W. Webb. Effect of body form and response threshold on the vulnerability of four species of teleost prey attacked by largemouth bass (Micropterus salmoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):763–771, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-094.

Campbell:1986:ELA

[1597] Alan Campbell and John Brattey. Egg loss from the American lobster, Homarus americanus, in relation to Nemertean, Pseudocarcinonemertes homari, infestation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):772–780, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-095.

Kenchington:1986:MCT

[1598] T. J. Kenchington. Morphological comparison of two Northwest Atlantic redfishes, Sebastes fasciatus and S. mentella, and techniques for their identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):781–787, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
096.

Yan:1986:EPC

[1599] N. D. Yan. Empirical prediction of crustacean zooplankton biomass in nutrient-poor Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):788–796, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-097.

Shortreed:1986:TSS

[1600] Ken S. Shortreed and John G. Stockner. Trophic status of 19 subarctic lakes in the Yukon Territory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):797–805, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-098.

Bendell-Young:1986:ACW

[1601] Leah I. Bendell-Young, Harold H. Harvey, and Jeffrey F. Young. Accumulation of cadmium by white suckers (Catostomus commersoni) in relation to fish growth and lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):806–811, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-099.

Loftus:1986:PLL

[1602] D. H. Loftus and P. F. Hulsman. Predation on larval lake whitefish (Coregonus clupeaformis) and lake herring (C. artedii) by adult rainbow smelt (Osmerus mordax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):812–818, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-100.

Madenjian:1986:IAP

[1603] Charles P. Madenjian, David J. Jude, and Frank J. Tesar. Intervention analysis of power plant impact on fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):819–829, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-101.

Walters:1986:IBP

[1604] Carl J. Walters, M. Stocker, A. V. Tyler, and S. J. Westrheim. Interaction between Pacific cod (Gadus macrocephalus) and herring (Clupea harengus pallasi) in the Hecate Strait, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):830–837, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-102.

Johnson:1986:WSC

[1605] David F. Johnson, Louis W. Botsford, Richard D. Methot, Jr., and Thomas C. Wainwright. Wind stress and cycles in Dungeness crab (Cancer magister) catch off California, Oregon, and Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):838–845, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-103.

Campbell:1986:EFR

[1606] C. E. Campbell and E. E. Prepas. Evaluation of factors related to the unusually low chlorophyll levels in prairie saline lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):846–854, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-104.

Haymes:1986:EAA

[1607] Gerard T. Haymes and Paul H. Patrick. Exclusion of adult alewife, Alosa pseudoharengus using low-frequency sound for application at water intakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):855–862, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-105.

Evans:1986:CSS

[1608] Geoffrey T. Evans and J. Brian Dempson. Calculating the sensitivity of a salmonid population model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):863–868, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-106.

Cunningham:1986:ILP

[1609] G. L. Cunningham and B. J. Shuter. Interaction of low pH and starvation on body weight and composition of young-of-year smallmouth bass (Micropterus dolomieui). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (4):869–876, April 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-107.

Cornett:1986:TBW

[1610] R. Jack Cornett and Ivan L. Ophel. Transport of ^{60}Co between water and sediments in a small shield lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):877–884, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-108.

Zaugg:1986:STS

[1611] W. S. Zaugg, J. E. Bodle, J. E. Manning, and E. Wold. Smolt transformation and seaward migration in 0-age progeny of adult spring Chinook salmon (Oncorhynchus tshawytscha) matured early with photoperiod control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):885–888, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
109.

Hannesson:1986:OTY

[1612] Rögnvaldur Hannesson. Optimal thinning of a year-class with densitydependent growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):889– 892, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-110.

Morantz:1986:SEW

[1613] D. L. Morantz, S. E. Barbour, and R. K. Sweeney. Source of error in water velocity measurement for aquatic studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):893–896, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-111.

Favro:1986:CRE

[1614] L. D. Favro, P. K. Kuo, and J. F. McDonald. Capture–recapture experiment with fly-caught brown (Salmo trutta) and rainbow trout (S. gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):896–899, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-112.

Minns:1986:LSR

[1615] C. K. Minns, J. R. M. Kelso, and M. G. Johnson. Large-scale risk assessment of acid rain impacts on fisheries: Models and lessons. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):900–921, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-113.

Anonymous:1986:BRCd

[1616] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(4):922–923, April 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-114.

Shenker:1986:LFG

[1617] Jonathan M. Shenker and Bori L. Olla. Laboratory feeding and growth.of juvenile sablefish, Anoplopoma fimbria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):930–937, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-115.

Chow-Fraser:1986:DCD

[1618] Patricia Chow-Fraser and C. Kim Wong. Dietary change during development in the freshwater calanoid copepod Epischura lacustris Forbes.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):938–944, May 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-116.

Reno:1986:EIP

[1619] Paul W. Reno, Katherine Kleftis, Stuart W. Sherburne, and Bruce L. Nicholson. Experimental infection and pathogenesis of viral erythrocytic necrosis (VEN) in Atlantic cod, Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):945–951, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-117.

Harding:1986:DVM

[1620] G. C. Harding, W. P. Vass, B. T. Hargrave, and S. Pearre, Jr. Diel vertical movements and feeding activity of zooplankton in St. Georges Bay, N.S., using net tows and a newly developed passive trap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):952–967, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-118.

Courtois:1986:RAP

[1621] R. Courtois and J. J. Dodson. Régime alimentaire et principaux facteurs influençant l’alimentation des larves de capelan (Mallotus villosus), d’éperlan (Osmerus mordax) et de hareng (Clupea harengus harengus) dans un estuaire partiellement mélangé. (French) [Diet and main factors influencing the diet of capelin (Mallotus villosus), perlan (Osmerus mordax) and herring (Clupea harengus harengus larvae) in a partially mixed estuary]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):968–979, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-119.

Jensen:1986:DAS

[1622] Arne Johan Jensen and Bjørn Ove Johnsen. Different adaptation strategies of Atlantic salmon (Salmo salar) populations to extreme climates with special reference to some cold Norwegian rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):980–984, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-120.

Riget:1986:PSE

[1623] F. F. Riget, K. H. Nygaard, and B. Christensen. Population structure, ecological segregation, and reproduction in a population of Arctic char (Salvelinus alpinus) from Lake Tasersuaq, Greenland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):985–992, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-121.

Lodge:1986:LTD

[1624] David M. Lodge, Timothy K. Kratz, and Gregory M. Capelli. Longterm dynamics of three crayfish species in Trout Lake, Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):993–998, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-122.

Furnell:1986:MMM

[1625] Donald J. Furnell and J. R. Brett. Model of monthly marine growth and natural mortality for Babine Lake sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):999–1004, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-123.

Williams:1986:FIM

[1626] D. Dudley Williams. Factors influencing the microdistribution of two sympatric species of Plecoptera: an experimental study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1005–1009, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-124.

Thomson:1986:CAL

[1627] C. A. Thomson and J. R. Geraci. Cortisol, aldosterone, and leucocytes in the stress response of bottlenose dolphins, Tursiops truncatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1010–1016, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-125.

Wood:1986:PRA

[1628] Chris M. Wood and Mary S. Rogano. Physiological responses to acid stress in crayfish (Orconectes): Haemolymph ions, acid–base status, and exchanges with the environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1017–1026, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-126.

McGlade:1986:SAF

[1629] J. M. McGlade. Selected articles on fish genetics: Introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1027, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-127.

Scott:1986:FAP

[1630] Gary K. Scott, Garth L. Fletcher, and Peter L. Davies. Fish antifreeze proteins: Recent gene evolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1028–1034, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-128.

Foster:1986:DCM

[1631] E. A. D. Foster. A distance constraint model for the prediction of protein tertiary structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (5):1035–1044, May 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-129.

Rehse:1986:ERF

[1632] Peter H. Rehse and William S. Davidson. Evolutionary relationship of a fish C type lactate dehydrogenase to other vertebrate lactate dehydrogenase isozymes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (5):1045–1051, May 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-130.

McGlade:1986:EME

[1633] Jacqueline McGlade and Peter Allen. Evolution of multifunctionalism in enzymes: Specialist versus generalist strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1052–1058, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-131.

Doyle:1986:ASG

[1634] Roger W. Doyle and André J. Talbot. Artificial selection on growth and correlated selection on competitive behaviour in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1059–1064, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-132.

Chow-Fraser:1986:EMP

[1635] Patricia Chow-Fraser. An empirical model to predict in situ grazing rates of Diaptomus minutus Lilljeborg on small algal particles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1065–1070, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-133.

Jude:1986:SDE

[1636] D. J. Jude, P. J. Mansfield, S. F. DeBoe, and F. J. Tesar. Spatial distribution of entrained fish larvae in a power plant discharge canal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1070–1074, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-134.

Verspoor:1986:SCT

[1637] Eric Verspoor. Spatial correlations of transferrin allele frequencies in Atlantic salmon (Salmo salar) populations from North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1074–1078, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-135.

Leslie:1986:CLF

[1638] J. K. Leslie and J. E. Moore. Changes in lengths of fixed and preserved young freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (5):1079–1081, May 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-136.

Anonymous:1986:ON

[1639] Anonymous. Obituary/nécrologie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1082–1083, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-137.

Anonymous:1986:LCLa

[1640] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1084–1094, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-138.

Anonymous:1986:BRCe

[1641] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(5):1095–1097, May 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f86-139.

Welch:1986:PCC

[1642] Harold E. Welch and John A. Legault. Precipitation chemistry and chemical limnology of fertilized and natural lakes at Saqvaqjuac, N.W.T. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1104–1134, June 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-140.

Reinertsen:1986:ACP

[1643] Helge Reinertsen, Arne Jensen, Arnfinn Langeland, and Yngvar Olsen. Algal competition for phosphorus: The influence of zooplankton and fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1135–1141, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-141.

Whyte:1986:ISS

[1644] J. N. C. Whyte, J. R. Englar, B. L. Carswell, and K. E. Medic. Influence of starvation and subsequent feeding on body composition and energy reserves in the prawn Pandalus platyceros. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1142–1148, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-142.

Hilton:1986:DUC

[1645] J. W. Hilton and S. J. Slinger. Digestibility and utilization of canola meal in practical-type diets for rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1149–1155, June 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-143.

Reiter:1986:CVS

[1646] Michael A. Reiter and Robert E. Carlson. Current velocity in streams and the composition of benthic algal mats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1156–1162, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-144.

Nalewajko:1986:EEV

[1647] C. Nalewajko and D. Voltolina. Effects of environmental variables on growth rates and physiological characteristics of Lake Superior phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1163–1170, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-145.

Hambright:1986:ESI

[1648] K. David Hambright, Robert J. Trebatoski, Ray W. Drenner, and Dean Kettle. Experimental study of the impacts of bluegill (Lepomis macrochirus) and largemouth bass (Micropterus salmoides) on pond community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6): 1171–1176, June 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-146.

Whalen:1986:SIC

[1649] S. C. Whalen and V. Alexander. Seasonal inorganic carbon and nitrogen transport by phytoplankton in an Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1177–1186, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-147.

Allen:1986:DDE

[1650] P. M. Allen and J. M. McGlade. Dynamics of discovery and exploitation: The case of the Scotian Shelf groundfish fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1187–1200, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-148.

Quigley:1986:PRP

[1651] Michael A. Quigley and John A. Robbins. Phosphorus release processes in nearshore Southern Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1201–1207, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-149.

Victor:1986:DMR

[1652] Benjamin C. Victor. Delayed metamorphosis with reduced larval growth in a coral reef fish (Thalassoma bifasciatum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1208–1213, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-150.

Richards:1986:ESA

[1653] Laura J. Richards and Jon T. Schnute. An experimental and statistical approach to the question: Is CPUE an index of abundance? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1214–1227, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-151.

Bird:1986:CHO

[1654] Jerry L. Bird, Duane T. Eppler, and David M. Checkley, Jr. Comparisons of herring otoliths using Fourier series shape analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1228–1234, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-152.

Morin:1986:MPC

[1655] Antoine Morin, Pierre-Paul Harper, and Robert Henry Peters. Microhabitat-preference curves of blackfly larvae (Diptera: Simuliidae): a comparison of three estimation methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1235–1241, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-153.

Myers:1986:VMP

[1656] Ransom A. Myers, Jeffrey A. Hutchings, and R. John Gibson. Variation in male parr maturation within and among populations of Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6): 1242–1248, June 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-154.

Thomas:1986:WHP

[1657] A. C. Thomas and W. J. Emery. Winter hydrography and plankton distribution on the Southern British Columbia Continental Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1249–1258, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-155.

Peterson:1986:CFT

[1658] Bruce J. Peterson, John E. Hobbie, and Teresa L. Corliss. Carbon flow in a tundra stream ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (6):1259–1270, June 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-156.

Findlay:1986:BBB

[1659] Stuart Findlay, Judy L. Meyer, and Rebecca Risley. Benthic bacterial biomass and production in two blackwater rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1271–1276, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-157.

Lasenby:1986:TPE

[1660] D. C. Lasenby, T. G. Northcote, and M. Fürst. Theory, practice, and effects of Mysis relicta introductions to North American and Scandinavian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1277–1284, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-158.

Burton:1986:IPA

[1661] Thomas M. Burton and Jon W. Allan. Influence of pH, aluminum, and organic matter on stream invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1285–1289, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-159.

Hawkins:1986:PRA

[1662] C. M. Hawkins, J. D. Castell, and V. Leroyer. Patterns and rates of ammonia excretion by juvenile American lobsters, Homarus americanus, fed casein- and crab protein-based diets. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1290–1294, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-160.

McIntyre:1986:EFL

[1663] D. B. McIntyre and F. J. Ward. Estimating fork lengths of fathead minnows, Pimephales promelas, from measurement of pharyngeal arches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1294–1297, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-161.

Wohlfarth:1986:DNF

[1664] Giora W. Wohlfarth. Decline in natural fisheries — a genetic analysis and suggestion for recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (6):1298–1306, June 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-162.

Anonymous:1986:BRCf

[1665] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(6):1307–1309, June 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f86-163.

McGladdery:1986:ASN

[1666] S. E. McGladdery. Anisakis simplex (Nematoda: Anisakidae) infection of the musculature and body cavity of Atlantic herring (Clupea harengus harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1312–1317, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
164.

Mansfield:1986:AAP

[1667] Pamela J. Mansfield and David J. Jude. Alewife (Alosa pseudoharengus) survival during the first growth season in Southeastern Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1318–1326, July 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-165.

Agellon:1986:RTS

[1668] L. B. Agellon, T. T. Chen, R. J. Van Beneden, R. A. Sonstegard, G. F. Wagner, and B. A. McKeown. Rainbow trout (Salmo gairdneri) growth hormone: in vitro translation of pituitary RNA and product analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1327–1331, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-166.

Walter:1986:RAE

[1669] Gilbert G. Walter. A robust approach to equilibrium yield curves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1332–1339, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-167.

Campbell:1986:EAO

[1670] P. J. Campbell and F. H. Rigler. Effect of ambient oxygen concentration on measurements of sediment oxygen consumption. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1340–1349, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-168.

Baron:1986:SDM

[1671] Jill Baron, Stephen A. Norton, David R. Beeson, and Raymond Herrmann. Sediment diatom and metal stratigraphy from Rocky Mountain lakes with special reference to atmospheric deposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1350–1362, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-169.

Hanson:1986:ECB

[1672] John Mark Hanson and William C. Leggett. Effect of competition between two freshwater fishes on prey consumption and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1363–1372, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-170.

Mulligan:1986:CAP

[1673] T. J. Mulligan and R. Kieser. Comparison of acoustic population estimates of salmon in a lake with a weir count. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1373–1385, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-171.

Levings:1986:DUC

[1674] C. D. Levings, C. D. McAllister, and B. D. Chang. Differential use of the Campbell River Estuary, British Columbia by wild and hatcheryreared juvenile Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1386–1397, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-172.

Leider:1986:CLH

[1675] Steven A. Leider, Mark W. Chilcote, and John J. Loch. Comparative life history characteristics of hatchery and wild steelhead trout (Salmo gairdneri) of summer and winter races in the Kalama River, Washington.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1398–1409, July 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-173.

Holdway:1986:IPE

[1676] D. A. Holdway and D. G. Dixon. Impact of pulse exposure to methoxychlor on flagfish (Jordanella floridae) over one reproductive cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1410–1415, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-174.

Hebert:1986:PCA

[1677] Paul D. N. Hebert and Brenda J. Hann. Patterns in the composition of Arctic tundra pond microcrustacean communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1416–1425, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-175.

Parks:1986:EPD

[1678] J. W. Parks, J. A. Sutton, and A. Lutz. Effect of point and diffuse source loadings on mercury concentrations in the Wabigoon River: Evidence of a seasonally varying sediment–water partition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1426–1444, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-176.

Peterson:1986:PVP

[1679] R. H. Peterson and D. J. Martin-Robichaud. Perivitelline and vitelline potentials in teleost eggs as influenced by ambient ionic strength, natal salinity, and electrode electrolyte; and the influence of these potentials on cadmium dynamics within the egg. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1445–1450, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-177.

Aiken:1986:OMS

[1680] D. E. Aiken and S. L. Waddy. Oocyte maturation and spawning in wild American lobsters (Homarus americanus): Lack of evidence for significant regulation by photoperiod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1451–1453, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-178.

Aiken:1986:GVD

[1681] D. E. Aiken and S. L. Waddy. Growth of the vasa deferentia of mature Homarus americanus: Conflicting results from field and laboratory studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1453–1457, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-179.

Strong:1986:ACP

[1682] M. B. Strong, J. D. Neilson, and J. J. Hunt. Aberrant crystallization of pollock (Pollachius virens) otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1457–1463, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-180.

Larsson:1986:ZFA

[1683] Per Larsson. Zooplankton and fish accumulate chlorinated hydrocarbons from contaminated sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1463–1466, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-181.

Anonymous:1986:LCLb

[1684] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1467–1468, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-182.

Anonymous:1986:BRCg

[1685] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(7):1469, July 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-183.

Suttle:1986:PUR

[1686] C. A. Suttle and P. J. Harrison. Phosphate uptake rates of phytoplankton assemblages grown at different dilution rates in semicontinuous culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1474–1481, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-184.

Currie:1986:DOU

[1687] David J. Currie. Does orthophosphate uptake supply sufficient phosphorus to phytoplankton to sustain their growth? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1482–1487, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-185.

Lauren:1986:IWH

[1688] Darrel Jon Laurén and D. G. McDonald. Influence of water hardness, pH, and alkalinity on the mechanisms of copper toxicity in juvenile rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8): 1488–1496, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-186.

Cusimano:1986:EPT

[1689] Robert F. Cusimano, David F. Brakke, and Gary A. Chapman. Effects of pH on the toxicities of cadmium, copper, and zinc to steelhead trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8): 1497–1503, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-187.

Hardy:1986:BPZ

[1690] F. Joan Hardy, Ken S. Shortreed, and John G. Stockner. Bacterioplankton, phytoplankton, and zooplankton communities in a British Columbia coastal lake before and after nutrient reduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1504–1514, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-188.

Dawson:1986:LLA

[1691] V. K. Dawson, D. A. Johnson, and J. L. Allen. Loss of lampricides by adsorption on bottom sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1515–1520, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-189.

Murphy:1986:ECC

[1692] Michael L. Murphy, Jonathan Heifetz, Scott W. Johnson, K. V. Koski, and John F. Thedinga. Effects of clear-cut logging with and without buffer strips on juvenile salmonids in Alaskan streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1521–1533, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-190.

Macdonald:1986:FRU

[1693] J. Stevenson Macdonald and Roger H. Green. Food resource utilization by five species of benthic feeding fish in Passamaquoddy Bay, New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1534–1546, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-191.

Schom:1986:GEM

[1694] C. B. Schom. Genetic, environmental, and maturational effects on Atlantic salmon (Salmo salar) survival in acute low pH trials. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1547–1555, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-192.

Scherer:1986:LAB

[1695] E. Scherer, S. E. Harrison, and S. B. Brown. Locomotor activity and blood plasma parameters of acid-exposed lake whitefish, Coregonus clupeaformis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1556–1561, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-193.

Suthers:1986:RHL

[1696] Iain M. Suthers and John H. Gee. Role of hypoxia in limiting diel spring and summer distribution of juvenile yellow perch (Perca flavescens) in a prairie marsh. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8): 1562–1570, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-194.

McQueen:1986:TRF

[1697] Donald J. McQueen, John R. Post, and Edward L. Mills. Trophic relationships in freshwater pelagic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1571–1581, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-195.

Feltmate:1986:DSN

[1698] Blair W. Feltmate, Robert L. Baker, and Philip J. Pointing. Distribution of the stonefly nymph Paragnetina media (Plecoptera: Perlidae): Influence of prey, predators, current speed, and substrate composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1582–1587, August 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-196.

Chilton:1986:NMA

[1699] D. E. Chilton and H. T. Bilton. New method for ageing Chinook salmon (Onchorhynchus tshawytscha) using dorsal fin rays, and evidence if its validity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1588–1594, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
197.

Helfman:1986:DDA

[1700] Gene S. Helfman. Diel distribution and activity of American eels (Anguilla rostrata) in a cave-spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1595–1605, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-198.

Kapuscinski:1986:CGF

[1701] Anne R. D. Kapuscinski and James E. Lannan. A conceptual genetic fitness model for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1606–1616, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-199.

Hamrin:1986:VDH

[1702] Stellan F. Hamrin. Vertical distribution and habitat partitioning between different size classes of vendace, Coregonus albula, in thermally stratified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1617–1625, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
200.

Berube:1986:TDD

[1703] Michel Bérubé and Mark A. Curtis. Transmission of Diphyllobothrium ditremum to Arctic char (Salvelinus alpinus) in two subarctic Quebec lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1626–1634, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
201.

Ward:1986:WDS

[1704] G. Milton Ward and Nicholas G. Aumen. Woody debris as a source of fine particulate organic matter in coniferous forest stream ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1635–1642, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-202.

Manzer:1986:FER

[1705] J. I. Manzer and I. Miki. Fecundity and egg retention of some sockeye salmon (Oncorhynchus nerka) stocks in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1643–1655, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-203.

Misra:1986:FCS

[1706] R. K. Misra. Fitting and comparing several growth curves of the generalized von Bertalanffy type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (8):1656–1659, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-204.

Cornett:1986:SMM

[1707] R. J. Cornett and F. H. Rigler. Simple method of measuring seston respiration in oligotrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (8):1660–1663, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-205.

McLeese:1986:TTS

[1708] J. M. McLeese and E. Don Stevens. Trypsin from two strains of rainbow trout, Salmo gairdneri, is influenced differently by assay and acclimation temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8): 1664–1667, August 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-206.

Fortin:1986:SBT

[1709] C. Fortin, D. Lapointe, and G. Charpentier. Susceptibility of brook trout (Salvelinus fontinalis) fry to a liquid formulation of Bacillus thuringiensis serovar. israelensis (Teknar(R)) used for blackfly control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1667–1670, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-207.

Longhurst:1986:IIC

[1710] Alan R. Longhurst. Instar increments in copepod growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1671–1674, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-208.

Anonymous:1986:LCLc

[1711] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1675–1677, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-209.

Anonymous:1986:BRCh

[1712] Anonymous. Book reviews/comptes rendus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1678–1679, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f86-210.

Anonymous:1986:CRPb

[1713] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(8):1680, August 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
211.

Schnute:1986:GMD

[1714] J. Schnute and J. O. T. Jensen. A general multivariate dose–response model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1684–1693, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-212.

Jensen:1986:AJS

[1715] J. O. T. Jensen, J. Schnute, and D. F. Alderdice. Assessing juvenile salmonid response to gas supersaturation using a general multivariate dose–response model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9): 1694–1709, September 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-213.

Conan:1986:FMT

[1716] Gérard Y. Conan and Michel Comeau. Functional maturity and terminal molt of male snow crab, Chionoecetes opilio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1710–1719, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-214.

Baker:1986:FLL

[1717] Robert L. Baker. Food limitation of larval dragonflies: a field test of spacing behaviour. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1720– 1725, September 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-215.

VanDuyn-Henderson:1986:ZCT

[1718] J. A. Van Duyn-Henderson and D. C. Lasenby. Zinc and cadmium transport by the vertically migrating opossum shrimp, Mysis relicta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1726–1732, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-216.

Donald:1986:DUL

[1719] D. B. Donald and D. J. Alger. Dynamics of unexploited and lightly exploited populations of rainbow trout (Salmo gairdneri) from coastal, montane, and subalpine lakes in Western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1733–1741, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-217.

Jensen:1986:FRC

[1720] A. L. Jensen. Functional regression and correlation analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1742–1745, September 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-218.

Hankin:1986:DER

[1721] David G. Hankin and M. C. Healey. Dependence of exploitation rates for maximum yield and stock collapse on age and sex structure of Chinook salmon (Oncorhynchus tshawytscha) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1746–1759, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-219.

Olson:1986:APY

[1722] Robert J. Olson and Christofer H. Boggs. Apex predation by Yellowfïn tuna (Thunnus albacares): Independent estimates from gastric evacuation and stomach contents, bioenergetics, and cesium concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1760–1775, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-220.

Burczynski:1986:ADB

[1723] Janusz J. Burczynski and Robert L. Johnson. Application of dualbeam acoustic survey techniques to limnetic populations of juvenile sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1776–1788, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-221.

Sprules:1986:PSS

[1724] W. Gary Sprules and M. Munawar. Plankton size spectra in relation to ecosystem productivity, size, and perturbation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1789–1794, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-222.

Stewart:1986:HGT

[1725] Kenton M. Stewart and Robert F. Platford. Hypersaline gradients in two Canadian High Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1795–1803, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-223.

Earle:1986:APC

[1726] J. Chris Earle, Hamish C. Duthie, and David A. Scruton. Analysis of the phytoplankton composition of 95 Labrador lakes, with special reference to natural and anthropogenic acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1804–1811, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-224.

Pugsley:1986:TDD

[1727] C. W. Pugsley and H. B. N. Hynes. Three-dimensional distribution of winter stonefly nymphs, Allocapnia pygmaea, within the substrate of a Southern Ontario river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9): 1812–1817, September 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-225.

Shepherd:1986:RBS

[1728] Bruce G. Shepherd, Gordon F. Hartman, and William J. Wilson. Relationships between stream and intragravel temperatures in coastal drainages, and some implications for fisheries workers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(9):1818–1822, September 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-226.

Heggberget:1986:GGV

[1729] Tor G. Heggberget, Roar A. Lund, Nils Ryman, and Gunnar Ståhl. Growth and genetic variation of Atlantic salmon (Salmo salar) from different sections of the River Alta, North Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1828–1835, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-227.

Dixit:1986:SVS

[1730] Sushil S. Dixit and R. Douglas Evans. Spatial variability in sedimentary algal microfossils and its bearing on diatom-inferred pH reconstructions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1836–1845, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-228.

Mulholland:1986:ESA

[1731] Patrick J. Mulholland, J. W. Elwood, A. V. Palumbo, and R. J. Stevenson. Effect of stream acidification on periphyton composition, chlorophyll, and productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):

1846–1858, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-229.

Guderley:1986:MCD

[1732] Helga Guderley, Pierre Blier, and Lorraine Richard. Metabolic changes during the reproductive migration of two sympatric coregonines, Coregonus artedii and Coregonus clupeaformis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1859–1865, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-230.

Gonzalez-Villasenor:1986:CCM

[1733] Lucia Irene González-Villaseñor, Amanda M. Burkhoff, Víctor Corces, and Dennis A. Powers. Characterization of cloned mitochondrial DNA from the teleost Fundulus heteroclitus and its usefulness as an interspecies hybridization probe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 1866–1872, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-231.

Jackson:1986:MML

[1734] Togwell A. Jackson. Methyl mercury levels in a polluted prairie river–lake system: Seasonal and site-specific variations, and the dominant influence of trophic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 1873–1887, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-232.

Rosseland:1986:LSR

[1735] B. O. Rosseland, O. K. Skogheim, H. Abrahamsen, and D. Matzow. Limestone slurry reduces physiological stress and increases survival of Atlantic salmon (Salmo salar) in an acidic Norwegian river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1888–1893, October 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-233.

OBrien:1986:NVP

[1736] W. John O’Brien, Barbara I. Evans, and Gregory L. Howick. A new view of the predation cycle of a planktivorous fish, white crappie (Pomoxis annularis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1894–1899, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
234.

Duarte:1986:PBC

[1737] Carlos M. Duarte, Jacob Kalff, and Robert H. Peters. Patterns in biomass and cover of aquatic macrophytes in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1900–1908, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-235.

Vollestad:1986:EFR

[1738] L. A. Vøllestad, B. Jonsson, N. A. Hvidsten, T. F. Næsje, Ø. Haraldstad, and J. Ruud-Hansen. Environmental factors regulating the seaward migration of European silver eels (Anguilla anguilla). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1909–1916, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-236.

Herman:1986:IAP

[1739] David Herman, Narinder K. Kaushik, and Keith R. Solomon. Impact of atrazine on periphyton in freshwater enclosures and some ecological consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1917–1925, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-237.

Portt:1986:BPF

[1740] Cameron B. Portt, Eugene K. Balon, and David L. G. Noakes. Biomass and production of fishes in natural and channelized streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1926–1934, October 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-238.

Drenner:1986:EAD

[1741] Ray W. Drenner, Stephen T. Threlkeld, and Michael D. McCracken. Experimental analysis of the direct and indirect effects of an omnivorous filter-feeding clupeid on plankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1935–1945, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-239.

Holtby:1986:SAS

[1742] L. B. Holtby and M. C. Healey. Selection for adult size in female Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1946–1959, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-240.

Jasper:1986:PCL

[1743] Stephen Jasper and Max L. Bothwell. Photosynthetic characteristics of lotic periphyton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 1960–1969, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-241.

Cunjak:1986:WHU

[1744] Richard A. Cunjak and Geoffrey Power. Winter habitat utilization by stream resident brook trout (Salvelinus fontinalis) and brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1970– 1981, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f86-242.

Smith:1986:PEW

[1745] T. G. Smith and M. O. Hammill. Population estimates of white whale, Delphinapterus leucas, in James Bay, Eastern Hudson Bay, and Ungava Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1982–1987, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f86-
243.

Hearn:1986:HUB

[1746] William E. Hearn and Boyd E. Kynard. Habitat utilization and behavioral interaction of juvenile Atlantic salmon (Salmo salar) and rainbow trout (S. gairdneri) in tributaries of the White River of Vermont. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):1988–1998, October 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-244.

Kadlec:1986:EFD

[1747] John A. Kadlec. Effects of flooding on dissolved and suspended nutrients in small diked marshes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 1999–2008, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-245.

Kadlec:1986:ION

[1748] John A. Kadlec. Input–output nutrient budgets for small diked marshes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2009–2016, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-246.

Raymond:1986:ITD

[1749] Pierre Raymond, Gérard Leduc, and Jack A. Kornblatt. Investigation sur la toxicodynamique du cyanure et sur sa biotransformation chez la truite arc-en-ciel (Salmo gairdneri). (French) [Investigation into the toxicodynamics of cyanide and its biotransformation in rainbow trout (Salmo gairdneri)]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2017–2024, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-247.

Wrona:1986:EAS

[1750] Frederick J. Wrona, Peter Calow, Ian Ford, Donald J. Baird, and Lorraine Maltby. Estimating the abundance of stone-dwelling organisms: a new method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2025– 2035, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f86-248.

Siddens:1986:CCE

[1751] Lisbeth K. Siddens, Wayne K. Seim, Lawrence R. Curtis, and Gary A. Chapman. Comparison of continuous and episodic exposure to acidic, aluminum-contaminated waters of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2036–2040, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-249.

Hollett:1986:VMI

[1752] Lois Hollett, Michael Berrill, and Locke Rowe. Variation in major ion concentration of Cambarus robustus and Orconectes rusticus following exposure to low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 2040–2044, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-250.

Roff:1986:HPM

[1753] John C. Roff and Russell R. Hopcroft. High precision microcomputer based measuring system for ecological research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2044–2048, October 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-251.

Tam:1986:RRG

[1754] W. H. Tam, P. D. Payson, and R. J. J. Roy. Retardation and recovery of growth in brook trout fry (Salvelinus fontinalis) exposed for various durations to acidified water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10): 2048–2050, October 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-252.

Anonymous:1986:LCLd

[1755] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(10):2051–2057, October 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-253.

Campbell:1986:IIW

[1756] Alan Campbell. Introduction to the International Workshop on Lobster Recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2064– 2065, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-254.

Stewart:1986:IA

[1757] James E. Stewart. Introductory address. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2068–2069, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-255.

Ennis:1986:SDR

[1758] G. P. Ennis. Stock definition, recruitment variability, and larval recruitment processes in the American lobster, Homarus americanus: a review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2072–2084, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-256.

Jamieson:1986:IFR

[1759] G. S. Jamieson. Implications of fluctuations in recruitment in selected crab populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2085– 2098, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-257.

Lyons:1986:PPR

[1760] William G. Lyons. Problems and perspectives regarding recruitment of spiny lobsters, Panulirus argus, to the South Florida fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2099–2106, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-258.

Pollock:1986:RFB

[1761] David E. Pollock. Review of the fishery for and biology of the Cape rock lobster Jasus lalandii with notes on larval recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2107–2117, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-259.

MacDonald:1986:RPS

[1762] Craig D. MacDonald. Recruitment of the puerulus of the spiny lobster, Panulirus marginatus, in Hawaii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2118–2125, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-260.

Phillips:1986:PCC

[1763] B. F. Phillips. Prediction of commercial catches of the western rock lobster Panulirus cygnus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11): 2126–2130, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-261.

Caputi:1986:RBI

[1764] N. Caputi and R. S. Brown. Relationship between indices of juvenile abundance and recruitment in the western rock lobster (Panulirus cygnus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2131– 2139, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-262.

Pringle:1986:CSL

[1765] J. D. Pringle. California spiny lobster (Panulirus interruptus) larval retention and recruitment: a review and synthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2142–2152, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-263.

Phillips:1986:PPS

[1766] B. F. Phillips and P. S. McWilliam. The pelagic phase of spiny lobster development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2153– 2163, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-264.

Hudon:1986:RHV

[1767] Christiane Hudon, Pierre Fradette, and Pierre Legendre. La répartition horizontale et verticale des larves de homard (Homarus americanus) autour des îles de la Madeleine, golfe du Saint-Laurent. (French) [The horizontal and vertical distribution of lobster larvae (Homarus americanus) around the Madeleine Islands, Gulf of St. Lawrence]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2164–2176, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-265.

Ennis:1986:SAL

[1768] G. P. Ennis. Swimming ability of larval American lobsters, Homarus americanus, in flowing water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (11):2177–2183, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f86-266.

Sulkin:1986:ALS

[1769] Stephen D. Sulkin. Application of laboratory studies of larval behavior to fisheries problems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11): 2184–2188, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-267.

Campbell:1986:REP

[1770] A. Campbell and D. S. Pezzack. Relative egg production and abundance of berried lobsters, Homarus americanus, in the Bay of Fundy and off Southwestern Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2190–2196, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-268.

Campbell:1986:MMO

[1771] Alan Campbell. Migratory movements of ovigerous lobsters, Homarus americanus, tagged off Grand Manan, Eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2197–2205, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-269.

Pezzack:1986:EMH

[1772] D. S. Pezzack and D. R. Duggan. Evidence of migration and homing of lobsters (Homarus americanus) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2206–2211, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-270.

Booth:1986:RPR

[1773] John D. Booth. Recruitment of packhorse rock lobster Jasus verreauxi in New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2212– 2220, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-271.

Marx:1986:SSL

[1774] J. M. Marx. Settlement of spiny lobster, Panulirus argus, pueruli in South Florida: an evaluation from two perspectives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2221–2227, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-272.

Gregory:1986:MSL

[1775] Douglas R. Gregory, Jr. and Ronald F. Labisky. Movements of the spiny lobster Panulirus argus in South Florida. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2228–2234, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-273.

Brown:1986:CRW

[1776] R. S. Brown and N. Caputi. Conservation of recruitment of the western rock lobster (Panulirus cygnus) by improving survival and growth of undersize rock lobsters captured and returned by fishermen to the sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2236–2242, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-274.

Hunt:1986:FAG

[1777] John H. Hunt and William G. Lyons. Factors affecting growth and maturation of spiny lobsters, Panulirus argus, in the Florida Keys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2243–2247, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-275.

Richards:1986:CSB

[1778] R. Anne Richards and J. Stanley Cobb. Competition for shelter between lobsters (Homarus americanus) and Jonah crabs (Cancer borealis): Effects of relative size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11): 2250–2255, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-276.

Aiken:1986:EIR

[1779] D. E. Aiken and S. L. Waddy. Environmental influence on recruitment of the American lobster Homarus americanus: a perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2258–2270, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-277.

Quackenbush:1986:CER

[1780] L. Scott Quackenbush. Crustacean endocrinology, a review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2271–2282, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-278.

Atema:1986:RSS

[1781] Jelle Atema. Review of sexual selection and chemical communication in the lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2283–2290, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-279.

Waddy:1986:MFC

[1782] S. L. Waddy and D. E. Aiken. Multiple fertilization and consecutive spawning in large American lobsters, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2291–2294, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-280.

Wickham:1986:EIN

[1783] Daniel E. Wickham. Epizootic infestations by Nemertean brood parasites on commercially important crustaceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2295–2302, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-281.

Eagles:1986:ILF

[1784] M. D. Eagles, D. E. Aiken, and S. L. Waddy. Influence of light and food on larval American lobsters, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2303–2310, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-282.

Sasaki:1986:NBC

[1785] Glenn C. Sasaki, Judith McDowell Capuzzo, and Patricia Biesiot. Nutritional and bioenergetic considerations in the development of the American lobster Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2311–2319, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-283.

Castell:1986:ERN

[1786] J. D. Castell and J. C. Kean. Evaluation of the role of nutrition in lobster recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11): 2320–2327, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-284.

Caddy:1986:MSR

[1787] J. F. Caddy. Modelling stock-recruitment processes in crustacea: Some practical and theoretical perspectives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2330–2344, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-285.

Botsford:1986:EEF

[1788] Louis W. Botsford. Effects of environmental forcing on age-structured populations: Northern California Dungeness crab (Cancer magister) as an example. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2345– 2352, November 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-286.

Bannister:1986:EAA

[1789] R. C. A. Bannister and J. T. Addison. Effect of assumptions about the stock–recruitment relationship on a lobster (Homarus gammarus) stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2353–2359, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-287.

Addison:1986:DDM

[1790] J. T. Addison. Density-dependent mortality and the relationship between size composition and fishing effort in lobster populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2360–2367, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-288.

Fogarty:1986:RDA

[1791] Michael J. Fogarty and Josef S. Idoine. Recruitment dynamics in an American lobster (Homarus americanus) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2368–2376, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-289.

Pringle:1986:SSL

[1792] J. D. Pringle. Summary of session 2: Lobster fisheries and recruitment variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2378–2379, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-290.

Phillips:1986:SSL

[1793] B. F. Phillips. Summary of session 3: Larval distribution dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2380–2382, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-291.

Herrnkind:1986:SSB

[1794] William F. Herrnkind. Summary of session 4: Behavior and movements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2383, November 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-292.

Conan:1986:SSR

[1795] Gérard Y. Conan. Summary of session 5: Recruitment enhancement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2384–2388, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-293.

Cobb:1986:SSE

[1796] J. Stanley Cobb. Summary of session 6: Ecology of population structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2389–2390, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-294.

Aiken:1986:SSF

[1797] D. E. Aiken. Summary of session 7: Factors affecting growth and reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2391, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-295.

Fogarty:1986:SSM

[1798] M. J. Fogarty. Summary of session 8: Models of stock–recruitment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(11):2392–2393, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-296.

Caddy:1986:SSS

[1799] J. F. Caddy and A. Campbell. Summary of session 9: Summary of research recommendations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43 (11):2394–2396, November 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f86-297.

MacCall:1986:VPA

[1800] Alec D. MacCall. Virtual population analysis (VPA) equations for nonhomogeneous populations, and a family of approximations including improvements on Pope’s cohort analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2406–2409, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-298.

Wildish:1986:BMP

[1801] D. J. Wildish, D. L. Peer, and D. A. Greenberg. Benthic macrofaunal production in the Bay of Fundy and the possible effects of a tidal power barrage at Economy Point–Cape Tenny. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2410–2417, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-299.

Rose:1986:ALT

[1802] K. A. Rose, J. K. Summers, R. A. Cummins, and D. G. Heimbuch. Analysis of long-term ecological data using categorical time series regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2418–2426, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-300.

Oliveira:1986:NIR

[1803] Luis Oliveira and Naval J. Antia. Nickel ion requirements for autotrophic growth of several marine microalgae with urea serving as nitrogen source. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2427–2433, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f86-301.

Simon:1986:FSE

[1804] Raymond C. Simon, John D. McIntyre, and A. R. Hemmingsen. Family size and effective population size in a hatchery stock of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12): 2434–2442, December 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-302.

Nickelson:1986:UHC

[1805] Thomas E. Nickelson, Mario F. Solazzi, and Steven L. Johnson. Use of hatchery Coho salmon (Oncorhynchus kisutch) presmolts to rebuild wild populations in Oregon coastal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2443–2449, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-303.

Hall:1986:CTS

[1806] Kenneth J. Hall and Thomas G. Northcote. Conductivity–temperature standardization and dissolved solids estimation in a meromictic saline lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2450–2454, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-304.

Lavin:1986:ADT

[1807] P. A. Lavin and J. D. McPhail. Adaptive divergence of trophic phenotype among freshwater populations of the threespine stickleback (Gasterosteus aculeatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12): 2455–2463, December 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-305.

Munn:1986:SDP

[1808] Nancy Munn and Ellie Prepas. Seasonal dynamics of phosphorus partitioning and export in two streams in Alberta, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2464–2471, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-306.

Stockner:1986:APM

[1809] John G. Stockner and Naval J. Antia. Algal picoplankton from marine and freshwater ecosystems: a multidisciplinary perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2472–2503, December 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-307.

Secchi:1986:EPU

[1810] François Secchi and Micheline Martoja. Existence de populations uricotéliques chez Abra alba et Dreissena polymorpha (bivalves eulamellibranches). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2504–2508, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-308.

Carter:1986:DVM

[1811] J. C. H. Carter and K. A. Goudie. Diel vertical migrations and horizontal distributions of Limnocalanus macrurus and Senecella calanoides (Copepoda, Calanoida) in lakes of Southern Ontario in relation to planktivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2508–2514, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-309.

Mulcahy:1986:STI

[1812] Dan M. Mulcahy and Ron J. Pascho. Sequential tests for infectious hematopoietic necrosis virus in individuals and populations of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2515–2519, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f86-310.

Smith:1986:PEF

[1813] Val H. Smith and Maud Wasllsten. Prediction of emergent and floatingleaved macrophyte cover in Central Swedish lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2519–2523, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-311.

Hershey:1986:SPP

[1814] Anne E. Hershey. Selective predation by Procladius in an Arctic Alaskan lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2523–2528, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f86-312.

Swink:1986:SSL

[1815] William D. Swink and Lee H. Hanson. Survival from sea lamprey (Petromyzon marinus) predation by two strains of lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2528– 2531, December 1986. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f86-313.

Ryan:1986:PAS

[1816] Patrick M. Ryan. Prediction of angler success in an Atlantic salmon, Salmo salar, fishery two fishing seasons in advance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2531–2534, December 1986. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f86-314.

Anonymous:1986:LCLe

[1817] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 43(12):2535–2537, December 1986. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f86-315.

Anonymous:1987:E

[1818] Anonymous. Editorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1): 2, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-001.

Johnson:1987:TEL

[1819] Murray G. Johnson. Trace element loadings to sediments of fourteen Ontario lakes and correlations with concentrations in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):3–13, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-002.

Buckley:1987:RGB

[1820] L. J. Buckley and R. G. Lough. Recent growth, biochemical composition and prey field of larval haddock (Melanogrammus aeglefinus) and Atlantic cod (Gadus morhua) on Georges Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):14–25, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-003.

Koslow:1987:RNA

[1821] J. Anthony Koslow, Keith R. Thompson, and William Silvert. Recruitment to Northwest Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) stocks: Influence of stock size and climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):26–39, January 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-004.

Hayes:1987:CSS

[1822] John W. Hayes. Competition for spawning space between brown (Salmo trutta) and rainbow trout (S. gairdneri) in a lake inlet tributary, New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):40–47, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-005.

Peterman:1987:DDG

[1823] Randall M. Peterman and Michael J. Bradford. Density-dependent growth of age 1 English sole (Parophrys vetuls) in Oregon and Washington coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (1):48–53, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-006.

deLafontaine:1987:ESE

[1824] Yves de Lafontaine and William C. Leggett. Evaluation of in situ enclosures for larval fish studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (1):54–65, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-007.

Blackett:1987:DPA

[1825] Roger F. Blackett. Development and performance of an Alaska steeppass fishway for sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):66–76, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-008.

Valett:1987:FQH

[1826] H. M. Valett and J. A. Stanford. Food quality and hydropsychid caddisfly density in a lake outlet stream in Glacier National Park, Montana, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):77–82, January 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-009.

Sanders:1987:NES

[1827] James G. Sanders, Stephen J. Cibik, Christopher F. D’Elia, and Walter R. Boynton. Nutrient enrichment studies in a coastal plain estuary: Changes in phytoplankton species composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):83–90, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-010.

Uthe:1987:CSS

[1828] J. F. Uthe and C. L. Chou. Cadmium in sea scallop (Placopecten magellanicus) tissues from clean and contaminated areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):91–98, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-011.

Lauren:1987:ACRa

[1829] Darrel Jon Laurén and D. G. McDonald. Acclimation to copper by rainbow trout, Salmo gairdneri: Physiology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):99–104, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-012.

Lauren:1987:ACRb

[1830] Darrel Jon Laurén and D. G. McDonald. Acclimation to copper by rainbow trout, Salmo gairdneri: Biochemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):105–111, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-013.

Ho:1987:DPL

[1831] Kay T. Y. Ho and Steven P. Gloss. Distribution and persistence of the lampricide Bayer 73 following a control application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):112–119, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-014.

Morantz:1987:SMS

[1832] D. L. Morantz, R. K. Sweeney, C. S. Shirvell, and D. A. Longard. Selection of microhabitat in summer by juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):120–129, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-015.

Axler:1987:SME

[1833] R. P. Axler and J. E. Reuter. A simple method for estimating the ^{15}N content of dissolved organic matter (DO^{15}N) in N-cycling studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):130–133, January 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-016.

Cochran:1987:ECL

[1834] Roger C. Cochran. Effects of coal leachates on fish spermatogenesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):134–139, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-017.

Donard:1987:RTM

[1835] Olivier F. X. Donard, Fredrick T. Short, and James H. Weber. Regulation of tin and methyltin compounds by the green alga Enteromorpha under simulated estuarine conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):140–145, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-018.

Cornett:1987:DSH

[1836] R. J. Cornett and F. H. Rigler. Decomposition of seston in the hyolimnion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):146–151, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-019.

MacDonald:1987:GRO

[1837] B. A. MacDonald and N. F. Bourne. Growth, reproductive output, and energy partitioning in weathervane scallops, Patinopecten caurinus, from British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1): 152–160, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-020.

Maule:1987:CIS

[1838] Alec G. Maule, Carl B. Schreck, and Stephen L. Kaattari. Changes in the immune system of Coho salmon (Oncorhynchus kisutch) during the parr-to-smolt transformation and after implantation of cortisol. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):161–166, January 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-021.

Couture:1987:SFP

[1839] Pierre Couture, Claude Thellen, Patsy A. Thompson, and Jean-Christian Auclair. Structure and function of phytoplanktonic and microbial communities in relation to industrial wastewater discharge: an ecotoxicological approach in a lotic system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):167–175, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-022.

McCauley:1987:ECZ

[1840] Edward McCauley and Jacob Kalff. Effect of changes in zooplankton on orthophosphate dynamics of natural phytoplankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):176–182, January 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-023.

Pascho:1987:ELI

[1841] R. J. Pascho and D. Mulcahy. Enzyme-linked immunosorbent assay for a soluble antigen of Renibacterium salmoninarum, the causative agent of salmonid bacterial kidney disease. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):183–191, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-024.

Davis:1987:AVG

[1842] Michael W. Davis and Bori L. Olla. Aggression and variation in growth of chum salmon (Oncorhynchus keta) juveniles in seawater: Effects of limited ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):192–197, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-025.

Vezina:1987:SSV

[1843] Alain F. Vézina and Trevor Platt. Small-scale variability of new production and particulate fluxes in the ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):198–205, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-026.

Elliott:1987:MFF

[1844] Diane G. Elliott and Theresa Y. Barila. Membrane filtration–fluorescent antibody staining procedure for detecting and quantifying Renibacterium salmoninarum in coelomic fluid of Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):206–210, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-027.

Kostecki:1987:SLS

[1845] Paul T. Kostecki, Patricia Clifford, Steven P. Gloss, and James C. Carlisle. Scale loss and survival in smolts of Atlantic salmon (Salmo salar) after turbine passage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (1):210–214, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-028.

McDowell:1987:SVA

[1846] William H. McDowell, Jonathan J. Cole, and Charles T. Driscoll. Simplified version of the ampoule–persulfate method for determination of dissolved organic carbon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (1):214–218, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-029.

Hardy:1987:DVB

[1847] Ronald W. Hardy, Edmundo Casillas, and Toshiro Masumoto. Determination of vitamin B_{6} deficiency in rainbow trout (Salmo gairdneri) by liver enzyme assay and HPLC analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):219–222, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-030.

Pepin:1987:IAP

[1848] P. Pepin. Influence of alternative prey abundance on pelagic fish predation of larval fish: a model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (1):222–227, January 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-031.

Holdway:1987:ATM

[1849] D. A. Holdway, E. J. Kempe, and D. G. Dixon. Acute toxicity of methoxychlor to larval white sucker (Catostomus commersoni) as modified by age, food availability, and egg preexposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(1):227–232, January 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-032.

Fresh:1987:IAS

[1850] Kurt L. Fresh and Steven L. Schroder. Influence of the abundance, size, and yolk reserves of juvenile chum salmon (Oncorhynchus keta) on predation by freshwater fishes in a small coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):236–243, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-033.

Beacham:1987:AVB

[1851] T. D. Beacham and C. B. Murray. Adaptive variation in body size, age, morphology, egg size, and developmental biology of chum salmon (Oncorhynchus keta) in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):244–261, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-034.

Hartman:1987:UST

[1852] G. F. Hartman and T. G. Brown. Use of small, temporary, floodplain tributaries by juvenile salmonids in a West Coast rain-forest drainage basin, Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):262–270, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-035.

Hume:1987:ESD

[1853] Jeremy M. B. Hume and Eric A. Parkinson. Effect of stocking density on the survival, growth, and dispersal of steelhead trout fry (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):271–281, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-036.

Bowen:1987:AED

[1854] W. D. Bowen, R. A. Myers, and K. Hay. Abundance estimation of a dispersed, dynamic population: Hooded seals (Cystophora cristata) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2): 282–295, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-037.

Cattaneo:1987:PLD

[1855] Antonella Cattaneo. Periphyton in lakes of different trophy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):296–303, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-038.

Drost:1987:RBA

[1856] Maarten R. Drost. Relation between aiming and catch success in larval fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):304–315, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-039.

Green:1987:AMR

[1857] Philip E. J. Green and P. D. M. Macdonald. Analysis of mark–recapture data from hatchery-raised salmon using log-linear models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):316–326, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-040.

Green:1987:NGT

[1858] Philip E. J. Green. New graphical techniques for analyzing salmon migration in mark–recapture experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):327–336, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-041.

Pepper:1987:SRA

[1859] V. A. Pepper, T. Nicholls, and N. P. Oliver. Seminatural rearing of Atlantic salmon (Salmo salar) in Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):337–347, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-042.

Williams:1987:CSS

[1860] I. V. Williams and J. R. Brett. Critical swimming speed of Fraser and Thompson River pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):348–356, February 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-043.

Collie:1987:FSY

[1861] Jeremy S. Collie. Food selection by yellowtail flounder (Limanda ferruginea) on Georges Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (2):357–367, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-044.

Menard:1987:LAV

[1862] Jacques Ménard and Bruno Myrand. L’incidence d’ Aerococcus viridans var. homari dans le stock naturel de homards (Homarus americanus) des îles de la Madeleine (Québec), suite à une épidémie de gaffkémie dans un élevage in situ. (French) [The incidence of Aerococcus viridans var. homari in the natural stock of lobsters (Homarus americanus) of the Madeleine Islands (Quebec), following an epidemic of gaffkemia in a farm in situ]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):368–372, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-045.

Jones:1987:BBS

[1863] Keith A. Jones, Scott B. Brown, and Toshiaki J. Hara. Behavioral and biochemical studies of onset and recovery from acid stress in Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2): 373–381, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-046.

Yan:1987:IED

[1864] N. D. Yan and G. L. Mackie. Improved estimation of the dry weight of Holopedium gibberum (Crustacea, Cladocera) using clutch size, a body fat index, and lake water total phosphorus concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):382–389, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-047.

McKee:1987:SRS

[1865] P. M. McKee, W. J. Snodgrass, D. R. Hart, H. C. Duthie, J. H. McAndrews, and W. Keller. Sedimentation rates and sediment core profiles of ^{238}U and ^{232}Th decay chain radionuclides in a lake affected by uranium mining and milling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (2):390–398, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-048.

Bernatchez:1987:RBB

[1866] Louis Bernatchez and Julian J. Dodson. Relationship between bioenergetics and behavior in anadromous fish migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):399–407, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-049.

Mendelssohn:1987:FFA

[1867] Roy Mendelssohn and Philippe Cury. Fluctuations of a fortnightly abundance index of the ivoirian coastal pelagic species and associated environmental conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2): 408–421, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-050.

Fournier:1987:LBS

[1868] David A. Fournier and Ian J. Doonan. A length-based stock assessment method utilizing a generalized delay-difference model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):422–437, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-051.

Taggart:1987:WFH

[1869] Christopher T. Taggart and William C. Leggett. Wind-forced hydrodynamics and their interaction with larval fish and plankton abundance: a time-series analysis of physical–biological data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):438–451, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-052.

Rodgers:1987:PCD

[1870] Jeffrey D. Rodgers, Richard D. Ewing, and James D. Hall. Physiological changes during seaward migration of wild juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2): 452–457, February 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-053.

Arts:1987:ERT

[1871] Michael T. Arts and W. Gary Sprules. Energy reserves of three zooplankton species from two lakes with different metal concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):458–466, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-054.

Rice:1987:EPS

[1872] James A. Rice, Larry B. Crowder, and Fred P. Binkowski. Evaluating potential sources of mortality for larval bloater (Coregonus hoyi): Starvation and vulnerability to predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):467–472, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-055.

Caraco:1987:NLP

[1873] Nina Caraco, Armando Tamse, Osiris Boutros, and Ivan Valiela. Nutrient limitation of phytoplankton growth in brackish coastal ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):473–476, February 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-056.

Russell:1987:LSP

[1874] J. E. Russell, F. W. H. Beamish, and R. J. Beamish. Lentic spawning by the Pacific lamprey, Lampetra tridentata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):476–478, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-057.

France:1987:TBA

[1875] R. L. France. Test of biotic and abiotic environmental determinants of amphipod (Hyalella azteca) preamplexus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):478–482, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-058.

Anonymous:1987:LCLa

[1876] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):483–484, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-059.

Anonymous:1987:CRP

[1877] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(2):486, February 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-060.

Grant:1987:LGS

[1878] W. Stewart Grant, Chang Ik Zhang, Tokimasa Kobayashi, and Gunnar Ståhl. Lack of genetic stock discretion in Pacific cod (Gadus macrocephalus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):490–498, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-061.

Fahnenstiel:1987:DLMa

[1879] Gary L. Fahnenstiel and Donald Scavia. Dynamics of Lake Michigan phytoplankton: Primary production and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):499–508, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-062.

Fahnenstiel:1987:DLMb

[1880] Gary L. Fahnenstiel and Donald Scavia. Dynamics of Lake Michigan phytoplankton: Recent changes in surface and deep communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):509–514, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-063.

Nalepa:1987:LTC

[1881] Thomas F. Nalepa. Long-term changes in the macrobenthos of Southern Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):515–524, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-064.

Beamish:1987:LHA

[1882] Richard J. Beamish and John H. Youson. Life history and abundance of young adult Lampetra ayresi in the Fraser River and their possible impact on salmon and herring stocks in the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):525–537, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-065.

Freedman:1987:IMB

[1883] B. Freedman and T. A. Clair. Ion mass balances and seasonal fluxes from four acidic brownwater streams in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):538–548, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-066.

Mills:1987:FCP

[1884] Edward L. Mills, Daniel V. Widzowski, and Suzanne R. Jones. Food conditioning and prey selection by young yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):549–555, March 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-067.

Seeb:1987:GVP

[1885] James E. Seeb, Lisa W. Seeb, David W. Oates, and Fred M. Utter. Genetic variation and postglacial dispersal of populations of northern pike (Esox lucius) in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):556–561, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-068.

Kieser:1987:ITE

[1886] R. Kieser, T. J. Mulligan, N. J. Williamson, and M. O. Nelson. Intercalibration of two echo integration systems based on acoustic backscattering measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):562–572, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-069.

Campton:1987:GSA

[1887] Donald E. Campton and Fred M. Utter. Genetic structure of anadromous cutthroat trout (Salmo clarki clarki) populations in the Puget Sound area: Evidence for restricted gene flow. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):573–582, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-070.

Millar:1987:MLE

[1888] Russell B. Millar. Maximum likelihood estimation of mixed stock fishery composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):583–590, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-071.

Lodge:1987:RSM

[1889] David M. Lodge and James G. Lorman. Reductions in submersed macrophyte biomass and species richness by the crayfish Orconectes rusticus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):591–597, March 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-072.

McQueen:1987:IWT

[1890] D. J. McQueen and D. R. S. Lean. Influence of water temperature and nitrogen to phosphorus ratios on the dominance of blue–green algae in Lake St. George, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (3):598–604, March 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-073.

Welch:1987:FDF

[1891] David W. Welch. Frequency domain filtering of age-structured population data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):605–618, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-074.

Pringle:1987:EWS

[1892] Catherine M. Pringle. Effects of water and substratum nutrient supplies on lotic periphyton growth: an integrated bioassay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):619–629, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-075.

Tam:1987:MCM

[1893] W. H. Tam, L. Birkett, R. Makaran, P. D. Payson, D. K. Whitney, and C. K.-C. Yu. Modification of carbohydrate metabolism and liver vitellogenic function in brook trout (Salvelinus fontinalis) by exposure to low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):630–635, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-076.

Cleland:1987:NKC

[1894] Glenn B. Cleland and Ronald A. Sonstegard. Natural killer cell activity in rainbow trout (Salmo gairdneri): Effect of dietary exposure to Aroclor 1254 and/or mirex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (3):636–638, March 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-077.

Earle:1987:FID

[1895] J. C. Earle, H. C. Duthie, and David A. Scruton. Factors influencing the distribution of phytoplankton in 97 headwater lakes in insular Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):639–649, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-078.

Paloheimo:1987:FIP

[1896] J. E. Paloheimo and R. R. Fulthorpe. Factors influencing plankton community structure and production in freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):650–657, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-079.

McLeay:1987:RAG

[1897] D. J. McLeay, I. K. Birtwell, G. F. Hartman, and G. L. Ennis. Responses of Arctic grayling (Thymallus arcticus) to acute and prolonged exposure to Yukon placer mining sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):658–673, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-080.

Cross:1987:SCE

[1898] Melvin L. Cross, Gary F. Newkirk, and Graham Johnson. Strategies for the cultivation of European oysters (Ostrea edulis) in the Maritime provinces of Canada: Preliminary analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):674–679, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-081.

Roesijadi:1987:ICC

[1899] G. Roesijadi and G. W. Fellingham. Influence of Cu, Cd, and Zn preexposure on Hg toxicity in the mussel Mytilus edulis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):680–684, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-082.

Leek:1987:VEI

[1900] Steve L. Leek. Viral erythrocytic inclusion body syndrome (EIBS) occurring in juvenile spring Chinook salmon (Oncorhynchus tshawytscha) reared in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (3):685–688, March 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-083.

Anonymous:1987:LCLb

[1901] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(3):689–694, March 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-084.

Johnston:1987:EPT

[1902] C. E. Johnston, R. W. Gray, A. McLennan, and A. Paterson. Effects of photoperiod, temperature, and diet on the reconditioning response, blood chemistry, and gonad maturation of Atlantic salmon kelts (Salmo salar) held in freshwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):702–711, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-085.

West:1987:ESS

[1903] Cameron J. West and P. A. Larkin. Evidence for size-selective mortality of juvenile sockeye salmon (Oncorhynchus nerka) in Babine Lake, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):712–721, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-086.

Parrish:1987:SAL

[1904] Christopher C. Parrish. Separation of aquatic lipid classes by chromarod thin-layer chromatography with measurement by latroscan flame ionization detection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):722–731, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
087.

Duarte:1987:UET

[1905] Carlos M. Duarte. Use of echosounder tracings to estimate the aboveground biomass of submerged plants in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):732–735, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-088.

Wehr:1987:HSN

[1906] John D. Wehr, Lewis M. Brown, and Kathryn O’Grady. Highly specialized nitrogen metabolism in a freshwater phytoplankter, Chrysochromulina breviturrita. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):736–742, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-089.

Knowles:1987:NSC

[1907] Roger Knowles and David R. S. Lean. Nitrification: a significant cause of oxygen depletion under winter ice. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):743–749, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-090.

Xun:1987:MSR

[1908] Luying Xun, N. E. R. Campbell, and John W. M. Rudd. Measurements of specific rates of net methyl mercury production in the water column and surface sediments of acidified and circumneutral lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):750–757, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-091.

Lyons:1987:PCA

[1909] John Lyons. Prey choice among piscivorous juvenile walleyes (Stizostedion vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):758–764, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
092.

Gharrett:1987:GRA

[1910] A. J. Gharrett, S. M. Shirley, and G. R. Tromble. Genetic relationships among populations of Alaskan Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):765–774, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-

Ostrofsky:1987:CPR

[1911] M. L. Ostrofsky and F. H. Rigler. Chlorophyll–phosphorus relationships for subarctic lakes in Western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):775–781, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-094.

Addison:1987:TOR

[1912] R. F. Addison and P. F. Brodie. Transfer of organochlorine residues from blubber through the circulatory system to milk in the lactating grey seal Halichoerus grypus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):782–786, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-095.

Gharrett:1987:GCP

[1913] A. J. Gharrett and M. A. Thomason. Genetic changes in pink salmon (Oncorhynchus gorbuscha) following their introduction into the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):787–792, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-096.

Richards:1987:CIA

[1914] Laura J. Richards. Comparing imprecise abundance indices with a symmetric model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):793–802, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
097.

Hightower:1987:OPR

[1915] Joseph E. Hightower and Gary D. Grossman. Optimal policies for rehabilitation of overexploited fish stocks using a deterministic model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):803–810, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-098.

Rudstam:1987:DSE

[1916] Lars G. Rudstam, Clarence S. Clay, and John J. Magnuson. Density and size estimates of cisco (Coregonus artedii) using analysis of echo peak PDF from a single-transducer sonar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):811–821, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-099.

Wehrhahn:1987:EVR

[1917] C. F. Wehrhahn and R. Powell. Electrophoretic variation, regional differences, and gene flow in the Coho salmon (Oncorhynchus kisutch) of Southern British Colombia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):822–831, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-100.

Ciborowski:1987:DDM

[1918] Jan J. H. Ciborowski. Dynamics of drift and microdistribution of two mayfly populations: a predictive model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):832–845, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-101.

Kemp:1987:TMB

[1919] Christopher J. Kemp and Lawrence R. Curtis. Thermally modulated biliary excretion of [^{14}C] Taurocholate in rainbow trout (Salmo gairdneri) and the Na^{+}, K^{+}-ATPase. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):846–851, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-102.

Cornett:1987:VTO

[1920] R. J. Cornett and F. H. Rigler. Vertical transport of oxygen into the hypolimnion of lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (4):852–858, April 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-103.

Misra:1987:MTT

[1921] R. K. Misra and J. F. Uthe. Methods of time trends analysis applied to contaminant levels in Canadian Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):859–865, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-104.

Wood:1987:SIM

[1922] C. C. Wood, S. McKinnell, T. J. Mulligan, and D. A. Fournier. Stock identification with the maximum-likelihood mixture model: Sensitivity analysis and application to complex problems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):866–881, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-105.

Winters:1987:RDS

[1923] G. H. Winters and J. P. Wheeler. Recruitment dynamics of springspawning herring in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):882–900, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-106.

Medland:1987:AVM

[1924] T. E. Medland and F. W. H. Beamish. Age validation for the mountain brook lamprey, Ichthyomyzon greeleyi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):901–904, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-107.

McFarlane:1987:SDO

[1925] G. A. McFarlane and R. J. Beamish. Selection of dosages of oxytetracycline for age validation studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):905–909, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-108.

Martin:1987:AMW

[1926] J. F. Martin, C. P. McCoy, W. Greenleaf, and L. Bennett. Analysis of 2-methylisoborneol in water, mud, and channel catfish (Ictalurus punctatus) from commercial culture ponds in Mississippi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):909–912, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-109.

Sissenwine:1987:APR

[1927] M. P. Sissenwine and J. G. Shepherd. An alternative perspective on recruitment overfishing and biological reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(4):913–918, April 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-110.

Schnute:1987:GFM

[1928] Jon Schnute. A general fishery model for a size-structured fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):924–940, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-111.

Wood:1987:PJPa

[1929] C. C. Wood. Predation of juvenile Pacific salmon by the common merganser (Mergus merganser) on Eastern Vancouver Island. i: Predation during the seaward migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):941–949, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-112.

Wood:1987:PJPb

[1930] C. C. Wood. Predation of juvenile Pacific salmon by the common merganser (Mergus merganser) on Eastern Vancouver Island. II: Predation of stream-resident juvenile salmon by merganser broods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):950–959, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-113.

Ostrofsky:1987:PSS

[1931] Milton L. Ostrofsky. Phosphorus species in the surficial sediments of lakes of Eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (5):960–966, May 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-114.

Rivard:1987:AEC

[1932] Denis Rivard and Malcolm G. Foy. An analysis of errors in catch projections for Canadian Atlantic fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):967–981, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-115.

Thompson:1987:DPD

[1933] A. B. Thompson and L. Margolis. Determination of population discreteness in two species of shrimp, Pandalus jordani and Pandalopsis dispar, from coastal British Columbia using parasite tags and other population characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):982–989, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
116.

Rasmussen:1987:EMZ

[1934] Joseph B. Rasmussen and Jacob Kalff. Empirical models for zoobenthic biomass in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):990–1001, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
117.

Walters:1987:EMS

[1935] Carl J. Walters, Edith Krause, William E. Neill, and Thomas G. Northcote. Equilibrium models for seasonal dynamics of plankton biomass in four oligotrophy lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (5):1002–1017, May 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-118.

Trippel:1987:RRF

[1936] Edward A. Trippel and Harold H. Harvey. Reproductive responses of five white sucker (Catostomus commersoni) populations in relation to lake acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1018–1023, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-119.

Mulligan:1987:MCF

[1937] T. J. Mulligan, R. Kieser, and T. Gjernes. A model for comparing fish density estimates from seine catches and from echo integration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1024–1030, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-120.

Shardlow:1987:CAI

[1938] Tom Shardlow, Ray Hilborn, and David Lightly. Components analysis of instream escapement methods for Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1031–1037, May 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-121.

Hamilton:1987:STE

[1939] Daniel T. Hamilton and William D. Taylor. Short-term effects of zooplankton manipulations on phosphate uptake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1038–1044, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-122.

Keller:1987:MFP

[1940] Aimee Keller. Modeling and forecasting primary production rates using Box–Jenkins transfer function models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1045–1052, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-123.

Swartzman:1987:BMP

[1941] Gordon L. Swartzman, Wayne M. Getz, and Robert C. Francis. Binational management of Pacific hake (Merluccius productus): a stochastic modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (5):1053–1063, May 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-124.

Palumbo:1987:MCL

[1942] Anthony V. Palumbo, Patrick J. Mulholland, and Jerry W. Elwood. Microbial communities on leaf material protected from macroinvertebrate grazing in acidic and circumneutral streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1064–1070, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-125.

Mulcahy:1987:IHN

[1943] D. Mulcahy and W. N. Batts. Infectious hematopoietic necrosis virus detected by separation and incubation of cells from salmonid cavity fluid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1071–1075, May 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-126.

Mulcahy:1987:TMS

[1944] D. Mulcahy, R. J. Pascho, and W. N. Batts. Testing of male sockeye salmon (Oncorhynchus nerka) and steel head trout (Salmo gairdneri) for infectious hematopoietic necrosis virus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1075–1078, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-127.

Jakobsen:1987:IGR

[1945] Per Johan Jakobsen, Geir Helge Johnsen, and Jens Christian Holm. Increased growth rate in Atlantic salmon parr (Salmo salar) by using a two-coloured diet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1079– 1082, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-128.

Kratz:1987:WSP

[1946] Timothy K. Kratz, Robert B. Cook, Carl J. Bowser, and Patrick L. Brezonik. Winter and spring pH depressions in Northern Wisconsin lakes caused by increases in pCO_{2}. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (5):1082–1088, May 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-129.

Tang:1987:PSR

[1947] Yong Tang and E. T. Garside. Preexposure and subsequent resistance to lead in yearling brook trout Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1089–1091, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-130.

Baird:1987:OCT

[1948] D. J. Baird, T. E. Gates, and Ronald W. Davies. Oxygen conditions in two prairie pothole lakes during winter ice cover. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1092–1095, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-131.

Anonymous:1987:LCLc

[1949] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(5):1096–1098, May 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-132.

France:1987:LSP

[1950] R. L. France and P. M. Stokes. Life stage and population variation in resistance and tolerance of Hyalella azteca (Amphipoda) to low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1102–1111, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-133.

France:1987:EHE

[1951] Robert L. France and Bruce D. LaZerte. Empirical hypothesis to explain the restricted distribution of Hyalella azteca (Amphipoda) in anthropogenically acidified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (6):1112–1121, June 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-134.

Kleiber:1987:APS

[1952] P. Kleiber, A. W. Argue, and R. E. Kearney. Assessment of Pacific skipjack tuna (Katsuwonus pelamis) resources by estimating standing stock and components of population turnover from tagging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1122–1134, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-135.

Amy:1987:MAU

[1953] Penny S. Amy, Bruce A. Caldwell, A. H. Soeldner, Richard Y. Morita, and L. J. Albright. Microbial activity and ultrastructure of mineralbased marine snow from Howe Sound, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1135–1142, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-136.

Incze:1987:ALT

[1954] Lewis S. Incze, David A. Armstrong, and Sharon L. Smith. Abundance of larval Tanner crabs (Chionoecetes spp.) in relation to adult females and regional oceanography of the Southeastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1143–1156, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-137.

Attard:1987:EDE

[1955] Jean Attard and Christiane Hudon. Embryonic development and energetic investment in egg production in relation to size of female lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6): 1157–1164, June 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-138.

Urban:1987:AIZ

[1956] Noel R. Urban, Steven J. Eisenreich, and Eville Gorham. Aluminum, iron, zinc, and lead in bog waters of Northeastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1165–1172, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-139.

Starodub:1987:ICP

[1957] M. E. Starodub, P. T. S. Wong, C. I. Mayfield, and Y. K. Chau. Influence of complexation and pH on individual and combined heavy metal toxicity to a freshwater green alga. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (6):1173–1180, June 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-140.

Hay:1987:FAD

[1958] D. E. Hay and A. R. Kronlund. Factors affecting the distribution, abundance, and measurement of Pacific herring (Clupea harengus pallasi) spawn. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1181–1194, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-141.

Lawton:1987:DAF

[1959] Peter Lawton. Diel activity and foraging behavior of juvenile American lobsters, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1195–1205, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-142.

Yamada:1987:RES

[1960] S. Behrens Yamada, T. J. Mulligan, and D. Fournier. Role of environment and stock on the elemental composition of sockeye salmon (Oncorhynchus nerka) vertebrae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (6):1206–1212, June 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-143.

Kope:1987:SVP

[1961] Robert G. Kope. Separable virtual population analysis of Pacific salmon with application to marked Chinook salmon, Oncorhynchus tshawytscha, from California’s Central Valley. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1213–1220, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-144.

Newsome:1987:EMC

[1962] G. E. (Buck) Newsome and S. K. Aalto. An egg-mass census method for tracking fluctuations in yellow perch (Perca flavescens) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1221–1232, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-145.

Macdonald:1987:FHU

[1963] J. Stevenson Macdonald, I. K. Birtwell, and G. M. Kruzynski. Food and habitat utilization by juvenile salmonids in the Campbell River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1233–1246, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-146.

Perrin:1987:EEC

[1964] Chris J. Perrin, Max L. Bothwell, and Pat A. Slaney. Experimental enrichment of a coastal stream in British Columbia: Effects of organic and inorganic additions on autotrophic periphyton production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1247–1256, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-147.

Wong:1987:BMT

[1965] P. T. S. Wong, Y. K. Chau, J. L. Yaromich, and O. Kramar. Bioaccumulation and metabolism of tri- and dialkyllead compounds by a freshwater alga. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1257–1260, June 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-148.

Jefferts:1987:AAS

[1966] Katharine Jefferts, Janusz Burczynski, and William G. Pearcy. Acoustical assessment of squid (Loligo opalescens) off the Central Oregon Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(6):1261–1267, June 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-149.

Stegeman:1987:ICP

[1967] John J. Stegeman, Frances Y. Teng, and Elisabeth A. Snowberger. Induced cytochrome P450 in winter flounder (Pseudopleuronectes americanus) from coastal Massachusetts evaluated by catalytic assay and monoclonal antibody probes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (7):1270–1277, July 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-150.

Morrison:1987:CSM

[1968] K. A. Morrison, N. Thérien, and B. Marcos. Comparison of six models for nutrient limitations on phytoplankton growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1278–1288, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-151.

Geraci:1987:TCA

[1969] J. R. Geraci, N. C. Palmer, and D. J. St. Aubin. Tumors in cetaceans: Analysis and new findings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (7):1289–1300, July 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-152.

Wilson:1987:MDA

[1970] Gary M. Wilson, W. Kelly Thomas, and Andrew T. Beckenbach. Mitochondrial DNA analysis of Pacific Northwest populations of Oncorhynchus tshawytscha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (7):1301–1305, July 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-153.

Engstrom:1987:IVH

[1971] Daniel R. Engstrom. Influence of vegetation and hydrology on the humus budgets of Labrador lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (7):1306–1314, July 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-154.

El-Shaarawi:1987:FSR

[1972] A. H. El-Shaarawi. Frequency of sampling required for monitoring the Niagara River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1315– 1319, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-155.

Chadwick:1987:ICB

[1973] E. Michael P. Chadwick, Ross R. Claytor, Claude E. Léger, and Richard L. Saunders. Inverse correlation between ovarian development of Atlantic salmon (Salmo salar) smolts and sea age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1320–1325, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-156.

Fortier:1987:PLE

[1974] Louis Fortier, William C. Leggett, and Serge Gosselin. Patterns of larval emergence and their potential impact on stock differentiation in beach spawning capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1326–1336, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-157.

Trimbee:1987:ETP

[1975] Annette M. Trimbee and E. E. Prepas. Evaluation of total phosphorus as a predictor of the relative biomass of blue-green algae with emphasis on Alberta lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (7):1337–1342, July 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-158.

Birtwell:1987:LAS

[1976] I. K. Birtwell and G. M. Kruzynski. Laboratory apparatus for studying the behaviour of organisms in vertically stratified waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1343–1350, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-159.

Fitzpatrick:1987:PTC

[1977] Martin S. Fitzpatrick, J. Michael Redding, Frank D. Ratti, and Carl B. Schreck. Plasma testosterone concentration predicts the ovulatory response of Coho salmon (Oncorhynchus kisutch) to gonadotropin-releasing hormone analog. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1351– 1357, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-160.

Kulakkattolickal:1987:ESA

[1978] Augusthy Thevasia Kulakkattolickal and Donald L. Kramer. Effect of surface access and oxygen concentration on the toxicity of Croton tiglium (Euphorbiaceae) seed extract to Brachydanio rerio (Cyprinidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1358–1361, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-161.

Tremblay:1987:IPS

[1979] M. J. Tremblay, L. D. Meade, and G. V. Hurley. Identification of planktonic sea scallop larvae (Placopecten magellanicus) (Gmelin). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1361–1366, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-162.

Hilborn:1987:GMS

[1980] Ray Hilborn and Carl J. Walters. A general model for simulation of stock and fleet dynamics in spatially heterogeneous fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1366–1369, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-163.

Getz:1987:MVM

[1981] Wayne M. Getz, Robert C. Francis, and Gordon L. Swartzman. On managing variable marine fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1370–1375, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-164.

Welch:1987:ITL

[1982] David W. Welch. Influence of three life history strategies on the information content of stock–recruitment data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1375–1379, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-165.

McGladdery:1987:PES

[1983] Sharon E. McGladdery. Potential of Eimeria sardinae (Apicomplexa: Eimeridae) oocysts for distinguishing between spawning groups and between first- and repeat-spawning Atlantic herring (Clupea harengus harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(7):1379–1385, July 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
166.

Grant:1987:EBU

[1984] James W. A. Grant and David L. G. Noakes. Escape behaviour and use of cover by young-of-the-year brook trout, Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1390–1396, August 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-167.

Bilby:1987:EPH

[1985] Robert E. Bilby and Peter A. Bisson. Emigration and production of hatchery Coho salmon (Oncorhynchus kisutch) stocked in streams draining an old-growth and a clear-cut watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1397–1407, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-168.

Guildford:1987:DPP

[1986] S. J. Guildford, F. P. Healey, and R. E. Hecky. Depression of primary production by humic matter and suspended sediment in limnocorral experiments at Southern Indian Lake, Northern Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1408–1417, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-169.

Gunn:1987:LEP

[1987] John M. Gunn and David L. G. Noakes. Latent effects of pulse exposure to aluminum and low pH on size, ionic composition, and feeding efficiency of lake trout (Salvelinus namaycush) alevins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1418–1424, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-170.

Kelly:1987:CAM

[1988] R. K. Kelly, J. F. Klaverkamp, R. V. Hunt, and O. Nielsen. Chemical analysis of muscle from walleye (Stizostedion vitreum vitreum) with myofibrogranuloma, a chronic myopathy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1425–1431, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-171.

Reckhow:1987:EMF

[1989] Kenneth H. Reckhow, Robert W. Black, Thomas B. Stockton, Jr., J. David Vogt, and Judith G. Wood. Empirical models of fish response to lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8): 1432–1442, August 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-172.

Chou:1987:EDC

[1990] C. L. Chou, J. F. Uthe, J. D. Castell, and J. C. Keanz. Effect of dietary cadmium on growth, survival, and tissue concentrations of cadmium, zinc, copper, and silver in juvenile American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1443–1450, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
173.

Welch:1987:ESI

[1991] Harold E. Welch, John A. Legault, and Martin A. Bergmann. Effects of snow and ice on the annual cycles of heat and light in Saqvaqjuac Lakes.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1451–1461, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-174.

McCormick:1987:PCP

[1992] Stephen D. McCormick, Richard L. Saunders, Eugene B. Henderson, and Paul R. Harmon. Photoperiod control of parr–smolt transformation in Atlantic salmon (Salmo salar): Changes in salinity tolerance, gill Na^{+}, K^{+}-ATPase activity, and plasma thyroid hormones. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1462–1468, August 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-175.

Stites:1987:PPD

[1993] David L. Stites. Population and production dynamics of an enchytraeid worm in a subtropical blackwater river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1469–1474, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-176.

Lacroix:1987:RJA

[1994] Gilles L. Lacroix and Donald R. Townsend. Responses of juvenile Atlantic salmon (Salmo salar) to episodic increases in acidity of Nova Scotia rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1475–1484, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
177.

Tanasichuk:1987:IIV

[1995] R. W. Tanasichuk and D. M. Ware. Influence of interannual variations in winter sea temperature on fecundity and egg size in Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (8):1485–1495, August 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-178.

Hay:1987:ODO

[1996] D. E. Hay, D. N. Outram, B. A. McKeown, and M. Hurlburt. Ovarian development and oocyte diameter as maturation criteria in Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (8):1496–1502, August 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-179.

Yamada:1987:MNS

[1997] S. Behrens Yamada and T. J. Mulligan. Marking nonfeeding salmonid fry with dissolved strontium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (8):1502–1506, August 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-180.

Bevan:1987:LDE

[1998] David J. Bevan and Donald L. Kramer. Lack of a depth effect on the growth of the air-breathing catfish, Clarias macrocephalus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(8):1507–1510, August 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-181.

France:1987:ALA

[1999] R. L. France. Aggregation in littoral amphipod populations: Transformation controversies revisited. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (8):1510–1515, August 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-182.

Doyle:1987:SIL

[2000] Roger W. Doyle, A. J. Talbot, and R. R. Nicholas. Statistical interrelation of length, growth, and scale circulus spacing: Appraisal of a growth rate estimator for fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9): 1520–1528, September 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-183.

Confer:1987:ISI

[2001] John L. Confer and Marianne V. Moore. Interpreting selectivity indices calculated from field data or conditions of prey replacement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1529–1533, September 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-184.

DeLafontaine:1987:ECS

[2002] Yves De Lafontaine and William C. Leggett. Effect of container size on estimates of mortality and predation rates in experiments with macrozooplankton and larval fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9): 1534–1543, September 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-185.

Thomas:1987:AMF

[2003] G. L. Thomas and Darrell R. Jackson. Acoustic measurement of fish schools using array phase information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1544–1550, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-186.

Collie:1987:ARM

[2004] Jeremy S. Collie and Carl J. Walters. Alternative recruitment models of Adams River sockeye salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1551–1561, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-187.

Noltie:1987:IES

[2005] Douglas B. Noltie. Incidence and effects of sea lamprey (Petromyzon marinus) parasitism on breeding pink salmon (Oncorhynchus gorbuscha) from the Carp River, Eastern Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1562–1567, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-188.

Drinkwater:1987:TPM

[2006] K. F. Drinkwater and R. A. Myers. Testing predictions of marine fish and shellfish landings from environmental variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1568–1573, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-189.

Klump:1987:DTS

[2007] J. Val Klump, John R. Krezoski, Michael E. Smith, and Jerry L. Kaster. Dual tracer studies of the assimilation of an organic contaminant from sediments by deposit feeding oligochaetes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1574–1583, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-190.

Gibson:1987:CMP

[2008] Katherine N. Gibson, John P. Smol, and Jesse Ford. Chrysophycean microfossils provide new insight into the recent history of a naturally acidic lake (Cone Pond, New Hampshire). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1584–1588, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-191.

Mallet:1987:MVN

[2009] A. L. Mallet, C. E. A. Carver, S. S. Coffen, and K. R. Freeman. Mortality variations in natural populations of the blue mussel, Mytilus edulis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1589–1594, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-192.

Galloway:1987:PCL

[2010] James N. Galloway, George R. Hendrey, Carl L. Schofield, Norman E. Peters, and Arland H. Johannes. Processes and causes of lake acidification during spring snowmelt in the West-Central Adirondack Mountains, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1595– 1602, September 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-193.

Reeves:1987:IBR

[2011] Gordon H. Reeves, Fred H. Everest, and James D. Hall. Interactions between the redside shiner (Richardsonius balteatus) and the steelhead trout (Salmo gairdneri) in Western Oregon: The influence of water temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1603–1613, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-194.

Houlihan:1987:EET

[2012] D. F. Houlihan and P. Laurent. Effects of exercise training on the performance, growth, and protein turnover of rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1614–1621, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-195.

Clark:1987:IVH

[2013] Karen L. Clark and Bruce D. LaZerte. Intraspecific variation in hydrogen ion and aluminum toxicity in Bufo americanus and Ambystoma maculatum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1622–1628, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-196.

Fischer:1987:PIC

[2014] Robert U. Fischer, Jr., Edward A. Standora, and James R. Spotila. Predator-induced changes in thermoregulation of bluegill, Lepomis macrochirus, from a thermally altered reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1629–1634, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-197.

Peterson:1987:PIA

[2015] R. H. Peterson and D. J. Martin-Robichaud. Permeability of the isolated Atlantic salmon (Salmo salar) chorion to ions as estimated by diffusion potentials. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1635– 1639, September 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-198.

Steinman:1987:EIC

[2016] Alan D. Steinman and C. David McIntire. Effects of irradiance on the community structure and biomass of algal assemblages in laboratory streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1640– 1648, September 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-199.

VanderWerf:1987:DTL

[2017] B. Van der Werf, J. Schrotenboer, A. F. Richter, J. R. Moed, H. L. Hoogveld, and H. De Haan. A durable and transportable limnetic enclosure system suitable for wind-exposed lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1649–1652, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-200.

Hall:1987:EAE

[2018] Ronald J. Hall and Fredric P. Ide. Evidence of acidification effects on stream insect communities in Central Ontario between 1937 and 1985. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1652–1657, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-201.

Baker:1987:DIV

[2019] Robert L. Baker and Blair W. Feltmate. Development of Ischnura verticalis (Coenagrionidae: Odonata): Effects of temperature and prey abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1658–1661, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-202.

Geraci:1987:SPW

[2020] J. R. Geraci and L. J. Bruce-Allen. Slow process of wound repair in beluga whales, Delphinapterus leucas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1661–1665, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-203.

Chambers:1987:NOS

[2021] Patricia A. Chambers. Nearshore occurrence of submersed aquatic macrophytes in relation to wave action. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1666–1669, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-204.

Pontasch:1987:PRG

[2022] Kurt W. Pontasch and Merlyn A. Brusven. Periphyton response to a gasoline spill in Wolf Lodge Creek, Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1669–1673, September 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-205.

Anonymous:1987:LCLd

[2023] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(9):1676–1680, September 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-206.

Wolff:1987:PDP

[2024] Matthias Wolff. Population dynamics of the Peruvian scallop Argopecten purpuratus during the El Niño phenomenon of 1983. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1684–1691, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-207.

MacIsaac:1987:APR

[2025] Hugh J. MacIsaac, Thomas C. Hutchinson, and W. Keller. Analysis of planktonic rotifer assemblages from Sudbury, Ontario, area lakes of varying chemical composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1692–1701, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-208.

Beacham:1987:BGS

[2026] T. D. Beacham, A. P. Gould, R. E. Withler, C. B. Murray, and L. W. Barner. Biochemical genetic survey and stock identification of chum salmon (Oncorhynchus keta) in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1702–1713, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-209.

Day:1987:IFE

[2027] Kristin E. Day, N. K. Kaushik, and K. R. Solomon. Impact of fenvalerate on enclosed freshwater planktonic communities and on in situ rates of filtration of zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10): 1714–1728, October 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-210.

Taggart:1987:CUO

[2028] Christopher T. Taggart and Kenneth T. Frank. Coastal upwelling and Oikopleura occurrence (“Slub”): a model and potential application to inshore fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1729– 1736, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f87-211.

Rempel:1987:MEA

[2029] Robert S. Rempel and John C. H. Carter. Modelling the effects of accelerated development and reduced fecundity on the population dynamics of aquatic Diptera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10): 1737–1742, October 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-212.

Rempel:1987:TIA

[2030] Robert S. Rempel and John C. H. Carter. Temperature influences on adult size, development, and reproductive potential of aquatic Diptera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1743–1752, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-213.

Shin:1987:PRP

[2031] C. N. Shin, G-Y. Rhee, and J. Chen. Phosphate requirement, photosynthesis, and diel cell cycle of Scenedesmus obliquas under fluctuating light.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1753–1758, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-214.

Duarte:1987:LID

[2032] Carlos M. Duarte and Jacob Kalff. Latitudinal influences on the depths of maximum colonization and maximum biomass of submerged angiosperms in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1759–1764, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f87-
215.

Zatta:1987:CTI

[2033] Paolo Zatta. Calcium and taurine interaction in the heart of the crab Carcinus maenas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1765– 1768, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f87-216.

Suttle:1987:ENP

[2034] C. A. Suttle, J. G. Stockner, and P. J. Harrison. Effects of nutrient pulses on community structure and cell size of a freshwater phytoplankton assemblage in culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10): 1768–1774, October 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-217.

Skaala:1987:FSB

[2035] Øystein Skaala and Knut E. Jørstad. Fine-spotted brown trout (Salmo trutta): Its phenotypic description and biochemical genetic variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1775–1780, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-218.

Beamish:1987:EPN

[2036] Richard J. Beamish. Evidence that parasitic and nonparasitic life history types are produced by one population of lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1779–1782, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-219.

Buerkle:1987:EFL

[2037] U. Buerkle. Estimation of fish length from acoustic target strengths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1782–1785, October 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-220.

Arts:1987:PMM

[2038] Michael T. Arts and D. O. Evans. Precision micrometer measurement of mouth gape of larval fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10): 1786–1791, October 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-221.

Thorne:1987:NVM

[2039] Richard E. Thorne and James J. Ames. A note on variability of marine survival of sockeye salmon (Oncorhynchus nerka) and effects of flooding on spawning success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10): 1791–1795, October 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-222.

Hilborn:1987:RIC

[2040] Ray Hilborn and Wilf Luedke. Rationalizing the irrational: a case study in user group participation in Pacific salmon management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1796–1805, October 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-223.

Anonymous:1987:LCLe

[2041] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(10):1806–1809, October 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-224.

Stottlemyer:1987:SIA

[2042] Robert Stottlemyer. Snowpack ion accumulation and loss in a basin draining to Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (11):1812–1819, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f87-225.

Lavin:1987:MDO

[2043] Patrick A. Lavin and J. D. McPhail. Morphological divergence and the organization of trophic characters among lacustrine populations of the threespine stickleback (Gasterosteus aculeatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1820–1829, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-226.

Marnell:1987:GIC

[2044] Leo F. Marnell, Behnke Robert J., and Fred W. Allendorf. Genetic identification of cutthroat trout, Salmo clarki, in Glacier National Park, Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1830–1839, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-227.

Post:1987:SSM

[2045] John R. Post and Andrew B. Prankevicius. Size-selective mortality in young-of-the-year yellow perch (Perca flavescens): Evidence from otolith microstructure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1840– 1847, November 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-228.

Anders:1987:FIG

[2046] K. Anders and H. Möller. Food-induced granulomatosis in European smelt, Osmerus eperlanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44 (11):1848–1854, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f87-229.

Hudon:1987:EGP

[2047] Christiane Hudon. Ecology and growth of postlarval and juvenile lobster, Homarus americanus, off Îles de la Madeleine (Quebec). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1855–1869, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-230.

Peterman:1987:SME

[2048] Randall M. Peterman, Michael J. Bradford, and Gordon H. Kruse. Simulation model of English sole (Parophrys vetulus) population dynamics in Washington and Oregon coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1870–1878, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-231.

Peterman:1987:SPT

[2049] Randall M. Peterman and Michael J. Bradford. Statistical power of trends in fish abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11): 1879–1889, November 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-232.

Cabilio:1987:FCL

[2050] Paul Cabilio, David L. DeWolfe, and Graham R. Daborn. Fish catches and long-term tidal cycles in Northwest Atlantic fisheries: a nonlinear regression approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11): 1890–1897, November 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-233.

Vanderploeg:1987:CCP

[2051] Henry A. Vanderploeg, Brian J. Eadie, James R. Liebig, Stephen J. Tarapchak, and Rebecca M. Glover. Contribution of calcite to the particle-size spectrum of Lake Michigan seston and its interactions with the plankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1898–1914, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-234.

Dickie:1987:CAR

[2052] L. M. Dickie and P. R. Boudreau. Comparison of acoustic reflections from spherical objects and fish using a dual-beam echosounder. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1915–1921, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-235.

Campana:1987:OML

[2053] Steven E. Campana, Jacques A. Gagné, and Jean Munro. Otolith microstructore of larval herring (Clupea harengus): Image or reality? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1922–1929, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-236.

VanDerKraak:1987:EMS

[2054] G. Van Der Kraak, E. M. Donaldson, H. M. Dye, G. A. Hunter, J. E. Rivier, and W. W. Vale. Effects of mammalian and salmon gonadotropinreleaslng hormones and analogues on plasma gonadotropin levels and ovulation in Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1930–1935, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-237.

Chambers:1987:SAM

[2055] R. Christopher Chambers and William C. Leggett. Size and age at metamorphosis in marine fishes: an analysis of laboratory-reared winter flounder (Pseudopleutonectes americanus) with a review of variation in other species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11): 1936–1947, November 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-238.

Hill:1987:ATS

[2056] A. R. Hill and J. Warwick. Ammonium transformations in springwater within the riparian zone of a small woodland stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1948–1956, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-239.

Fleming:1987:ETF

[2057] Ian A. Fleming and Susan Ng. Evaluation of techniques for fixing, preserving, and measuring salmon eggs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1957–1962, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-240.

Quinn:1987:HWS

[2058] T. P. Quinn, C. C. Wood, L. Margolis, B. E. Riddell, and K. D. Hyatt. Homing in wild sockeye salmon (Oncorhynchus nerka) populations as inferred from differences in parasite prevalance and allozyme allele frequencies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1963–1971, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-241.

Dwyer:1987:FHD

[2059] Deborah A. Dwyer, Kevin M. Bailey, and Patricia A. Livingston. Feeding habits and daily ration of walleye pollock (Theragra chalcogramma) in the Eastern Bering Sea, with special reference to cannibalism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1972–1984, November 1987. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-242.

Machado:1987:HPE

[2060] J. P. Machado, D. L. Garling, Jr., N. R. Kevern, A. L. Trapp, and T. G. Bell. Histopathoiogy and the pathogenesis of embolism (gas bubble disease) in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1985–1994, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-243.

Bowen:1987:DPR

[2061] Stephen H. Bowen. Dietary protein requirements of fishes — a reassessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):1995–2001, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-244.

Mohn:1987:SCB

[2062] R. K. Mohn and R. W. Elner. A simulation of the Cape Breton snow crab, Chionoecetes opilio, fishery for testing the robustness of the Leslie method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2002–2008, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-245.

Runge:1987:MEP

[2063] Jeffrey A. Runge. Measurement of egg production rate of Calanus finmarchicus from preserved samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2009–2012, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-246.

Pepin:1987:PLF

[2064] P. Pepin, S. Pearre, Jr., and J. A. Koslow. Predation on larval fish by Atlantic mackerel Scomber scombrus, with a comparison of predation by zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11): 2012–2018, November 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-247.

Bowman:1987:TSF

[2065] G. T. Bowman, B. A. Schuknecht, J. E. Wilken, and W. C. Sonzogni. A time-saving filtration system for nutrient analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2018–2021, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-248.

Nash:1987:SVG

[2066] Richard D. M. Nash, John J. Magnuson, Clarence S. Clay, and Timothy K. Stanton. A synoptic view of the Gulf Stream front with 70-kHz sonar: Taking advantage of a closer look. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2022–2024, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-249.

Cattaneo:1987:SDP

[2067] Antonella Cattaneo. Size distribution in periphyton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2025–2028, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-250.

Confer:1987:IPT

[2068] John L. Confer and Gary J. Lake. Influence of prey type on growth of young yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2028–2033, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-251.

Getchell:1987:EVN

[2069] R. G. Getchell. Effects of V-notching on the lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(11):2033–2037, November 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-252.

Lean:1987:OSN

[2070] D. R. S. Lean. Overview of studies on the nutrient status of Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2042–2046, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-253.

Simons:1987:SCU

[2071] T. J. Simons and W. M. Schertzer. Stratification, currents, and upwelling in Lake Ontario, summer 1982. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2047–2058, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-254.

Stevens:1987:RLO

[2072] R. J. J. Stevens and M. A. Neilson. Response of Lake Ontario to reductions in phosphorus load, 1967–82. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2059–2068, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-255.

Lean:1987:PDL

[2073] D. R. S. Lean, A. A. Abbott, and F. R. Pick. Phosphorus deficiency of Lake Ontario plankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12): 2069–2076, December 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-256.

Lean:1987:SPK

[2074] D. R. S. Lean and R. L. Cuhel. Subcellular phosphorus kinetics for Lake Ontario plankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12): 2077–2086, December 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-257.

Pick:1987:IAP

[2075] F. R. Pick. Interpretations of alkaline phosphatase activity in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2087–2094, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-258.

Pick:1987:CPC

[2076] F. R. Pick. Carbohydrate and protein content of Lake Seston in relation to plankton nutrient deficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2095–2101, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-259.

Cuhel:1987:PSL

[2077] Russell L. Cuhel and David R. S. Lean. Protein synthesis by lake plankton measured using in situ carbon dioxide and sulfate assimilation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2102–2117, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f87-260.

Cuhel:1987:ILI

[2078] Russell L. Cuhel and David R. S. Lean. Influence of light intensity, light quality, temperature, and daylength on uptake and assimilation of carbon dioxide and sulfate by lake plankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2118–2132, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-261.

Lean:1987:NTL

[2079] D. R. S. Lean and R. Knowles. Nitrogen transformations in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2133–2143, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-262.

Putt:1987:PDE

[2080] M. Putt, G. P. Harris, and R. L. Cuhel. Photoinhibition of DCMUenhanced fluorescence in Lake Ontario phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2144–2154, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-263.

Gray:1987:DBN

[2081] I. M. Gray. Differences between nearshore and offshore phytoplankton communities in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2155–2163, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-264.

Pick:1987:PNB

[2082] F. R. Pick and D. A. Caron. Picoplankton and nanoplankton biomass in Lake Ontario: Relative contribution of phototrophic and heterotrophic communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2164– 2172, December 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-265.

Leppard:1987:CCP

[2083] Gary G. Leppard, Dina Urciuoli, and F. R. Pick. Characterization of cyanobacterial picoplankton in Lake Ontario by transmission electron microscopy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2173–2177, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f87-266.

Taylor:1987:ZSS

[2084] W. D. Taylor, H-J. Fricker, and D. R. S. Lean. Zooplankton seasonal succession in Lake Ontario at northshore, midlake, and southshore stations in 1982, and a comparison with 1970. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2178–2184, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-267.

Taylor:1987:SVD

[2085] William D. Taylor and Miriam L. Heynen. Seasonal and vertical distribution of ciliophora in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2185–2191, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-268.

Neilson:1987:SHN

[2086] M. A. Neilson and R. J. J. Stevens. Spatial heterogeneity of nutrients and organic matter in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2192–2203, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-269.

Manning:1987:PII

[2087] P. C. Manning. Phosphate ion interactions at the sediment–water interface in Lake Ontario: Relationship to sediment adsorption capacities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2204–2211, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f87-270.

Painter:1987:RCB

[2088] D. S. Painter and G. Kamaitis. Reduction of Cladophora biomass and tissue phosphorus in Lake Ontario, 1972–83. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2212–2215, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-271.

Zimmermann:1987:DTT

[2089] G. M. Zimmermann, D. R. S. Lean, and M. N. Charlton. Differential thermal and thermogravimetric analysis of sediment-forming materials from Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12): 2216–2224, December 1987. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f87-272.

Rodgers:1987:TOF

[2090] G. Keith Rodgers. Time of onset of full thermal stratification in Lake Ontario in relation to lake temperatures in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2225–2229, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f87-273.

Lean:1987:LOL

[2091] D. R. S. Lean, H-J. Fricker, M. N. Charlton, R. L. Cuhel, and F. R. Pick. The Lake Ontario life support system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 44(12):2230–2240, December 1987. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f87-274.

Fevolden:1988:GPS

[2092] Svein Erik Fevolden and Tore Haug. Genetic population structure of Atlantic halibut, Hippoglossus hippoglossus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):2–7, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-001.

Peterman:1988:CEL

[2093] Randall M. Peterman, Michael J. Bradford, Nancy C. H. Lo, and Richard D. Methot, Jr. Contribution of early life stages to interannual variability in recruitment of northern anchovy (Engraulis mordax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):8–16, January 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-002.

Osenberg:1988:GPB

[2094] Craig W. Osenberg, Earl E. Werner, Gary G. Mittelbach, and Donald J. Hall. Growth patterns in bluegill (Lepomis macrochirus) and pumpkinseed (L. gibbosus) sunfish: Environmental variation and the importance of ontogenetic niche shifts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (1):17–26, January 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-003.

Naesje:1988:ISC

[2095] Tor F. Næsje and Bror Jonsson. Impacted stress: a causal agent of reduced whitefish (Coregonus iavaretus) egg incubation time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):27–31, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-004.

Spry:1988:RCD

[2096] D. J. Spry, P. V. Hodson, and C. M. Wood. Relative contributions of dietary and waterborne zinc in the rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):32–41, January 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-005.

Clarke:1988:RRM

[2097] W. Craig Clarke and Richard J. Beamish. Response of recently metamorphosed anadromous parasitic lamprey (Lampetra tridentata) to confinement in fresh water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (1):42–47, January 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-006.

Frank:1988:IDF

[2098] Kenneth T. Frank. Independent distributions of fish larvae and their prey: Natural paradox or sampling artifact? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):48–59, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-007.

Saunders:1988:ECD

[2099] R. L. Saunders and E. B. Henderson. Effects of constant day length on sexual maturation and growth of Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):60–64, January 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-008.

Jansen:1988:EWA

[2100] Wolfgang A. Jansen and John H. Gee. Effects of water acidity on swimbladder function and swimming in the fathead minnow, Pimephales promelas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):65–77, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-009.

Seeb:1988:GVP

[2101] Lisa Wishard Seeb and Donald R. Gunderson. Genetic variation and population structure of Pacific Ocean perch (Sebastes alutus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):78–88, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-010.

Hisden:1988:SFR

[2102] Mskael Hisdén. Significance of the functional response of predators to changes in prey abundance in multispecies virtual population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):89–96, January 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-011.

Jackson:1988:MPR

[2103] Togwell A. Jackson. The mercury problem in recently formed reservoirs of Northern Manitoba (Canada): Effects of impoundment and other factors on the production of methyl mercury by microorganisms in sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):97–121, January 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-012.

Berg:1988:GFG

[2104] W. J. Berg and G. A. E. Gall. Gene flow and genetic differentiation among California coastal rainbow trout populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):122–131, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-013.

Ankley:1988:EDP

[2105] Gerald T. Ankley and Vicki S. Blazer. Effects of diet on PCB-induced changes in xenobiotic metabolism in the liver of channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):132–137, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-014.

Boisclair:1988:SEE

[2106] D. Boisclair and W. C. Leggett. An in situ experimental evaluation of the Elliott and Persson and the eggers models for estimating fish daily ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):138–145, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-015.

Agellon:1988:PRG

[2107] L. B. Agellon, C. J. Emery, J. M. Jones, S. L. Davies, A. D. Dingle, and T. T. Chen. Promotion of rapid growth of rainbow trout (Salmo gairdneri) by a recombinant fish growth hormone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):146–151, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-016.

Beamish:1988:RDB

[2108] R. J. Beamish and C. A. McFarlane. Resident and dispersal behavior of adult sablefish (Anaplopoma fimbria) in the slope waters off Canada’s West Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):152–164, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
017.

Scavia:1988:DLM

[2109] Donald Scavia, Gregory A. Lang, and James F. Kitchell. Dynamics of Lake Michigan plankton: a model evaluation of nutrient loading, competition, and predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1): 165–177, January 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-018.

Caputi:1988:FAT

[2110] N. Caputi. Factors affecting the time series bias in stock–recruitment relationships and the interaction between time series and measurement error bias. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):178–184, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-019.

Kope:1988:EIC

[2111] Robert G. Kope. Effects of initial conditions on Monte Carlo estimates of bias in estimating functional relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):185–187, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-020.

Evans:1988:RSC

[2112] Barbara I. Evans and W. John O’Brien. A reevaluation of the search cycle of planktivorous Arctic graylings Thymallus arcticus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):187–192, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-021.

Wallace:1988:IEA

[2113] Jeff C. Wallace and Tor G. Heggberget. Incubation of eggs of Atlantic salmon (Salmo salar) from different Norwegian streams at temperatures below 1^{o}C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(1):193–196, January 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
022.

Prairie:1988:DPD

[2114] Yves T. Prairie and Jacob Kalff. Dissolved phosphorus dynamics in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2): 200–209, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-023.

Prairie:1988:PPD

[2115] Yves T. Prairie and Jacob Kalff. Particulate phosphorus dynamics in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2): 210–215, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-024.

Duarte:1988:ILM

[2116] Carlos M. Duarte and Jacob Kalff. Influence of lake morphometry on the response of submerged macrophytes to sediment fertilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):216–221, February 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-025.

Niimi:1988:IMW

[2117] A. J. Niimi and B. G. Oliver. Influence of molecular weight and molecular volume on dietary absorption efficiency of chemicals by fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):222–227, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-026.

Evans:1988:CFB

[2118] R. Douglas Evans, Daniel Andrews, and R. Jack Cornell. Chemical fractionation and bioavailability of cobalt-60 to benthic deposit-feeders. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):228–236, February 1988.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-027.

Taylor:1988:AVR

[2119] Eric B. Taylor. Adaptive variation in rheotactic and agonistic behavior in newly emerged fry of Chinook salmon, Oncorhynchus tshawytscha, from ocean- and stream-type populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):237–243, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-028.

Reid:1988:ECC

[2120] S. D. Reid and D. G. McDonald. Effects of cadmium, copper, and low pH on ion fluxes in the rainbow trout, Salmo gairdneri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):244–253, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-029.

Nalewajko:1988:EAP

[2121] C. Nalewajko and M. A. O’Mahony. Effects of acid pH shock on phosphate concentrations and microbial phosphate uptake in an acidifying and a circumneutral lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (2):254–260, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-030.

Bothwell:1988:GRR

[2122] Max L. Bothwell. Growth rate responses of lotic periphytic diatoms to experimental phosphorus enrichment: The influence of temperature and light. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):261–270, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-031.

Carrick:1988:RLM

[2123] Hunter J. Carrick and Rex L. Lowe. Response of Lake Michigan benthic algae to in situ enrichment with Si, N, and P. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):271–279, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-032.

Jackson:1988:SDP

[2124] John K. Jackson and Vincent H. Resh. Sequential decision plans in monitoring benthic macroinvertebrates: Cost savings, classification accuracy, and development of plans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (2):280–286, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-033.

Doe:1988:IPA

[2125] K. G. Doe, W. R. Ernst, W. R. Parker, G. R. J. Julien, and P. A. Hennigar. Influence of pH on the acute lethality of fenitrothion, 2,4-D, and aminocarb and some pH-altered sublethal effects of aminocarb on rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (2):287–293, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-034.

Brenner:1988:RBC

[2126] Mark Brenner and Michael W. Binford. Relationships between concentrations of sedimentary variables and trophic state in Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):294–300, February 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-035.

Christie:1988:MOT

[2127] Gavin C. Christie and Henry A. Regier. Measures of optimal thermal habitat and their relationship to yields for four commercial fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):301–314, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-036.

Morin:1988:AMS

[2128] Antoine Morin, Michel Constantin, and Robert Henry Peters. Allometric models of simuliid growth rates and their use for estimation of production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):315–324, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-037.

Alberts:1988:DCD

[2129] James J. Alberts, John W. Bowling, and Mark Emmons. Dissolved carbohydrate distribution and dynamics in two Southeastern United States reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):325–332, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-038.

Welch:1988:MLM

[2130] D. W. Welch and R. P. Foucher. A maximum likelihood methodology for estimating length-at-maturity with application to Pacific cod (Gadus macrocephalus) population dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):333–343, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-039.

Gibson:1988:ISR

[2131] R. J. Gibson and R. A. Myers. Influence of seasonal river discharge on survival of juvenile Atlantic salmon Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):344–348, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-040.

Levy:1988:PCN

[2132] E. M. Levy and K. Lee. Potential contribution of natural hydrocarbon seepage to benthic productivity and the fisheries of Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):349–352, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-041.

Fletcher:1988:EAP

[2133] Garth L. Fletcher, Margaret A. Shears, Madonna J. King, Peter L. Davies, and Choy L. Hew. Evidence for antifreeze protein gene transfer in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):352–357, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-042.

Knight:1988:BHU

[2134] Ivor T. Knight, D. Jay Grimes, and Rita R. Colwell. Bacterial hydrolysis of urea in the tissues of carcharhinid sharks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):357–360, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-043.

Northcote:1988:FSF

[2135] T. G. Northcote. Fish in the structure and function of freshwater ecosystems: a “Top-down” view. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (2):361–379, February 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-044.

Anonymous:1988:LCLa

[2136] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):380–382, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-045.

Anonymous:1988:CRPa

[2137] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(2):383, February 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-046.

Hay:1988:EIP

[2138] D. E. Hay, J. R. Brett, E. Bilinski, D. T. Smith, E. M. Donaldson, G. A. Hunter, and A. V. Solmie. Experimental impoundments of prespawning Pacific herring (Clupea harengus pallasi): Effects of feeding and density on maturation, growth, and proximate analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):388–398, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-047.

Hay:1988:MFP

[2139] D. E. Hay and J. R. Brett. Maturation and fecundity of Pacific herring (Clupea harengus pallasi): an experimental study with comparisons to natural populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (3):399–406, March 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-048.

Cornett:1988:PRT

[2140] R. J. Cornett and L. Chant. Pu residence times in freshwaters and accumulation in Shield lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):407–415, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-049.

Harding:1988:DHA

[2141] Gareth C. Harding and Ronald W. Trites. Dispersal of Homarus americanus larvae in the Gulf of Maine from Browns Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):416–425, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-050.

Jakobsen:1988:EPR

[2142] Per Johan Jakobsen, Geir Helge Johnsen, and Petter Larsson. Effects of predation risk and parasitism on the feeding ecology, habitat use, and abondance of lacustrine threespine stickleback (Gasterosteus aculeatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):426–431, March 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-051.

Mulligan:1988:USC

[2143] T. J. Mulligan, S. McKinnell, and C. C. Wood. Uncertainty in stock composition estimates of oceanic steelhead trout using electrophoretic characteristics: Comments on a recent study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):432–442, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-052.

Cunjak:1988:PCO

[2144] Richard A. Cunjak. Physiological consequences of overwintering in streams: The cost of acclimitization? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):443–452, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-053.

Nurnberg:1988:PPR

[2145] Gertrud K. Nürnberg. Prediction of phosphorus release rates from total and reductant-soluble phosphorus in anoxic lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):453–462, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-054.

Dempson:1988:IVE

[2146] J. B. Dempson, E. Verspoor, and J. Hammar. Intrapopulation variation of the esterase-2 polymorphism in the serum of anadromous Arctic charr, Salvelinus alpinus, from a Northern Labrador river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):463–468, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-055.

Earle:1988:FAS

[2147] J. C. Earle, H. C. Duthie, W. A. Glooschenko, and P. B. Hamilton. Factors affecting the spatial distribution of diatoms on the surface sediments of three Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):469–478, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-056.

Doyon:1988:VCN

[2148] Jean-François Doyon, John A. Downing, and Etienne Magnin. Variation in the condition of northern pike, Esox lncius. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):479–483, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-057.

Stobo:1988:MAH

[2149] Wayne T. Stobo, John D. Neilson, and Patricia G. Simpson. Movements of Atlantic halibut (Hippoglossus hippoglossus) in the Canadian North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):484–491, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-058.

Steedman:1988:MAI

[2150] Robert J. Steedman. Modification and assessment of an index of biotic integrity to quantify stream quality in Southern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):492–501, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-059.

Holtby:1988:ELS

[2151] L. Blair Holtby. Effects of logging on stream temperatures in Carnation Creek, British Columbia, and associated impacts on the Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (3):502–515, March 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-060.

Winters:1988:MCS

[2152] G. H. Winters and E. L. Dalley. Meristic composition of sand lance (Ammodytes spp.) in Newfoundland waters with a review of species designations in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):516–529, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-061.

Walters:1988:EDE

[2153] Carl J. Walters, Jeremy S. Collie, and Timothy Webb. Experimental designs for estimating transient responses to management disturbances. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):530–538, March 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-062.

Prager:1988:SVV

[2154] Michael H. Prager and Alec D. MacCall. Sensitivities and variances of virtual population analysis as applied to the mackerel, Scomber japonicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):539–547, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-063.

Beyer:1988:DTE

[2155] Jean M. Beyer, Gino Lucchetti, and Gerry Gray. Digestive tract evacuation in northern squawfish (Ptychocheilus oregonensis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):548–553, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-064.

Armstrong:1988:BES

[2156] Michael J. Armstrong and Peter A. Shelton. Bias in estimation of stockrecruit function parameters caused by nonrandom environmental variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):554–557, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-065.

Smith:1988:CPE

[2157] Barry D. Smith. Comparison of productivity estimates for Laminaria in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):557–562, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-066.

Welch:1988:SSC

[2158] Harold E. Welch, Martin A. Bergmann, John K. Jorgenson, and William Burton. A subice suction corer for sampling epontic ice algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):562–568, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-067.

Montevecchi:1988:MPA

[2159] W. A. Montevecchi, D. K. Cairns, and V. L. Birt. Migration of postsmolt Atlantic salmon, Salmo salar, off Northeastern Newfoundland, as inferred by tag recoveries in a seabird colony. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):568–571, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-068.

Trimbee:1988:DLO

[2160] Annette M. Trimbee and E. E. Prepas. Dependence of lake oxygen depletion rates on maximum oxygen storage in a partially meromictic lake in Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(3):571–576, March 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-069.

Bowering:1988:AMC

[2161] W. R. Bowering. An analysis of morphometric characters of Greenland halibut (Reinhardtius hippoglossoides) in the Northwest Atlantic using a multivariate analysis of covariance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):580–585, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-070.

Bailey:1988:ESC

[2162] R. E. Bailey, L. Margolis, and C. Groot. Estimating stock composition of migrating juvenile Fraser River (British Columbia) sockeye salmon, Oncorhynchus nerka, using parasites as natural tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):586–591, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-071.

Magnhagen:1988:PRF

[2163] Carin Magnhagen. Predation risk and foraging in juvenile pink (Oncorhynchus gorbuscha) and chum salmon (O. keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):592–596, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-072.

Rose:1988:HSC

[2164] G. A. Rose and W. C. Leggett. Hydroacoustic signal classification of fish schools by species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (4):597–604, April 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-073.

MacLennan:1988:PAF

[2165] D. N. MacLennan and I. G. MacKenzie. Precision of acoustic fish stock estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):605–616, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-074.

Johnson:1988:PAP

[2166] M. G. Johnson. Production by the amphipod Pontoporeia hoyi in South Bay, Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):617–624, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
075.

deMarch:1988:ATB

[2167] B. G. E. de March. Acute toxicity of binary mixtures of five cations (Cu^{2+}, Cd^{2+}, Zn^{2+}, Mg^{2+}, and K^{+}) to the freshwater amphipod Gammarus lacustris (Sars): Alternative descriptive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):625–633, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-076.

Ramamoorthy:1988:EPS

[2168] S. Ramamoorthy. Effect of pH on speciation and toxicity of aluminum to rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):634–642, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-077.

Billington:1988:MDV

[2169] Neil Billington and Paul D. N. Hebert. Mitochondrial DNA variation in Great Lakes walleye (Stizostedion vitreum) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):643–654, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-078.

Gurtz:1988:HRI

[2170] Martin E. Gurtz, G. Richard Marzolf, Keith T. Killingbeck, David L. Smith, and J. Vaun McArthur. Hydrologic and riparian influences on the import and storage of coarse particulate organic matter in a prairie stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):655–665, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-079.

Krise:1988:ELL

[2171] William F. Krise and James W. Meade. Effects of low-level gas supersaturation on lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):666–674, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-080.

Majkowski:1988:TRR

[2172] Jacek Majkowski, William S. Hearn, and Ronald L. Sandland. A tag– release/recovery method for predicting the effect of changing the catch of one component of a fishery upon the remaining components. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):675–684, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-081.

Vinyard:1988:ESP

[2173] Gary L. Vinyard, Ray W. Drenner, Moshe Gophen, Utsa Pollingher, Dana L. Winkelman, and K. David Hambright. An experimental study of the plankton community impacts of two omnivorous filter-feeding cichlids, Tilapia galilaea and Tilapia aurea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):685–690, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-082.

Deutschman:1988:SPT

[2174] Mark R. Deutschman and John J. Peterka. Secondary production of tiger salamanders (Ambystoma tigrinum) in three North Dakota prairie lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):691–697, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-083.

Bigler:1988:FSD

[2175] Brian Bigler. Focal scale damage among chum salmon (Oncorhynchus keta) of Hokkaido, Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (4):698–704, April 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-084.

Morin:1988:EBF

[2176] Antoine Morin, Christian Back, Alain Chalifour, Jacques Boisvert, and Robert H. Peters. Effect of black fly ingestion and assimilation on seston transport in a Quebec lake outlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):705–714, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-085.

Stenson:1988:APC

[2177] G. B. Stenson and R. A. Myers. Accuracy of pup classifications and its effect on population estimates in the hooded seal (Cystophora cristata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):715–719, April 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-086.

Groeger:1988:PCM

[2178] Alan W. Groeger and Bruce L. Kimmel. Photosynthetic carbon metabolism by phytoplankton in a nitrogen-limited reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):720–730, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-087.

Welch:1988:ECD

[2179] Harold E. Welch, John K. Jorgenson, and Martin F. Curtis. Emergence of Chironomidae (Diptera) in fertilized and natural lakes at Saqvaqjuac, N.W.T. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):731–737, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-088.

Welch:1988:MAE

[2180] Harold E. Welch, John K. Jorgenson, and Martin F. Curtis. Measuring abundance of emerging Chironomidae (Diptera): Experiments on trap size and design, set duration, and transparency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):738–741, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-089.

Ringler:1988:VDS

[2181] Neil H. Ringler and James D. Hall. Vertical distribution of sediment and organic debris in Coho salmon (Oncorhynchus kisutch) redds in three small Oregon streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (4):742–747, April 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-090.

Barrett:1988:SEI

[2182] Bruce A. Barrett and Brian A. McKeown. Sustained exercise increases plasma growth hormone concentrations in two anadromous salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):747–749, April 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-091.

Jones:1988:BAA

[2183] Keith A. Jones and Toshiaki J. Hara. Behavioral alterations in Arctic char (Salvelinus alpinus) briefly exposed to sublethal chlorine levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(4):749–752, April 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-092.

Chopin:1988:RCD

[2184] T. Chopin, J. D. Pringle, and R. E. Semple. Reproductive capacity of dragraked and non-dragraked Irish moss (Chondrus crispus Stackhouse) beds in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):758–766, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-093.

Pringle:1988:IHI

[2185] J. D. Pringle and R. E. Semple. Impact of harvesting on Irish moss (Chondrus crispus Stackhouse) frond size-class structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):767–773, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-094.

Heise:1988:PHL

[2186] Brian A. Heise, John F. Flannagan, and Terry D. Galloway. Production of Hexagenia limbata (Serville) and Ephemera simulans Walker (Ephemeroptera) in Dauphin Lake, Manitoba, with a note on weight loss due to preservatives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (5):774–781, May 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-095.

Lane:1988:IDM

[2187] Daniel E. Lane. Investment decision making by fishermen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):782–796, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-096.

Nelson:1988:FIE

[2188] K. Nelson, D. Hedgecock, and W. Borgeson. Factors influencing egg extrusion in the American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):797–804, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-097.

Nelson:1988:ERU

[2189] K. Nelson, D. Hedgecock, and W. Borgeson. Effects of reproduction upon molting and growth in female American lobsters (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):805–821, May 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-098.

Hardy:1988:HLV

[2190] John T. Hardy, James A. Coley, Liam D. Antrim, and Steven L. Kiesser. A hydrophobic large-volume sampler for collecting aquatic surface microlayers: Characterization and comparison with the glass plate method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):822–826, May 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-099.

Peterson:1988:CPE

[2191] Mark S. Peterson. Comparative physiological ecology of centrarchids in hyposaline environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (5):827–833, May 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-100.

Hankin:1988:ETF

[2192] David G. Hankin and Gordon H. Reeves. Estimating total fish abundance and total habitat area in small streams based on visual estimation methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):834–844, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-101.

Heggberget:1988:TSN

[2193] Tor G. Heggberget. Timing of spawning in Norwegian Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):845–849, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
102.

Fried:1988:IFB

[2194] Stephen M. Fried and Ray Hilborn. Inseason forecasting of Bristol Bay, Alaska, sockeye salmon (Oncorhynchus nerka) abundance using Bayesian probability theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):850– 855, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-103.

Kyle:1988:DDT

[2195] G. B. Kyle, J. P. Koenings, and B. M. Barrett. Density-dependent, trophic level responses to an introduced run of sockeye salmon (Oncorhynchus nerka) at Frazer Lake, Kodiak Island, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):856–867, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-104.

Hudon:1988:PGL

[2196] Christiane Hudon and Pierre Fradette. Planktonic growth of larval lobster (Homarus americanus) off Îles de la Madeleine (Quebec), Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):868–878, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-105.

Pepin:1988:LSF

[2197] P. Pepin, J. A. Koslow, and S. Pearre, Jr. Laboratory study of foraging by Atlantic mackerel, Scomber scombrus, on natural zooplankton assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):879–887, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-106.

Hall:1988:AHS

[2198] D. L. Hall, R. Hilborn, M. Stocker, and C. J. Walters. Alternative harvest strategies for Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):888–897, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-107.

Ulvund:1988:FAC

[2199] Kjell Arne Ulvund and Otto Grahl-Nielsen. Fatty acid composition in eggs of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):898–901, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-108.

Kao:1988:JAC

[2200] Ming H. Kao and Garth L. Fletcher. Juvenile Atlantic cod (Gadus morhua) can be more freeze resistant than adults. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):902–905, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-109.

Fechhelm:1988:WAR

[2201] Robert G. Fechhelm and David B. Fissel. Wind-aided recruitment of Canadian Arctic cisco (Coregonus autumnalis) into Alaskan waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):906–910, May 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-110.

Prepas:1988:CSE

[2202] E. E. Prepas, M. E. Dunnigan, and A. M. Trimbee. Comparison of in situ estimates of chlorophyll a obtained with whatman GF/F and GF/C glass-fiber filters in mesotrophic to hypereutrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):910–914, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-111.

Watson:1988:CPN

[2203] Susan Watson and Edward McCauley. Contrasting patterns of netand nanoplankton production and biomass among lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):915–920, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-112.

Holzkamm:1988:PFL

[2204] Tim Holzkamm and Michael McCarthy. Potential fishery for lake sturgeon (Acipenser fulvescens) as indicated by the returns of the Hudson’s Bay Company Lac la Pluie District. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(5):921–923, May 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-113.

Stocker:1988:EFP

[2205] M. Stocker and D. J. Noakes. Evaluating forecasting procedures for predicting Pacific herring (Clupea harengus pallasi) recruitment in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):928–935, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-114.

Francis:1988:GPE

[2206] R. I. C. C. Francis. Are growth parameters estimated from tagging and age–length data comparable? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (6):936–942, June 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-115.

Wieser:1988:GRG

[2207] W. Wieser, H. Forstner, F. Schiemer, and W. Mark. Growth rates and growth efficiencies in larvae and juveniles of Rutilus rutilus and other cyprinid species: Effects of temperature and food in the laboratory and in the field. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):943–950, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-116.

Houlihan:1988:GRP

[2208] D. F. Houlihan, S. J. Hall, C. Gray, and B. S. Noble. Growth rates and protein turnover in Atlantic cod, Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):951–964, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-117.

Nemec:1988:UBA

[2209] A. F. L. Nemec and R. O. Brinkhurst. Using the bootstrap to assess statistical significance in the cluster analysis of species abundance data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):965–970, June 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-118.

Nemec:1988:FMS

[2210] A. F. L. Nemec and R. O. Brinkhurst. The fowlkes–Mallows statistic and the comparison of two independently determined dendrograms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):971–975, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-119.

Cyr:1988:ERP

[2211] Hélène Cyr and John A. Downing. Empirical relationships of phytomacrofaunal abundance to plant biomass and macrophyte bed characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):976–984, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-120.

Ryg:1988:TST

[2212] Morten Ryg, Thomas G. Smith, and Nils Are Øritsland. Thermal significance of the topographical distribution of blubber in ringed seals (Phoca hispida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):985–992, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-121.

Gregory:1988:RSM

[2213] R. S. Gregory and P. M. Powles. Relative selectivities of Miller highspeed samplers and light traps for collecting ichthyoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):993–998, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-122.

Magnan:1988:IBB

[2214] Pierre Magnan. Interactions between brook charr, Salvelinus fontinalis, and nonsalmonid species: Ecological shift, morphological shift, and their impact on zooplankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):999–1009, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-123.

Chambers:1988:USA

[2215] P. A. Chambers and E. E. Prepas. Underwater spectral attenuation and its effect on the maximum depth of angiosperm colonization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1010–1017, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-124.

Ihssen:1988:BGD

[2216] P. E. Ihssen, J. M. Casselman, G. W. Martin, and R. B. Phillips. Biochemical genetic differentiation of lake trout (Salvelinus namaycush) stocks of the Great Lakes region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1018–1029, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-125.

Conley:1988:SPF

[2217] Daniel J. Conley, Michael A. Quigley, and Claire L. Schelske. Silica and phosphorus flux from sediments: Importance of internal recycling in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1030–1035, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
126.

Fisher:1988:GJC

[2218] J. P. Fisher and W. G. Pearcy. Growth of juvenile Coho salmon (Oncorhynchus kisutch) off Oregon and Washington, USA, in years of differing coastal upwelling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (6):1036–1044, June 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-127.

Giles:1988:ACR

[2219] Michael A. Giles. Accumulation of cadmium by rainbow trout, Salmo gairdneri, during extended exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1045–1053, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-128.

Deriso:1988:DAS

[2220] Richard B. Deriso and Ana M. Parma. Dynamics of age and size for a stochastic population model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (6):1054–1068, June 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-129.

Greene:1988:CSF

[2221] Charles H. Greene and Michael R. Landry. Carnivorous suspension feeding by the subarctic calanoid copepod Neocalanus cristatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1069–1074, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-130.

Hallerman:1988:DLP

[2222] Eric M. Hallerman and Jacques S. Beckmann. DNA-level polymorphism as a tool in fisheries science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (6):1075–1087, June 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-131.

Eckmann:1988:ECD

[2223] Reiner Eckmann, Ursula Gaedke, and Hans Johst Wetzlar. Effects of climatic and density-dependent factors on year-class strength of Coregonus lavaretus in Lake Constance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (6):1088–1093, June 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-132.

Taylor:1988:CDP

[2224] W. D. Taylor, J. W. Barko, and W. F. James. Contrasting diel patterns of vertical migration in the dinoflagellate Ceratium hirundinella in relation to phosphorus supply in a North temperate reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1093–1098, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-133.

Cury:1988:PSN

[2225] Philippe Cury. Pressions sélectives et nouveautés évolutives : une hypothèse pour comprendre certains aspects des fluctuations à long terme des poissons pélagiques côtiers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(6):1099–1107, June 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-134.

Ward:1988:LHS

[2226] B. R. Ward and P. A. Slaney. Life history and smolt-to-adult survival of Keogh River steelhead trout (Salmo gairdneri) and the relationship to smolt size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1110–1122, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
135.

Warwick:1988:MDC

[2227] W. F. Warwick and N. A. Tisdale. Morphological deformities in Chironomus, Cryptochironomus, and Procladius larvae (Diptera: Chironomidae) from two differentially stressed sites in Tobin Lake, Saskatchewan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1123–1144, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-136.

Fogarty:1988:TSM

[2228] Michael J. Fogarty. Time series models of the Maine lobster fishery: The effect of temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (7):1145–1153, July 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-137.

Andersson:1988:ELD

[2229] Paul Andersson and Hans Borg. Effects of liming on the distribution of cadmium in water, sediment, and organisms in a Swedish lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1154–1162, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-138.

Heggenes:1988:EEI

[2230] Jan Heggenes. Effect of experimentally increased intraspecific competition on sedentary adult brown trout (Salmo trutta) movement and stream habitat choice. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (7):1163–1172, July 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-139.

deLafontaine:1988:PJL

[2231] Yves de Lafontaine and W. C. Leggett. Predation by jellyfish on larval fish: an experimental evaluation employing in situ enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1173–1190, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-140.

Dunbrack:1988:FJC

[2232] Robert L. Dunbrack. Feeding of juvenile Coho salmon (Oncorhynchus kisutch): Maximum appetite, sustained feeding rate, appetite return, and body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1191–1196, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-141.

Tam:1988:GIG

[2233] W. H. Tam, J. N. Fryer, I. Ali, M. R. Dallaire, and B. Valentine. Growth inhibition, gluconeogenesis, and morphometric studies of the pituitary and interrenal cells of acid-stressed brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1197–1211, July 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-142.

Fryer:1988:PCC

[2234] J. N. Fryer, W. H. Tam, B. Valentine, and R. E. Tikkala. Prolactin cell cytology, plasma electrolytes, and whole-body sodium efflux in acid-stressed brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1212–1221, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-143.

Rose:1988:AOCa

[2235] G. A. Rose and W. C. Leggett. Atmosphere–ocean coupling in the Northern Gulf of St. Lawrence: Frequency-dependent wind-forced variations in nearshore sea temperatures and currents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1222–1233, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-144.

Rose:1988:AOCb

[2236] G. A. Rose and W. C. Leggett. Atmosphere–ocean coupling and Atlantic cod migrations: Effect of wind-forced variations in sea temperatures and currents or nearshore distributions and catch rates of Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1234–1243, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-145.

Thorne:1988:EED

[2237] Richard E. Thorne. An empirical evaluation of the duration-in-beam technique for hydroacoustic estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1244–1248, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-146.

Crittenden:1988:WDB

[2238] R. N. Crittenden, G. L. Thomas, D. A. Marino, and R. E. Thorne. A weighted duration-in-beam estimator for the volume sampled by a quantitative echo sounder. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (7):1249–1256, July 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-147.

Roesijadi:1988:IQM

[2239] G. Roesijadi, M. E. Unger, and J. E. Morris. Immunochemical quantification of metallothioneins of a marine mollusc. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1257–1263, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-148.

Whittier:1988:CBE

[2240] Thomas R. Whittier, Robert M. Hughes, and David P. Larsen. Correspondence between ecoregions and spatial patterns in stream ecosystems in Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1264–1278, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
149.

Swain:1988:ESV

[2241] D. P. Swain. Evidence of selection for vertebral number of fry in peamouth Mylocheilus caurinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1279–1290, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-150.

Schelske:1988:SRB

[2242] Claire L. Schelske, John A. Robbins, Wayne S. Gardner, Daniel J. Conley, and Richard A. Bourbonniere. Sediment record of biogeochemical responses to anthropogenic perturbations of nutrient cycles in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1291–1303, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-151.

Taugbol:1988:ESD

[2243] Trend Taugbøl, Jostein Skurdal, and Ragnvald Andersen. Ecological significance of differences in frequency of white fin margins among four brown trout (Salmo trutta) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1304–1309, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-152.

Fisher:1988:EIA

[2244] William S. Fisher and Mark Tamplin. Environmental influence on activities and foreign-particle binding by hemocytes of American oysters, Crassostrea virginica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (7):1309–1315, July 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-153.

Irvine:1988:PML

[2245] J. R. Irvine. Poststocking movements of large Coho salmon (Oncorhynchus kisutch) fingerlings introduced in February into a small coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1316– 1321, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-154.

Henderson:1988:RSS

[2246] M. A. Henderson and T. G. Northcote. Retinal structure of sympatric and allopatric populations of cutthroat trout (Salmo clarki clarki) and Dolly Varden char (Salvelinus malma) in relation to their spatial distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(7):1321–1326, July 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-155.

Blouw:1988:TVM

[2247] D. M. Blouw, S. D. Saxon, and E. M. P. Chadwick. Temporal variation of meristic traits within an Atlantic salmon (Salmo salar) stock, and implications for stock identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1330–1339, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-156.

Gunnerod:1988:OSR

[2248] T. B. Gunnerød, N. A. Hvidsten, and T. G. Heggberget. Open sea releases of Atlantic salmon smolts, Salmo salar, in Central Norway, 1973– 83. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1340–1345, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-157.

Bocking:1988:PFS

[2249] Robert C. Bocking and Randall M. Peterman. Preseason forecasts of sockeye salmon (Oncorhynchus nerka): Comparison of methods and economic considerations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8): 1346–1354, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-158.

Munkittrick:1988:GFE

[2250] K. R. Munkittrick and D. G. Dixon. Growth, fecundity, and energy stores of white sucker (Catostomus commersoni) from lakes containing elevated levels of copper and zinc. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (8):1355–1365, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-159.

Macdonald:1988:FET

[2251] J. Stevenson Macdonald, Colin D. Levings, Carey D. McAllister, U. H. M. Fagerlund, and J. R. McBride. A field experiment to test the importance of estuaries for Chinook salmon (Oncorhynchus tshawytscha) survival: Short-term results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (8):1366–1377, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-160.

Tan:1988:OIS

[2252] F. C. Tan, D. Cai, and D. L. Roddick. Oxygen isotope studies on sea scallops, Placopecten magellanicus, from Browns Bank, Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1378–1386, August 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-161.

Audet:1988:LTS

[2253] Céline Audet, R. Stephen Munger, and Chris M. Wood. Long-term sublethal acid exposure in rainbow trout (Salmo gairdneri) in soft water: Effects on ion exchanges and blood chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1387–1398, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-162.

Audet:1988:DRT

[2254] Céline Audet and Chris M. Wood. Do rainbow trout (Salmo gairdneri) acclimate to low pH? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8): 1399–1405, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-163.

Bergstedt:1988:ASL

[2255] Roger A. Bergstedt and Clifford P. Schneider. Assessment of sea lamprey (Petromyzon marinus) predation by recovery of dead lake trout (Salvelinus namaycush) from Lake Ontario, 1982–85. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1406–1410, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-164.

Dixit:1988:SCC

[2256] Sushil S. Dixit, Aruna S. Dixit, and R. Douglas Evans. Scaled chrysophytes (chrysophyceae) as indicators of pH in Sudbury, Ontario, lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1411–1421, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-165.

Robinson:1988:ODG

[2257] S. M. C. Robinson and D. M. Ware. Ontogenetic development of growth rates in larval Pacific herrings, Clupea harengus pallasi, measured with RNA–DNA ratios in the Strait of Georgia, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1422–1429, August 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-166.

Beckman:1988:CIC

[2258] Brian R. Beckman and Waldo S. Zaugg. Copper intoxication in Chinook salmon (Oncorhynchus tshawytscha) induced by natural springwater: Effects on gill Na^{+}, K^{+}-ATPase, hematocrit, and plasma glucose. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1430–1435, August 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-167.

Rasmussen:1988:LZB

[2259] Joseph B. Rasmussen. Littoral zoobenthic biomass in lakes, and its relationship to physical, chemical, and trophic factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1436–1447, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-168.

Kope:1988:DEI

[2260] Robert G. Kope and Louis W. Botsford. Detection of environmental influence on recruitment using abundance data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1448–1458, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-169.

Laird:1988:DLM

[2261] Gwenyth A. Laird, Donald Scavia, Gary L. Fahnenstiel, Lisa A. Strong, and Gregory A. Lang. Dynamics of Lake Michigan phytoplankton; relationship to nitrogen and silica fluxes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1459–1466, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-170.

Canfield:1988:INE

[2262] Daniel E. Canfield, Jr. and Mark V. Hoyer. Influence of nutrient enrichment and light availability on the abundance of aquatic macrophytes in Florida streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8): 1467–1472, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-171.

Loftus:1988:ELT

[2263] Andrew J. Loftus, William W. Taylor, and Myrl Keller. An evaluation of lake trout (Salvelinus namaycush) hooking mortality in the Upper Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1473–1479, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
172.

Heath:1988:CPD

[2264] Robert T. Heath and David A. Francko. Comparison of phosphorus dynamics in two Oklahoma reservoirs and a natural lake varying in abiogenic turbidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8): 1480–1486, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-173.

Johnson:1988:OAW

[2265] Scott W. Johnson and Jonathan Heifetz. Osmoregulatory ability of wild Coho salmon (Oncorhynchus kisutch) and Dolly Varden char (Salvelinus malma) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8): 1487–1490, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-174.

Parkinson:1988:GSH

[2266] Eric A. Parkinson and Kanji Tsumura. Growth and survival of hormonesterilized Coho (Oncorhynchus kisutch) and Kokanee salmon (O. nerka) in a lacustrine environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (8):1490–1494, August 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-175.

Cochran:1988:EEF

[2267] P. A. Cochran and K. J. Knutsen. Error in estimation of feeding rates from changes in mean body mass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1494–1498, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-176.

Anonymous:1988:LCLb

[2268] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1499–1503, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-177.

Anonymous:1988:CRPb

[2269] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(8):1504, August 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-

Vandevelde:1988:IBV

[2270] Thierry Vandevelde, Louis Legendre, Serge Demers, and Jean-Claude Therriault. Interrelationship between in vivo fluorescence of phytoplankton and light beam transmission with reference to fluorescence yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1508–1513, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-179.

Mosegaard:1988:USO

[2271] Henrik Mosegaard, Henrik Svedäng, and Kjell Taberman. Uncoupling of somatic and otolith growth rates in Arctic char (Salvelinus alpinus) as an effect of differences in temperature response. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1514–1524, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-180.

Andersson:1988:PDF

[2272] Tommy Andersson, Lars Förlin, Jan Härdig, and Åke Larsson. Physiological disturbances in fish living in coastal water polluted with bleached kraft pulp mill effluents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9): 1525–1536, September 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-181.

Jonsson:1988:LHV

[2273] B. Jonsson, S. Skúlason, S. S. Snorrason, O. T. Sandlund, H. J. Malmquist, P. M. Jónasson, R. Cydemo, and T. Lindem. Life history variation of polymorphic Arctic charr (Salvelinus alpinus) in Thingvallavatn, Iceland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1537– 1547, September 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-182.

Corti:1988:MMA

[2274] M. Corti, R. S. Thorpe, L. Sola, V. Sbordoni, and S. Cataudella. Multivariate morphometrics in aquaculture: a case study of six stocks of the common carp (Cyprinus carpio) from Italy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1548–1554, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-183.

Wissmar:1988:ECJ

[2275] Robert C. Wissmar and Charles A. Simenstad. Energetic constraints of juvenile chum salmon (Oncorhynchus keta) migrating in estuaries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1555–1560, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-184.

Bergman:1988:LAF

[2276] H. L. Bergman, J. S. Mattice, and D. J. A. Brown. Lake acidification and fisheries project: Adult brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1561–1562, September 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-185.

Booth:1988:EAL

[2277] C. E. Booth, D. G. McDonald, B. P. Simons, and C. M. Wood. Effects of aluminum and low pH on net ion fluxes and ion balance in the brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9): 1563–1574, September 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-186.

Wood:1988:BGA

[2278] C. M. Wood, R. C. Playle, B. P. Simons, G. G. Goss, and D. G. McDonald. Blood gases, acid–base status, ions, and hematology in adult brook trout (Salvelinus fontinalis) under acid/aluminum exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1575–1586, September 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-187.

Wood:1988:PEAa

[2279] C. M. Wood, D. G. McDonald, C. E. Booth, B. P. Simons, C. G. Ingersoll, and H. L. Bergman. Physiological evidence of acclimation to acid/ aluminum stress in adult brook trout (Salvelinus fontinalis). 1. blood composition and net sodium fluxes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1587–1596, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-188.

Wood:1988:PEAb

[2280] C. M. Wood, B. P. Simons, D. R. Mount, and H. L. Bergman. Physiological evidence of acclimation to acid/aluminum stress in adult brook trout (Salvelinus fontinalis). 2. blood parameters by cannulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1597–1605, September 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-189.

McDonald:1988:STB

[2281] D. G. McDonald and C. L. Millsgan. Sodium transport in the brook trout, Salvelinus fontinalis: Effects of prolonged low pH exposure in the presence and absence of aluminum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1606–1613, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-190.

Walker:1988:ELP

[2282] R. L. Walker, C. M. Wood, and H. L. Bergman. Effects of low pH and aluminum on ventilation in the brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1614–1622, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-191.

Mount:1988:ELTa

[2283] D. R. Mount, C. G. Ingersoll, D. D. Gulley, J. D. Fernandez, T. W. LaPoint, and H. L. Bergman. Effect of long-term exposure to acid, aluminum, and low calcium on adult brook trout (Salvelinus fontinalis).

1. survival, growth, fecundity, and progeny survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1623–1632, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-192.

Mount:1988:ELTb

[2284] D. R. Mount, J. R. Hockett, and W. A. Gern. Effect of long-term exposure to acid aluminum, and low calcium on adult brook trout (Salvelinus fontinalis). 2. Vitellogenesss and osmoregulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1633–1642, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-193.

Tietge:1988:MCG

[2285] J. E. Tietge, R. D. Johnson, and H. L. Bergman. Morphometric changes in gill secondary lamellae of brook trout (Salvelinus fontinalis) after longterm exposure to acid and aluminum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1643–1648, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-194.

Rowe:1988:HSM

[2286] Locke Rowe, Jeffrey Hudson, and Michael Berrill. Hatching success of mayfly eggs at low pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9): 1649–1652, September 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-195.

Hiltz:1988:CDM

[2287] Kenneth S. Hiltz, Jonathan G. Dean, Peter Schwinghamer, and Donald J. Belliveau. Computer-driven motor control with positional memory for mechanical microscope stage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9):1652–1656, September 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-196.

Miller:1988:LSR

[2288] Thomas J. Miller, Larry B. Crowder, James A. Rice, and Elizabeth A. Marschall. Larval size and recruitment mechanisms in fishes: Toward a conceptual framework. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(9): 1657–1670, September 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-197.

Neilson:1988:SSS

[2289] J. D. Neilson, E. M. DeBlois, and P. C. F. Hurley. Stock structure of Scotian Shelf flatfish as inferred from ichthyoplankton survey data and the geographic distribution of mature females. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1674–1685, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-198.

Verspoor:1988:RGV

[2290] Eric Verspoor. Reduced genetic variability in first-generation hatchery populations of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1686–1690, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-199.

Heggberget:1988:WRS

[2291] Tor G. Heggberget, Lars P. Hansen, and Tor F. Næsje. Within-river spawning migration of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1691–1698, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-200.

Nelson:1988:BOC

[2292] Jay A. Nelson, John J. Magnuson, and Weerewan Chulakasem. Blood oxygen capacity differences in yellow perch (Perca flavescens) from Northern Wisconsin lakes differing in pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1699–1704, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f88-201.

Stephenson:1988:MAC

[2293] M. Stephenson and G. L. Mackie. Multivariate analysis of correlations between environmental parameters and cadmium concentrations in Hyalella azteca (Crustacea: Amphipoda) from Central Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1705–1710, October 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-202.

Morin:1988:EMP

[2294] Antoine Morin, Christian Back, Alain Chalifour, Jacques Boisvert, and Robert H. Peters. Empirical models predicting ingestion rates of black fly larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1711–1719, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
203.

Murawski:1988:BBM

[2295] S. A. Murawski and J. T. Finn. Biological bases for mixed-species fisheries: Species co-distribution in relation to environmental and biotic variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1720–1735, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
204.

Mullins:1988:BVT

[2296] Julia Mullins, Hal Whitehead, and Linda S. Weilgart. Behaviour and vocalizations of two single sperm whales, Physeter macrocephalus, off Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1736–1743, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f88-
205.

Jackson:1988:AMP

[2297] Togwell A. Jackson. Accumulation of mercury by plankton and benthic invertebrates in riverine lakes of Northern Manitoba (Canada): Importance of regionally and seasonally varying environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1744–1757, October 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-206.

Vanni:1988:FZC

[2298] Michael J. Vanni. Freshwater zooplankton community structure: Introduction of large invertebrate predators and large herbivores to a small species community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10): 1758–1770, October 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-207.

Godbout:1988:PDS

[2299] Lyse Godbout and Robert H. Peters. Potential determinants of stable catch in the brook trout (Salvelinus fontinalis) sport fishery in Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1771–1778, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-208.

Burr:1988:IDP

[2300] Robert L. Burr. Inferring the distribution of the parameters of the von Bertalanffy growth model from length moments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1779–1788, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-209.

Lovejoy:1988:ESO

[2301] William S. Lovejoy. Effect of stochasticity on optimal harvesting strategies in some lumped-parameter fishery models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1789–1800, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-210.

Heggenes:1988:SPB

[2302] Jan Heggenes. Substrate preferences of brown trout fry (Salmo trutta) m artificial stream channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (10):1801–1806, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f88-211.

Blanck:1988:VET

[2303] Hans Blanck and Sten-Åke Wängberg. Validity of an ecotoxicological test system: Short-term and long-term effects of arsenate on marine periphyton communities in laboratory systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1807–1815, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-212.

Blanck:1988:ICT

[2304] Hans Blanck and Sten-Åke Wängberg. Induced community tolerance in marine periphyton established under arsenate stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1816–1819, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-213.

Post:1988:OCD

[2305] John R. Post and Donald J. McQueen. Ontogenetic changes in the distribution of larval and juvenile yellow perch (Perca flavescens): a response to prey or predators? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10): 1820–1826, October 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-214.

Martoja:1988:ECC

[2306] Roger Martoja, Christiane Ballan-Dufrançais, Anne-Yvonne Jeantet, Pierre Gouzerh, Jean-Claude Amiard, Claude Amiard-Triquet, Brigitte Berthet, and Jean-Pierre Baud. Effets chimiques et cytologiques de la contamination expérimentale de l’huître Crassostrea gigas Thunberg par l’argent administré sous forme dissoute et par voie alimentaire. (French) [Chemical and cytological effects of experimental contamination of oysters Crassostrea gigas Thunberg by silver administered in dissolved form and through food]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10): 1827–1841, October 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-215.

Madenjian:1988:PND

[2307] Charles M. Madenjian, Gary L. Rogers, and Arlo W. Fast. Predicting nighttime dissolved oxygen loss in aquaculture ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1842–1847, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-216.

Walters:1988:REF

[2308] Carl J. Walters and Jeremy S. Collie. Is research on environmental factors useful to fisheries management? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1848–1854, October 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-217.

Anonymous:1988:LCLc

[2309] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(10):1855–1856, October 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-218.

Loder:1988:STS

[2310] John W. Loder, Charles K. Ross, and Peter C. Smith. A space- and timescale characterization of circulation and mixing over submarine banks, with application to the Northwestern Atlantic Continental Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1860–1885, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-219.

Dermott:1988:SIR

[2311] Ronald Dermott and Karl Corning. Seasonal ingestion rates of Pontopareia hoyi (Amphipoda) in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1886–1895, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-220.

Nilsson:1988:PCS

[2312] Christer Nilsson and Paul A. Keddy. Predictability of change in shoreline vegetation in a hydroelectric reservoir, Northern Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1896–1904, November 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-221.

Kelso:1988:RHL

[2313] J. R. M. Kelso and D. S. Jeffries. Response of headwater lakes to varying atmospheric deposition in North-Central Ontario, 1979–85. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1905–1911, November 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-222.

Riessen:1988:CPP

[2314] Howard P. Riessen, Jonathan W. Sommerville, Christopher Chiappari, and Daniel Gustafson. Chaoborus predation, prey vulnerability, and their effect in zooplankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1912–1920, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-223.

Moore:1988:SHU

[2315] Kelly M. S. Moore and Stanley V. Gregory. Summer habitat utilization and ecology of cutthroat trout fry (Salmo clarki) in Cascade Mountain streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1921–1930, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-224.

Matuszek:1988:FSR

[2316] J. E. Matuszek and G. L. Beggs. Fish species richness in relation to lake area, pH, and other abiotic factors in Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1931–1941, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-225.

Malison:1988:GFR

[2317] Jeffrey A. Malison, Terrence B. Kayes, Bernard C. Wentworth, and Clyde H. Amundson. Growth and feeding responses of male versus female yellow perch (Perca flavescens) treated with Estradiol-17β. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1942–1948, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-226.

Freeman:1988:FAS

[2318] M. C. Freeman, M. K. Crawford, J. C. Barrett, D. E. Facey, M. G. Flood, J. Hill, D. J. Stouder, and G. D. Grossman. Fish assemblage stability in a Southern Appalachian stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1949–1958, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-227.

Galat:1988:MBG

[2319] David L. Galat and James P. Verdin. Magnitude of blue-green algal blooms in a saline desert lake evaluated by remote sensing: Evidence for nitrogen control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11): 1959–1967, November 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-228.

Martin:1988:DAT

[2320] Jennifer L. Martin and Alan W. White. Distribution and abundance of the toxic dinoflagellate Gonyaulax excavata in the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1968–1975, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-229.

Canfield:1988:PBC

[2321] Daniel E. Canfield, Jr. and Carlos M. Duarte. Patterns in biomass and cover of aquatic macrophytes in lakes: Test with Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1976–1982, November 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-230.

Payne:1988:WSL

[2322] Jerry F. Payne, J. Kiceniuk, L. L. Fancey, U. Williams, G. L. Fletcher, A. Rahimtula, and B. Fowler. What is a safe level of polycyclic aromatic hydrocarbons for fish: Subchronic toxicity study on winter flounder (Pseudopleuronectes americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1983–1993, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-231.

Giberson:1988:SVF

[2323] Donna J. Giberson and Ronald J. Hall. Seasonal variation in faunal distribution within the sediments of a Canadian Shield stream, with emphasis on responses to spring floods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):1994–2002, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-232.

Pouliot:1988:ADR

[2324] T. Pouliot and J. de la Noüe. Apparent digestibility in rainbow trout (Salmo gairdneri): Influence of hypoxia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):2003–2009, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-233.

Cuffney:1988:POM

[2325] Thomas F. Cuffney and J. Bruce Wallace. Particulate organic matter export from three headwater streams: Discrete versus continuous measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):2010–2016, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f88-234.

Clements:1988:IHM

[2326] Williams H. Clements, Donald S. Cherry, and John Cairns, Jr. Impact of heavy metals on insect communities in streams: a comparison of observational and experimental results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):2017–2025, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-235.

Klotz:1988:SCS

[2327] R. L. Klotz. Sediment control of soluble reactive phosphorus in Hoxie Gorge Creek, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11): 2026–2034, November 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-236.

Billington:1988:EIH

[2328] Neil Billington, Paul D. N. Hebert, and Robert D. Ward. Evidence of introgressive hybridization in the genus Stizostedion: Interspecific transfer of mitochondrial DNA between sauger and walleye. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):2035–2041, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-237.

Scarnecchia:1988:SMS

[2329] Dennis L. Scarnecchia. Salmon management and the search for values. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(11):2042–2050, November 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-238.

McQuinn:1988:AAE

[2330] I. H. McQuinn, L. Gendron, and J. H. Himmelman. Area of attraction and effective area fished by a whelk (Buccinum undatum) trap under variable conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12): 2054–2060, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-239.

Lathrop:1988:EWL

[2331] Richard C. Lathrop. Evaluation of whole-lake nitrogen fertilization for controlling blue-green algal blooms in a hypereutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2061–2075, December 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-240.

Rao:1988:DAN

[2332] D. V. Subba Rao, M. A. Quilliam, and R. Pocklington. Domoic acid — a neurotoxic amino acid produced by the marine diatom Nitzschia pungens in culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12): 2076–2079, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-241.

Andrus:1988:WDC

[2333] C. W. Andrus, B. A. Long, and H. A. Froehlich. Woody debris and its contribution to pool formation in a coastal stream 50 years after logging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2080–2086, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-242.

Petrosky:1988:RWR

[2334] C. E. Petrosky and T. C. Bjornn. Response of wild rainbow (Salmo gairdneri) and cutthroat trout (S. clarki) to stocked rainbow trout in fertile and infertile streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12): 2087–2105, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-243.

Ennis:1988:FMS

[2335] G. P. Ennis, R. G. Hooper, and D. M. Taylor. Functional maturity in small male snow crabs (Chionoecetes opilio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2106–2109, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-244.

Hinch:1988:SEC

[2336] Scott G. Hinch and Roger H. Green. Shell etching on clams from lowalkalinity Ontario lakes: a physical or chemical process? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2110–2113, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-245.

Grewe:1988:MDD

[2337] Peter M. Grewe and Paul D. N. Hebert. Mitochondrial DNA diversity among broodstocks of the lake trout, Salvelinus namaycush. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2114–2122, December 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-246.

Hall:1988:FAS

[2338] Ronald J. Hall, R. C. Bailey, and Joseph Findeis. Factors affecting survival and cation concentration in the blackflies Prosimulium fuscum/ mixtum and the mayfly Leptophlebia cupida during spring snowmelt. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2123–2132, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f88-247.

Johnson:1988:CEO

[2339] Lyndal L. Johnson, Edmundo Casillas, Tracy K. Collier, Bruce B. Mc-Cain, and Usha Varanasi. Contaminant effects on ovarian development in English sole (Parophrys vetulus) from Puget Sound, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2133–2146, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-248.

Arts:1988:EIE

[2340] Michael T. Arts and W. Gary Sprules. Evidence for indirect effects of fish predation on maternal lipid investment in Holopedium gibberum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2147–2155, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-249.

Cunjak:1988:BMY

[2341] Richard A. Cunjak. Behaviour and microhabitat of young Atlantic salmon (Salmo salar) during winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2156–2160, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f88-250.

Metcalfe:1988:EIL

[2342] C. D. Metcalfe, V. W. Cairns, and J. D. Fitzsimons. Experimental induction of liver tumours in rainbow trout (Salmo gairdneri) by contaminated sediment from Hamilton Harbour, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2161–2167, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-251.

King:1988:MCE

[2343] J. M. King, J. A. Day, P. R. Hurly, M-P. Henshall-Howard, and B. R. Davies. Macroinvertebrate communities and environment in a Southern African mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45 (12):2168–2181, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f88-252.

Yamamori:1988:GRT

[2344] Kunio Yamamori, Moritaka Nakamura, Takashi Matsui, and Toshiaki j. Hara. Gustatory responses to tetrodotoxin and saxitoxin in fish: a possible mechanism for avoiding marine toxins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2182–2186, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-253.

Forster:1988:EDC

[2345] Ian Forster, David A. Higgs, Gordon R. Bell, B. S. Dosanjh, and B. E. March. Effect of diets containing herring oil oxidized to different degrees on growth and immunocompetence of juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12): 2187–2194, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-254.

Erman:1988:EWF

[2346] Don C. Erman, Edmund D. Andrews, and Michael Yoder-Williams. Effects of winter floods on fishes in the Sierra Nevada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2195–2200, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-255.

Mclsaac:1988:EHC

[2347] Donald O. Mclsaac and Thomas P. Quinn. Evidence for a hereditary component in homing behavior of Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2201– 2205, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-256.

Spry:1988:ZIA

[2348] Douglas J. Spry and Chris M. Wood. Zinc influx across the isolated, perfused head preparation of the rainbow trout (Salmo gairdneri) in hard and soft water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12): 2206–2215, December 1988. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f88-257.

Starr:1988:RHR

[2349] Paul Starr and Ray Hilborn. Reconstruction of harvest rates and stock contribution in gauntlet salmon fisheries: Application to British Columbia and Washington sockeye (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2216–2229, December 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-258.

Tarapchak:1988:OCL

[2350] Stephen J. Tarapchak and Lynn R. Herche. Orthophosphate concentrations in lake water: Analysis of Rigler’s radiobioassay method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2230–2237, December 1988. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f88-259.

Fausch:1988:TCB

[2351] Kurt D. Fausch. Tests of competition between native and introduced salmonids in streams: What have we learned? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 45(12):2238–2246, December 1988. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f88-260.

Levine:1989:PNC

[2352] S. N. Levine and D. W. Schindler. Phosphorus, nitrogen, and carbon dynamics of Experimental Lake 303 during recovery from eutrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):2–10, January 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-001.

Collins:1989:DEF

[2353] Nicholas C. Collins. Daytime exposure to fish predation for littoral benthic organisms in unproductive lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):11–15, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-002.

Scarnecchia:1989:EOV

[2354] Dennis L. Scarnecchia, Árni Ísaksson, and S. E. White. Effects of oceanic variations and the West Greenland fishery on age at maturity of Icelandic West Coast stocks of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):16–27, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-003.

Hilborn:1989:TPS

[2355] Ray Hilborn and Paul Medley. Tuna purse-seine fishing with fishaggregating devices (FAD): Models of tuna FAD interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):28–32, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-004.

Curtis:1989:RED

[2356] Lawrence R. Curtis, Wayne K. Seim, Lisbeth K. Siddens, Debra A. Meager, Richard A. Carchman, Walter H. Carter, and Gary A. Chapman. Role of exposure duration in hydrogen ion toxicity to brook (Salvelinus fontinalis) and rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):33–40, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-005.

Ford:1989:GSW

[2357] T. E. Ford and Robert J. Naiman. Groundwater–surface water relationships in boreal forest watersheds: Dissolved organic carbon and inorganic nutrient dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (1):41–49, January 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-006.

Edmonds:1989:TEA

[2358] J. S. Edmonds, M. J. Moran, N. Caputi, and M. Morita. Trace element analysis of fish sagittae as an aid to stock identifications: Pink snapper (Chrysophrys auratus) in Western Australian waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):50–54, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-007.

Latham:1989:OAP

[2359] Keith E. Latham and John J. Just. Oxygen availability provides a signal for hatching in the rainbow trout (Salmo gairdneri) embryo. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):55–58, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-008.

Lucotte:1989:PRS

[2360] Marc Lucotte. Phosphorus reservoirs in the St. Lawrence Upper Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):59–65, January 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-009.

Reisenbichler:1989:GVS

[2361] R. R. Reisenbichler and S. R. Phelps. Genetic variation in steelhead (Salmo gairdneri) from the North Coast of Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):66–73, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-010.

Machado:1989:ECM

[2362] J. P. Machado, T. G. Bell, A. L. Trapp, D. L. Garling, Jr., and N. R. Kevern. Effect of carbon monoxide exposure on gas bubble disease in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (1):74–80, January 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-011.

Robinson:1989:IEF

[2363] Clifford L. K. Robinson and William M. Tonn. Influence of environmental factors and piscivory in structuring fish assemblages of small Alberta lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):81–89, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-012.

Welch:1989:PNP

[2364] Harold E. Welch, John A. Legauit, and Hedy J. Kling. Phytoplankton, nutrients, and primary production in fertilized and natural lakes at Saqvaqjuac, N.W.T. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (1):90–107, January 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-013.

Reznick:1989:SGR

[2365] David Reznick, Eric Lindbeck, and Heather Bryga. Slower growth results in larger otoliths: an experimental test with guppies (Poecilia reticulata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):108–112, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-014.

Secor:1989:SGE

[2366] David H. Secor and John Mark Dean. Somatic growth effects on the otolith–fish size relationship in young pond-reared striped bass, Morone saxatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):113–121, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-015.

Morin:1989:CRN

[2367] Pierre-Philippe Morin, Julian J. Dodson, and François Y. Doré. Cardiac responses to a natural odorant as evidence of a sensitive period for olfactory imprinting in young Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):122–130, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-016.

Morin:1989:TAC

[2368] Pierre-Philippe Morin, Julian J. Dodson, and François Y. Doré. Thyroid activity concomitant with olfactory learning and heart rate changes in Atlantic salmon Salmo salar, during smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):131–136, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-017.

Ludwig:1989:RMP

[2369] D. Ludwig and C. J. Walters. A robust method for parameter estimation from catch and effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (1):137–144, January 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-018.

Aiken:1989:ITP

[2370] D. E. Aiken and S. L. Waddy. Interaction of temperature and photoperiod in the regulation of spawning by American lobsters (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):145–148, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-019.

Foote:1989:BGC

[2371] Chris J. Foote, Chris C. Wood, and Ruth E. Withler. Biochemical genetic comparison of sockeye salmon and Kokanee, the anadromous and nonanadromous forms of Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):149–158, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-020.

Cravedi:1989:DER

[2372] J. P. Cravedi, G. Delous, and D. Rao. Disposition and elimination routes of 17α-methyltestosterone in rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):159–165, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-021.

Henault:1989:CMM

[2373] Michel Hénault and Réjean Fortin. Comparison of meristic and morphometric characters among spring- and fall-spawning ecotypes of cisco (Coregonus artedii) in Southern Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):166–173, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-022.

Anonymous:1989:LCLa

[2374] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(1):174, January 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-023.

Down:1989:PAR

[2375] N. E. (Ted) Down, Edward M. Donaldson, Helen M. Dye, Thomas C. Boone, Keith E. Langley, and Lawrence M. Souza. A potent analog of recombinant bovine somatotropin accelerates growth in juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):178–183, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-024.

Aalto:1989:EDS

[2376] S. K. Aalto and G. E. (Buck) Newsome. Evidence of demic structure for a population of yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):184–190, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-025.

Scott:1989:MAC

[2377] J. S. Scott. Matrix analysis of co-occurrences of fishes of the Scotian Shelf and Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (2):191–197, February 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-026.

Hansell:1989:MEU

[2378] Dennis A. Hansell and John J. Goering. A method for estimating uptake and production rates for urea in seawater using [^{14}C] urea and [^{15}N] urea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):198–202, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-027.

Newman:1989:DBT

[2379] Raymond M. Newman and Thomas F. Waters. Differences in brown trout (Salmo trutta) production among contiguous sections of an entire stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):203–213, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-028.

Mather:1989:CMT

[2380] Martha E. Mather and David H. Wahl. Comparative mortality of three esocids due to stocking stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):214–217, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-029.

Burtle:1989:LRC

[2381] G. J. Burtle and R. T. Lovell. Lack of response of channel catfish (Ictalurus punctatus) to dietary myo-inositol. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):218–222, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-030.

Robbins:1989:RAT

[2382] J. A. Robbins, T. Keilty, D. S. White, and D. N. Edgington. Relationships among tubificid abundances, sediment composition, and accumulation rates in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (2):223–231, February 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-031.

Swales:1989:RCP

[2383] S. Swales and C. D. Levings. Role of off-channel ponds in the life cycle of Coho salmon (Oncorhynchus kisutch) and other juvenile salmonids in the Coldwater River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):232–242, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-032.

Small:1989:ETS

[2384] Susan A. Small and D. J. Randall. Effects of triploidy on the swimming performance of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):243–245, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-033.

Porcella:1989:LAM

[2385] D. B. Porcella. Lake Acidification Mitigation Project (LAMP): an overview of an ecosystem perturbation experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):246–248, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-034.

Driscoll:1989:STC

[2386] Charles T. Driscoll, Gerald F. Fordham, William A. Ayling, and Leah M. Oliver. Short-term changes in the chemistry of trace metals following calcium carbonate treatment of acidic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):249–257, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-035.

Driscoll:1989:CRL

[2387] Charles T. Driscoll, William A. Ayling, G. F. Fordham, and Leah M. Oliver. Chemical response of lakes treated with CaCO_{3} to reacidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):258–267, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-036.

Staubitz:1989:HTH

[2388] Ward W. Staubitz and Phillip J. Zarriello. Hydrology of two headwater lakes in the Adirondack Mountains of New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):268–276, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-037.

Gloss:1989:SGR

[2389] Steven P. Gloss, Carl L. Schofield, Robert L. Spateholts, and Barbara A. Plonski. Survival, growth, reproduction, and diet of brook trout (Salvelinus fontinalis) stocked into lakes after liming to mitigate acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):277–286, February 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-038.

Roberts:1989:ELE

[2390] Deborah A. Roberts and Charles W. Boylen. Effects of liming on the epipelic algal community of Woods Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):287–294, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-039.

Schaffner:1989:ENA

[2391] William R. Schaffner. Effects of neutralization and addition of brook trout (Salvelinus fontinalis) on the limnetic zooplankton communities of two acidic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):295–305, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-040.

Fordham:1989:STC

[2392] G. F. Fordham and C. T. Driscoll. Short-term changes in the acid/ base chemistry of two acidic lakes following calcium carbonate treatment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):306–314, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-041.

Young:1989:CDS

[2393] Thomas C. Young, Joseph V. DePinto, James R. Rhea, and Richard D. Scheffe. Calcite dose selection, treatment efficiency, and residual calcite fate after whole-lake neutralization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):315–322, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-042.

DePinto:1989:PRC

[2394] Joseph V. DePinto, Richard D. Scheffe, William G. Booty, and Thomas C. Young. Predicting reacidification of calcite treated acid lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):323–332, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-043.

Schofield:1989:PGE

[2395] Carl L. Schofield, Steven P. Gloss, Barbara Plonski, and Robert Spateholts. Production and growth efficiency of brook trout (Salvelinus fontinalis) in two Adirondack mountain (New York) lakes following liming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):333–341, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-044.

Evans:1989:RLI

[2396] Richard A. Evans. Response of limnetic insect populations of two acidic, fishless lakes to liming and brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):342–351, February 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-045.

Bukaveckas:1989:ECT

[2397] Paul A. Bukaveckas. Effects of calcite treatment on primary producers in acidified Adirondack lakes. II. Short-term response by phytoplankton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):352–359, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-046.

Addison:1989:OMM

[2398] R. F. Addison. Organochlorines and marine mammal reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):360–368, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-047.

Anonymous:1989:CRPa

[2399] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(2):369, February 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-048.

Hyatt:1989:RNR

[2400] Kim D. Hyatt and Neil H. Ringler. Role of nest raiding and egg predation in regulating population density of threespine sticklebacks (Gasterosteus aculeatus) in a coastal British Columbia lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):372–383, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-049.

Beamish:1989:IDM

[2401] F. W. H. Beamish, J. C. Howlett, and T. E. Medland. Impact of diet on metabolism and swimming performance in juvenile lake trout, Salvelinus namaycush. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):384–388, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-050.

Krantzberg:1989:MRT

[2402] G. Krantzberg and P. M. Stokes. Metal regulation, tolerance, and body burdens in the larvae of the genus Chironomus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):389–398, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-051.

Borgmann:1989:ECS

[2403] U. Borgmann, E. S. Millard, and C. C. Charlton. Effect of cadmium on a stable, large volume, laboratory ecosystem containing Daphnia and phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):399–405, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-052.

Vuorinen:1989:GDA

[2404] Jukka Vuorinen and Ole Kristian Berg. Genetic divergence of anadromous and nonanadromous Atlantic salmon (Salmo salar) in the River Namsen, Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):406– 409, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-053.

Hansen:1989:IPM

[2405] L. P. Hansen, B. Jonsson, R. I. G. Morgan, and J. E. Thorpe. Influence of parr maturity on emigration of smolting Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):410–415, March 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-054.

Ostrofsky:1989:RBA

[2406] M. L. Ostrofsky, D. A. Osborne, and T. J. Zebulske. Relationship between anaerobic sediment phosphorus release rates and sedimentary phosphorus species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (3):416–419, March 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-055.

Beamish:1989:EPA

[2407] R. J. Beamish and T. G. Northcote. Extinction of a population of anadromous parasitic lamprey, Lampetra tridentata, upstream of an impassable dam. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):420–425, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-056.

McMurtry:1989:RMC

[2408] Michael J. McMurtry, Donna L. Wales, Wolfgang A. Scheider, Gail L. Beggs, and Patricia E. Dimond. Relationship of mercury concentrations in lake trout (Salvelinus namaycush) and smallmouth bass (Micropterus dolomieui) to the physical and chemical characteristics of Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):426–434, March 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-057.

Chambers:1989:RVS

[2409] P. A. Chambers, E. E. Prepas, M. L. Bothwell, and H. R. Hamilton. Roots versus shoots in nutrient uptake by aquatic macrophytes in flowing waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):435–439, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-058.

Hamilton:1989:BCM

[2410] Steven J. Hamilton and Terry A. Haines. Bone characteristics and metal concentrations in white suckers (Catostomus commersoni) from one neutral and three acidified lakes in Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):440–446, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-059.

Young:1989:EVP

[2411] Michael K. Young, Wayne A. Hubert, and Thomas A. Wesche. Evaluation of variation in permeability measurements when using the MARK VI standpipe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):447–450, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-060.

Hare:1989:GSB

[2412] Landis Hare, Peter G. C. Campbell, André Tessier, and Nelson Belzile. Gut sediments in a burrowing mayfly (Ephemeroptera, Hexagenia limbata): Their contribution to animal trace element burdens, their removal, and the efficacy of a correction for their presence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):451–456, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-061.

Boisclair:1989:APVa

[2413] D. Boisclair and W. C. Leggett. Among-population variability of fish growth: I. Influence of the quantity of food consumed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):457–467, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-062.

Boisclair:1989:APVb

[2414] D. Boisclair and W. C. Leggett. Among-population variability of fish growth: II. Influence of prey type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):468–482, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-063.

Allen:1989:SPU

[2415] K. Radway Allen and W. S. Hearn. Some procedures for use in cohort analysis and other population simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):483–488, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-064.

Smith:1989:GEW

[2416] R. L. Smith, J. M. Paul, and A. J. Paul. Gastric evacuation in walleye pollock, Theragra chalcogramma. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):489–493, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-065.

Korver:1989:ZAF

[2417] R. M. Korver and J. B. Sprague. Zinc avoidance by fathead minnows (Pimephales promelas): Computerized tracking and greater ecological relevance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):494–502, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-066.

Schneider:1989:PLA

[2418] David C. Schneider and Richard L. Haedrich. Prediction limits of allometric equations: a reanalysis of Ryder’s morphoedaphic index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):503–508, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-067.

Shaffer:1989:TLC

[2419] Paul W. Shaffer and M. Robbins Church. Terrestrial and in-lake contributions to alkalinity budgets of drainage lakes: an assessment of regional differences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):509–515, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-068.

Moodie:1989:IEC

[2420] G. E. E. Moodie, N. L. Loadman, M. D. Wiegand, and J. A. Mathias. Influence of egg characteristics on survival, growth and feeding in larval walleye (Stizostedion vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(3):516–521, March 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-069.

Giguere:1989:CWE

[2421] Louis A. Giguère, J.-F. St-Pierre, B. Bernier, A. Vézina, and j. G. Rondeau. Can we estimate the true weight of zooplankton samples after chemical preservation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (3):522–527, March 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-070.

Kerr:1989:CAM

[2422] S. R. Kerr and R. A. Ryder. Current approaches to multispecies analyses of marine fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (3):528–534, March 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-071.

Dickie:1989:FEP

[2423] L. M. Dickie and P. C. Smith. Fisheries Ecology Program (FEP): General introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):538, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-072.

Smith:1989:CDB

[2424] Peter C. Smith. Circulation and dispersion on Browns Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):539–559, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-073.

McLaren:1989:CPS

[2425] Ian A. McLaren, M. John Tremblay, C. J. Corkett, and John C. Roff. Copepod production on the Scotian Shelf based on life-history analyses and laboratory bearings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (4):560–583, April 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-074.

Wildish:1989:BMP

[2426] D. J. Wildish, A. J. Wilson, and B. Frost. Benthic macrofaunal production of Browns Bank, Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):584–590, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-075.

Suthers:1989:IAD

[2427] Iain M. Suthers and Kenneth T. Frank. Inter-annual distributions of larval and pelagic juvenile cod (Gadus morhua) in Southwestern Nova Scotia determined with two different gear types. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):591–602, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-076.

Campana:1989:ATM

[2428] Steven E. Campana and Peter C. F. Hurley. An age- and temperaturemediated growth model for cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) larvae in the Gulf of Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):603–613, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-077.

Boudreau:1989:BMF

[2429] P. R. Boudreau and L. M. Dickie. Biological model of fisheries production based on physiological and ecological scalings of body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):614–623, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-078.

Chenoweth:1989:ODL

[2430] S. B. Chenoweth, D. A. Libby, R. L. Stephenson, and M. J. Power. Origin and dispersion of larval herring (Clupea harengus) in coastal waters of Eastern Maine and Southwestern New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):624–632, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-079.

Heifetz:1989:MTS

[2431] Jonathan Heifetz, Scott W. Johnson, K. V. Koski, and Michael L. Murphy. Migration timing, size, and salinity tolerance of sea-type sockeye salmon (Oncorhynchus nerka) in an Alaska estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):633–637, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-080.

Hanna:1989:BAP

[2432] Micheline Hanna. Biologically available phosphorus: Estimation and prediction using an anion-exchange resin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):638–643, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-081.

Darwish:1989:PAF

[2433] G. S. Darwish, F. R. van de Voort, and J. P. Smith. Proximate analysis of fish tissue by mid-infrared transmission spectroscopy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):644–649, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-082.

Jeffries:1989:PWF

[2434] Perry Jeffries, Aimee Keller, and Stephen Hale. Predicting winter flounder (Pseudopleuronectes americanus) catches by time series analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):650–659, April 1989.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-083.

Polovina:1989:DDS

[2435] Jeffrey J. Polovina. Density dependence in spiny lobster, Panulirus marginatus, in the Northwestern Hawaiian Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):660–665, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-084.

Chapman:1989:FFN

[2436] Lauren J. Chapman, William C. Mackay, and Craig W. Wilkinson. Feeding flexibility in northern pike (Esox lucius): Fish versus invertebrate prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):666–669, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-085.

Cury:1989:OEW

[2437] Philippe Cury and Claude Roy. Optimal environmental window and pelagic fish recruitment success in upwelling areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):670–680, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-086.

Scrivener:1989:EFH

[2438] J. C. Scrivener and M. J. Brownlee. Effects of forest harvesting on spawning gravel and incubation survival of chum (Oncorhynchus keta) and Coho salmon (O. kisutch) in Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):681–696, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-087.

Holm:1989:MDJ

[2439] Jens Chr Holm. Mono- and duoculture of juvenile Atlantic salmon (Salmo salar) and Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):697–704, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-088.

Dauble:1989:RBE

[2440] D. D. Dauble and L. R. Curtis. Rapid branchial excretion of dietary quinoline by rainbow trout (Salmo gairdneri). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):705–713, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-089.

Foster:1989:IHA

[2441] J. R. Foster and C. B. Schom. Imprinting and homing of Atlantic salmon (Salmo salar) kelts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (4):714–719, April 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-090.

LeFoll:1989:ADS

[2442] Didier Le Foll, Evelyne Brichet, Jean Louis Reyss, Claude Lalou, and Daniel Latrouite. Age determination of the spider crab Maja squinado and the European lobster Homarus gammarus by ^{228}Th/^{228}Ra chronology: Possible extension to other crustaceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):720–724, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-091.

Anonymous:1989:LCLb

[2443] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(4):725–727, April 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-092.

Schnute:1989:FGI

[2444] Jon T. Schnute, Laura J. Richards, and Alan J. Cass. Fish growth: Investigations based on a size-structured model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):730–742, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-093.

Schnute:1989:FSR

[2445] Jon T. Schnute, Laura J. Richards, and Alan J. Cass. Fish survival and recruitment: Investigations based on a size-structured model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):743–769, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-094.

Nicholls:1989:RCP

[2446] K. H. Nicholls and D. A. Hurley. Recent changes in the phytoplankton of the Bay of Quinte, Lake Ontario: the relative importance of fish, nutrients, and other factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (5):770–779, May 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-095.

Speare:1989:FAR

[2447] D. J. Speare and H. W. Ferguson. Fixation artifacts in rainbow trout (Salmo gairdneri) gills: a morphometric evaluation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):780–785, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-096.

Jensen:1989:TRA

[2448] Arne J. Jensen, Bjørn O. Johnsen, and Laila Saksgård. Temperature requirements in Atlantic salmon (Salmo salar), brown trout (Salmo trutta), and Arctic char (Salvelinus alpinus) from hatching to initial feeding compared with geographic distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):786–789, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-097.

Wiggs:1989:ARE

[2449] A. J. Wiggs, E. B. Henderson, R. L. Saunders, and M. N. Kutty. Activity, respiration, and excretion of ammonia by Atlantic salmon (Salmo salar) smolt and postsmolt. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (5):790–795, May 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-098.

StAubin:1989:ACH

[2450] D. J. St. Aubin and J. R. Geraci. Adaptive changes in hematologic and plasma chemical constituents in captive beluga whales, Delphinapterus leucas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):796–803, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-099.

Giles:1989:OVM

[2451] M. A. Giles and D. M. Rystephanuk. Ontogenetic variation in the multiple hemoglobins of Arctic charr, Salvelinus alpinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):804–809, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-100.

Johannes:1989:GSN

[2452] Mark R. S. Johannes, Donald J. McQueen, Tom J. Stewart, and John R. Post. Golden shiner (Notemigonus crysoleucas) population abundance correlations with food and predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):810–817, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-101.

Crittenden:1989:CGL

[2453] R. N. Crittenden and G. L. Thomas. A conditional generalized least squares method for estimating the size of a closed population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):818–823, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-102.

Gauthier:1989:RSA

[2454] D. Gauthier, L. Desjardins, J. A. Robitaille, and Y. Vigneault. River spawning of artificially reconditioned Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):824–826, May 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-103.

Planas:1989:SOS

[2455] Dolors Planas, Louise Lapierre, Guy Moreau, and Martine Allard. Structural organization and species composition of a lotic periphyton community in response to experimental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):827–835, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-104.

Castonguay:1989:DBA

[2456] Martin Castonguay, Jean-Denis Dutil, and Claude Desjardins. Distinction between American eels (Anguilla rostrata) of different geographic origins on the basis of their organochlorine contaminant levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):836–843, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-105.

Hogan:1989:HGS

[2457] Daniel L. Hogan and Michael Church. Hydraulic geometry in small, coastal streams: Progress toward quantification of salmonid habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):844–852, May 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-106.

Craig:1989:CBC

[2458] J. F. Craig, K. Smiley, and J. A. Babaluk. Changes in the body composition with age of goldeye, Hiodon alosoides. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):853–858, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-107.

Chiles:1989:MTB

[2459] Thomas C. Chiles, Peter D. Pendoley, and Roy B. Laughlin, Jr. Mechanisms of tri-n-butyltin bioaccumulation by marine phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):859–862, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-108.

Shaw:1989:RSA

[2460] Margo A. Shaw and Gerald L. Mackie. Reproductive success of Amnicola limosa (Gastropoda) in low alkalinity lakes in South-Central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):863–869, May 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-109.

Graves:1989:MDS

[2461] John E. Graves and Andrew E. Dizon. Mitochondrial DNA sequence similarity of Atlantic and Pacific albacore tuna (Thunnus alalunga). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):870–873, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-110.

Hayward:1989:GBF

[2462] Robert S. Hayward, F. Joseph Margraf, Carey T. Knight, and Daniel J. Glomski. Gear bias in field estimation of the amount of food consumed by fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):874–876, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-111.

Hennigan:1989:CAS

[2463] Martha Hennigan, Laurence M. Vaughan, Timothy J. Foster, Peter Smith, and Frank Gannon. Characterization of Aeromonas salmonicida strains using DNA probe technology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):877–879, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-112.

Mills:1989:EZA

[2464] Edward L. Mills, Ruth Sherman, and Douglas S. Robson. Effect of zooplankton abundance and body size on growth of age-0 yellow perch (Perca flavescens) in Oneida Lake, New York, 1975–86. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):880–886, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-113.

Neilson:1989:SAH

[2465] John D. Neilson, Kenneth G. Waiwood, and Stephen J. Smith. Survival of Atlantic halibut (Hippoglossus hippoglossus) caught by longline and otter trawl gear. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):887–897, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-114.

Evans:1989:PPP

[2466] Geoffrey T. Evans and Pierre Pepin. Potential for predicting plankton populations (and fish recruitment) from environmental data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):898–903, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-115.

Bird:1989:BOM

[2467] D. F. Bird and C. M. Duarte. Bacteria–organic matter relationship in sediments: a case of spurious correlation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(5):904–908, May 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-116.

Davies:1989:PCC

[2468] I. J. Davies. Population collapse of the crayfish Orconectes virilis in response to experimental whole-lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):910–922, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-117.

Davies:1989:DOS

[2469] I. J. Davies and D. J. Ramsey. A diver operated suction gun and collection bucket for sampling crayfish and other aquatic macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):923–927, June 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-118.

Labee-Lund:1989:SMM

[2470] Jan Hennsng L’abée-Lund. Significance of mature male parr in a small population of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):928–931, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-119.

Jensen:1989:BRC

[2471] Gregory C. Jensen and David A. Armstrong. Biennial reproductive cycle of blue king crab, Paralithodes platypus, at the Pribilof Island, Alaska and comparison to a congener, P. camtschatica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):932–940, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-120.

Kimura:1989:VTS

[2472] Daniel K. Kimura. Variability, tuning, and simulation for the Doubleday– DerisoCatch-at-Age model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):941–949, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-121.

Ballantyne:1989:ETO

[2473] J. S. Ballantyne, D. Flannigan, and T. B. White. Effects of temperature on the oxidation of fatty acids, acyl carnitines, and ketone bodies by mitochondria isolated from the liver of the lake charr, Salvelinus namaycush. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):950–954, June 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-122.

Berg:1989:SGT

[2474] Ole Kristian Berg and Magnus Berg. Sea growth and time of migration of anadromous Arctic char (Salvelinus alpinus) from the Vardnes River, in Northern Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):955–960, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-123.

Polovina:1989:SSD

[2475] Jeffrey J. Polovina. A system of simultaneous dynamic production and forecast models for mulispecies or multiarea applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):961–963, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-124.

Batts:1989:CIH

[2476] W. N. Batts and J. R. Winton. Concentration of infectious hematopoietic necrosis virus from water samples by tangential flow filtration and polyethylene glycol precipitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):964–968, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f89-125.

McQuinn:1989:ISA

[2477] Ian H. McQuinn. Identification of spring- and autumn-spawning herring (Clupea harengus harengus) using maturity stages assigned from a gonadosomatic index model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):969–980, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-126.

Shaw:1989:TSP

[2478] J. F. H. Shaw and E. E. Prepas. Temporal and spatial patterns of porewater phosphorus in shallow sediments, and its potential transport into Narrow Lake, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):981–988, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-127.

Tanasschuk:1989:QQC

[2479] R. W. Tanasschuk and W. C. Mackay. Quantitative and qualitative characteristics of somatic and gonadal growth of yellow perch (Perca flavescens) from Lac Ste. Anne, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):989–994, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-128.

Wanzenbock:1989:PDC

[2480] Josef Wanzenböck and Fritz Schiemer. Prey detection in cyprinids during early development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):995–1001, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-129.

Fournier:1989:EFM

[2481] David A. Fournier and A. R. Warburton. Evaluating fisheries management models by simulated adaptive control — introducing the composite model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1002–1012, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-130.

Agbeti:1989:ULF

[2482] M. Agbeti and M. Dickman. Use of lake fossil diatom assemblages to determine historical changes in trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1013–1021, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-131.

Ryan:1989:CRF

[2483] P. M. Ryan and J. J. Kerekes. Correction of relative fish abundance estimates from catch data for variable fishing intensity during lake surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1022–1025, June 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-132.

Confer:1989:CPR

[2484] John L. Confer and Linda M. O’Bryan. Changes in prey rank and preference by young planktivores for short-term and long-term ingestion periods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1026–1032, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-133.

El-Shaarawi:1989:MLE

[2485] A. H. El-Shaarawi and D. M. Dolan. Maximum likelihood estimation of water quality concentrations from censored data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1033–1039, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-134.

Allard:1989:MCS

[2486] Martine Allard and Pamela M. Stokes. Mercury in crayfish species from thirteen Ontario lakes in relation to water chemistry and smallmouth bass (Micropterus dolomieui) mercury. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1040–1046, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-135.

Sebastien:1989:IMS

[2487] Robert J. Sebastien, Reinhart A. Brust, and David M. Rosenberg. Impact of methoxychlor on selected nontarget organisms in a riffle of the Souris River, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (6):1047–1061, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-136.

Hiramatsu:1989:RMO

[2488] Kazuhiko Hiramatsu and Yukimasa Ishida. Random movement and orientation in pink salmon (Oncothynchus gorbuscha) migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1062–1066, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-137.

Preall:1989:CAP

[2489] Richard J. Preall and Neil H. Ringler. Comparison of actual and potential growth rates of brown trout (Salmo trutta) in natural streams based on bioenergetic models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6): 1067–1076, June 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-138.

Anonymous:1989:LCLc

[2490] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(6):1077–1082, June 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-139.

Irvine:1989:PTS

[2491] J. R. Irvine and B. R. Ward. Patterns of timing and size of wild Coho salmon (Oncorhynchus kisutch) smolts migrating from the Keogh River watershed on Northern Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1086–1094, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-140.

Likely:1989:CPD

[2492] C. G. Likely and M. D. B. Burt. Cultivation of Pseudoterranova decipiens (sealworm) from third-stage larvae to egg-laying adults in vitro. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1095–1096, July 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-141.

Stout:1989:ECT

[2493] R. Jean Stout. Effects of condensed tannins on leaf processing in mid-latitude and tropical streams: a theoretical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1097–1106, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-142.

Hard:1989:GVO

[2494] Jeffrey J. Hard, Alex C. Wertheimer, and William F. Johnson. Geographic variation in the occurrence of red- and white-fleshed Chinook salmon (Oncorhynchus tshawytscha) in Western North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1107–1113, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-143.

Waltman:1989:DTP

[2495] W. Douglas Waltman II, Emmett B. Shotts, and Richard E. Wooley. Development and transfer of plasmid-mediated antimicrobial resistance in Edwardsiella ictaluri. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (7):1114–1117, July 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-144.

Gardner:1989:MFN

[2496] Wayne S. Gardner, Brian J. Eadie, Joann F. Chandler, Christopher C. Parrish, and John M. Malczyk. Mass flux and “Nutritional composition” of settling epilimnetic particles in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1118–1124, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-145.

Gauvin:1989:EFR

[2497] Jeffery M. Gauvin, Wayne S. Gardner, and Michael A. Quigley. Effects of food removal on nutrient release rates and lipid content of Lake Michigan Pontoporeia hoyi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (7):1125–1130, July 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-146.

Bloom:1989:DPL

[2498] Nicolas Bloom. Determination of picogram levels of methylmercury by aqueous phase ethylation, followed by cryogenic gas chromatography with cold vapour atomic fluorescence detection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1131–1140, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-147.

Ankley:1989:DLF

[2499] Gerald T. Ankley, Vicki S. Blazer, Steven M. Plakas, and Robert E. Reinert. Dietary lipid as a factor modulating xenobiotic metabolism in channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1141–1146, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-148.

Bigler:1989:MOF

[2500] Brian Bigler. Mechanism and occurrence of focal scale resorption among chum salmon (Oncorhynchus keta) of the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1147–1153, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-149.

Mallet:1989:GMS

[2501] A. L. Mallet and C. E. A. Carver. Growth, mortality, and secondary production in natural populations of the blue mussel, Mytilus edulis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1154–1159, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-150.

Bacher:1989:CRT

[2502] C. Bacher and P. Goulletquer. Comparaison des relations trophiques de Ruditapes philippinarum en milieux estuarien et océanique à partir d’un modèle de croissance. (French) [Comparison of the trophic relationships of Ruditapes philippinarum in estuarine and oceanic environments using a growth model]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (7):1160–1170, July 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-151.

McCauley:1989:SRB

[2503] Edward McCauley, John A. Downing, and Susan Watson. Sigmoid relationships between nutrients and chlorophyll among lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1171–1175, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-152.

Prairie:1989:UNC

[2504] Yves T. Prairie, Carlos M. Duarte, and Jacob Kalff. Unifying nutrient– chlorophyll relationships in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1176–1182, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-153.

Peterman:1989:ASP

[2505] Randall M. Peterman. Application of statistical power analysis to the Oregon Coho salmon (Oncorhynchus kisutch) problem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1183–1187, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-154.

Holtze:1989:LLP

[2506] K. E. Holtze and N. J. Hutchinson. Lethality of low pH and Al to early life stages of six fish species inhabiting PreCambrian Shield waters in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1188–1202, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-155.

Bates:1989:PDN

[2507] S. S. Bates, C. J. Bird, A. S. W. de Freitas, R. Foxall, M. Gilgan, L. A. Hanic, G. R. Johnson, A. W. McCulloch, P. Odense, R. Pocklington, M. A. Quilliam, P. G. Sim, J. C. Smith, D. V. Subba Rao, E. C. D. Todd, J. A. Walter, and J. L. C. Wright. Pennate diatom Nitzschia pungens as the primary source of domoic acid, a toxin in shellfish from Eastern Prince Edward Island, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1203–1215, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-156.

Mathews:1989:SOG

[2508] Stephen B. Mathews and Yukimasa Ishida. Survival, ocean growth, and ocean distribution of differentially timed releases of hatchery Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1216–1226, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-157.

Cooper:1989:STP

[2509] W. J. Cooper, D. R. S. Lean, and J. H. Carey. Spatial and temporal patterns of hydrogen peroxide in lake waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1227–1231, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-158.

Canfield:1989:FIA

[2510] Daniel E. Canfield, Jr., Edward Phlips, and Carlos M. Duarte. Factors influencing the abundance of blue-green algae in Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1232–1237, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-159.

Simard:1989:MAE

[2511] Yvan Simard and David L. Mackas. Mesoscale aggregations of euphausiid sound scattering layers on the Continental Shelf of Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1238–1249, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-160.

Ehlinger:1989:FMS

[2512] Timothy J. Ehlinger. Foraging mode switches in the golden shiner (Notemigonus crysoleucas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (7):1250–1254, July 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-161.

Warwick:1989:MDL

[2513] W. F. Warwick. Morphological deformities in larvae of Procladius Skuse (Diptera: Chironomidae) and their biomonitoring potential. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(7):1255–1270, July 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-162.

Evans:1989:RAS

[2514] Geoffrey T. Evans. Rational approximations to solutions of the VPA equations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1274–1276, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-163.

Horne:1989:EFI

[2515] John K. Horne and Steven E. Campana. Environmental factors influencing the distribution of juvenile groundfish in nearshore habitats of Southwest Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8): 1277–1286, August 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-164.

Xie:1989:PRM

[2516] Liusen Xie and William W. Hsieh. Predicting the return migration routes of the Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1287–1292, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-165.

Bothwell:1989:PLG

[2517] Max L. Bothwell. Phosphorus–limited growth dynamics of lotic periphytic diatom communities: Areal biomass and cellular growth rate responses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1293–1301, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-
166.

Koslow:1989:MNV

[2518] J. Anthony Koslow. Managing nonrandomly varying fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1302–1308, August 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-167.

Dixit:1989:LAR

[2519] Sushil S. Dixit, Aruna S. Dixit, and John P. Smol. Lake acidification recovery can be monitored using chrysophycean microfossils. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1309–1312, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-168.

Charles:1989:BSE

[2520] Anthony T. Charles. Bio-socio-economic fishery models: Labour dynamics and multi-objective management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1313–1322, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-169.

Fox:1989:EPD

[2521] Michael G. Fox. Effect of prey density and prey size on growth and survival of juvenile walleye (Stizostedion vitreum vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1323–1328, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-170.

Ciborowski:1989:FID

[2522] Jan J. H. Ciborowski and Douglas A. Craig. Factors influencing dispersion of larval black flies (Diptera: Simuliidae): Effects of current velocity and food concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8): 1329–1341, August 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-171.

Thorp:1989:FPN

[2523] James H. Thorp, Lawrence D. Goldsmith, John A. Polgreen, and Lynn M. Mayer. Foraging patterns of nesting and nonnesting sunfish (Centrarchidae: Lepomis auritus and L. gibbosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1342–1346, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-172.

Cederholm:1989:FCS

[2524] C. J. Cederholm, D. B. Houston, D. L. Cole, and W. J. Scarlett. Fate of Coho salmon (Oncorhynchus kisutch) carcasses in spawning streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1347–1355, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-173.

Hilborn:1989:MTD

[2525] Ray Hilborn. Models of tag dynamics with exchange between available and unavailable populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (8):1356–1366, August 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-174.

Lleonart:1989:DAC

[2526] Jordi Lleonart and Jordi Salat. A dynamic approach to catch–effort data analysis using a new concept: the inertia of the stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1367–1374, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-175.

Reiter:1989:DBA

[2527] Michael A. Reiter. Development of benthic algal assemblages subjected to differing near-substrate hydrodynamic regimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1375–1382, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-176.

Thedinga:1989:ELS

[2528] John F. Thedinga, Michael L. Murphy, Jonathan Heifetz, K. V. Koski, and Scott W. Johnson. Effects of logging on size and age composition of juvenile Coho salmon (Oncorhynchus kisutch) and density of presmolts in Southeast Alaska streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (8):1383–1391, August 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-177.

Bailey:1989:STV

[2529] Robert C. Bailey and Roger H. Green. Spatial and temporal variation in a population of freshwater mussels in Shell Lake, N.W.T. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1392–1395, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-178.

Holtby:1989:STI

[2530] L. Blair Holtby, Thomas E. McMahon, and J. Charles Scrivener. Stream temperatures and inter-annual variability in the emigration timing of Coho salmon (Oncorhynchus kisutch) smolts and fry and chum salmon (O. Keta) fry from Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1396–1405, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-179.

Swain:1989:DMB

[2531] Douglas P. Swain and L. Blair Holtby. Differences in morphology and behavior between juvenile Coho salmon (Oncorhynchus kisutch) rearing in a lake and in its tributary stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1406–1414, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-180.

Purcell:1989:PFL

[2532] Jennifer E. Purcell. Predation on fish larvae and eggs by the hydromedusa Aequorea victoria at a herring spawning ground in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1415–1427, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-181.

Trites:1989:EJS

[2533] Andrew W. Trites. Estimating the juvenile survival rate of male northern fur seals (Callorhinus ursinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1428–1436, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-182.

Trites:1989:DFP

[2534] Andrew W. Trites and Peter A. Larkin. The decline and fall of the Pribilof fur seal (Callorhinus ursinus): a simulation study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1437–1445, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-183.

Bentzen:1989:MDP

[2535] Paul Bentzen, Gregory C. Brown, and William C. Leggett. Mitochondrial DNA polymorphism, population structure, and life history variation in American shad (Alosa sapidissima). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1446–1454, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-184.

Munkittrick:1989:UWS

[2536] K. R. Munkittrick and D. G. Dixon. Use of white sucker (Catostomus commersoni) populations to assess the health of aquatic ecosystems exposed to low-level contaminant stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1455–1462, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-185.

Anonymous:1989:CRPb

[2537] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(8):1463, August 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-
186.

Cunjak:1989:DME

[2538] R. A. Cunjak, E. M. P. Chadwick, and M. Shears. Downstream movements and estuarine residence by Atlantic salmon parr (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1466–1471, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-187.

Nielsen:1989:SPF

[2539] O. Nielsen, R. K. Kelly, W. R. Lillie, J. W. Clayton, R. S. Fujioka, and B. S. Yoneyama. Some properties of a finite cell line from beluga whale (Delphinapterus leucas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9): 1472–1477, September 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-188.

Lokkeborg:1989:RCG

[2540] Svein Løkkeborg, Åsmund Bjordal, and Anders Fernö. Responses of cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) to baited hooks in the natural environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1478–1483, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-189.

Scrudato:1989:UTM

[2541] Ronald J. Scrudato and William H. McDowell. Upstream transport of mirex by migrating salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1484–1488, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-190.

Plante:1989:PFI

[2542] Céline Plante and John A. Downing. Production of freshwater invertebrate populations in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (9):1489–1498, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f89-191.

Haefner:1989:ELP

[2543] James W. Haefner and Linda C. Abbott. Extrapolation of laboratory pH dose–response data to fluctuating environments: Comparisons with a null model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1499– 1509, September 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-192.

Hanson:1989:SDM

[2544] John Mark Hanson, E. E. Prepas, and William C. Mackay. Size distribution of the macroinvertebrate community in a freshwater lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1510–1519, September 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-193.

Williams:1989:ECD

[2545] D. Dudley Williams and Kathryn A. Moore. Environmental complexity and the drifting behaviour of a running water amphipod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1520–1530, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-194.

Trippel:1989:LTS

[2546] Edward A. Trippel and F. William H. Beamish. Lake trout (Salvelinus namaycush) growth potential predicted from cisco (Coregonus artedii) population structure and conductivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1531–1538, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-195.

Boisclair:1989:APVc

[2547] D. Boisclair and W. C. Leggett. Among-population variability of fish growth: III. Influence of fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1539–1550, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-196.

McMahon:1989:ICC

[2548] Thomas E. McMahon and Gordon F. Hartman. Influence of cover complexity and current velocity on winter habitat use by juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1551–1557, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-197.

Wagemann:1989:CHM

[2549] R. Wagemann. Comparison of heavy metals in two groups of ringed seals (Phoca hispida) from the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1558–1563, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-198.

Jessop:1989:EHS

[2550] B. M. Jessop and W. E. Anderson. Effects of heterogeneity in the spatial and temporal pattern of juvenile alewife (Alosa pseudoharengus) and blueback herring (A. aestivalis) density on estimation of an index of abondance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1564– 1574, September 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-199.

Feltmate:1989:IRT

[2551] Blair W. Feltmate and D. Dudley Williams. Influence of rainbow trout (Oncothynchus mykiss) on density and feeding behaviour of a perlid stonefly. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1575–1580, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-200.

Craig:1989:RCW

[2552] J. F. Craig and J. A. Babaluk. Relationship of condition of walleye (Stizostedion vitreum) and northern pike (Esox lucius) to water clarity, with special reference to Dauphin Lake, Manitoba. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1581–1586, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-201.

Hebert:1989:EGS

[2553] Paul D. N. Hebert, B. W. Muncaster, and G. L. Mackie. Ecological and genetic studies on Dreissena polymorpha (Pallas): a new mollusc in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1587– 1591, September 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-202.

Bernard:1989:UEC

[2554] F. R. Bernard. Uptake and elimination of coliform bacteria by four marine bivalve mollusks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9): 1592–1599, September 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-203.

Smith:1989:MSD

[2555] Barry D. Smith and Glen S. Jamieson. A model for standardizing Dungeness crab (Cancer magister) catch rates among traps which experienced different soak times. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (9):1600–1608, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f89-204.

Smith:1989:EMM

[2556] Barry D. Smith and Glen S. Jamieson. Exploitation and mortality of male Dungeness crabs (Cancer magister) near Tofino, British Colombia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1609–1614, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-205.

Thorrold:1989:AOM

[2557] Simon R. Thorrold and David McB. Williams. Analysis of otolith microstructure to determine growth histories in larval cohorts of a tropical herring (Herklotsichthys castelnaui). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1615–1624, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-206.

Karnofsky:1989:BRL

[2558] Elisa B. Karnofsky and Holly J. Price. Behavioural response of the lobster Homarus americanus to traps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1625–1632, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-207.

Chambers:1989:EAA

[2559] R. Christopher Chambers and William C. Leggett. Event analysis applied to timing in marine fish ontogeny. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1633–1641, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-208.

Helgen:1989:LMV

[2560] J. C. Helgen. Larval mosquitoes as vulnerable prey: Chaoborus predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(9):1642–1650, September 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-209.

Miller:1989:CAL

[2561] Robert J. Miller. Catchability of American lobsters (Homarus americanus) and rock crabs (Cancer irroratus) by traps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1652–1657, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-210.

Molot:1989:FAA

[2562] Lewis A. Molot, P. J. Dillon, and B. D. LaZerte. Factors affecting alkalinity concentrations of streamwater during snowmelt in Central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1658–1666, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-211.

Dixit:1989:RBC

[2563] Sushil S. Dixit, Aruna S. Dixit, and John P. Smol. Relationship between chrysophyte assemblages and environmental variables in seventytwo Sudbury lakes as examined by canonical correspondence analysis (CCA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1667–1676, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-
212.

Murphy:1989:HUJ

[2564] Michael L. Murphy, Jonathan Heifetz, John F. Thedinga, Scott W. Johnson, and K. V. Koski. Habitat utilization by juvenile Pacific salmon (Onchorynchus) in the glacial Taku River, Southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1677–1685, October 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-213.

Mendelssohn:1989:TSD

[2565] Roy Mendelssohn and Philippe Cury. Temporal and spatial dynamics of a coastal pelagic species, Sardinella maderensis off the Ivory Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1686–1697, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-214.

Swamikannu:1989:ESG

[2566] Xavier Swamikannu and Kyle D. Hoagland. Effects of snail grazing on the diversity and structure of a periphyton community in a eutrophic pond. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1698–1704, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-215.

Vainola:1989:MDE

[2567] Risto Väinölä and Sirkka-Liisa Varvio. Molecular divergence and evolutionary relationships in Pontoporeia (Crustacea: Amphipoda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1705–1713, October 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-216.

Bousfield:1989:RMR

[2568] E. L. Bousfield. Revised morphological relationships within the amphipod genera Pontoporeia and Gammaracanthus and the “Glacial Relict” significance of their postglacial distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1714–1725, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-217.

Williams:1989:ENE

[2569] E. M. Williams and F. B. Eddy. Effect of nitrite on the embryonic development of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1726–1729, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-218.

Yamashita:1989:SGC

[2570] S. Yamashita, R. E. Evans, and Toshiaki J. Hara. Specificity of the gustatory chemoreceptors for CO_{2} and H^{+} in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1730–1734, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-

Mazumder:1989:EFP

[2571] Asit Mazumder, W. D. Taylor, D. J. McQueen, and D. R. S. Lean. Effects of fertilization and planktivorous fish on epilimnetic phosphorus and phosphorus sedimentation in large enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1735–1742, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-220.

Carscadden:1989:CMV

[2572] J. E. Carscadden, K. T. Frank, and D. S. Miller. Capelin (Mallotus villosus) spawning on the southeast shoal: Influence of physical factors past and present. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10): 1743–1754, October 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-221.

Krantzberg:1989:AEN

[2573] G. Krantzberg. Accumulation of essential and nonessential metals by chironomid larvae in relation to physical and chemical properties of the elements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1755–1761, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-
222.

Koeller:1989:FEJ

[2574] P. A. Koeller, L. Coates-Markle, and J. D. Neilson. Feeding ecology of juvenile (age-0) silver hake (Merluccius bilinearis) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1762–1768, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-223.

Bunn:1989:SIA

[2575] Stuart E. Bunn, David R. Barton, H. B. Noel Hynes, Geoffrey Power, and Mark A. Pope. Stable isotope analysis of carbon flow in a tundra river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1769–1775, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-

Ware:1989:BBM

[2576] D. M. Ware and R. W. Tanasichuk. Biological basis of maturation and spawning waves in Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1776–1784, October 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-225.

Morin:1989:CME

[2577] Antoine Morin, Christian Back, Jacques Boisvert, and Robert H. Peters. A conceptual model for the estimation of the sensitivity of black fly larvae to Bacillus thuringiensis var. israelensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1785–1792, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-226.

Welch:1989:SDI

[2578] Harold E. Welch and Martin A. Bergmann. Seasonal development of ice algae and its prediction from environmental factors near Resolute, N.W.T., Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1793– 1804, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f89-227.

Poff:1989:ISV

[2579] N. LeRoy Poff and J. V. Ward. Implications of streamflow variability and predictability for lotic community structure: a regional analysis of streamflow patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10): 1805–1818, October 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-228.

Mohr:1989:ESS

[2580] Michael S. Mohr and David G. Hankin. Estimation of size-specific molting probabilities in adult decapod crustaceans based on postmolt indicator data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1819–1830, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f89-

Thompson:1989:DSR

[2581] Keith R. Thompson and Fred H. Page. Detecting synchrony of recruitment using short, autocorrelated time series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(10):1831–1838, October 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-230.

Campbell:1989:DAL

[2582] Alan Campbell. Dispersal of American lobsters, Homarus americanus, tagged off Southern Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1842–1844, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-231.

Cochran:1989:MFP

[2583] Philip A. Cochran and James F. Kitchell. A model of feeding by parasitic lampreys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1845– 1852, November 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-232.

Ward:1989:SBS

[2584] B. R. Ward, P. A. Slaney, A. R. Facchin, and R. W. Land. Size-biased survival in steelhead trout (Oncorhynchus mykiss): Back-calculated lengths from adults’ scales compared to migrating smolts at the Keogh River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1853– 1858, November 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-233.

Boisclair:1989:IAB

[2585] D. Boisclair and W. C. Leggett. The importance of activity in bioenergetics models applied to actively foraging fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1859–1867, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-234.

Fraser:1989:TCC

[2586] Angus J. Fraser. Triacylglycerol content as a condition index for fish, bivalve, and crustacean larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46 (11):1868–1873, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f89-235.

Stock:1989:EBE

[2587] Marsha S. Stock and Amelia K. Ward. Establishment of a bedrock epilithic community in a small stream: Microbial (algal and bacterial) metabolism and physical structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1874–1883, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-236.

Radtke:1989:LFA

[2588] Richard L. Radtke. Larval fish age, growth, and body shrinkage: Information available from otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1884–1894, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-237.

Geraci:1989:HWM

[2589] Joseph R. Geraci, Donald M. Anderson, Ralph J. Timperi, David J. St. Aubin, Gregory A. Early, John H. Prescott, and Charles A. Mayo. Humpback whales (Megaptera novaeangliae) fatally poisoned by dinoflagellate toxin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1895– 1898, November 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-238.

Neilson:1989:MSR

[2590] John D. Neilson and W. R. Bowering. Minimum size regulations and the implications for yield and value in the Canadian Atlantic halibut (Hippoglossus hippoglossus) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1899–1903, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-239.

Rose:1989:IEG

[2591] G. A. Rose and W. C. Leggett. Interactive effects of geophysically-forced sea temperatures and prey abundance on mesoscale coastal distributions of a marine predator, Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1904–1913, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-240.

Buck:1989:MCW

[2592] John D. Buck, L. Louise Shepard, Patricia M. Bubucis, Stephen Spotte, Kate McClave, and Robert A. Cook. Microbiological characteristics of white whale (Delphinapterus leucas) from capture through extended captivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1914–1921, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-241.

Carrick:1989:BSS

[2593] Hunter J. Carrick and Gary L. Fahnenstiel. Biomass, size structure, and composition of phototrophic and heterotrophic nanoflagellate communities in Lakes Huron and Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1922–1928, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-242.

Eilers:1989:CCA

[2594] J. M. Eilers, G. E. Glass, A. K. Pollack, and J. A. Sorensen. Changes in conductivity, alkalinity, calcium, and pH during a 50-year period in selected Northern Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1929–1944, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-243.

Okamoto:1989:GCS

[2595] Hiroaki Okamoto. A genetic comparison of sympatric populations of sand lance (genus Ammodytes) from the region east of Cape Soya, Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1945–1951, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-244.

Dabrowski:1989:AAA

[2596] K. Dabrowski and G. Köck. Absorption of ascorbic acid and ascorbic sulfate and their interaction with minerals in the digestive tract of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1952–1957, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-245.

Post:1989:SDO

[2597] John R. Post and David O. Evans. Size-dependent overwinter mortality of young-of-the-year yellow perch (Perca flavescens): Laboratory, in situ enclosure, and field experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1958–1968, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-246.

Walters:1989:VST

[2598] Carl J. Walters. Value of short-term forecasts of recruitment variation for harvest management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11): 1969–1976, November 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-247.

Sandheinrich:1989:SCE

[2599] Mark B. Sandheinrich and Gary J. Atchison. Sublethal copper effects on bluegill, Lepomis macrochirus, foraging behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1977–1985, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-248.

Saenger:1989:BNS

[2600] R. Alfred Saenger. Bivariate normal swimbladder size allometry models and allometric exponents for 38 mesopelagic swimbladdered fish species commonly found in the North Sargasso Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):1986–2002, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-249.

Lacroix:1989:PJA

[2601] Gilles L. Lacroix. Production of juvenile Atlantic salmon (Salmo salar) in two acidic rivers of Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):2003–2018, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-250.

Parrish:1989:MSS

[2602] R. H. Parrish, R. Serra, and W. S. Grant. The monotypic sardines, Sardina and Sardinops: Their taxonomy, distribution, stock structure, and zoogeography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):2019– 2036, November 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-251.

Anonymous:1989:LCLd

[2603] Anonymous. Letter and comment/lettre et commentaire. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(11):2037–2044, November 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-252.

Noakes:1989:NAG

[2604] D. J. Noakes. A nonparametric approach to generating inseason forecasts of salmon returns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12): 2046–2055, December 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-253.

Nero:1989:CPA

[2605] R. W. Nero and J. J. Magnuson. Characterization of patches along transects using high-resolution 70-kHz integrated acoustic data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2056–2064, December 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-254.

Spotte:1989:PCC

[2606] Stephen Spotte and Gary Anderson. Plasma cortisol changes in seawateradapted mummichogs (Fundulus heteroclitus) exposed to ammonia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2065–2069, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-255.

Kugel:1989:PFP

[2607] B. Kügel and R. H. Peterson. Perivitelline fluid pH of rainbow trout (Oncorhynchus mykiss) eggs in relation to ambient pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2070–2073, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-256.

Ward:1989:CAM

[2608] Robert D. Ward, Neil Billington, and Paul D. N. Hebert. Comparison of allozyme and mitochondrial DNA variation in populations of walleye, Stizostedion vitreum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12): 2074–2084, December 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-257.

Lambert:1989:SDG

[2609] Timothy C. Lambert and Shoukry N. Messieh. Spawning dynamics of Gulf of St. Lawrence herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2085–2094, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-258.

Wahl:1989:CVT

[2610] David H. Wahl and Roy A. Stein. Comparative vulnerability of three esocids to largemouth bass (Micropterus salmoides) predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2095–2103, December 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-259.

Raymond:1989:FRS

[2611] James A. Raymond. Freezing resistance in some northern populations of Pacific herring, Clupea harengus pallasi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2104–2107, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-260.

Wood:1989:ISS

[2612] Chris C. Wood, Dennis T. Rutherford, and Skip McKinnell. Identification of sockeye salmon (Oncorhynehus nerka) stocks in mixed-stock fisheries in British Columbia and Southeast Alaska using biological markers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2108–2120, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-261.

Wood:1989:USD

[2613] Chris C. Wood. Utility of similarity dendrograms in stock composition analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2121–2128, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-262.

Lapointe:1989:TFM

[2614] Michael F. Lapointe, Randall M. Peterman, and Alec D. MacCall. Trends in fishing mortality rate along with errors in natural mortality rate can cause spurious time trends in fish stock abundances estimated by virtual population analysis (VPA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2129–2139, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-263.

Leatherland:1989:THC

[2615] J. F. Leatherland, L. Lin, N. E. Down, and E. M. Donaldson. Thyroid hormone content of eggs and early developmental stages of five Oncorhynchus species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12): 2140–2145, December 1989. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f89-264.

Leatheland:1989:THC

[2616] J. F. Leatheland, L. Lin, N. E. Down, and E. M. Donaldson. Thyroid hormone content of eggs and early developmental stages of three stocks of goitred Coho salmon (Oncorhynchus kisutch) from the Great Lakes of North America, and a comparison with a stock from British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2146–2152, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-265.

McDonald:1989:SSL

[2617] Michael E. McDonald and Anne E. Hershey. Size structure of a lake trout (Salvelinus namaycush) population an Arctic lake: Influence of angling and implications for fish community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2153–2156, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-266.

Ferraro:1989:PCE

[2618] Steven P. Ferraro, Faith A. Cole, Waldemar A. DeBen, and Richard C. Swartz. Power-cost efficiency of eight macrobenthic sampling schemes in Puget Sound, Washington, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2157–2165, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-267.

Kessler:1989:PPTa

[2619] T. A. Kessler and T. R. Parsons. Primary production in a tidally energetic fjord: Evidence of seasonal and inter-annual tidal forcing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2166–2172, December 1989. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f89-268.

Kessler:1989:PPTb

[2620] T. A. Kessler and T. R. Parsons. Primary production in a tidally energetic fjord: Environmental forcing of vertical phytoplankton distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2173–2183, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f89-269.

Parks:1989:WCM

[2621] J. W. Parks, A. Lutz, and J. A. Sutton. Water column methylmercury in the Wabigoon/English River–lake system: Factors controlling concentrations, speciation, and net production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2184–2202, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f89-270.

Hail:1989:PPP

[2622] Donald J. Hail and Timothy J. Ehlinger. Perturbation planktivory, and pelagic community structure: The consequence of winterkill in a small lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2203–2209, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f89-271.

Randall:1989:ESA

[2623] R. G. Randall. Effect of sea-age on the reproductive potential of Atlantic salmon (Salmo salar) in Eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2210–2218, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-272.

Mitchell:1989:NCL

[2624] Edward D. Mitchell. A new cetacean from the Late Eocene La Meseta Formation Seymour Island, Antarctic Peninsula. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 46(12):2219–2235, December 1989. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f89-273.